### v1

# TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models

Abstract
Data collection is often difficult in critical fields such as medicine, physics, and chemistry. As a result, classification methods usually perform poorly with these small datasets, leading to weak predictive performance. Increasing the training set with additional synthetic data, similar to data augmentation in images, is commonly believed to improve downstream classification performance. However, current tabular generative methods that learn either the joint distribution $p({\mathbf{x}},y)$ or the class-conditional distribution $p({\mathbf{x}}\mid y)$ often overfit on small datasets, resulting in poor-quality synthetic data, usually worsening classification performance compared to using real data alone. To solve these challenges, we introduce TabEBM, a novel class-conditional generative method using Energy-Based Models (EBMs). Unlike existing methods that use a shared model to approximate all class-conditional densities, our key innovation is to create distinct EBM generative models for each class, each modelling its class-specific data distribution individually. This approach creates robust energy landscapes, even in ambiguous class distributions. Our experiments show that TabEBM generates synthetic data with higher quality and better statistical fidelity than existing methods. When used for data augmentation, our synthetic data consistently improves the classification performance across diverse datasets of various sizes, especially small ones.
1 Introduction
Figure 1: Evaluation of TabEBM and other state-of-the-art tabular generative methods across six key metrics (larger area indicates better performance). The results demonstrate that TabEBM excels in data augmentation (utility), leading to downstream classification improvements by a larger margin than all other methods.
In scientific fields such as medicine, physics, and chemistry, collecting tabular data is often challenging due to the experimental nature of data acquisition [5, 50, 4, 32, 77, 12]. Due to the small size of such datasets [5, 46], training machine learning models that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], often suffer from poor performance [75, 52, 37]. In the case of vision and language tasks, which focus on image and text data, a standard remedy to address model performance due to data scarcity is employing data augmentation techniques [73, 74, 60, 72] that generate additional synthetic samples from existing data. This increases the diversity of the training set and often leads to improved model generalisation [73]. For instance, techniques such as rotating, flipping, or altering the colour of images [73], or generating paraphrased sentences for text [74], have proven effective for many downstream tasks. However, applying data augmentation to tabular data remains understudied, as tabular datasets are very diverse and lack explicit symmetries [8], such as rotations or translations seen in images. Additionally, feature correlations in tabular data are weaker compared to the strong spatial or semantic relationships in image or text data [8]. Consequently, existing tabular data augmentation methods often yield mixed results and can even degrade model performance [51, 72, 48], hindering their widespread adoption.
Tabular augmentation typically involves training generative models to approximate either the joint distribution $p({\mathbf{x}},y)$ [86, 24] or the class-conditional distribution $p({\mathbf{x}}|y)$ [86, 42, 84, 47, 48]. A key challenge of joint distribution methods is maintaining the original training label distribution, as sampling from such generators can produce label distributions that deviate from the original and even fail to generate data for specific classes (see Appendix B for an example). These issues compromise the effectiveness of data augmentation [51] by undermining the label accuracy and distribution. On the other hand, while class-conditional models learning $p({\mathbf{x}}|y)$ preserve the stratification of the original data, they often employ a shared model to represent all class-conditional densities. This, however, can lead to overfitting, particularly in imbalanced datasets where the model may prioritise more frequent classes [21], ignoring unique features needed for generating label-invariant samples. Additionally, in datasets with limited data, this can lead to mode collapse [69, 71], where the model does not effectively capture the diversity of each class [71], and thus tends to perform poorly in a multi-class setting.
Figure 2: An overview of the proposed TabEBM. We learn distinct class-specific Energy-Based Models (EBMs) $E_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}blue}}}({\mathbf{x}})$ and $E_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{red}}}({\mathbf{x}})$ exclusively on the points of their respective class. Each EBM approximates a class-conditional distribution $p({\mathbf{x}}|y)$. TabEBM allows synthetic data generation by sampling from the estimated distributions for each class $p({\mathbf{x}}|y={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{blue}})$ and $p({\mathbf{x}}|y={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{red}})$.
To address the challenges of class-conditional tabular generation, we introduce TabEBM (Figure 2), a new method for tabular data augmentation utilising Energy-Based Models (EBMs). Our method introduces two innovations: (i) Distinct class-specific models: TabEBM constructs a collection of individual models – one for each class – which, by design, enables learning distinct marginal distributions for the inputs associated with each class. This, in turn, enables performing data augmentation while maintaining the original label distribution. (ii) Generative models: we build novel class-specific generators that produce high-quality synthetic data even from extremely few samples. Specifically, we create a surrogate binary classification task for each class and fit it with a pre-trained tabular in-context classifier. We then convert the binary classifier into an EBM, a generative model, without additional training. Using class-specific EBMs makes the energy landscape more robust to class overlaps, compared to using a single shared EBM to approximate the class-conditional distribution.
Our contributions can be summarised as:
  • Technical: We propose TabEBM, which is the first generative method to create class-specific EBMs, learning the marginal distribution for each class separately.
  • Empirical: We present the first comprehensive analysis of tabular data augmentation across different dataset sizes and use cases beyond predictive performance. Our analysis compares TabEBM with eight leading tabular generative models across various datasets, demonstrating that TabEBM consistently improves data augmentation performance on small datasets, while our generated data demonstrates better statistical fidelity and privacy-preserving properties (Figure 1).
  • Library: We release TabEBM as an open-source library (included as a zip file with this submission and to be made publicly available after publication, details in Section A.5). Our library enables off-the-shelf data generation and data augmentation on any tabular dataset without requiring training.
2 TabEBM
Notation. We address classification problems with $C$ classes, denoted by ${\mathcal{Y}}=\{1,2,\ldots,C\}$. Let $\{({\mathbf{x}}^{(i)},y_{i})\}_{i=1}^{N}$ represent a dataset of $N$ samples, each being a $D$-dimensional vector ${\mathbf{x}}^{(i)}\in{\mathbb{R}}^{D}$, with a corresponding label $y_{i}\in{\mathcal{Y}}$. For each class $c\in{\mathcal{Y}}$, we define ${\mathcal{X}}_{c}=\{{\mathbf{x}}^{(i)}\mid y_{i}=c\}$ as the subset of samples labelled with class $c$. Let $f_{\theta}(\cdot)$ denote a classifier. The expression $f_{\theta}({\mathbf{x}})[y]$ represents the (unnormalised) logit assigned to the class $y$ for the input ${\mathbf{x}}$.
2.1 Preliminaries on Energy-Based Models
An Energy-Based Model (EBM) [43] defines a probability density function $p_{\theta}({\mathbf{x}})$ through an energy function $E({\mathbf{x}})$. Specifically, the model posits that $p({\mathbf{x}})\propto e^{-E({\mathbf{x}})}$, where $E({\mathbf{x}})$ represents the unnormalised negative log-density of the input ${\mathbf{x}}$. In this framework, lower energy values correspond to higher probability densities. This relationship allows EBMs to model distributions by learning to assign lower energy to more probable configurations of ${\mathbf{x}}$ and higher energy to less probable ones.
An important observation is that energy-based models can utilise the same model architectures as standard classification models [29]. Typically, the logits $f_{\theta}({\mathbf{x}})[y]$ from a classification model define a discriminative distribution through the softmax function, expressed as $p_{\theta}(y|{\mathbf{x}})=\mathrm{softmax}(f_{\theta}({\mathbf{x}})[y])$. Intriguingly, these same logits can be reinterpreted to define an energy-based model for the joint distribution $p({\mathbf{x}},y)$. This is achieved by setting the energy function to $E({\mathbf{x}},y)=-f_{\theta}({\mathbf{x}})$. Furthermore, the energy function for the marginal distribution $p({\mathbf{x}})$ is obtained by marginalising over $p({\mathbf{x}},y)$, resulting in $E({\mathbf{x}})=-\text{LogSumExp}_{y^{\prime}}f_{\theta}({\mathbf{x}})[y^{\prime}]$.
Such an energy-based model, trained with EBM-specific protocols on multiple classes, is typically used as a classifier, as demonstrated on several computer vision tasks in [29]. In contrast, in this work our focus is the opposite: We propose employing trained classifiers, one for each specific class, as a generative energy-based model for the class-conditional distributions $p({\mathbf{x}}|y)$. We apply our TabEBM method for generative tasks on tabular data.
2.2 Distinct Class-Specific Energy-Based Models
TabEBM is a class-conditional generative model $p({\mathbf{x}}|y)$ implemented using a set of EBMs, $\{E_{1}({\mathbf{x}}),E_{2}({\mathbf{x}}),\ldots,E_{C}({\mathbf{x}})\}$. Our approach assumes that the class-conditional density $p({\mathbf{x}}|y=c)$ is best modelled using its class-specific data $\mathcal{X}_{c}$. Thus, for each class $c$, we construct a class-specific EBM, $E_{c}({\mathbf{x}})$, using only the data from that class, $\mathcal{X}_{c}$, such that $p({\mathbf{x}}|y=c)\propto\exp(-E_{c}({\mathbf{x}}))$.
Figure 3: The class-specific energy function $E_{c}({\mathbf{x}})$ from the surrogate binary task, where the blue region represents low energy (i.e., high data density). Placing the negative samples in a hypercube distant from the data results in an accurate energy function.
We derive each class-specific EBM $E_{c}({\mathbf{x}})$ by training a classifier on a novel task and reinterpreting its logits. Specifically, for each class $c$, we propose a surrogate binary classification task to determine if a sample belongs to class $c$ by comparing $\mathcal{X}_{c}$ against a set of surrogate negative samples $\mathcal{X}_{c}^{\text{neg}}$, which we show in Figure 3. Specifically, we generate the negative samples at the corners of a hypercube in $R^{D}$. For each dimension $d$, the coordinates of a negative sample are either $\alpha^{\text{neg}}_{\text{dist}}\sigma_{d}$ or $-\alpha^{\text{neg}}_{\text{dist}}\sigma_{d}$, where $\alpha^{\text{neg}}_{\text{dist}}$ is a fixed constant and $\sigma_{d}$ is the standard deviation of dimension $d$. For example, in $R^{3}$, a negative sample might have coordinates $[\alpha^{\text{neg}}_{\text{dist}}\sigma_{1},\alpha^{\text{neg}}_{\text{dist}}\sigma_{2},-\alpha^{\text{neg}}_{\text{dist}}\sigma_{3}]$. Placing the negative samples at the corners of a hypercube ensures they are easily distinguishable from the real data, which is crucial for an accurate energy function (see Section C.1.1). This placement is also robust to variations in the number and distance of the negative samples (see Sections C.1.2 and C.1.3).
We create the combined dataset $\mathcal{D}_{c}$ for the surrogate binary classification task by labelling $\mathcal{X}_{c}$ as 1 and $\mathcal{X}_{c}^{\text{neg}}$ as 0:
$\mathcal{D}_{c}=(\mathcal{X}_{c}\cup\mathcal{X}_{c}^{\text{neg}},\{1\}^{|\mathcal{X}_{c}|}\cup\{0\}^{|\mathcal{X}_{c}^{\text{neg}}|})$ (1)
We then train a binary classifier $f^{c}_{\theta}(\cdot)$ on $\mathcal{D}_{c}$ and use it to construct the class-specific energy $E_{c}({\mathbf{x}})$ for class $c$. To do this, we reinterpret the logits $\{f_{\theta}^{c}({\mathbf{x}})[0],f_{\theta}^{c}({\mathbf{x}})[1]\}$ of the trained binary classifier as components of an approximated joint distribution for the surrogate binary task:
$p_{c}({\mathbf{x}},0)=\frac{\exp(f_{\theta}^{c}({\mathbf{x}})[0])}{Z},\quad p_{c}({\mathbf{x}},1)=\frac{\exp(f_{\theta}^{c}({\mathbf{x}})[1])}{Z}\quad\text{\small{($Z$ is the normalisation constant)}}$) (2)
Next, we derive the approximated distribution $p_{c}({\mathbf{x}})$ by marginalisation:
$p_{c}({\mathbf{x}})$ $=p_{c}({\mathbf{x}},0)+p_{c}({\mathbf{x}},1)$
$=\frac{\exp(f_{\theta}^{c}({\mathbf{x}})[0])+\exp(f_{\theta}^{c}({\mathbf{x}})[1])}{Z}$
$=\frac{\exp\left(\log\left(\exp(f_{\theta}^{c}({\mathbf{x}})[0])+\exp(f_{\theta}^{c}({\mathbf{x}})[1])\right)\right)}{Z}$
$\rightarrow E_{c}({\mathbf{x}})$ $=-\log\left(\exp(f_{\theta}^{c}({\mathbf{x}})[0])+\exp(f_{\theta}^{c}({\mathbf{x}})[1])\right)$ (TabEBM class-specific energy) (3)
For the binary classifier $f^{c}_{\theta}(\cdot)$ in the surrogate binary classification, we use TabPFN [33], a pre-trained tabular in-context model. Note that TabPFN is intended for inference only, with no updates to its parameters (see Section 4 for more details about TabPFN). In this context, “training” the TabPFN classifier is analogous to the K Nearest Neighbour algorithm, which simply performs inference based on a training dataset provided to the model. We apply TabPFN multiple times on separate datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{C}\}$ to obtain multiple classifiers $\{f_{\theta}^{1},f_{\theta}^{2},\ldots,f_{\theta}^{C}\}$. In Section 3.4, we explore why reinterpreting TabPFN’s logits, trained on our surrogate binary tasks, can be useful for estimating an energy function. We emphasise that TabEBM is a general method, capable of using any gradient-based classifier that computes logits (using Equation 3), and is not limited to TabPFN.
Generating data with TabEBM involves two steps. First, we sample a class $c$ from the empirical distribution $c\sim p(y)$. Then, we sample a data point ${\mathbf{x}}$ from the conditional distribution ${\mathbf{x}}\sim p({\mathbf{x}}|y=c)$ approximated by the class-specific energy-based model $E_{c}({\mathbf{x}})$, as outlined in Algorithm 1. We employ Stochastic Gradient Langevin Dynamics (SGLD) [85] to perform this sampling. SGLD is an efficient method for high-dimensional data, combining stochastic gradient descent (SGLD) with Langevin dynamics. The update rule for SGLD at each iteration is:
${\mathbf{x}}_{t+1}={\mathbf{x}}_{t}-\frac{\eta}{2}\nabla E({\mathbf{x}}_{t})+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,\eta\mathbf{I})$ (4)
where a Gaussian noise term $\epsilon_{t}$ introduces randomness into the sampling process, enhancing the exploration of the distribution. In practice, the step size and the noise standard deviations are often chosen separately, resulting in a biased sampler that allows for faster training. Section C.2 further shows that TabEBM is stable to hyperparameters for the sampling process.
In our method, SGLD performs iterative augmentation. We start by sampling close to real data and iteratively adjust these synthetic data points, steering them towards regions of higher probability under the learned energy model. TabEBM enables sampling from any specified class distribution, including the original class distribution, which is crucial for data augmentation.
Algorithm 1 TabEBM sampling from Class-Specific EBM $E_{c}({\mathbf{x}})$
Input: Training data $\mathcal{X}_{c}$ for class $c$, step size $\alpha_{\text{step}}$, noise scale $\alpha_{\text{noise}}$, initial perturbation $\sigma_{\text{start}}$, number of steps $T$
Output: Set of synthetic samples for class $c$
1:Initialise the surrogate binary classification task and train the model
2:Assign new labels to the samples $\mathcal{X}_{c}$ from class $c$, setting them to class 1
3:Generate a set of surrogate negative samples $\mathcal{X}_{c}^{\text{neg}}$ and assign them class 0 labels
4:Train a binary classifier $f_{\theta}^{c}$ on the dataset $\mathcal{D}_{c}=(\mathcal{X}_{c}\cup\mathcal{X}_{c}^{\text{neg}},\{1\}^{|\mathcal{X}_{c}|}\cup\{0\}^{|\mathcal{X}_{c}^{\text{neg}}|})$
5:Synthesise samples using Stochastic Gradient Langevin Dynamics (SGLD)
6:Initialise synthetic data points ${\mathbf{x}}_{0}^{\text{synth}}$ by sampling from $\mathcal{N}(\mathcal{X}_{c},\sigma_{\text{start}}^{2}\mathbf{I})$
7:for each iteration $t=0,1,\dots,T-1$ do
8:    $E_{c}({\mathbf{x}}_{t}^{\text{synth}})=-\log\left(\exp(f_{\theta}^{c}({\mathbf{x}}_{t}^{\text{synth}})[0])+\exp(f_{\theta}^{c}({\mathbf{x}}_{t}^{\text{synth}})[1])\right)$
9:    ${\mathbf{x}}_{t+1}^{\text{synth}}={\mathbf{x}}_{t}^{\text{synth}}-\alpha_{\text{step}}\nabla E_{c}({\mathbf{x}}_{t}^{\text{synth}})+\mathcal{N}(0,\alpha_{\text{noise}}^{2}\mathbf{I})$
10:end for
11:return ${\mathbf{x}}_{T}^{\text{synth}}$ as the generated synthetic data for class $c$
3 Experiments
We evaluate TabEBM by focusing on four research questions:
  • Data Augmentation Improvement (Q1, Section 3.1): Can TabEBM generate synthetic data that improves the accuracy of downstream predictors via data augmentation?
  • Statistical Fidelity (Q2, Section 3.2): Can TabEBM generate synthetic data with high statistical fidelity (i.e., with similar distributions to those of real data)?
  • Privacy Preservation (Q3, Section 3.3): Can TabEBM generate synthetic data that finds a competitive trade-off between downstream performance and privacy preservation?
  • Understanding TabEBM’s energy formulation (Q4, Section 3.4): Why is TabEBM’s class-specific energy effective, and how do the proposed surrogate tasks influence this?
Datasets. We utilise 14 open-source tabular datasets–eight from OpenML [7] and six from UCI [22] – across five domains: Medicine, Chemistry, Engineering, Language and Economics. These diverse datasets contain 7 to 77 features and 698 to 5500 samples across 2 to 26 classes. Five datasets contain both numerical and categorical features, while the remaining are numerical only. To provide comprehensive and fine-grained conclusions on model performance, we further enlarge the evaluation scope by varying the degrees of data availability (i.e., $N_{\text{real}}$), leading up to 33 different test cases for the eight OpenML datasets. Section A.1 provide detailed dataset descriptions.
Benchmark generators. We compare TabEBM against eight existing tabular data generation methods of eight different categories: (i) a standard interpolation method SMOTE [13]; (ii) a Variational Autoencoders (VAE) based method TVAE [86]; (iii) a Generative Adversarial Networks (GAN) method CTGAN [86]; (iv) a normalising flow model Neural Spine Flows (NFLOW) [24]; (v) a diffusion model TabDDPM [42]; (vi) a tree-based method Adversarial Random Forests (ARF) [84]; (vii) a Graph Neural Network (GNN) based method GOGGLE [47]; and (viii) a Prior-Data Fitted Networks (PFN) based method TabPFGen [48]. Furthermore, we also include a “Baseline” model, where no data augmentation is applied (i.e., only real data is used to train downstream predictors). In Section A.6, we detail the settings used for TabEBM and all other generators. We have attached the implementation code for TabEBM and will make it publicly available post-publication.
Downstream predictors. We select six representative downstream predictors, including three standard baselines: Logistic Regression (LR) [16], KNN [27] and MLP [28]; two tree-based methods: Random Forest (RF) [10] and XGBoost [14]; and a PFN method: TabPFN [33].
General experimental setup. For each dataset of $N$ samples, we first split it into stratified train and test sets. We create large test sets to reduce the likelihood that the model’s performance is accidentally inflated due to a small, unrepresentative set of samples [70], and thus the test size is computed via $N_{\text{test}}=\min\left(\frac{N}{2},500\right)$. The full train set approximates the upper bound of the quality of synthetic data, and we call this set “oracle”. We subsample the full train set to simulate different levels of data availability, thus the subset size $N_{\text{real}}$ varies over $\{20,50,100,200,500\}$. We split each subset into stratified training and validation sets with a ratio of 4:1. We provide detailed descriptions of data splitting in Section A.2 and preprocessing in Section A.3. We repeat the splitting ten times, summing up to 10 runs per subset size. The reported results are averaged by default over ten runs on the test sets. When aggregating results across datasets, we use the average distance to the minimum (ADTM) metric via affine renormalisation between the top-performing and worse-performing models [30, 54]. We provide the evaluation results averaged over six downstream predictors for a general conclusion, and the fine-grained numerical results for each predictor are in Appendix C.
Data augmentation setup. Given $N_{\text{real}}$ real samples, we first train generators on the real training data and then generate $N_{\text{syn}}$ synthetic samples. For training the downstream predictors, we expand the real training split by adding the synthetic samples. The real validation data is used for early stopping, and the real test set is used for evaluating the predictor’s performance. The optimal $N_{\text{syn}}$ remains an open problem for tabular data [51, 72, 31]. Prior work [47, 48] mainly use synthetic sets with equivalent sizes to the real sets (i.e., $N_{\text{real}}=N_{\text{syn}}$). However, we observe that $N_{\text{real}}=N_{\text{syn}}$ can lead to highly unstable results, especially on small datasets we investigate. Recent work has used different $N_{\text{syn}}$ for various generators, such as by applying post-processing [31, 72]. In this work, we want to provide a head-to-head comparison of the effect of data augmentation across subsampled datasets of varying size $N_{\text{real}}\in\{20,50,100,200,500\}$. Therefore, we perform data augmentation with a large synthetic set ($N_{\text{syn}}=500$) across all splits, and the synthetic data has the same class distribution as the real training data. We provide an illustrative figure of the data splitting setup in Section A.2.
Table 1: Classification accuracy (%) aggregated over six downstream predictors, comparing data augmentation on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean balanced accuracy of other methods. We bold the highest accuracy for each dataset of different sample size. Our method, TabEBM, consistently outperforms training on real data alone, and achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ Baseline (Real data) SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 28.14±${}_{\pm\text{6.83}}$ N/A 21.18±${}_{\pm\text{1.48}}$ 22.00±${}_{\pm\text{3.43}}$ 21.30±${}_{\pm\text{2.84}}$ 22.12±${}_{\pm\text{5.30}}$ 24.82±${}_{\pm\text{2.88}}$ 22.40±${}_{\pm\text{9.28}}$ 33.25±${}_{\pm\text{5.01}}$ 33.84±${}_{\pm\text{4.92}}$
50 50.72±${}_{\pm\text{10.53}}$ 54.52±${}_{\pm\text{8.59}}$ 39.54±${}_{\pm\text{5.19}}$ 36.32±${}_{\pm\text{7.17}}$ 35.37±${}_{\pm\text{8.00}}$ 35.11±${}_{\pm\text{11.78}}$ 41.99±${}_{\pm\text{5.24}}$ 37.53±${}_{\pm\text{14.72}}$ 54.45±${}_{\pm\text{7.96}}$ 55.91±${}_{\pm\text{6.41}}$
100 67.83±${}_{\pm\text{11.72}}$ 73.25±${}_{\pm\text{7.48}}$ 59.28±${}_{\pm\text{7.20}}$ 57.64±${}_{\pm\text{9.95}}$ 52.57±${}_{\pm\text{9.55}}$ 56.37±${}_{\pm\text{9.64}}$ 57.01±${}_{\pm\text{8.56}}$ 51.69±${}_{\pm\text{16.68}}$ 71.53±${}_{\pm\text{9.87}}$ 73.31±${}_{\pm\text{6.77}}$
200 81.66±${}_{\pm\text{10.18}}$ 85.65±${}_{\pm\text{6.24}}$ 76.42±${}_{\pm\text{7.71}}$ 74.88±${}_{\pm\text{8.20}}$ 72.10±${}_{\pm\text{10.04}}$ 75.86±${}_{\pm\text{9.30}}$ 74.07±${}_{\pm\text{8.74}}$ 73.57±${}_{\pm\text{6.74}}$ 84.95±${}_{\pm\text{7.47}}$ 86.14±${}_{\pm\text{5.50}}$
500 93.49±${}_{\pm\text{5.28}}$ 94.73±${}_{\pm\text{3.67}}$ 92.24±${}_{\pm\text{3.73}}$ 91.48±${}_{\pm\text{4.43}}$ 90.44±${}_{\pm\text{5.54}}$ 90.62±${}_{\pm\text{5.63}}$ 91.79±${}_{\pm\text{4.53}}$ 91.31±${}_{\pm\text{5.20}}$ 94.87±${}_{\pm\text{3.70}}$ 95.18±${}_{\pm\text{3.10}}$
fourier 20 28.30±${}_{\pm\text{12.09}}$ N/A 21.32±${}_{\pm\text{4.06}}$ 18.19±${}_{\pm\text{3.90}}$ 17.30±${}_{\pm\text{3.03}}$ 15.35±${}_{\pm\text{3.26}}$ 21.75±${}_{\pm\text{2.76}}$ 16.70±${}_{\pm\text{2.91}}$ 36.72±${}_{\pm\text{7.30}}$ 37.13±${}_{\pm\text{6.01}}$
50 53.69±${}_{\pm\text{8.04}}$ 55.51±${}_{\pm\text{7.43}}$ 37.96±${}_{\pm\text{4.48}}$ 35.09±${}_{\pm\text{7.46}}$ 31.94±${}_{\pm\text{8.99}}$ 35.99±${}_{\pm\text{13.06}}$ 40.32±${}_{\pm\text{6.70}}$ 33.56±${}_{\pm\text{14.02}}$ 55.11±${}_{\pm\text{10.66}}$ 56.57±${}_{\pm\text{7.12}}$
100 63.70±${}_{\pm\text{6.76}}$ 64.10±${}_{\pm\text{6.89}}$ 50.46±${}_{\pm\text{8.61}}$ 49.26±${}_{\pm\text{9.15}}$ 44.58±${}_{\pm\text{8.40}}$ 52.79±${}_{\pm\text{10.04}}$ 51.13±${}_{\pm\text{6.35}}$ 41.93±${}_{\pm\text{15.60}}$ 63.86±${}_{\pm\text{7.76}}$ 65.21±${}_{\pm\text{6.42}}$
200 70.99±${}_{\pm\text{4.88}}$ 71.43±${}_{\pm\text{4.47}}$ 62.17±${}_{\pm\text{7.29}}$ 62.92±${}_{\pm\text{7.87}}$ 59.15±${}_{\pm\text{8.33}}$ 68.05±${}_{\pm\text{6.91}}$ 62.53±${}_{\pm\text{6.97}}$ 56.44±${}_{\pm\text{10.13}}$ 71.81±${}_{\pm\text{5.35}}$ 72.36±${}_{\pm\text{3.77}}$
500 77.72±${}_{\pm\text{2.36}}$ 77.51±${}_{\pm\text{2.60}}$ 73.29±${}_{\pm\text{4.97}}$ 74.61±${}_{\pm\text{4.89}}$ 71.74±${}_{\pm\text{6.54}}$ 77.04±${}_{\pm\text{3.64}}$ 74.31±${}_{\pm\text{4.40}}$ 70.61±${}_{\pm\text{6.01}}$ 77.15±${}_{\pm\text{2.57}}$ ${}_{\pm\text{2.87}}$
biodeg 20 66.20±${}_{\pm\text{4.26}}$ 68.59±${}_{\pm\text{1.17}}$ 66.77±${}_{\pm\text{2.64}}$ 58.03±${}_{\pm\text{2.47}}$ 59.37±${}_{\pm\text{1.74}}$ 52.72±${}_{\pm\text{2.38}}$ 61.17±${}_{\pm\text{2.00}}$ 61.39±${}_{\pm\text{6.39}}$ 68.99±${}_{\pm\text{2.54}}$ 69.79±${}_{\pm\text{2.15}}$
50 72.66±${}_{\pm\text{3.98}}$ 72.80±${}_{\pm\text{3.08}}$ 71.31±${}_{\pm\text{2.71}}$ 67.99±${}_{\pm\text{3.63}}$ 62.40±${}_{\pm\text{4.28}}$ 60.72±${}_{\pm\text{10.11}}$ 71.62±${}_{\pm\text{2.43}}$ 66.68±${}_{\pm\text{6.00}}$ 73.29±${}_{\pm\text{3.53}}$ 73.78±${}_{\pm\text{3.42}}$
100 76.69±${}_{\pm\text{2.70}}$ 76.31±${}_{\pm\text{2.42}}$ 75.38±${}_{\pm\text{2.06}}$ 74.82±${}_{\pm\text{2.89}}$ 69.50±${}_{\pm\text{4.59}}$ 68.28±${}_{\pm\text{9.54}}$ 74.42±${}_{\pm\text{2.38}}$ 71.68±${}_{\pm\text{3.72}}$ 76.22±${}_{\pm\text{2.31}}$ 76.45±${}_{\pm\text{3.08}}$
200 80.01±${}_{\pm\text{2.66}}$ 79.67±${}_{\pm\text{2.56}}$ 78.11±${}_{\pm\text{2.68}}$ ${}_{\pm\text{1.78}}$ 75.05±${}_{\pm\text{4.68}}$ 74.43±${}_{\pm\text{8.09}}$ 77.97±${}_{\pm\text{2.32}}$ 77.13±${}_{\pm\text{3.01}}$ 79.76±${}_{\pm\text{2.63}}$ 80.11±${}_{\pm\text{2.33}}$
500 82.63±${}_{\pm\text{2.43}}$ 82.85±${}_{\pm\text{1.93}}$ 82.13±${}_{\pm\text{1.94}}$ 82.42±${}_{\pm\text{1.58}}$ 81.11±${}_{\pm\text{3.23}}$ 79.19±${}_{\pm\text{6.60}}$ 81.92±${}_{\pm\text{2.28}}$ 81.24±${}_{\pm\text{2.30}}$ 82.35±${}_{\pm\text{2.21}}$ 82.29±${}_{\pm\text{2.15}}$
steel 20 57.51±${}_{\pm\text{4.58}}$ 58.32±${}_{\pm\text{3.27}}$ 57.99±${}_{\pm\text{3.06}}$ 56.61±${}_{\pm\text{1.70}}$ 53.89±${}_{\pm\text{1.73}}$ 55.74±${}_{\pm\text{6.02}}$ 54.24±${}_{\pm\text{2.08}}$ 53.04±${}_{\pm\text{2.36}}$ 63.21±${}_{\pm\text{5.86}}$ 63.27±${}_{\pm\text{5.45}}$
50 75.06±${}_{\pm\text{10.43}}$ 65.63±${}_{\pm\text{4.00}}$ 64.18±${}_{\pm\text{3.95}}$ 63.70±${}_{\pm\text{6.10}}$ 58.90±${}_{\pm\text{6.39}}$ 65.85±${}_{\pm\text{14.84}}$ 61.72±${}_{\pm\text{3.39}}$ 56.72±${}_{\pm\text{3.47}}$ 78.67±${}_{\pm\text{11.79}}$ 80.50±${}_{\pm\text{8.67}}$
100 86.87±${}_{\pm\text{12.49}}$ 74.61±${}_{\pm\text{5.99}}$ 70.12±${}_{\pm\text{5.76}}$ 69.89±${}_{\pm\text{5.58}}$ 65.67±${}_{\pm\text{9.10}}$ 76.01±${}_{\pm\text{17.54}}$ 67.33±${}_{\pm\text{5.15}}$ 60.56±${}_{\pm\text{5.37}}$ ${}_{\pm\text{9.50}}$ 92.71±${}_{\pm\text{7.57}}$
200 92.90±${}_{\pm\text{9.14}}$ 81.97±${}_{\pm\text{4.12}}$ 78.73±${}_{\pm\text{5.06}}$ 78.36±${}_{\pm\text{6.98}}$ ${}_{\pm\text{9.57}}$ 85.45±${}_{\pm\text{15.03}}$ 78.65±${}_{\pm\text{6.70}}$ 68.20±${}_{\pm\text{5.30}}$ 95.56±${}_{\pm\text{5.85}}$ 96.29±${}_{\pm\text{4.64}}$
500 97.52±${}_{\pm\text{3.76}}$ 92.44±${}_{\pm\text{4.46}}$ 92.47±${}_{\pm\text{3.66}}$ 92.42±${}_{\pm\text{4.76}}$ 88.20±${}_{\pm\text{8.36}}$ 96.34±${}_{\pm\text{4.67}}$ 90.41±${}_{\pm\text{5.35}}$ 84.23±${}_{\pm\text{10.90}}$ 98.14±${}_{\pm\text{2.67}}$ 98.47±${}_{\pm\text{2.15}}$
stock 20 78.75±${}_{\pm\text{4.39}}$ 82.18±${}_{\pm\text{2.15}}$ 74.11±${}_{\pm\text{3.71}}$ 64.25±${}_{\pm\text{6.29}}$ 72.64±${}_{\pm\text{2.01}}$ 78.61±${}_{\pm\text{3.57}}$ 69.54±${}_{\pm\text{1.65}}$ 76.35±${}_{\pm\text{5.08}}$ 82.42±${}_{\pm\text{2.17}}$ 83.49±${}_{\pm\text{1.60}}$
50 86.10±${}_{\pm\text{3.62}}$ 87.82±${}_{\pm\text{3.41}}$ 82.81±${}_{\pm\text{3.51}}$ 79.63±${}_{\pm\text{3.93}}$ 80.14±${}_{\pm\text{3.90}}$ 86.72±${}_{\pm\text{4.29}}$ 82.48±${}_{\pm\text{2.95}}$ 83.36±${}_{\pm\text{5.23}}$ 88.14±${}_{\pm\text{3.01}}$ 88.44±${}_{\pm\text{3.14}}$
100 89.07±${}_{\pm\text{3.71}}$ 89.99±${}_{\pm\text{3.22}}$ 87.55±${}_{\pm\text{4.25}}$ 86.44±${}_{\pm\text{4.40}}$ 84.64±${}_{\pm\text{4.79}}$ 89.40±${}_{\pm\text{4.26}}$ 87.32±${}_{\pm\text{4.42}}$ 87.44±${}_{\pm\text{5.46}}$ 90.27±${}_{\pm\text{3.33}}$ 90.36±${}_{\pm\text{3.51}}$
200 90.85±${}_{\pm\text{4.39}}$ 91.75±${}_{\pm\text{3.73}}$ 90.12±${}_{\pm\text{5.44}}$ 89.44±${}_{\pm\text{4.94}}$ 88.47±${}_{\pm\text{6.06}}$ 90.76±${}_{\pm\text{5.27}}$ 89.59±${}_{\pm\text{5.37}}$ 89.62±${}_{\pm\text{6.29}}$ 91.56±${}_{\pm\text{3.91}}$ 91.71±${}_{\pm\text{3.77}}$
More than 10 classes energy 50 17.77±${}_{\pm\text{6.15}}$ N/A 12.30±${}_{\pm\text{2.59}}$ 12.11±${}_{\pm\text{3.16}}$ 10.14±${}_{\pm\text{2.87}}$ 10.55±${}_{\pm\text{2.44}}$ 11.99±${}_{\pm\text{2.27}}$ 15.46±${}_{\pm\text{3.54}}$ N/A 23.98±${}_{\pm\text{2.73}}$
100 25.94±${}_{\pm\text{4.86}}$ N/A 17.78±${}_{\pm\text{4.73}}$ 18.60±${}_{\pm\text{6.09}}$ 18.56±${}_{\pm\text{6.39}}$ 18.84±${}_{\pm\text{6.23}}$ 19.91±${}_{\pm\text{5.21}}$ 17.65±${}_{\pm\text{5.88}}$ N/A 31.24±${}_{\pm\text{5.53}}$
200 35.99±${}_{\pm\text{8.92}}$ N/A 27.65±${}_{\pm\text{11.12}}$ 27.77±${}_{\pm\text{10.55}}$ 28.37±${}_{\pm\text{10.82}}$ 29.50±${}_{\pm\text{10.33}}$ 29.57±${}_{\pm\text{9.18}}$ 28.95±${}_{\pm\text{10.40}}$ N/A 41.28±${}_{\pm\text{7.66}}$
collins 100 11.44±${}_{\pm\text{2.77}}$ N/A 8.38±${}_{\pm\text{1.52}}$ 8.11±${}_{\pm\text{1.00}}$ 7.93±${}_{\pm\text{1.40}}$ ${}_{\pm\text{2.16}}$ 7.53±${}_{\pm\text{1.10}}$ 9.21±${}_{\pm\text{2.35}}$ N/A 13.07±${}_{\pm\text{2.51}}$
200 15.74±${}_{\pm\text{3.73}}$ 17.45±${}_{\pm\text{3.46}}$ 12.08±${}_{\pm\text{3.03}}$ 11.37±${}_{\pm\text{1.20}}$ 10.74±${}_{\pm\text{1.72}}$ 15.39±${}_{\pm\text{3.37}}$ 10.71±${}_{\pm\text{1.37}}$ 14.30±${}_{\pm\text{3.42}}$ N/A 17.03±${}_{\pm\text{3.20}}$
texture 50 72.40±${}_{\pm\text{13.07}}$ 76.40±${}_{\pm\text{10.50}}$ 55.32±${}_{\pm\text{6.20}}$ 54.80±${}_{\pm\text{12.97}}$ 55.39±${}_{\pm\text{10.65}}$ 62.27±${}_{\pm\text{8.01}}$ 55.65±${}_{\pm\text{10.58}}$ 62.94±${}_{\pm\text{12.06}}$ N/A 78.90±${}_{\pm\text{7.96}}$
100 82.42±${}_{\pm\text{10.38}}$ 84.35±${}_{\pm\text{9.67}}$ 66.00±${}_{\pm\text{7.21}}$ 69.49±${}_{\pm\text{10.93}}$ 71.78±${}_{\pm\text{9.06}}$ 76.25±${}_{\pm\text{7.40}}$ 70.93±${}_{\pm\text{9.71}}$ 76.34±${}_{\pm\text{9.55}}$ N/A 86.01±${}_{\pm\text{7.36}}$
200 87.54±${}_{\pm\text{7.62}}$ 89.29±${}_{\pm\text{6.20}}$ 78.37±${}_{\pm\text{6.03}}$ 82.44±${}_{\pm\text{7.15}}$ 81.94±${}_{\pm\text{6.30}}$ 84.67±${}_{\pm\text{4.79}}$ 83.29±${}_{\pm\text{6.32}}$ 82.53±${}_{\pm\text{7.99}}$ N/A 89.77±${}_{\pm\text{5.77}}$
500 92.96±${}_{\pm\text{4.07}}$ 93.69±${}_{\pm\text{3.83}}$ 90.09±${}_{\pm\text{3.56}}$ 91.48±${}_{\pm\text{3.50}}$ 90.50±${}_{\pm\text{2.71}}$ 91.53±${}_{\pm\text{3.29}}$ 91.76±${}_{\pm\text{3.98}}$ 91.24±${}_{\pm\text{3.56}}$ N/A 93.76±${}_{\pm\text{3.64}}$
Average rank 3.30±${}_{\pm\text{1.02}}$ 3.03±${}_{\pm\text{1.25}}$ 6.79±${}_{\pm\text{1.80}}$ 7.48±${}_{\pm\text{1.50}}$ 8.94±${}_{\pm\text{0.70}}$ 6.39±${}_{\pm\text{2.41}}$ 6.94±${}_{\pm\text{1.50}}$ 7.76±${}_{\pm\text{2.03}}$ 3.15±${}_{\pm\text{1.27}}$ 1.21±${}_{\pm\text{0.74}}$
Figure 4: Mean normalised balanced accuracy improvement (%) across different sample sizes (Left) and across datasets with varying numbers of classes (Right). Because TabPFGen is not applicable for datasets with more than ten classes, we plot short bars at zeros for visual clearance. Positive values indicate that the generator improves downstream classification performance. TabEBM generally outperforms benchmark generators across varying sample sizes and number of classes.
3.1 Data Augmentation Improvement (Q1)
We evaluate the effect of using synthetic data for data augmentation by comparing the balanced accuracy of downstream predictors before and after augmentation. Typically, higher classification accuracy (i.e., $\text{ACC}_{\text{Generator}})$ and accuracy improvements (i.e., $\text{ACC}_{\text{Generator}}-\text{ACC}_{\text{Baseline}}>0$) demonstrate the effectiveness of the synthetic data for data augmentation.
TabEBM effectively improves downstream performance across sample sizes, especially for very low-sample-size regimes. Table 1 and Figure 4 (Left) show that TabEBM exhibits competitive performance in data augmentation, generally achieving the highest downstream accuracy and average rank across most datasets and sample sizes. Notably, TabEBM is the only generator that consistently improves performance across sample sizes. A key observation is that most modern benchmark generators even underperform the Baseline, indicating poor approximated distributions in the low-sample-size regime. Moreover, TabEBM achieves the largest overall performance improvement on six additional UCI datasets, further supporting its effectiveness (see Section C.5.2 for detailed results).
Furthermore, TabEBM is the most widely applicable method among the top three competitive generators on the considered datasets: (i) SMOTE requires at least two samples per class for interpolation, and thus it is not applicable for some datasets, such as the “protein” dataset ($N_{\text{real}}=20$); (ii) TabPFGen cannot scale up to more than ten classes, such as the “collins” dataset. In addition, TabEBM can stabilise downstream performance, especially when real data is very scarce ($N_{\text{real}}=20$): TabEBM leads to smaller standard deviations than Baseline on seven out of eight datasets.
TabEBM effectively improves downstream performance across the number of classes, especially for more than ten classes. Figure 4 (Right) shows that TabEBM consistently outperforms the Baseline with notable improvements, particularly in datasets with more than ten classes. In contrast, an increased number of classes tends to cause a performance degradation in the benchmark generators.
Figure 5: Mean normalised balanced accuracy improvement (%) on imbalanced datasets. TabEBM consistently outperforms the Baseline and other generators across different levels of data imbalance.
Figure 6: Median data augmentation time vs. mean normalised balanced accuracy. TabEBM achieves higher downstream accuracy while typically operating 3-30 times faster than most other methods.
TabEBM is robust on imbalanced datasets. For the three binary OpenML datasets (i.e., “biodeg”, “steel” and “stock”), we adjust the class distribution in the training data to vary the class imbalance, while keeping the test data fixed. Figure 6 shows that TabEBM consistently outperforms Baseline, while the other generators exhibit performance degradation as data imbalance increases.
TabEBM is computationally efficient. Figure 6 shows the trade-off between accuracy and the time needed for generating stratified synthetic data (for data augmentation). We measure the total duration of (i) training the model and (ii) generating 500 synthetic samples. The results show that TabEBM is practical, as it achieves higher downstream accuracy with lower time costs.
3.2 Statistical Fidelity (Q2)
We evaluate the fidelity of synthetic data by focusing on two aspects of synthetic data: similarity to real train data and real test data. We evaluate the similarity between real and synthetic data via (i) average inverse of the Kullback–Leibler Divergence (inverse KL) [17], (ii) p-value of Kolmogorov–Smirnov test (KS test) [39] and (iii) p-value of Chi-squared test ($\chi^{2}$ test) [55]. The full results, including $\chi^{2}$ test and the similarity between real test data and synthetic data, are in Section C.6. For all three metrics, a bigger value denotes that synthetic data is more likely to have the same distribution as real data.
In Figure 7 (a1&a2), TabEBM consistently achieves the highest accuracy and distribution similarity between real train data and synthetic data, showing that TabEBM learns the distributions of real train data better than benchmark generators. In Section C.6, we further show that TabEBM remains the most competitive method in similarity between real test data and synthetic data. This indicates that TabEBM can extrapolate beyond real train data and thus generate synthetic data from a more general distribution that aligns with both train and test data. This extrapolation ability also explains why TabEBM can outperform Baseline via data augmentation (Section 3.1).
3.3 Privacy Preservation (Q3)
More broadly, data privacy is a critical concern for organisations and governments handling sensitive data [76]. Privacy-preserving synthetic data also allows researchers and practitioners to bypass ethical and logistical issues while enabling model training and testing [38]. We further explore the use of TabEBM-generated data for data sharing, where only synthetic data is accessible for downstream users [76, 90, 87, 23, 42]. In this case, downstream models are trained exclusively on synthetic data.
Specifically, we evaluate synthetic data via three metrics: (i) balanced accuracy of downstream predictors trained with only synthetic data (i.e., train-on-synthetic, test-on-real [86, 42, 88]); (ii) median Distance to Closest Record (DCR) [89], where a greater DCR denotes synthetic data is less likely to be copied from real data; and (iii) ${\delta}$-presense [62], where a smaller value denotes a lower re-identification risk for real data from synthetic data. Full numerical results are in Section C.7.
Figure 7 (b1&b2) shows that TabEBM consistently finds a better trade-off between accuracy and privacy preservation. Notably, the “train-on-synthetic, test-on-real” scenario poses a greater challenge for generators in achieving high accuracy because real data is inaccessible for model training and data augmentation. Despite this difficulty, TabEBM is the only generator that surpasses the overall performance of training on real data (i.e., Baseline). The relatively high DCR for TabEBM indicates that it can extrapolate beyond real train data, aligning with the finding that TabEBM’s synthetic data is statistically similar to real test data (Section 3.2). These results further suggest that TabEBM learns the general distribution of real data, and TabEBM can generate high-quality synthetic data suitable for various purposes, including privacy preservation.
Figure 7: (a1&a2): Median inverse KL and KS test vs. mean normalised balanced accuracy improvement (%) between real train data and synthetic data. (b1&b2): Median DCR and $\delta$-presence vs. mean normalised balanced accuracy change (%) between real train data and synthetic data. Note that “accuracy improvement” is for data augmentation, and “accuracy change” is for data sharing. The figures with with standard deviations are in Section C.4. TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
3.4 Why is TabEBM effective for estimating Energy-Based Models? (Q4)
Figure 8: (Left) Logit distribution of TabPFN trained on our surrogate binary tasks at increasing distances from the real data (on “steel”). (Right) The corresponding unnormalised density approximated by TabEBM. TabEBM assigns higher density closer to the real data.
Having established that TabEBM excels in data augmentation, we explore why classifier logits can be useful when reinterpreted as a class-conditional energy function. Figure 8 shows the logit distribution of TabPFN trained on surrogate binary tasks and the corresponding energy function of TabEBM (with TabPFN as the binary classifier) as the Euclidean distance from the real data increases.
We found it essential to place the negative samples far from the real data, since TabPFN, which pre-trained to approximate Bayesian inference [33], has its confidence influenced by the distance from the training data [53]. Figure 8 (left) shows that TabPFN outputs high logit values near the real data. As the distance from the real data increases, the logit $f({\mathbf{x}})[1]$ decreases smoothly until the two logits become similar, making the classifier uncertain (because the class probabilities become equal). The right figure shows that TabEBM’s inferred density drops significantly as the maximum logit decreases, because $p_{c}({\mathbf{x}})\propto(\exp(f({\mathbf{x}})[0])+\exp(f({\mathbf{x}})[1]))$ from Equation 3. Since SGLD sampling performs gradient ascent on the density, the TabEBM-generated samples will be close to the real data. These findings are consistent across datasets (see Section C.3), where TabPFN’s logits remain positive, with similar ranges and a relatively constant sum as distance increases, warranting further investigation. Overall, TabPFN’s distance-based uncertainty is useful for inferring accurate energy functions within our TabEBM framework. Since TabEBM can be paired with any other gradient-based classifier that produces logits, we leave these extensions for future work.
4 Discussion & Related Work
Section 3 showed that TabEBM efficiently generates high-fidelity data that can effectively improve the downstream performance via data augmentation. In Table 2, we further provide a summary of tabular data generative models analysed from three important perspectives: (i) Training: the type of distribution that the generators learn (crucial for preserving the original training label distribution), and the training costs associated with learning; (ii) Generation: do the generators employ class-specific models (reflecting their capability to capture unique features essential for label-invariant generation), and do models support stratified generation (crucial for effective data augmentation); (iii) Practicability: the scalability of the generators with respect to the number of classes (a common requirement in real-world multi-class tasks), and consistent downstream performance improvement across different class sizes.
Generative Models for Tabular Data. The common paradigm for tabular data generation is to adapt Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) [86, 63]. For instance, TableGAN employs a convolutional neural network to optimise the label quality [63], and TVAE is introduced in [86] as a variant of VAE for tabular data. However, these methods learn the joint distribution and thus cannot preserve the stratification of the original data (Appendix B). CTGAN [86] refines the generation to be class-conditional. The recent ARF [84] is an adversarial variant of random forest for density estimation, and GOGGLE [47] enhances VAE by learning relational structure with a Graph Neural Network (GNN). Some recent work focuses on generation with denoising diffusion models [42, 88, 40, 44]. For instance, TabDDPM [42] demonstrates that diffusion models can approximate typical distributions of tabular data. Although these class-conditional models can preserve the label distribution, they struggle to outperform Baseline and standard SMOTE in data augmentation [72, 48].
We attribute the performance degradation in current class-conditional models to their reliance on a single shared model to approximate all class-conditional densities. For instance, another promising generative approach uses pre-trained models like Prior-Data Fitted Networks (PFNs), and the recent TabPFGen [48] adapts such models into one shared class-conditional generator. However, TabPFGen’s shared generator can lead to inaccurate density estimates, particularly in high-noise and class-imbalance situations (see examples in Appendix B). As noise increases, TabPFGen’s inferred densities fluctuate significantly and diverge from the true data distributions. In contrast, TabEBM uses class-specific EBMs to model each class’s marginal distributions, and the results in Appendix B reveal that our design choice reduces the impact of noise and data imbalance. TabEBM focuses on approximating and generating for one class at a time, remaining unaffected by noise from other classes. Overall, our results demonstrate that TabEBM consistently improves performance across different datasets and sample sizes, outperforming TabPFGen. Moreover, TabPFGen is limited in usability (e.g., it supports only up to ten classes), while TabEBM scales to any number of classes.
In a broader context, some recent work attempts to adapt Large Language Models (LLMs) for tabular data generation [25, 72, 9]. However, data contamination is an inherent issue with such LLM-based models [19, 36, 18, 49]. As the pre-training data is not typically open-source, these models can have unfair advantages in downstream tasks (i.e., the full real dataset, including the real test data, may have been used for pre-training). Therefore, in this paper, we focus on models without support from LLMs, thus avoiding potential biases from data contamination.
Table 2: Comparison of the properties between TabEBM and prior tabular generative methods. TabEBM has novel design rationales of training-free class-specific models, and TabEBM is highly practicable with wide applicability and consistent accuracy improvement.
Methods Category Training Generation Practicability
Learned distribution Training-free Class-specific models Stratified generation Unlimited classes ACC improve ($\leq$ 10 classes) ACC improve ($>$ 10 classes)
SMOTE [13] Interpolation N/A N/A
TVAE [86] VAE $p({\mathbf{x}},y)$
CTGAN [86] GAN $p({\mathbf{x}}\mid y)$
NFLOW [24] Normal. Flows $p({\mathbf{x}},y)$
TabDDPM [42] Diffusion $p({\mathbf{x}}\mid y)$
ARF [84] Random Forest $p({\mathbf{x}}\mid y)$
GOGGLE [47] VAE $p({\mathbf{x}}\mid y)$
TabPFGen [48] PFN $p({\mathbf{x}}\mid y)$
TabEBM (Ours) PFN $p({\mathbf{x}}\mid y)$
Data Augmentation (DA) for Tabular Data. DA is an omnipresent technique in computer vision and natural language processing [83, 74, 73, 60, 26, 2]. However, DA for tabular data remains underexplored, and existing methods often perform poorly in real-world tasks, sometimes even reduce performance [51]. Recent studies show that using the same transformations across all classes leads to varied performance impacts [3, 41], indicating that data augmentation effects are class-specific and suggesting that different classes may require distinct augmentations. Given the lack of symmetries in tabular data, we believe this class-dependent effect is even more pronounced. Therefore, we propose TabEBM as a class-specific generative model to produce tailored augmentations for each class.
Prior-fitted Networks (PFNs) for Tabular Data. Recent work proposes to approximate the posterior predictive distribution with transformers [59, 33, 61, 80, 20]. PFNs can be adapted for various purposes by pre-training the transformer with corresponding “prior data”, and then it can make in-context predictions with unseen downstream data. For instance, TabPFN is a variant that is pre-trained on a prior designed for tabular data [33]. We note that prior data is different to synthetic data in this paper. Specifically, prior data refers to manually crafted fake data (e.g., $y=2x$) with no real-world semantics. In contrast, synthetic data from generators is expected to have the same semantics as real data. Inspired by TabPFN’s success in small-size classification tasks, TabEBM converts TabPFN into multiple EBMs that learn the marginal distribution for each class. The training-free nature of TabPFN enables TabEBM to generate high-quality tabular data without introducing extra training costs. Additionally, our class-specific design lets TabEBM surpass TabPFN’s limits and scale to more than ten classes.
Limitations and Future Work. TabEBM is a general method that relies on an underlying binary classifier, and as such, its strengths and weaknesses are directly tied to this classifier. We used TabPFN because it is a well-established open-source pre-trained model for tabular data. Therefore, TabEBM inherits some of TabPFN’s limitations, particularly in scaling to a larger number of features. TabEBM can handle datasets with over 1000 samples, overcoming TabPFN’s limitation, as it processes one class at a time. In Section C.5.3, we show that TabEBM outperforms other generators on larger datasets, though the performance gains decrease as the sample size increases. Although TabPFN has a 10-class limit, TabEBM can handle an unlimited number of classes by training TabPFN on binary surrogate tasks instead of the original multi-class problem. Our results also show that TabEBM can handle categorical data by encoding categorical features with leave-one-out target statistics. We stress that TabEBM is compatible with any classifier that can be adapted into EBMs, as described in Section 2. As foundational models for tabular data evolve [82], new models capable of handling more features and samples are expected. Integrating them into TabEBM will enhance its ability to manage high-dimensional datasets, increasing its versatility and utility.
5 Conclusion
We introduce TabEBM, the first tabular data augmentation method that creates class-specific EBM generators, learning the marginal distribution for each class separately. We also provide the first comprehensive analysis of tabular data augmentation across various dataset sizes. Our results demonstrate that TabEBM improves downstream performance through data augmentation on real-world datasets, outperforming other benchmark generators. The statistical evaluation confirms that TabEBM generates high-fidelity synthetic data, particularly for small datasets. We release our method as an open-source library, allowing users to generate data immediately without additional training.
Broader Impact Statement
This paper introduces a novel data augmentation approach, TabEBM, that aims to advance the field of machine learning by addressing challenges in the low-sample-size regime. Furthermore, TabEBM offers an elegant solution to learning the unique features in generating samples for each class, leading to high-fidelity synthetic data that can effectively improve downstream performance. These characteristics can be particularly useful in data-scarce domains like healthcare (e.g., pre-clinical drug evaluation in early-stage clinical trials [6, 58]). Moreover, we also demonstrate that TabEBM is readily applicable for privacy-preserving data sharing in high-stake tasks [90, 78].
TabEBM’s impact further extends to enabling broader machine learning applications in data-scarce domains, for instance, facilitating data analysis in clinical scenarios with limited access to data collection techniques. Improving the performance of machine learning models in such applications can further foster the uptake of more sophisticated ML approaches and, ultimately, help improve the quality of healthcare [1, 15, 57]. TabEBM can further facilitate research and enhance machine learning accessibility in various communities across societal and scientific domains. To this end, our work has only been evaluated in a strictly research setting. Further applications of our work in scenarios with sensitive data bear some risks. As TabEBM is a generative model, training models with the resulting generated samples can bias the downstream model. Therefore, this risk, together with other data privacy risks during downstream deployment, must be carefully managed.
Appendix: TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models
[…]
Appendix A Reproducibility
A.1 Datasets
All eight datasets are publicly available on OpenML [7], and their details are listed in Table 3. To ensure consistent stratified data-splitting across all datasets, we remove classes with fewer than 10 samples. For example, the original “energy” dataset contains 14 classes with fewer than 10 samples, which could result in a validation set lacking samples from these classes, leading to unstratified data splitting.
Table 3: Details of the eight real-world tabular datasets.
Dataset OpenML ID Not evaluated in TabPFN [33] # Samples ($N$) # Features ($D$) # Classes $N/D$ Samples per class (Min) # Samples per class (Max)
At most 10 classes
protein 40966 1,080 77 8 14.03 105 150
fourier 14 2,000 76 10 26.32 200 200
biodeg 1494 1,055 41 2 25.73 356 699
steel 1504 1,941 33 2 58.82 673 1,268
stock 841 950 9 2 105.56 462 488
More than 10 classes
energy 1472 698 9 23 77.56 10 74
collins 40971 970 19 26 51.05 17 80
texture 40499 5,500 40 11 137.5 500 500
A.2 Data Splitting
Figure 9 shows the data splitting setup used across all datasets. Note that data sharing (Section 3.3) shares the same data splitting as data augmentation, except that the “Training set” and “Validation set” containing real data are no longer used for training the downstream predictors.
Figure 9: Data splitting strategies for data augmentation for all datasets.
A.3 Data Preprocessing
Following the procedures presented in prior work [54, 30], we perform preprocessing in two steps. We first compute the required statistics with training data and then transform it. Firstly, we impute the missing values with the mean value for numerical features and the most mode value for categorical features. Secondly, we convert the categorical features into numerical features equal to Leave-one-out Target Statistic [66, 56]. Next, we perform Z-score normalisation for each feature. Specifically, we compute each feature’s mean and standard deviation in the training data and then transform the training samples to have a mean of zero and a variance of one for each feature. Finally, we apply the same transformation to the validation and test data before conducting evaluations.
A.4 Software and Computing Resources
Software implementation. (i) For generators: We implemented TabEBM using PyTorch 1.13 [64], an open-source deep learning library with a BSD licence. We implemented SMOTE with Imbalanced-learn [45], an open-source Python library for imbalanced datasets with an MIT licence. For other benchmark generators, we used their open-source implementations in Synthcity [67], a library for generating and evaluating synthetic tabular data with an Apache-2.0 license. (ii) For downstream predictors: We implemented TabPFN with its open-source implementation (https://github.com/automl/TabPFN). We implemented the other five downstream predictors (i.e., Logistic Regression, KNN, MLP, Random Forest and XGBoost) with their open-source implementation in scikit-learn [65], an open-source Python library under the 3-Clause BSD license. (iii) For result analysis and visualisation: All numerical plots and graphics have been generated using Matplotlib 3.7 [34], a Python-based plotting library with a BSD licence. The model architecture was generated using draw.io (https://github.com/jgraph/drawio), a free drawing software under Apache License 2.0.
We ensure the consistency and reproducibility of experimental results by implementing a uniform pipeline using PyTorch Lightning, an open-source library under an Apache-2.0 licence. We further fixed the random seeds for data loading and evaluation throughout the training and evaluation process. This ensured that TabEBM and all benchmark models were trained and evaluated on the same set of samples. We also attach our code to this submission and will release it under the MIT licence upon publication. The experimental environment settings, including library dependencies, are specified in the associated code for reference and reproduction purposes.
Computing Resources. We trained 140,000 models for evaluations (including over 35,000 of generators and over 10,500 for downstream predictors). All our experiments are run on a single machine from an internal cluster with a GPU Nvidia Quadro RTX 8000 with 48GB memory and an Intel(R) Xeon(R) Gold 5218 CPU with 16 cores (at 2.30GHz). The operating system was Ubuntu 20.4.4 LTS.
A.5 TabEBM open-source library
We implemented TabEBM as an extensible library, included as a zip file in the Supplementary material. The library will be made publicly available after publication. For practitioners, it offers an easy-to-use, domain-agnostic tool that requires no training, making it particularly suitable for data augmentation, especially in small datasets. For researchers, the library includes the complete implementation of TabEBM, facilitating future extensions and investigations into class-specific energy-based models.
The library has two core functionalities:
  1. 1.
    Generate synthetic data: The library can generate data for augmentation.
    from TabEBM import TabEBM
    tabebm = TabEBM()
    augmented_data = tabebm.generate(X_train, y_train, num_samples=100)
    % augmented_data[class_id] = numpy.ndarray of generated data
    %                            for a specific ’’class_id‘‘
  2. 2.
    Compute and visualise the energy function: The library allows computation of TabEBM’s energy function and the unnormalised data density. The demo notebook, TabEBM_approximated_density.ipynb, shows the TabEBM-inferred densities under conditions of data noise and class imbalance (thus recreating the plots from Appendix B).
A.6 Implementation of Generators
TabEBM. In all our experiments, the surrogate binary classifier in TabEBM is a pretrained in-context model, TabPFN [33], using the official model weights released by the authors (https://github.com/automl/TabPFN/raw/main/tabpfn/models_diff/prior_diff_real_checkpoint_n_0_epoch_42.cpkt). We use TabPFN with three ensembles. We use four surrogate negative samples, $\mathcal{X}_{c}^{\text{neg}}$, positioned at $\alpha^{\text{neg}}_{\text{dist}}=5$ standard deviations from zero, in random corners of a hypercube in ${\mathbb{R}}^{D}$ (as explained in Section 2.2), distant from any real data. In Section C.1, we show that TabEBM is robust to the distribution of the negative samples.
We use SGLD [85] for sampling from TabEBM, where the starting points ${\mathbf{x}}_{0}^{\text{synth}}$ are initialised by adding Gaussian noise with zero mean and standard deviation $\sigma_{\text{start}}=0.01$ to a randomly selected sample of the specific class, i.e., ${\mathbf{x}}_{0}^{\text{synth}}\sim\mathcal{N}(\mathcal{X}_{c},\sigma_{\text{start}}^{2}\mathbf{I})$. For SGLD, we used the following parameters: step size $\alpha_{\text{step}}=0.1$, noise scale $\alpha_{\text{noise}}=0.01$ and number of steps $T=200$. We found TabEBM to be robust to the SGLD settings (see Section C.2). We attach the implementation code for TabEBM and will release it as a public open-source library available after publication (see Section A.5 for details on our library).
TabPFGen. We re-implemented TabPFGen [48] by closely following the original paper since no official implementation is available. As recommended in [48], the starting points are initialised by adding Gaussian noise with zero mean and standard deviation of 0.01 to the training points.
SMOTE. We use the open-source implementation of SMOTE from Imbalanced-learn [45], and the number neighbours $k$ is set within the range of $\{1,3,5\}$. When applicable, we always set the maximum value for nearest neighbours (i.e., $k=5$). However, very low-sample-size datasets may not contain sufficient samples for large $k$. For instance, the “fourier” dataset ($N_{\text{real}}=20$) only has two samples per class. We set $k=1$ to generate synthetic data with SMOTE in these cases.
For the other six benchmark generators, we use their open-source implementations in Synthcity [68]. Following prior studies [88, 81, 72, 48], we use the default settings for all generators.
A.7 Implementation of Downstream Predictors
We implemented TabPFN with its official implementation [33] and the other five downstream predictors with the scikit-learn library [65]. Following prior studies [81, 72], we use the default settings for all downstream predictors.
Appendix B Limitations of Existing Generative Methods
We showcase three limitations of current generative models: (1) Figure 10 shows that models approximating the joint distribution $p({\mathbf{x}},y)$ may fail to preserve the stratification of the real data and even fail to generate samples from specific classes. (2) Figure 11(e) evaluates the approximated class-conditional distributions $p({\mathbf{x}}\mid y)$ on data with increasing noise levels, and (3) Figure 12(d) evaluates the approximated class-conditional distributions $p({\mathbf{x}}\mid y)$ on data with increasing class imbalance.
Figure 10: Comparison of class distribution between real data and synthetic data from TVAE. We first train TVAE on the “energy-efficiency” dataset and then randomly generate 10,000 samples with it. We highlight the classes where no synthetic samples are generated. TVAE fails to generate samples for 4 of 23 classes, showing the impracticability to preserve stratification by generative methods that learn joint distribution $p({\mathbf{x}},y)$.
Figure 11: Evaluating the approximated class-conditional distributions on data with increasing noise levels. Darker blue indicates a higher assigned probability. TabPFGen uses a single shared energy-based model to infer the class-conditional distribution $p({\mathbf{x}}|y)$. As noise increases, TabPFGen’s probability assignments vary significantly and end up assigning very high probabilities that are far from the real data. For instance, the areas of assigned probability for $p({\mathbf{x}}|y=1)$ completely flip when noise increases from 0.5 to 1. In contrast, our TabEBM uses class-specific energy models, resulting in robust inferred conditionals. TabEBM performs well even under very high noise (see $p({\mathbf{x}}|y=0)$ for noise level 2), while TabPFGen struggles.
Figure 12: Evaluating the approximated class-conditional distributions on a toy dataset of 300 samples with varying class imbalances. The two clusters maintain their positions. Darker blue indicates a higher assigned probability. TabPFGen uses a single shared energy-based model to infer the class-conditional distribution $p({\mathbf{x}}|y)$. As class imbalance increases, TabPFGen starts assigning high probability in areas far from the real data, for instance, in the case of $p({\mathbf{x}}|y=1)$ for class ratio 10:290. In contrast, our TabEBM fits class-specific energy models only on the class-wise data $\mathcal{X}_{c}=\{{\mathbf{x}}^{(i)}\mid y_{i}=c\}$. This results in very robust inferred conditional distributions even under heavy class imbalance (e.g., see that $p({\mathbf{x}}|y=1)$ remains relatively constant).
Appendix C Extended Experimental Results
C.1 Ablations on the distribution of the surrogate negative samples
C.1.1 Ablations on placing the negative samples
Figure 13: TabEBM energy $E_{c}(x)$ for different choices of negative samples. The blue region represents low energy, indicating high data density. In (A), TabEBM, with the proposed negative samples placed in a hypercube far from the data, infers an accurate energy surface, resulting in generated data close to the real points. In (B), labelling a random subset of the real data as negative samples leads to a completely inaccurate energy surface. In (C), labelling half of the real points as negative samples reduces density near the decision boundary, as TabPFN assigns low maximal logit due to the high uncertainty. In conclusion, placing negative samples far from the real data results in a robust energy surface.
Figure 13 shows TabEBM’s energy $E_{c}(x)$ when varying the selection of the negative samples. TabEBM infers an accurate energy surface with distant negative samples, and the energy surface becomes inaccurate when negative samples resemble real samples. This occurs because TabPFN is uncertain when points of different classes are close, affecting its logits magnitude and making them unsuitable for density estimation.
C.1.2 Varying the number of negative samples
We evaluate the impact of the ratio $|\mathcal{X}^{\text{neg}}_{c}|:|\mathcal{X}_{c}|$ between the negative samples $\mathcal{X}^{\text{neg}}_{c}$ and the real samples $|\mathcal{X}_{c}|$. We vary $|\mathcal{X}^{\text{neg}}_{c}|$ while keeping $|\mathcal{X}_{c}|$ fixed, simulating both balanced and highly imbalanced scenarios. The negative samples are placed in random corners of the hypercube (as described in Section 2), at five standard deviations in each direction (i.e., $\alpha^{\text{neg}}_{\text{dist}}=5$). To ensure reliable outcomes, we maintained a consistent ratio across all classes, keeping the same proportion of negative samples for each class.
Table 4 shows the results across six datasets with $N_{\text{real}}=100$ real samples, demonstrating that TabEBM is robust to imbalances in the surrogate binary tasks. The column with $|\mathcal{X}^{\text{neg}}_{c}|=4$ represents the TabEBM results from the main paper, where four negative samples were placed in the corners (as described in Section 2). There are negligible differences in performance, and TabEBM consistently outperforms both the baseline and other generators (as shown in Table 1).
Table 4: Evaluating the impact of varying the ratio $|\mathcal{X}^{\text{neg}}_{c}|:|\mathcal{X}_{c}|$. We show the test classification accuracy performance (%) of TabEBM on data augmentation averaged over six datasets and ten repeats. TabEBM shows consistent performance and outperforms the baseline, regardless of the number of negative samples.
TabEBM Baseline (Real data)
Ratio $|\mathcal{X}^{\text{neg}}_{c}|:|\mathcal{X}_{c}|$ 0.1 0.2 0.5 1 Fixed $|\mathcal{X}^{\text{neg}}_{c}|=4$
biodeg 76.59±${}_{\pm\text{3.95}}$ 76.54±${}_{\pm\text{3.95}}$ 76.47±${}_{\pm\text{4.05}}$ 76.81±${}_{\pm\text{3.58}}$ 76.45±${}_{\pm\text{3.08}}$ 76.69±${}_{\pm\text{2.70}}$
steel 92.71±${}_{\pm\text{7.46}}$ 92.60±${}_{\pm\text{7.45}}$ 92.79±${}_{\pm\text{7.50}}$ 92.63±${}_{\pm\text{7.59}}$ 92.71±${}_{\pm\text{7.57}}$ 86.87±${}_{\pm\text{12.4}}$
stock 90.46±${}_{\pm\text{3.49}}$ 90.41±${}_{\pm\text{3.65}}$ 90.52±${}_{\pm\text{3.52}}$ 90.31±${}_{\pm\text{3.63}}$ 90.36±${}_{\pm\text{3.14}}$ 89.07±${}_{\pm\text{3.71}}$
energy 31.20±${}_{\pm\text{6.22}}$ 31.20±${}_{\pm\text{6.22}}$ 30.89±${}_{\pm\text{5.83}}$ 30.90±${}_{\pm\text{6.09}}$ 31.24±${}_{\pm\text{5.53}}$ 25.94±${}_{\pm\text{4.86}}$
collins 13.06±${}_{\pm\text{2.88}}$ 13.02±${}_{\pm\text{2.85}}$ 13.05±${}_{\pm\text{2.89}}$ 12.97±${}_{\pm\text{2.79}}$ 13.07±${}_{\pm\text{2.51}}$ 11.44±${}_{\pm\text{2.77}}$
texture 85.91±${}_{\pm\text{6.92}}$ 85.91±${}_{\pm\text{6.92}}$ 85.94±${}_{\pm\text{6.76}}$ 86.26±${}_{\pm\text{6.72}}$ 86.01±${}_{\pm\text{7.36}}$ 82.42±${}_{\pm\text{10.38}}$
Average accuracy 64.99 64.95 64.94 64.98 64.97 62.07
C.1.3 Varying the distance of the negative samples
We assess the effect of varying the distance of negative samples. We use TabEBM with four negative samples positioned randomly at the corners of the hypercube, as outlined in Section 2 (this corresponds to the experimental setup from the main paper). The distance of the negative samples, denoted as $\alpha^{\text{neg}}_{\text{dist}}$, is varied. Table 5 demonstrates that TabEBM remains generally robust to changes in this distance, with only small performance variations across different datasets. Importantly, using TabEBM for data augmentation consistently improves performance by approximately 3% compared to the Baseline, regardless of the distance used.
Table 5: Evaluating the impact of varying the distance of the negative samples $\alpha^{\text{neg}}_{\text{dist}}$ across various datasets. We show the test classification accuracy performance (%) of TabEBM on data augmentation averaged over six datasets and ten repeats. TabEBM is robust, and optional tuning of the negative samples could slightly improve performance.
TabEBM Baseline (Real data)
Per-dimension distance $\alpha_{d}$ of the negative samples 0.1 0.2 0.5 1 2 5 -
biodeg 76.72±${}_{\pm\text{3.33}}$ 76.62±${}_{\pm\text{3.40}}$ 77.12±${}_{\pm\text{2.60}}$ 76.85±${}_{\pm\text{3.14}}$ 76.50±${}_{\pm\text{3.93}}$ 76.45±${}_{\pm\text{3.08}}$ 76.69±${}_{\pm\text{2.70}}$
steel 93.97±${}_{\pm\text{5.76}}$ 93.46±${}_{\pm\text{6.24}}$ 93.00±${}_{\pm\text{6.92}}$ 92.60±${}_{\pm\text{7.31}}$ 92.68±${}_{\pm\text{7.38}}$ 92.71±${}_{\pm\text{7.57}}$ 86.87±${}_{\pm\text{12.4}}$
stock 90.42±${}_{\pm\text{3.46}}$ 90.29±${}_{\pm\text{3.61}}$ 90.56±${}_{\pm\text{3.46}}$ 90.38±${}_{\pm\text{3.64}}$ 90.43±${}_{\pm\text{3.56}}$ 90.36±${}_{\pm\text{3.14}}$ 89.07±${}_{\pm\text{3.71}}$
energy 31.73±${}_{\pm\text{6.21}}$ 31.42±${}_{\pm\text{6.08}}$ 31.86±${}_{\pm\text{6.12}}$ 32.53±${}_{\pm\text{5.96}}$ 31.65±${}_{\pm\text{6.06}}$ 31.24±${}_{\pm\text{5.53}}$ 25.94±${}_{\pm\text{4.86}}$
collins 13.03±${}_{\pm\text{2.59}}$ 12.92±${}_{\pm\text{2.60}}$ 12.97±${}_{\pm\text{2.69}}$ 13.03±${}_{\pm\text{2.84}}$ 13.08±${}_{\pm\text{2.93}}$ 13.07±${}_{\pm\text{2.51}}$ 11.44±${}_{\pm\text{2.77}}$
texture 85.62±${}_{\pm\text{7.41}}$ 85.58±${}_{\pm\text{7.49}}$ 85.50±${}_{\pm\text{7.65}}$ 85.05±${}_{\pm\text{8.21}}$ 85.20±${}_{\pm\text{7.95}}$ 86.01±${}_{\pm\text{7.36}}$ 82.42±${}_{\pm\text{10.38}}$
Average accuracy 65.25 65.05 65.17 65.07 64.92 64.97 62.07
C.2 Ablations on the sensitivity to the hyperparameters of SGLD sampling
We vary two key hyperparameters of SGLD on the “biodeg” binary dataset with $N_{\text{real}}=100$: the step size $\alpha_{\text{step}}$ and the noise scale $\alpha_{\text{noise}}$. Table 6 shows that TabEBM remains stable with respect to these hyperparameters. Note that smaller values of $\alpha_{\text{noise}}$ are expected to perform better because SGLD sampling adds noise at each iteration (see Line 7 in Algorithm 1), thus larger values of $\alpha_{\text{noise}}$ will hinder convergence of the SGLD sampler.
Table 6: Test classification accuracy (%) of TabEBM (averaged over six downstream predictors) with different SGLD settings. Increasing $\alpha_{\text{noise}}$ (added at each SGLD step) is expected to degrade performance, as it causes the sampling to diverge further from the real data.
$\alpha_{\text{noise}}$ $\alpha_{\text{step}}$
0.1 0.3 0.5 1.0
0.01 76.45 77.09 77.04 76.58
0.02 76.86 76.96 76.77 76.26
0.05 75.93 75.89 75.94 75.70
C.3 Distribution of Logits and Unnormalized Density in TabEBM
Figure 14: Additional results for Section 3.4. The logit distribution of TabPFN trained on our surrogate binary tasks across four datasets. Starting from the real samples, random points are selected at increasing distances (shown on the x-axis). The top row shows the logit distributions for the surrogate task. Close to the real data, TabPFN outputs a high logit value. As the distance increases, the logits converge due to increased predictive uncertainty, leading to equal class probabilities after applying softmax. Notably, across datasets, TabPFN’s logits are always positive, have similar ranges, and maintain a relatively constant sum as distance increases. The bottom row TabEBM’s unnormalized density, $p_{c}(x)\propto\exp(-E_{c}(x))\rightarrow p_{c}(x)\propto(\exp(f(x)[0])+\exp(f(x)[1]))$. The density decreases significantly far from the data, becoming negligible. Because sampling using SGLD perform gradient ascent on the density, the TabEBM-generated samples will be similar when using one or both logits.
C.4 Complete Trade-off Figures with Error Bars
Figure 15: (a1&a2): Median inverse KL and KS test vs. mean normalised balanced accuracy improvement (%) between real train data and synthetic data. (b1&b2): Median DCR and $\delta-presence$ vs. mean normalised balanced accuracy change (%) between real train data and synthetic data. Note that “accuracy improvement” is for data augmentation, and “accuracy change” is for data sharing. TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
C.5 Results on Data Augmentation
C.5.1 Results on eight OpenML datasets.
Table 7: Classification accuracy (%) of LR, comparing data augmentation on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes that a specific generator was not applicable or the downstream predictor failed to converge, and the rank is computed with the mean balanced accuracy of other methods. We bold the highest accuracy for each dataset of different sample sizes. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ Baseline (Real data) SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 36.33±${}_{\pm\text{3.04}}$ N/A 22.02±${}_{\pm\text{2.91}}$ 21.04±${}_{\pm\text{4.76}}$ 18.40±${}_{\pm\text{4.82}}$ 18.77±${}_{\pm\text{3.84}}$ 25.92±${}_{\pm\text{4.30}}$ 36.61±${}_{\pm\text{2.53}}$ 38.07±${}_{\pm\text{1.25}}$ 38.01±${}_{\pm\text{2.38}}$
50 62.14±${}_{\pm\text{3.77}}$ 61.43±${}_{\pm\text{4.34}}$ 37.04±${}_{\pm\text{2.79}}$ 33.10±${}_{\pm\text{5.99}}$ 31.25±${}_{\pm\text{4.21}}$ 23.98±${}_{\pm\text{2.75}}$ 43.64±${}_{\pm\text{5.07}}$ 54.95±${}_{\pm\text{3.28}}$ 63.00±${}_{\pm\text{3.69}}$ 63.05±${}_{\pm\text{3.84}}$
100 79.97±${}_{\pm\text{3.24}}$ 79.53±${}_{\pm\text{3.37}}$ 61.07±${}_{\pm\text{5.06}}$ 55.44±${}_{\pm\text{1.92}}$ 46.37±${}_{\pm\text{4.10}}$ 45.55±${}_{\pm\text{4.24}}$ 56.77±${}_{\pm\text{3.06}}$ 67.25±${}_{\pm\text{4.50}}$ 80.54±${}_{\pm\text{3.27}}$ 80.32±${}_{\pm\text{3.12}}$
200 91.53±${}_{\pm\text{1.58}}$ 90.92±${}_{\pm\text{1.81}}$ 77.43±${}_{\pm\text{2.75}}$ 71.27±${}_{\pm\text{3.07}}$ 66.16±${}_{\pm\text{4.31}}$ 66.37±${}_{\pm\text{3.42}}$ 70.52±${}_{\pm\text{2.17}}$ 76.30±${}_{\pm\text{3.70}}$ 91.69±${}_{\pm\text{1.66}}$ 91.34±${}_{\pm\text{1.77}}$
500 97.86±${}_{\pm\text{0.83}}$ 97.69±${}_{\pm\text{0.80}}$ 90.77±${}_{\pm\text{0.93}}$ 89.05±${}_{\pm\text{1.52}}$ 85.09±${}_{\pm\text{1.99}}$ 83.58±${}_{\pm\text{2.22}}$ 88.55±${}_{\pm\text{1.54}}$ 90.64±${}_{\pm\text{0.81}}$ 97.97±${}_{\pm\text{0.61}}$ 97.88±${}_{\pm\text{0.86}}$
fourier 20 42.90±${}_{\pm\text{5.30}}$ N/A 22.46±${}_{\pm\text{5.88}}$ 16.00±${}_{\pm\text{4.70}}$ 15.48±${}_{\pm\text{3.79}}$ 13.58±${}_{\pm\text{4.30}}$ 22.04±${}_{\pm\text{4.42}}$ 15.80±${}_{\pm\text{4.15}}$ 44.67±${}_{\pm\text{8.85}}$ 43.02±${}_{\pm\text{5.14}}$
50 60.62±${}_{\pm\text{1.64}}$ 58.40±${}_{\pm\text{1.95}}$ 33.42±${}_{\pm\text{2.98}}$ 31.18±${}_{\pm\text{5.47}}$ 28.70±${}_{\pm\text{3.74}}$ 26.18±${}_{\pm\text{3.80}}$ 39.04±${}_{\pm\text{3.12}}$ 40.00±${}_{\pm\text{4.97}}$ 60.07±${}_{\pm\text{2.14}}$ 60.36±${}_{\pm\text{1.55}}$
100 67.76±${}_{\pm\text{2.49}}$ 65.84±${}_{\pm\text{2.35}}$ 41.36±${}_{\pm\text{2.85}}$ 40.32±${}_{\pm\text{3.49}}$ 40.32±${}_{\pm\text{5.82}}$ 41.44±${}_{\pm\text{5.02}}$ 47.90±${}_{\pm\text{3.74}}$ 39.78±${}_{\pm\text{3.99}}$ 67.40±${}_{\pm\text{1.51}}$ 67.44±${}_{\pm\text{2.46}}$
200 73.13±${}_{\pm\text{2.41}}$ 71.56±${}_{\pm\text{2.67}}$ 54.76±${}_{\pm\text{3.46}}$ 55.00±${}_{\pm\text{3.72}}$ 52.40±${}_{\pm\text{3.18}}$ 58.08±${}_{\pm\text{3.52}}$ 58.48±${}_{\pm\text{2.08}}$ 50.98±${}_{\pm\text{2.68}}$ 70.30±${}_{\pm\text{2.91}}$ 72.38±${}_{\pm\text{3.01}}$
500 77.44±${}_{\pm\text{1.20}}$ 76.42±${}_{\pm\text{1.28}}$ 68.28±${}_{\pm\text{2.12}}$ 70.18±${}_{\pm\text{1.89}}$ 68.12±${}_{\pm\text{1.62}}$ 72.36±${}_{\pm\text{1.65}}$ 71.54±${}_{\pm\text{1.95}}$ 69.48±${}_{\pm\text{1.71}}$ 76.52±${}_{\pm\text{1.69}}$ 77.50±${}_{\pm\text{2.14}}$
biodeg 20 71.34±${}_{\pm\text{5.63}}$ 70.10±${}_{\pm\text{5.49}}$ 70.16±${}_{\pm\text{5.75}}$ 58.17±${}_{\pm\text{8.00}}$ 58.05±${}_{\pm\text{9.91}}$ 49.99±${}_{\pm\text{5.88}}$ 62.61±${}_{\pm\text{6.45}}$ 69.47±${}_{\pm\text{6.00}}$ 70.76±${}_{\pm\text{3.95}}$ 71.24±${}_{\pm\text{4.85}}$
50 76.35±${}_{\pm\text{2.88}}$ 75.69±${}_{\pm\text{3.03}}$ 73.63±${}_{\pm\text{2.64}}$ 67.44±${}_{\pm\text{3.83}}$ 62.87±${}_{\pm\text{7.30}}$ 49.44±${}_{\pm\text{2.63}}$ 74.44±${}_{\pm\text{2.77}}$ 71.75±${}_{\pm\text{5.27}}$ 75.68±${}_{\pm\text{2.31}}$ 76.41±${}_{\pm\text{2.93}}$
100 78.91±${}_{\pm\text{1.40}}$ 78.39±${}_{\pm\text{1.53}}$ 77.09±${}_{\pm\text{2.80}}$ 74.89±${}_{\pm\text{2.54}}$ 68.62±${}_{\pm\text{5.21}}$ 55.61±${}_{\pm\text{3.56}}$ 75.62±${}_{\pm\text{2.77}}$ 72.45±${}_{\pm\text{3.31}}$ 77.92±${}_{\pm\text{2.41}}$ 78.34±${}_{\pm\text{2.18}}$
200 82.00±${}_{\pm\text{1.47}}$ 81.42±${}_{\pm\text{1.39}}$ 80.07±${}_{\pm\text{1.82}}$ 78.56±${}_{\pm\text{3.43}}$ 72.35±${}_{\pm\text{1.72}}$ 59.06±${}_{\pm\text{4.65}}$ 78.03±${}_{\pm\text{1.95}}$ 73.73±${}_{\pm\text{2.09}}$ 81.24±${}_{\pm\text{1.71}}$ 81.43±${}_{\pm\text{1.78}}$
500 83.83±${}_{\pm\text{0.57}}$ 83.74±${}_{\pm\text{0.90}}$ 81.69±${}_{\pm\text{0.82}}$ 82.12±${}_{\pm\text{1.17}}$ 78.06±${}_{\pm\text{2.13}}$ 66.86±${}_{\pm\text{5.43}}$ 81.47±${}_{\pm\text{0.93}}$ 77.98±${}_{\pm\text{1.27}}$ 83.43±${}_{\pm\text{0.82}}$ 83.10±${}_{\pm\text{0.98}}$
steel 20 63.66±${}_{\pm\text{8.98}}$ 57.88±${}_{\pm\text{5.72}}$ 60.27±${}_{\pm\text{7.47}}$ 57.90±${}_{\pm\text{4.45}}$ 53.10±${}_{\pm\text{7.28}}$ 54.20±${}_{\pm\text{6.99}}$ 55.41±${}_{\pm\text{4.92}}$ 53.29±${}_{\pm\text{4.31}}$ 66.81±${}_{\pm\text{9.74}}$ 67.03±${}_{\pm\text{9.35}}$
50 87.91±${}_{\pm\text{5.88}}$ 69.01±${}_{\pm\text{6.60}}$ 66.22±${}_{\pm\text{3.63}}$ 66.22±${}_{\pm\text{5.77}}$ 57.05±${}_{\pm\text{5.51}}$ 57.46±${}_{\pm\text{8.48}}$ 64.81±${}_{\pm\text{4.64}}$ 57.20±${}_{\pm\text{5.19}}$ 93.63±${}_{\pm\text{4.78}}$ 92.20±${}_{\pm\text{4.81}}$
100 98.85±${}_{\pm\text{1.20}}$ 82.67±${}_{\pm\text{4.30}}$ 74.33±${}_{\pm\text{3.85}}$ 70.49±${}_{\pm\text{5.35}}$ 65.09±${}_{\pm\text{7.30}}$ 52.77±${}_{\pm\text{7.06}}$ 67.85±${}_{\pm\text{4.94}}$ 61.62±${}_{\pm\text{4.05}}$ 99.24±${}_{\pm\text{0.82}}$ 99.21±${}_{\pm\text{0.86}}$
200 99.43±${}_{\pm\text{0.58}}$ 87.18±${}_{\pm\text{3.06}}$ 82.77±${}_{\pm\text{3.21}}$ 80.34±${}_{\pm\text{2.93}}$ 70.49±${}_{\pm\text{5.27}}$ 72.99±${}_{\pm\text{13.98}}$ 80.27±${}_{\pm\text{7.32}}$ 64.52±${}_{\pm\text{2.16}}$ 99.45±${}_{\pm\text{0.69}}$ 99.51±${}_{\pm\text{0.69}}$
500 99.75±${}_{\pm\text{0.29}}$ 96.63±${}_{\pm\text{2.11}}$ 94.59±${}_{\pm\text{2.98}}$ 96.32±${}_{\pm\text{1.52}}$ 84.15±${}_{\pm\text{2.69}}$ 98.07±${}_{\pm\text{1.37}}$ 95.35±${}_{\pm\text{2.06}}$ 70.11±${}_{\pm\text{2.58}}$ 99.84±${}_{\pm\text{0.20}}$ 99.84±${}_{\pm\text{0.20}}$
stock 20 77.99±${}_{\pm\text{4.40}}$ 80.45±${}_{\pm\text{3.98}}$ 74.21±${}_{\pm\text{6.36}}$ 59.20±${}_{\pm\text{12.69}}$ 72.50±${}_{\pm\text{7.92}}$ 72.09±${}_{\pm\text{9.75}}$ 69.04±${}_{\pm\text{6.25}}$ 80.59±${}_{\pm\text{3.59}}$ 79.54±${}_{\pm\text{4.46}}$ 80.39±${}_{\pm\text{3.42}}$
50 80.68±${}_{\pm\text{2.65}}$ 81.49±${}_{\pm\text{2.95}}$ 76.41±${}_{\pm\text{3.95}}$ 72.95±${}_{\pm\text{2.17}}$ 75.41±${}_{\pm\text{6.00}}$ 78.44±${}_{\pm\text{4.40}}$ 76.91±${}_{\pm\text{2.36}}$ 75.49±${}_{\pm\text{5.31}}$ 82.37±${}_{\pm\text{3.20}}$ 82.21±${}_{\pm\text{2.60}}$
100 82.11±${}_{\pm\text{1.11}}$ 83.86±${}_{\pm\text{1.97}}$ 79.85±${}_{\pm\text{2.79}}$ 78.47±${}_{\pm\text{2.71}}$ 76.99±${}_{\pm\text{3.49}}$ 80.82±${}_{\pm\text{3.57}}$ 78.89±${}_{\pm\text{2.36}}$ 77.65±${}_{\pm\text{2.60}}$ 83.67±${}_{\pm\text{1.60}}$ 83.52±${}_{\pm\text{1.76}}$
200 82.18±${}_{\pm\text{0.81}}$ 84.29±${}_{\pm\text{1.19}}$ 79.24±${}_{\pm\text{2.82}}$ 79.86±${}_{\pm\text{2.42}}$ 76.49±${}_{\pm\text{1.37}}$ 80.21±${}_{\pm\text{2.13}}$ 78.87±${}_{\pm\text{2.46}}$ 76.91±${}_{\pm\text{1.04}}$ 83.75±${}_{\pm\text{1.53}}$ 84.17±${}_{\pm\text{1.42}}$
More than 10 classes energy 50 22.22±${}_{\pm\text{2.36}}$ N/A 10.11±${}_{\pm\text{2.20}}$ 9.58±${}_{\pm\text{3.15}}$ 7.70±${}_{\pm\text{1.83}}$ 8.20±${}_{\pm\text{2.01}}$ 10.51±${}_{\pm\text{1.28}}$ 17.10±${}_{\pm\text{5.03}}$ N/A 21.66±${}_{\pm\text{1.54}}$
100 24.00±${}_{\pm\text{2.30}}$ N/A 13.80±${}_{\pm\text{2.23}}$ 13.01±${}_{\pm\text{1.71}}$ 12.14±${}_{\pm\text{1.87}}$ 10.79±${}_{\pm\text{3.19}}$ 15.65±${}_{\pm\text{2.40}}$ 14.45±${}_{\pm\text{2.90}}$ N/A 28.10±${}_{\pm\text{2.19}}$
200 29.37±${}_{\pm\text{2.63}}$ N/A 16.39±${}_{\pm\text{2.68}}$ 16.56±${}_{\pm\text{3.58}}$ 16.78±${}_{\pm\text{3.15}}$ 18.11±${}_{\pm\text{1.71}}$ 20.10±${}_{\pm\text{2.48}}$ 20.92±${}_{\pm\text{2.79}}$ N/A 34.38±${}_{\pm\text{2.60}}$
collins 100 14.28±${}_{\pm\text{1.63}}$ N/A 10.57±${}_{\pm\text{1.72}}$ 8.69±${}_{\pm\text{1.17}}$ 9.59±${}_{\pm\text{1.35}}$ 13.31±${}_{\pm\text{1.67}}$ 8.69±${}_{\pm\text{1.80}}$ 12.08±${}_{\pm\text{1.56}}$ N/A 14.01±${}_{\pm\text{2.55}}$
200 19.20±${}_{\pm\text{1.71}}$ 19.39±${}_{\pm\text{1.88}}$ 16.03±${}_{\pm\text{1.74}}$ 11.64±${}_{\pm\text{1.76}}$ 10.97±${}_{\pm\text{1.46}}$ 17.06±${}_{\pm\text{1.51}}$ 11.31±${}_{\pm\text{1.58}}$ 17.80±${}_{\pm\text{1.21}}$ N/A 19.33±${}_{\pm\text{1.55}}$
texture 50 86.56±${}_{\pm\text{2.96}}$ 86.93±${}_{\pm\text{2.77}}$ 55.01±${}_{\pm\text{5.77}}$ 42.17±${}_{\pm\text{6.36}}$ 44.63±${}_{\pm\text{5.41}}$ 60.07±${}_{\pm\text{10.11}}$ 44.46±${}_{\pm\text{6.63}}$ 77.68±${}_{\pm\text{4.33}}$ N/A 88.54±${}_{\pm\text{2.88}}$
100 94.07±${}_{\pm\text{1.70}}$ 93.87±${}_{\pm\text{1.82}}$ 65.36±${}_{\pm\text{4.49}}$ 60.07±${}_{\pm\text{6.81}}$ 60.76±${}_{\pm\text{5.18}}$ 73.16±${}_{\pm\text{5.11}}$ 64.69±${}_{\pm\text{4.79}}$ 84.13±${}_{\pm\text{1.97}}$ N/A 94.38±${}_{\pm\text{1.24}}$
200 96.65±${}_{\pm\text{1.24}}$ 96.53±${}_{\pm\text{1.33}}$ 75.91±${}_{\pm\text{5.58}}$ 80.02±${}_{\pm\text{5.13}}$ 77.07±${}_{\pm\text{3.89}}$ 86.24±${}_{\pm\text{3.62}}$ 85.90±${}_{\pm\text{2.78}}$ 85.94±${}_{\pm\text{2.88}}$ N/A 96.53±${}_{\pm\text{1.27}}$
500 98.03±${}_{\pm\text{0.36}}$ 98.05±${}_{\pm\text{0.23}}$ 91.87±${}_{\pm\text{0.93}}$ 92.93±${}_{\pm\text{1.78}}$ 90.01±${}_{\pm\text{1.80}}$ 93.92±${}_{\pm\text{0.81}}$ 94.83±${}_{\pm\text{0.89}}$ 91.72±${}_{\pm\text{1.49}}$ N/A 97.75±${}_{\pm\text{0.42}}$
Average rank 2.36±${}_{\pm\text{1.14}}$ 3.45±${}_{\pm\text{1.35}}$ 6.52±${}_{\pm\text{1.48}}$ 7.53±${}_{\pm\text{1.42}}$ 9.08±${}_{\pm\text{0.77}}$ 7.61±${}_{\pm\text{2.33}}$ 6.70±${}_{\pm\text{1.47}}$ 6.67±${}_{\pm\text{2.53}}$ 3.17±${}_{\pm\text{1.81}}$ 1.92±${}_{\pm\text{0.75}}$
Table 8: Classification accuracy (%) of KNN, comparing data augmentation on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes that a specific generator was not applicable or the downstream predictor failed to converge, and the rank is computed with the mean balanced accuracy of other methods. We bold the highest accuracy for each dataset of different sample sizes. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ Baseline (Real data) SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 21.34±${}_{\pm\text{2.93}}$ N/A 21.78±${}_{\pm\text{2.06}}$ 21.18±${}_{\pm\text{4.22}}$ 21.30±${}_{\pm\text{1.90}}$ 22.00±${}_{\pm\text{2.70}}$ 22.69±${}_{\pm\text{3.86}}$ 16.99±${}_{\pm\text{3.45}}$ 35.78±${}_{\pm\text{4.46}}$ 35.76±${}_{\pm\text{4.37}}$
50 36.41±${}_{\pm\text{4.33}}$ 55.24±${}_{\pm\text{3.81}}$ 35.85±${}_{\pm\text{2.50}}$ 36.13±${}_{\pm\text{4.24}}$ 35.40±${}_{\pm\text{4.27}}$ 36.77±${}_{\pm\text{4.06}}$ 36.84±${}_{\pm\text{4.05}}$ 31.02±${}_{\pm\text{4.11}}$ ${}_{\pm\text{3.53}}$ 53.49±${}_{\pm\text{3.30}}$
100 50.17±${}_{\pm\text{3.11}}$ 70.11±${}_{\pm\text{2.82}}$ 51.97±${}_{\pm\text{2.84}}$ 50.61±${}_{\pm\text{3.15}}$ 50.62±${}_{\pm\text{3.27}}$ 50.63±${}_{\pm\text{3.55}}$ 50.36±${}_{\pm\text{3.44}}$ 44.70±${}_{\pm\text{2.22}}$ 67.99±${}_{\pm\text{2.43}}$ 68.27±${}_{\pm\text{2.51}}$
200 65.84±${}_{\pm\text{2.78}}$ 80.43±${}_{\pm\text{2.44}}$ 65.52±${}_{\pm\text{2.96}}$ 66.05±${}_{\pm\text{2.74}}$ 66.14±${}_{\pm\text{2.42}}$ 67.50±${}_{\pm\text{2.57}}$ 66.52±${}_{\pm\text{3.16}}$ 63.92±${}_{\pm\text{3.26}}$ 79.94±${}_{\pm\text{2.26}}$ 80.55±${}_{\pm\text{2.02}}$
500 85.63±${}_{\pm\text{1.41}}$ 90.92±${}_{\pm\text{1.42}}$ 87.08±${}_{\pm\text{1.86}}$ 85.77±${}_{\pm\text{1.43}}$ 85.51±${}_{\pm\text{1.50}}$ 86.47±${}_{\pm\text{1.40}}$ 85.87±${}_{\pm\text{1.63}}$ 85.64±${}_{\pm\text{1.94}}$ 91.32±${}_{\pm\text{1.09}}$ 91.67±${}_{\pm\text{1.11}}$
fourier 20 18.06±${}_{\pm\text{3.30}}$ N/A 26.56±${}_{\pm\text{4.92}}$ 24.88±${}_{\pm\text{3.66}}$ 19.80±${}_{\pm\text{3.77}}$ 19.30±${}_{\pm\text{3.53}}$ 23.42±${}_{\pm\text{3.45}}$ 18.78±${}_{\pm\text{2.17}}$ 41.08±${}_{\pm\text{6.56}}$ 42.78±${}_{\pm\text{5.83}}$
50 48.00±${}_{\pm\text{2.47}}$ 60.38±${}_{\pm\text{1.67}}$ 39.86±${}_{\pm\text{3.73}}$ 46.82±${}_{\pm\text{3.52}}$ ${}_{\pm\text{3.45}}$ 49.54±${}_{\pm\text{2.78}}$ 42.98±${}_{\pm\text{2.80}}$ 28.12±${}_{\pm\text{2.75}}$ 59.50±${}_{\pm\text{1.99}}$ 58.54±${}_{\pm\text{1.86}}$
100 58.36±${}_{\pm\text{3.26}}$ 66.96±${}_{\pm\text{2.47}}$ 48.44±${}_{\pm\text{4.14}}$ 53.94±${}_{\pm\text{3.47}}$ 53.50±${}_{\pm\text{2.54}}$ 60.80±${}_{\pm\text{4.28}}$ 52.74±${}_{\pm\text{3.15}}$ 35.70±${}_{\pm\text{2.40}}$ 63.88±${}_{\pm\text{2.53}}$ 65.08±${}_{\pm\text{2.47}}$
200 68.60±${}_{\pm\text{2.55}}$ 71.90±${}_{\pm\text{2.02}}$ 59.66±${}_{\pm\text{3.31}}$ 66.54±${}_{\pm\text{2.75}}$ 65.16±${}_{\pm\text{2.71}}$ 70.22±${}_{\pm\text{2.26}}$ 64.52±${}_{\pm\text{2.44}}$ 51.24±${}_{\pm\text{7.29}}$ 70.32±${}_{\pm\text{1.94}}$ 71.08±${}_{\pm\text{1.87}}$
500 76.90±${}_{\pm\text{1.30}}$ 77.64±${}_{\pm\text{1.07}}$ 73.20±${}_{\pm\text{1.68}}$ 76.22±${}_{\pm\text{1.62}}$ 75.72±${}_{\pm\text{1.40}}$ 78.88±${}_{\pm\text{1.58}}$ 76.54±${}_{\pm\text{0.77}}$ 63.66±${}_{\pm\text{2.49}}$ 74.30±${}_{\pm\text{1.51}}$ 75.35±${}_{\pm\text{1.34}}$
biodeg 20 65.23±${}_{\pm\text{5.01}}$ 68.99±${}_{\pm\text{3.31}}$ 66.63±${}_{\pm\text{7.83}}$ 56.99±${}_{\pm\text{5.55}}$ 59.91±${}_{\pm\text{6.09}}$ 55.85±${}_{\pm\text{4.94}}$ 58.77±${}_{\pm\text{5.93}}$ 56.62±${}_{\pm\text{7.29}}$ 67.79±${}_{\pm\text{4.64}}$ 69.76±${}_{\pm\text{4.43}}$
50 71.26±${}_{\pm\text{3.13}}$ 73.19±${}_{\pm\text{2.46}}$ 70.80±${}_{\pm\text{2.14}}$ 70.00±${}_{\pm\text{5.92}}$ 65.90±${}_{\pm\text{3.57}}$ 73.50±${}_{\pm\text{4.43}}$ 70.23±${}_{\pm\text{3.35}}$ 65.29±${}_{\pm\text{4.57}}$ 72.08±${}_{\pm\text{3.84}}$ ${}_{\pm\text{3.57}}$
100 76.12±${}_{\pm\text{1.98}}$ 76.07±${}_{\pm\text{1.74}}$ 74.02±${}_{\pm\text{2.78}}$ 75.36±${}_{\pm\text{2.18}}$ 73.24±${}_{\pm\text{2.61}}$ 77.34±${}_{\pm\text{2.19}}$ 74.28±${}_{\pm\text{2.02}}$ 72.26±${}_{\pm\text{2.46}}$ 74.56±${}_{\pm\text{1.58}}$ 75.60±${}_{\pm\text{1.55}}$
200 78.86±${}_{\pm\text{2.19}}$ 79.67±${}_{\pm\text{1.68}}$ 77.31±${}_{\pm\text{2.93}}$ 78.05±${}_{\pm\text{3.07}}$ 77.64±${}_{\pm\text{2.71}}$ 77.84±${}_{\pm\text{2.62}}$ 78.81±${}_{\pm\text{2.66}}$ 76.82±${}_{\pm\text{2.29}}$ 77.46±${}_{\pm\text{1.68}}$ 78.46±${}_{\pm\text{1.69}}$
500 82.59±${}_{\pm\text{1.17}}$ 83.07±${}_{\pm\text{1.50}}$ 82.13±${}_{\pm\text{1.21}}$ 82.17±${}_{\pm\text{1.32}}$ 82.80±${}_{\pm\text{1.28}}$ 81.06±${}_{\pm\text{1.22}}$ 82.15±${}_{\pm\text{1.33}}$ 82.10±${}_{\pm\text{0.79}}$ 79.99±${}_{\pm\text{1.76}}$ 81.01±${}_{\pm\text{1.66}}$
steel 20 56.40±${}_{\pm\text{4.48}}$ 63.95±${}_{\pm\text{3.14}}$ 59.45±${}_{\pm\text{8.27}}$ 57.04±${}_{\pm\text{5.05}}$ 54.59±${}_{\pm\text{5.81}}$ 65.46±${}_{\pm\text{6.10}}$ 56.97±${}_{\pm\text{5.43}}$ 52.90±${}_{\pm\text{3.76}}$ 70.68±${}_{\pm\text{3.87}}$ 69.31±${}_{\pm\text{4.02}}$
50 73.95±${}_{\pm\text{4.76}}$ 70.24±${}_{\pm\text{3.44}}$ 67.60±${}_{\pm\text{4.10}}$ 68.77±${}_{\pm\text{2.85}}$ 67.00±${}_{\pm\text{4.58}}$ 85.14±${}_{\pm\text{8.76}}$ 64.02±${}_{\pm\text{3.71}}$ 57.54±${}_{\pm\text{2.34}}$ 82.09±${}_{\pm\text{3.09}}$ 80.47±${}_{\pm\text{3.48}}$
100 84.70±${}_{\pm\text{5.57}}$ 77.46±${}_{\pm\text{3.67}}$ 71.87±${}_{\pm\text{2.98}}$ 72.94±${}_{\pm\text{4.62}}$ 77.09±${}_{\pm\text{2.63}}$ 94.05±${}_{\pm\text{3.84}}$ 72.62±${}_{\pm\text{5.21}}$ 61.08±${}_{\pm\text{1.93}}$ 87.77±${}_{\pm\text{3.13}}$ 87.67±${}_{\pm\text{3.22}}$
200 90.44±${}_{\pm\text{2.80}}$ 82.46±${}_{\pm\text{1.43}}$ 80.83±${}_{\pm\text{2.65}}$ 82.73±${}_{\pm\text{3.88}}$ 85.49±${}_{\pm\text{3.59}}$ 98.99±${}_{\pm\text{0.75}}$ 83.38±${}_{\pm\text{2.67}}$ 69.12±${}_{\pm\text{2.56}}$ 92.01±${}_{\pm\text{1.73}}$ 92.06±${}_{\pm\text{1.48}}$
500 94.99±${}_{\pm\text{1.09}}$ 89.97±${}_{\pm\text{0.88}}$ 91.34±${}_{\pm\text{1.69}}$ 92.42±${}_{\pm\text{1.36}}$ 93.37±${}_{\pm\text{1.13}}$ 99.71±${}_{\pm\text{0.21}}$ 92.02±${}_{\pm\text{2.03}}$ 80.79±${}_{\pm\text{1.93}}$ 95.08±${}_{\pm\text{1.30}}$ 95.50±${}_{\pm\text{1.49}}$
stock 20 71.89±${}_{\pm\text{4.37}}$ 84.41±${}_{\pm\text{5.28}}$ 73.80±${}_{\pm\text{4.68}}$ 66.38±${}_{\pm\text{9.10}}$ 68.93±${}_{\pm\text{10.49}}$ 81.82±${}_{\pm\text{8.38}}$ 67.53±${}_{\pm\text{8.58}}$ 71.80±${}_{\pm\text{4.99}}$ 84.41±${}_{\pm\text{4.22}}$ 84.69±${}_{\pm\text{4.16}}$
50 85.03±${}_{\pm\text{3.39}}$ 89.77±${}_{\pm\text{1.99}}$ 84.32±${}_{\pm\text{3.97}}$ 83.49±${}_{\pm\text{3.67}}$ 84.43±${}_{\pm\text{2.04}}$ 89.34±${}_{\pm\text{1.59}}$ 84.33±${}_{\pm\text{3.22}}$ 83.64±${}_{\pm\text{2.53}}$ 89.67±${}_{\pm\text{1.88}}$ 89.68±${}_{\pm\text{1.87}}$
100 89.66±${}_{\pm\text{1.39}}$ 92.32±${}_{\pm\text{0.99}}$ 89.58±${}_{\pm\text{1.22}}$ 89.61±${}_{\pm\text{1.36}}$ 89.66±${}_{\pm\text{1.01}}$ 91.40±${}_{\pm\text{1.41}}$ 89.66±${}_{\pm\text{2.10}}$ 89.44±${}_{\pm\text{1.41}}$ 92.02±${}_{\pm\text{0.81}}$ 92.47±${}_{\pm\text{0.83}}$
200 91.65±${}_{\pm\text{1.08}}$ 93.46±${}_{\pm\text{0.82}}$ 92.37±${}_{\pm\text{1.18}}$ 91.55±${}_{\pm\text{1.19}}$ 91.43±${}_{\pm\text{1.34}}$ 92.92±${}_{\pm\text{1.00}}$ 91.14±${}_{\pm\text{1.58}}$ 91.53±${}_{\pm\text{1.05}}$ 93.15±${}_{\pm\text{0.72}}$ 93.62±${}_{\pm\text{1.14}}$
More than 10 classes energy 50 10.85±${}_{\pm\text{1.76}}$ N/A 10.64±${}_{\pm\text{2.36}}$ 8.22±${}_{\pm\text{2.03}}$ 8.83±${}_{\pm\text{1.53}}$ 8.92±${}_{\pm\text{2.51}}$ 9.14±${}_{\pm\text{1.95}}$ 11.86±${}_{\pm\text{2.33}}$ N/A 25.36±${}_{\pm\text{2.27}}$
100 18.60±${}_{\pm\text{1.83}}$ N/A 13.71±${}_{\pm\text{1.66}}$ 15.81±${}_{\pm\text{1.50}}$ 14.67±${}_{\pm\text{1.55}}$ 16.18±${}_{\pm\text{1.75}}$ 15.71±${}_{\pm\text{2.79}}$ 17.64±${}_{\pm\text{2.68}}$ N/A 29.82±${}_{\pm\text{2.74}}$
200 26.45±${}_{\pm\text{1.49}}$ N/A 20.71±${}_{\pm\text{1.02}}$ 21.71±${}_{\pm\text{3.23}}$ 23.40±${}_{\pm\text{2.15}}$ 23.95±${}_{\pm\text{2.94}}$ 23.09±${}_{\pm\text{2.56}}$ 27.35±${}_{\pm\text{2.28}}$ N/A 35.93±${}_{\pm\text{2.85}}$
collins 100 10.59±${}_{\pm\text{1.48}}$ N/A 7.58±${}_{\pm\text{0.74}}$ 7.95±${}_{\pm\text{1.12}}$ 7.55±${}_{\pm\text{1.32}}$ 14.24±${}_{\pm\text{1.48}}$ 7.42±${}_{\pm\text{1.17}}$ 8.79±${}_{\pm\text{0.93}}$ N/A 15.16±${}_{\pm\text{1.92}}$
200 15.84±${}_{\pm\text{1.74}}$ 19.81±${}_{\pm\text{1.73}}$ 9.79±${}_{\pm\text{1.14}}$ 11.21±${}_{\pm\text{1.45}}$ 12.24±${}_{\pm\text{1.65}}$ 16.30±${}_{\pm\text{1.54}}$ 10.96±${}_{\pm\text{1.43}}$ 12.86±${}_{\pm\text{1.50}}$ N/A 18.05±${}_{\pm\text{1.65}}$
texture 50 62.96±${}_{\pm\text{2.49}}$ 78.80±${}_{\pm\text{2.75}}$ 55.51±${}_{\pm\text{3.69}}$ 61.86±${}_{\pm\text{4.48}}$ 62.08±${}_{\pm\text{3.17}}$ 61.91±${}_{\pm\text{2.24}}$ 62.67±${}_{\pm\text{2.29}}$ 56.81±${}_{\pm\text{2.98}}$ N/A 75.57±${}_{\pm\text{2.67}}$
100 77.16±${}_{\pm\text{1.25}}$ 86.15±${}_{\pm\text{2.62}}$ 69.54±${}_{\pm\text{2.66}}$ 76.53±${}_{\pm\text{2.22}}$ 76.85±${}_{\pm\text{1.56}}$ 77.77±${}_{\pm\text{1.80}}$ 76.70±${}_{\pm\text{2.05}}$ 72.64±${}_{\pm\text{1.81}}$ N/A 84.83±${}_{\pm\text{1.67}}$
200 85.34±${}_{\pm\text{1.18}}$ 89.07±${}_{\pm\text{1.74}}$ 81.70±${}_{\pm\text{1.32}}$ 85.46±${}_{\pm\text{1.28}}$ 84.62±${}_{\pm\text{1.09}}$ 85.94±${}_{\pm\text{1.35}}$ 85.11±${}_{\pm\text{1.20}}$ 84.72±${}_{\pm\text{0.80}}$ N/A 89.48±${}_{\pm\text{2.01}}$
500 91.40±${}_{\pm\text{1.60}}$ 93.14±${}_{\pm\text{1.28}}$ 89.88±${}_{\pm\text{1.44}}$ 91.40±${}_{\pm\text{1.55}}$ 91.34±${}_{\pm\text{1.60}}$ 92.31±${}_{\pm\text{1.60}}$ 91.46±${}_{\pm\text{1.51}}$ 91.91±${}_{\pm\text{1.63}}$ N/A 93.46±${}_{\pm\text{0.66}}$
Average rank 5.15±${}_{\pm\text{2.06}}$ 2.70±${}_{\pm\text{1.97}}$ 7.67±${}_{\pm\text{2.10}}$ 7.03±${}_{\pm\text{1.55}}$ 7.27±${}_{\pm\text{1.68}}$ 4.12±${}_{\pm\text{2.34}}$ 6.82±${}_{\pm\text{1.76}}$ 8.42±${}_{\pm\text{2.33}}$ 3.67±${}_{\pm\text{1.96}}$ 2.15±${}_{\pm\text{1.75}}$
Table 9: Classification accuracy (%) of MLP, comparing data augmentation on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes that a specific generator was not applicable or the downstream predictor failed to converge, and the rank is computed with the mean balanced accuracy of other methods. We bold the highest accuracy for each dataset of different sample sizes. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ Baseline (Real data) SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 35.12±${}_{\pm\text{2.59}}$ N/A 21.89±${}_{\pm\text{3.62}}$ 26.95±${}_{\pm\text{4.13}}$ 24.56±${}_{\pm\text{5.06}}$ 27.30±${}_{\pm\text{3.23}}$ 27.96±${}_{\pm\text{4.51}}$ 27.09±${}_{\pm\text{3.47}}$ 36.19±${}_{\pm\text{2.84}}$ 36.26±${}_{\pm\text{2.65}}$
50 58.11±${}_{\pm\text{4.13}}$ 57.24±${}_{\pm\text{4.60}}$ 40.77±${}_{\pm\text{3.59}}$ 43.04±${}_{\pm\text{5.43}}$ 44.78±${}_{\pm\text{3.92}}$ 49.43±${}_{\pm\text{2.51}}$ 46.84±${}_{\pm\text{5.08}}$ 44.78±${}_{\pm\text{2.67}}$ 58.62±${}_{\pm\text{4.41}}$ 58.75±${}_{\pm\text{4.48}}$
100 76.82±${}_{\pm\text{3.33}}$ 76.78±${}_{\pm\text{3.10}}$ 62.00±${}_{\pm\text{3.21}}$ 64.14±${}_{\pm\text{3.19}}$ 63.24±${}_{\pm\text{4.05}}$ 69.20±${}_{\pm\text{2.75}}$ 65.08±${}_{\pm\text{2.67}}$ 62.45±${}_{\pm\text{3.22}}$ 77.84±${}_{\pm\text{3.49}}$ 77.63±${}_{\pm\text{3.69}}$
200 89.53±${}_{\pm\text{2.34}}$ 90.28±${}_{\pm\text{2.13}}$ 80.28±${}_{\pm\text{3.49}}$ 81.94±${}_{\pm\text{2.56}}$ 81.85±${}_{\pm\text{2.46}}$ 85.48±${}_{\pm\text{2.07}}$ 82.57±${}_{\pm\text{2.19}}$ 78.04±${}_{\pm\text{2.66}}$ 90.74±${}_{\pm\text{2.13}}$ 90.48±${}_{\pm\text{2.06}}$
500 98.23±${}_{\pm\text{0.91}}$ 98.25±${}_{\pm\text{0.78}}$ 95.08±${}_{\pm\text{1.34}}$ 95.74±${}_{\pm\text{0.95}}$ 96.15±${}_{\pm\text{1.09}}$ 96.01±${}_{\pm\text{0.86}}$ 96.50±${}_{\pm\text{0.87}}$ 96.23±${}_{\pm\text{1.67}}$ 98.52±${}_{\pm\text{0.80}}$ 98.50±${}_{\pm\text{0.70}}$
fourier 20 33.66±${}_{\pm\text{3.92}}$ N/A 23.20±${}_{\pm\text{5.54}}$ 17.08±${}_{\pm\text{2.75}}$ 19.40±${}_{\pm\text{4.03}}$ 18.32±${}_{\pm\text{3.82}}$ 23.26±${}_{\pm\text{4.21}}$ 19.64±${}_{\pm\text{2.40}}$ 37.00±${}_{\pm\text{2.85}}$ 35.02±${}_{\pm\text{3.77}}$
50 53.72±${}_{\pm\text{1.67}}$ 53.02±${}_{\pm\text{1.96}}$ 37.16±${}_{\pm\text{3.08}}$ 37.60±${}_{\pm\text{4.52}}$ 35.14±${}_{\pm\text{2.44}}$ 40.90±${}_{\pm\text{2.80}}$ 42.82±${}_{\pm\text{2.83}}$ 32.66±${}_{\pm\text{5.19}}$ 55.40±${}_{\pm\text{2.23}}$ 55.34±${}_{\pm\text{1.40}}$
100 62.78±${}_{\pm\text{1.60}}$ 61.44±${}_{\pm\text{2.74}}$ 43.68±${}_{\pm\text{3.15}}$ 48.80±${}_{\pm\text{2.66}}$ 46.18±${}_{\pm\text{3.96}}$ 56.52±${}_{\pm\text{5.04}}$ 52.50±${}_{\pm\text{2.66}}$ 37.74±${}_{\pm\text{2.99}}$ 63.00±${}_{\pm\text{1.95}}$ 63.54±${}_{\pm\text{1.83}}$
200 70.18±${}_{\pm\text{1.85}}$ 70.06±${}_{\pm\text{2.10}}$ 58.90±${}_{\pm\text{2.56}}$ 62.36±${}_{\pm\text{2.86}}$ 58.40±${}_{\pm\text{2.53}}$ 70.08±${}_{\pm\text{1.90}}$ 62.14±${}_{\pm\text{2.00}}$ 50.92±${}_{\pm\text{5.13}}$ 71.49±${}_{\pm\text{1.41}}$ 71.36±${}_{\pm\text{1.36}}$
500 77.94±${}_{\pm\text{1.65}}$ 77.18±${}_{\pm\text{1.35}}$ 72.14±${}_{\pm\text{1.79}}$ 74.30±${}_{\pm\text{1.65}}$ 71.38±${}_{\pm\text{1.54}}$ 77.78±${}_{\pm\text{1.26}}$ 74.32±${}_{\pm\text{1.56}}$ 67.28±${}_{\pm\text{2.91}}$ 78.34±${}_{\pm\text{1.72}}$ 79.30±${}_{\pm\text{0.99}}$
biodeg 20 71.31±${}_{\pm\text{5.13}}$ 68.84±${}_{\pm\text{5.95}}$ 66.64±${}_{\pm\text{8.27}}$ 62.11±${}_{\pm\text{4.95}}$ 62.61±${}_{\pm\text{6.78}}$ 52.96±${}_{\pm\text{4.22}}$ 62.06±${}_{\pm\text{3.69}}$ 65.81±${}_{\pm\text{7.24}}$ 72.04±${}_{\pm\text{5.12}}$ 72.09±${}_{\pm\text{4.81}}$
50 76.73±${}_{\pm\text{3.16}}$ 74.97±${}_{\pm\text{2.51}}$ 72.02±${}_{\pm\text{4.74}}$ 71.83±${}_{\pm\text{3.17}}$ 67.86±${}_{\pm\text{6.02}}$ 69.92±${}_{\pm\text{4.83}}$ 74.03±${}_{\pm\text{3.05}}$ 71.01±${}_{\pm\text{2.78}}$ 77.17±${}_{\pm\text{2.93}}$ 77.11±${}_{\pm\text{3.20}}$
100 79.13±${}_{\pm\text{1.91}}$ 78.20±${}_{\pm\text{1.68}}$ 76.78±${}_{\pm\text{2.79}}$ 77.85±${}_{\pm\text{2.73}}$ 76.01±${}_{\pm\text{2.88}}$ 76.74±${}_{\pm\text{3.62}}$ 76.08±${}_{\pm\text{2.39}}$ 76.24±${}_{\pm\text{2.45}}$ 78.23±${}_{\pm\text{2.29}}$ 79.08±${}_{\pm\text{2.03}}$
200 82.39±${}_{\pm\text{1.48}}$ 81.70±${}_{\pm\text{1.22}}$ 80.43±${}_{\pm\text{2.02}}$ 79.96±${}_{\pm\text{2.35}}$ 79.92±${}_{\pm\text{1.55}}$ 80.51±${}_{\pm\text{1.26}}$ 79.59±${}_{\pm\text{1.72}}$ 80.34±${}_{\pm\text{1.93}}$ 81.74±${}_{\pm\text{1.36}}$ 82.24±${}_{\pm\text{1.54}}$
500 84.50±${}_{\pm\text{0.61}}$ 84.50±${}_{\pm\text{0.81}}$ 83.78±${}_{\pm\text{1.51}}$ 83.67±${}_{\pm\text{0.81}}$ 84.13±${}_{\pm\text{1.20}}$ ${}_{\pm\text{0.84}}$ 83.76±${}_{\pm\text{1.27}}$ 82.97±${}_{\pm\text{1.21}}$ 84.37±${}_{\pm\text{0.48}}$ 84.14±${}_{\pm\text{0.59}}$
steel 20 62.35±${}_{\pm\text{6.30}}$ 60.34±${}_{\pm\text{5.73}}$ 61.63±${}_{\pm\text{8.82}}$ 59.09±${}_{\pm\text{4.25}}$ 56.99±${}_{\pm\text{7.13}}$ 60.67±${}_{\pm\text{9.18}}$ 55.23±${}_{\pm\text{3.92}}$ 55.78±${}_{\pm\text{3.01}}$ 64.49±${}_{\pm\text{6.03}}$ 64.22±${}_{\pm\text{5.89}}$
50 79.65±${}_{\pm\text{5.53}}$ 68.18±${}_{\pm\text{3.16}}$ 69.01±${}_{\pm\text{3.48}}$ 70.30±${}_{\pm\text{4.77}}$ 66.96±${}_{\pm\text{5.12}}$ 84.04±${}_{\pm\text{8.03}}$ 64.79±${}_{\pm\text{4.07}}$ 58.95±${}_{\pm\text{1.66}}$ 82.72±${}_{\pm\text{6.02}}$ 82.15±${}_{\pm\text{5.78}}$
100 92.18±${}_{\pm\text{2.93}}$ 78.44±${}_{\pm\text{3.67}}$ 76.37±${}_{\pm\text{3.06}}$ 76.92±${}_{\pm\text{3.63}}$ 76.50±${}_{\pm\text{3.18}}$ 95.83±${}_{\pm\text{1.90}}$ 71.85±${}_{\pm\text{3.20}}$ 67.35±${}_{\pm\text{2.51}}$ 95.16±${}_{\pm\text{3.13}}$ 95.41±${}_{\pm\text{3.23}}$
200 97.31±${}_{\pm\text{1.63}}$ 83.93±${}_{\pm\text{1.83}}$ 82.42±${}_{\pm\text{2.75}}$ 84.70±${}_{\pm\text{2.54}}$ 84.06±${}_{\pm\text{4.46}}$ 98.75±${}_{\pm\text{0.68}}$ 79.66±${}_{\pm\text{3.26}}$ 78.36±${}_{\pm\text{3.86}}$ 98.83±${}_{\pm\text{0.80}}$ 98.84±${}_{\pm\text{0.67}}$
500 99.78±${}_{\pm\text{0.30}}$ 91.37±${}_{\pm\text{2.26}}$ 93.08±${}_{\pm\text{1.82}}$ 94.82±${}_{\pm\text{1.54}}$ 93.99±${}_{\pm\text{1.83}}$ 99.47±${}_{\pm\text{0.44}}$ 90.34±${}_{\pm\text{2.19}}$ 96.06±${}_{\pm\text{1.03}}$ 99.81±${}_{\pm\text{0.24}}$ 99.81±${}_{\pm\text{0.24}}$
stock 20 83.56±${}_{\pm\text{3.89}}$ 83.62±${}_{\pm\text{4.06}}$ 77.25±${}_{\pm\text{5.02}}$ 69.90±${}_{\pm\text{9.27}}$ 72.85±${}_{\pm\text{7.77}}$ 80.60±${}_{\pm\text{5.73}}$ 69.30±${}_{\pm\text{6.76}}$ ${}_{\pm\text{3.38}}$ 83.81±${}_{\pm\text{3.92}}$ 83.89±${}_{\pm\text{4.05}}$
50 89.57±${}_{\pm\text{2.01}}$ 89.71±${}_{\pm\text{2.21}}$ 82.62±${}_{\pm\text{3.49}}$ 79.35±${}_{\pm\text{2.71}}$ 81.52±${}_{\pm\text{1.99}}$ 88.48±${}_{\pm\text{2.18}}$ 83.36±${}_{\pm\text{2.30}}$ 88.38±${}_{\pm\text{2.41}}$ 90.23±${}_{\pm\text{2.02}}$ 90.38±${}_{\pm\text{2.17}}$
100 90.63±${}_{\pm\text{0.83}}$ 91.17±${}_{\pm\text{0.83}}$ 88.37±${}_{\pm\text{2.40}}$ 86.60±${}_{\pm\text{3.27}}$ 83.19±${}_{\pm\text{3.77}}$ 90.65±${}_{\pm\text{1.39}}$ 88.64±${}_{\pm\text{1.15}}$ 91.61±${}_{\pm\text{0.70}}$ 91.70±${}_{\pm\text{1.17}}$ 91.75±${}_{\pm\text{0.97}}$
200 91.25±${}_{\pm\text{0.74}}$ 92.58±${}_{\pm\text{0.91}}$ 91.27±${}_{\pm\text{0.95}}$ 90.34±${}_{\pm\text{1.60}}$ 89.32±${}_{\pm\text{2.45}}$ 92.19±${}_{\pm\text{0.78}}$ 90.37±${}_{\pm\text{1.32}}$ 91.89±${}_{\pm\text{1.32}}$ 92.91±${}_{\pm\text{0.62}}$ 92.47±${}_{\pm\text{0.63}}$
More than 10 classes energy 50 24.79±${}_{\pm\text{1.76}}$ N/A 12.51±${}_{\pm\text{2.89}}$ 12.61±${}_{\pm\text{3.45}}$ 8.43±${}_{\pm\text{2.11}}$ 10.91±${}_{\pm\text{2.11}}$ 12.45±${}_{\pm\text{1.87}}$ 20.42±${}_{\pm\text{4.41}}$ N/A 24.04±${}_{\pm\text{1.39}}$
100 26.86±${}_{\pm\text{1.51}}$ N/A 16.20±${}_{\pm\text{2.12}}$ 15.78±${}_{\pm\text{2.50}}$ 15.70±${}_{\pm\text{2.44}}$ 17.75±${}_{\pm\text{3.18}}$ 18.16±${}_{\pm\text{2.46}}$ 20.72±${}_{\pm\text{2.99}}$ N/A 29.30±${}_{\pm\text{2.32}}$
200 33.36±${}_{\pm\text{2.98}}$ N/A 22.30±${}_{\pm\text{2.44}}$ 23.00±${}_{\pm\text{4.24}}$ 23.53±${}_{\pm\text{1.84}}$ 26.12±${}_{\pm\text{1.78}}$ 26.28±${}_{\pm\text{3.03}}$ 33.03±${}_{\pm\text{3.38}}$ N/A 41.27±${}_{\pm\text{2.93}}$
collins 100 14.16±${}_{\pm\text{1.31}}$ N/A 9.24±${}_{\pm\text{1.71}}$ 9.16±${}_{\pm\text{1.57}}$ 9.04±${}_{\pm\text{1.79}}$ 14.03±${}_{\pm\text{1.24}}$ 8.59±${}_{\pm\text{1.84}}$ 10.81±${}_{\pm\text{1.68}}$ N/A 14.07±${}_{\pm\text{1.58}}$
200 19.35±${}_{\pm\text{1.24}}$ 19.06±${}_{\pm\text{1.49}}$ 14.62±${}_{\pm\text{2.00}}$ 13.17±${}_{\pm\text{0.94}}$ 12.38±${}_{\pm\text{1.56}}$ 18.63±${}_{\pm\text{1.56}}$ 12.48±${}_{\pm\text{1.91}}$ ${}_{\pm\text{1.76}}$ N/A 19.53±${}_{\pm\text{1.44}}$
texture 50 84.50±${}_{\pm\text{2.81}}$ 84.12±${}_{\pm\text{3.02}}$ 62.92±${}_{\pm\text{4.09}}$ 67.69±${}_{\pm\text{5.49}}$ 61.99±${}_{\pm\text{3.58}}$ 69.69±${}_{\pm\text{4.62}}$ 64.45±${}_{\pm\text{7.84}}$ 69.68±${}_{\pm\text{3.53}}$ N/A 85.51±${}_{\pm\text{2.89}}$
100 91.50±${}_{\pm\text{1.34}}$ ${}_{\pm\text{1.59}}$ 74.53±${}_{\pm\text{3.39}}$ 79.96±${}_{\pm\text{4.37}}$ 80.36±${}_{\pm\text{4.58}}$ 85.42±${}_{\pm\text{2.74}}$ 80.23±${}_{\pm\text{2.18}}$ 85.59±${}_{\pm\text{1.49}}$ N/A 92.17±${}_{\pm\text{1.31}}$
200 93.81±${}_{\pm\text{1.35}}$ 94.18±${}_{\pm\text{1.26}}$ 86.57±${}_{\pm\text{2.33}}$ 90.68±${}_{\pm\text{1.55}}$ 88.97±${}_{\pm\text{2.12}}$ 90.10±${}_{\pm\text{2.26}}$ 89.14±${}_{\pm\text{1.85}}$ 91.66±${}_{\pm\text{1.43}}$ N/A 94.35±${}_{\pm\text{1.57}}$
500 96.55±${}_{\pm\text{0.63}}$ 97.21±${}_{\pm\text{0.40}}$ 94.66±${}_{\pm\text{1.17}}$ 96.27±${}_{\pm\text{0.74}}$ 94.34±${}_{\pm\text{1.36}}$ 94.72±${}_{\pm\text{0.61}}$ 95.83±${}_{\pm\text{1.03}}$ 96.49±${}_{\pm\text{0.48}}$ N/A 97.13±${}_{\pm\text{0.53}}$
Average rank 3.00±${}_{\pm\text{1.32}}$ 4.06±${}_{\pm\text{1.62}}$ 7.82±${}_{\pm\text{1.63}}$ 7.48±${}_{\pm\text{1.50}}$ 8.45±${}_{\pm\text{1.33}}$ 5.55±${}_{\pm\text{2.14}}$ 7.48±${}_{\pm\text{1.72}}$ 6.82±${}_{\pm\text{2.57}}$ 2.67±${}_{\pm\text{1.69}}$ 1.67±${}_{\pm\text{0.74}}$
Table 10: Classification accuracy (%) of RF, comparing data augmentation on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes that a specific generator was not applicable or the downstream predictor failed to converge, and the rank is computed with the mean balanced accuracy of other methods. We bold the highest accuracy for each dataset of different sample sizes. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ Baseline (Real data) SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 28.52±${}_{\pm\text{2.19}}$ N/A 22.74±${}_{\pm\text{4.23}}$ 24.94±${}_{\pm\text{5.15}}$ ${}_{\pm\text{2.46}}$ 29.62±${}_{\pm\text{4.06}}$ 27.69±${}_{\pm\text{3.73}}$ 25.76±${}_{\pm\text{1.62}}$ 32.04±${}_{\pm\text{2.40}}$ 34.19±${}_{\pm\text{2.21}}$
50 53.40±${}_{\pm\text{3.26}}$ 55.69±${}_{\pm\text{2.61}}$ 46.95±${}_{\pm\text{3.13}}$ 43.91±${}_{\pm\text{4.98}}$ 43.28±${}_{\pm\text{4.07}}$ 47.93±${}_{\pm\text{4.49}}$ 44.48±${}_{\pm\text{3.35}}$ 47.25±${}_{\pm\text{4.81}}$ 54.29±${}_{\pm\text{2.57}}$ 56.85±${}_{\pm\text{2.49}}$
100 68.13±${}_{\pm\text{3.19}}$ 72.89±${}_{\pm\text{2.60}}$ 63.24±${}_{\pm\text{1.78}}$ 61.19±${}_{\pm\text{2.10}}$ 59.64±${}_{\pm\text{3.48}}$ 65.19±${}_{\pm\text{3.22}}$ 60.05±${}_{\pm\text{2.88}}$ 65.05±${}_{\pm\text{3.00}}$ 71.47±${}_{\pm\text{3.61}}$ 72.57±${}_{\pm\text{2.50}}$
200 80.34±${}_{\pm\text{2.35}}$ 83.60±${}_{\pm\text{2.71}}$ 78.51±${}_{\pm\text{2.58}}$ 75.84±${}_{\pm\text{1.61}}$ 76.84±${}_{\pm\text{2.35}}$ 78.74±${}_{\pm\text{2.24}}$ 75.61±${}_{\pm\text{2.90}}$ 79.44±${}_{\pm\text{2.73}}$ 83.36±${}_{\pm\text{2.40}}$ 84.30±${}_{\pm\text{1.97}}$
500 93.01±${}_{\pm\text{1.12}}$ 93.82±${}_{\pm\text{0.67}}$ 92.86±${}_{\pm\text{1.66}}$ 91.37±${}_{\pm\text{1.25}}$ 93.00±${}_{\pm\text{1.08}}$ 92.93±${}_{\pm\text{0.97}}$ 92.38±${}_{\pm\text{0.90}}$ 92.95±${}_{\pm\text{0.92}}$ 94.49±${}_{\pm\text{1.16}}$ 93.94±${}_{\pm\text{1.23}}$
fourier 20 35.10±${}_{\pm\text{4.56}}$ N/A 19.06±${}_{\pm\text{3.91}}$ 17.52±${}_{\pm\text{2.84}}$ 20.78±${}_{\pm\text{2.54}}$ 16.98±${}_{\pm\text{2.31}}$ 23.78±${}_{\pm\text{3.12}}$ 19.00±${}_{\pm\text{2.93}}$ 34.88±${}_{\pm\text{5.93}}$ 38.60±${}_{\pm\text{5.66}}$
50 64.10±${}_{\pm\text{3.80}}$ 64.76±${}_{\pm\text{4.00}}$ 37.20±${}_{\pm\text{3.35}}$ 32.82±${}_{\pm\text{4.56}}$ 37.78±${}_{\pm\text{3.11}}$ 51.76±${}_{\pm\text{3.50}}$ 47.22±${}_{\pm\text{4.35}}$ 53.86±${}_{\pm\text{3.61}}$ 66.92±${}_{\pm\text{3.05}}$ 66.26±${}_{\pm\text{3.16}}$
100 73.86±${}_{\pm\text{3.06}}$ 73.78±${}_{\pm\text{3.22}}$ 64.40±${}_{\pm\text{2.51}}$ 60.82±${}_{\pm\text{3.71}}$ 51.64±${}_{\pm\text{4.16}}$ 66.14±${}_{\pm\text{1.91}}$ 58.62±${}_{\pm\text{3.73}}$ 68.16±${}_{\pm\text{3.12}}$ 73.13±${}_{\pm\text{2.70}}$ 74.84±${}_{\pm\text{3.10}}$
200 78.54±${}_{\pm\text{2.15}}$ 79.18±${}_{\pm\text{1.92}}$ 74.86±${}_{\pm\text{1.60}}$ 74.26±${}_{\pm\text{2.20}}$ 69.36±${}_{\pm\text{2.61}}$ 76.42±${}_{\pm\text{1.95}}$ 72.88±${}_{\pm\text{1.22}}$ 76.64±${}_{\pm\text{1.99}}$ 82.20±${}_{\pm\text{0.85}}$ 79.18±${}_{\pm\text{2.08}}$
500 81.84±${}_{\pm\text{1.01}}$ 82.14±${}_{\pm\text{1.49}}$ 81.02±${}_{\pm\text{1.59}}$ 81.18±${}_{\pm\text{1.43}}$ 80.08±${}_{\pm\text{1.62}}$ 81.26±${}_{\pm\text{1.40}}$ 80.28±${}_{\pm\text{1.54}}$ 80.62±${}_{\pm\text{1.52}}$ 81.45±${}_{\pm\text{1.45}}$ 83.40±${}_{\pm\text{1.24}}$
biodeg 20 61.11±${}_{\pm\text{7.87}}$ 68.38±${}_{\pm\text{5.90}}$ 65.44±${}_{\pm\text{8.89}}$ 56.29±${}_{\pm\text{7.96}}$ 58.19±${}_{\pm\text{6.60}}$ 52.90±${}_{\pm\text{4.74}}$ 62.33±${}_{\pm\text{6.14}}$ 63.52±${}_{\pm\text{7.29}}$ 67.15±${}_{\pm\text{5.74}}$ 67.82±${}_{\pm\text{5.13}}$
50 68.38±${}_{\pm\text{4.82}}$ 70.64±${}_{\pm\text{3.44}}$ 71.77±${}_{\pm\text{2.99}}$ 66.78±${}_{\pm\text{4.89}}$ 61.39±${}_{\pm\text{4.94}}$ 63.98±${}_{\pm\text{3.65}}$ 68.78±${}_{\pm\text{5.22}}$ 70.34±${}_{\pm\text{3.39}}$ 71.38±${}_{\pm\text{3.60}}$ 72.12±${}_{\pm\text{3.29}}$
100 73.19±${}_{\pm\text{2.46}}$ 75.36±${}_{\pm\text{2.56}}$ 74.98±${}_{\pm\text{2.58}}$ 72.68±${}_{\pm\text{2.98}}$ 69.62±${}_{\pm\text{3.53}}$ 73.11±${}_{\pm\text{2.39}}$ 72.16±${}_{\pm\text{2.58}}$ 74.22±${}_{\pm\text{2.32}}$ 75.85±${}_{\pm\text{1.56}}$ 75.65±${}_{\pm\text{1.53}}$
200 77.85±${}_{\pm\text{2.72}}$ 78.86±${}_{\pm\text{1.97}}$ 76.42±${}_{\pm\text{2.25}}$ 76.68±${}_{\pm\text{2.77}}$ 73.43±${}_{\pm\text{3.01}}$ 76.16±${}_{\pm\text{2.00}}$ 75.79±${}_{\pm\text{2.49}}$ 77.42±${}_{\pm\text{2.24}}$ 79.68±${}_{\pm\text{1.74}}$ 79.22±${}_{\pm\text{1.70}}$
500 81.42±${}_{\pm\text{0.73}}$ 82.03±${}_{\pm\text{1.02}}$ 81.88±${}_{\pm\text{0.87}}$ 81.71±${}_{\pm\text{1.54}}$ 80.50±${}_{\pm\text{1.21}}$ 81.43±${}_{\pm\text{1.26}}$ 81.34±${}_{\pm\text{1.58}}$ 81.94±${}_{\pm\text{0.85}}$ 82.38±${}_{\pm\text{1.35}}$ 82.10±${}_{\pm\text{1.31}}$
steel 20 52.77±${}_{\pm\text{1.60}}$ 56.16±${}_{\pm\text{4.50}}$ 57.23±${}_{\pm\text{3.97}}$ 54.65±${}_{\pm\text{3.40}}$ 53.75±${}_{\pm\text{3.49}}$ 51.70±${}_{\pm\text{1.66}}$ 54.09±${}_{\pm\text{4.36}}$ 55.50±${}_{\pm\text{2.97}}$ 57.04±${}_{\pm\text{3.07}}$ 57.41±${}_{\pm\text{2.67}}$
50 59.75±${}_{\pm\text{3.11}}$ 62.12±${}_{\pm\text{2.46}}$ 60.65±${}_{\pm\text{1.96}}$ 58.09±${}_{\pm\text{1.75}}$ 54.69±${}_{\pm\text{2.44}}$ 58.14±${}_{\pm\text{4.21}}$ 57.67±${}_{\pm\text{2.52}}$ 60.34±${}_{\pm\text{2.90}}$ 65.07±${}_{\pm\text{3.11}}$ 67.74±${}_{\pm\text{3.36}}$
100 64.97±${}_{\pm\text{2.05}}$ 69.08±${}_{\pm\text{3.62}}$ 64.46±${}_{\pm\text{4.17}}$ 61.62±${}_{\pm\text{1.98}}$ 58.43±${}_{\pm\text{2.46}}$ 60.53±${}_{\pm\text{3.64}}$ 62.71±${}_{\pm\text{3.43}}$ 63.07±${}_{\pm\text{2.23}}$ 73.28±${}_{\pm\text{3.39}}$ 79.63±${}_{\pm\text{3.41}}$
200 75.45±${}_{\pm\text{3.26}}$ 74.71±${}_{\pm\text{3.79}}$ 71.45±${}_{\pm\text{2.18}}$ 68.52±${}_{\pm\text{3.80}}$ 62.15±${}_{\pm\text{2.80}}$ 68.10±${}_{\pm\text{3.59}}$ 67.61±${}_{\pm\text{1.81}}$ 67.36±${}_{\pm\text{1.63}}$ 85.12±${}_{\pm\text{4.44}}$ 88.85±${}_{\pm\text{5.10}}$
500 90.93±${}_{\pm\text{2.83}}$ 85.37±${}_{\pm\text{2.36}}$ 85.63±${}_{\pm\text{3.14}}$ 84.51±${}_{\pm\text{3.22}}$ 76.12±${}_{\pm\text{2.70}}$ 89.19±${}_{\pm\text{3.20}}$ 81.44±${}_{\pm\text{2.64}}$ 80.35±${}_{\pm\text{3.54}}$ 94.35±${}_{\pm\text{1.34}}$ 95.90±${}_{\pm\text{1.06}}$
stock 20 79.47±${}_{\pm\text{5.83}}$ 81.99±${}_{\pm\text{4.49}}$ 77.94±${}_{\pm\text{5.21}}$ 72.53±${}_{\pm\text{7.16}}$ 73.20±${}_{\pm\text{9.98}}$ 80.99±${}_{\pm\text{7.01}}$ 72.57±${}_{\pm\text{8.76}}$ 78.10±${}_{\pm\text{5.91}}$ 83.96±${}_{\pm\text{5.57}}$ 84.73±${}_{\pm\text{3.46}}$
50 87.57±${}_{\pm\text{2.60}}$ 89.69±${}_{\pm\text{1.99}}$ 86.62±${}_{\pm\text{3.44}}$ 83.75±${}_{\pm\text{4.32}}$ 84.28±${}_{\pm\text{2.98}}$ 88.69±${}_{\pm\text{2.11}}$ 84.92±${}_{\pm\text{2.09}}$ 88.65±${}_{\pm\text{2.55}}$ 89.35±${}_{\pm\text{2.18}}$ 89.99±${}_{\pm\text{2.63}}$
100 91.44±${}_{\pm\text{1.59}}$ 91.47±${}_{\pm\text{2.16}}$ 91.07±${}_{\pm\text{2.08}}$ 89.82±${}_{\pm\text{2.69}}$ 89.33±${}_{\pm\text{1.92}}$ 91.33±${}_{\pm\text{2.07}}$ 90.48±${}_{\pm\text{2.38}}$ 92.00±${}_{\pm\text{2.36}}$ 92.07±${}_{\pm\text{1.22}}$ 92.17±${}_{\pm\text{1.24}}$
200 93.52±${}_{\pm\text{0.80}}$ 93.94±${}_{\pm\text{1.09}}$ 93.35±${}_{\pm\text{1.05}}$ 92.62±${}_{\pm\text{1.02}}$ 92.77±${}_{\pm\text{1.25}}$ 93.65±${}_{\pm\text{1.08}}$ 93.08±${}_{\pm\text{0.53}}$ 93.87±${}_{\pm\text{1.25}}$ 93.65±${}_{\pm\text{1.02}}$ 93.67±${}_{\pm\text{1.07}}$
More than 10 classes energy 50 18.96±${}_{\pm\text{1.40}}$ N/A 16.63±${}_{\pm\text{2.27}}$ 15.66±${}_{\pm\text{2.43}}$ 14.81±${}_{\pm\text{3.16}}$ 14.49±${}_{\pm\text{1.26}}$ 15.05±${}_{\pm\text{3.06}}$ 15.58±${}_{\pm\text{3.26}}$ N/A 27.74±${}_{\pm\text{3.71}}$
100 30.85±${}_{\pm\text{2.19}}$ N/A 24.59±${}_{\pm\text{2.27}}$ 28.59±${}_{\pm\text{2.63}}$ 27.59±${}_{\pm\text{2.86}}$ 27.23±${}_{\pm\text{2.39}}$ 27.99±${}_{\pm\text{2.18}}$ 25.43±${}_{\pm\text{2.46}}$ N/A 41.03±${}_{\pm\text{2.24}}$
200 45.80±${}_{\pm\text{2.32}}$ N/A 42.10±${}_{\pm\text{2.57}}$ 41.69±${}_{\pm\text{3.84}}$ 44.41±${}_{\pm\text{2.51}}$ 44.58±${}_{\pm\text{1.37}}$ 41.33±${}_{\pm\text{3.90}}$ 44.64±${}_{\pm\text{2.54}}$ N/A 53.87±${}_{\pm\text{2.81}}$
collins 100 10.41±${}_{\pm\text{1.61}}$ N/A 6.75±${}_{\pm\text{0.69}}$ 8.23±${}_{\pm\text{1.76}}$ 7.34±${}_{\pm\text{1.46}}$ 12.84±${}_{\pm\text{1.61}}$ 6.73±${}_{\pm\text{1.36}}$ 8.43±${}_{\pm\text{0.94}}$ N/A 13.35±${}_{\pm\text{1.49}}$
200 13.75±${}_{\pm\text{1.12}}$ 17.56±${}_{\pm\text{1.79}}$ 10.51±${}_{\pm\text{1.41}}$ 11.00±${}_{\pm\text{1.37}}$ 9.85±${}_{\pm\text{1.38}}$ 15.15±${}_{\pm\text{1.22}}$ 9.90±${}_{\pm\text{0.72}}$ 13.40±${}_{\pm\text{1.09}}$ N/A 16.51±${}_{\pm\text{1.53}}$
texture 50 71.27±${}_{\pm\text{1.99}}$ 71.17±${}_{\pm\text{3.89}}$ 57.41±${}_{\pm\text{3.33}}$ 62.78±${}_{\pm\text{4.21}}$ 65.24±${}_{\pm\text{4.52}}$ 69.45±${}_{\pm\text{2.15}}$ 62.93±${}_{\pm\text{4.84}}$ 64.33±${}_{\pm\text{3.57}}$ N/A 75.79±${}_{\pm\text{3.07}}$
100 80.40±${}_{\pm\text{2.45}}$ 80.38±${}_{\pm\text{2.67}}$ 65.63±${}_{\pm\text{4.21}}$ 75.38±${}_{\pm\text{3.99}}$ 77.67±${}_{\pm\text{2.62}}$ 79.31±${}_{\pm\text{1.88}}$ 75.98±${}_{\pm\text{2.56}}$ 77.30±${}_{\pm\text{2.44}}$ N/A 82.30±${}_{\pm\text{2.21}}$
200 84.00±${}_{\pm\text{1.56}}$ 85.12±${}_{\pm\text{3.07}}$ 76.98±${}_{\pm\text{2.25}}$ 84.44±${}_{\pm\text{2.41}}$ 85.30±${}_{\pm\text{1.96}}$ 84.00±${}_{\pm\text{1.20}}$ 83.70±${}_{\pm\text{2.05}}$ 80.02±${}_{\pm\text{1.60}}$ N/A 85.92±${}_{\pm\text{2.18}}$
500 89.43±${}_{\pm\text{0.80}}$ 90.17±${}_{\pm\text{1.25}}$ 88.97±${}_{\pm\text{1.44}}$ 90.00±${}_{\pm\text{1.66}}$ 89.99±${}_{\pm\text{1.06}}$ 90.17±${}_{\pm\text{1.32}}$ 91.01±${}_{\pm\text{1.32}}$ 88.98±${}_{\pm\text{1.26}}$ N/A 90.77±${}_{\pm\text{1.10}}$
Average rank 4.36±${}_{\pm\text{1.95}}$ 3.02±${}_{\pm\text{1.14}}$ 6.88±${}_{\pm\text{2.25}}$ 7.82±${}_{\pm\text{1.61}}$ 8.45±${}_{\pm\text{1.95}}$ 6.12±${}_{\pm\text{2.23}}$ 7.85±${}_{\pm\text{1.77}}$ 6.00±${}_{\pm\text{1.75}}$ 3.12±${}_{\pm\text{1.75}}$ 1.38±${}_{\pm\text{0.57}}$
Table 11: Classification accuracy (%) of XGBoost, comparing data augmentation on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes that a specific generator was not applicable or the downstream predictor failed to converge, and the rank is computed with the mean balanced accuracy of other methods. We bold the highest accuracy for each dataset of different sample sizes. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ Baseline (Real data) SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 19.70±${}_{\pm\text{6.33}}$ N/A 19.44±${}_{\pm\text{4.11}}$ 17.32±${}_{\pm\text{2.75}}$ 18.11±${}_{\pm\text{3.07}}$ 16.15±${}_{\pm\text{3.80}}$ 20.71±${}_{\pm\text{5.24}}$ 17.40±${}_{\pm\text{4.89}}$ 24.00±${}_{\pm\text{3.64}}$ 24.18±${}_{\pm\text{3.05}}$
50 39.01±${}_{\pm\text{4.92}}$ 37.68±${}_{\pm\text{5.40}}$ 33.07±${}_{\pm\text{4.18}}$ 24.38±${}_{\pm\text{3.45}}$ 23.09±${}_{\pm\text{4.55}}$ 30.87±${}_{\pm\text{5.70}}$ 34.13±${}_{\pm\text{6.45}}$ 33.62±${}_{\pm\text{3.67}}$ 39.78±${}_{\pm\text{6.03}}$ 44.46±${}_{\pm\text{4.97}}$
100 57.59±${}_{\pm\text{3.69}}$ 60.16±${}_{\pm\text{5.75}}$ 49.23±${}_{\pm\text{5.51}}$ 43.33±${}_{\pm\text{7.92}}$ 37.69±${}_{\pm\text{5.96}}$ 48.36±${}_{\pm\text{4.08}}$ 43.97±${}_{\pm\text{5.45}}$ 47.00±${}_{\pm\text{3.29}}$ 53.74±${}_{\pm\text{7.94}}$ 62.77±${}_{\pm\text{5.85}}$
200 74.05±${}_{\pm\text{2.92}}$ 76.90±${}_{\pm\text{4.96}}$ 69.71±${}_{\pm\text{4.28}}$ 67.46±${}_{\pm\text{4.39}}$ 58.29±${}_{\pm\text{7.96}}$ 69.68±${}_{\pm\text{4.23}}$ 63.69±${}_{\pm\text{6.32}}$ 66.09±${}_{\pm\text{4.78}}$ 73.19±${}_{\pm\text{6.06}}$ 79.25±${}_{\pm\text{3.83}}$
500 88.89±${}_{\pm\text{1.71}}$ 90.02±${}_{\pm\text{1.51}}$ 90.10±${}_{\pm\text{1.80}}$ 89.37±${}_{\pm\text{1.81}}$ 86.03±${}_{\pm\text{2.31}}$ 87.29±${}_{\pm\text{2.08}}$ 90.05±${}_{\pm\text{2.70}}$ 85.04±${}_{\pm\text{2.07}}$ 89.66±${}_{\pm\text{1.17}}$ 91.81±${}_{\pm\text{1.44}}$
fourier 20 10.00±${}_{\pm\text{0.00}}$ N/A 14.64±${}_{\pm\text{3.13}}$ 13.58±${}_{\pm\text{2.57}}$ 13.82±${}_{\pm\text{4.14}}$ 11.72±${}_{\pm\text{4.19}}$ 16.38±${}_{\pm\text{3.86}}$ 12.34±${}_{\pm\text{3.59}}$ 23.50±${}_{\pm\text{1.56}}$ 26.78±${}_{\pm\text{4.82}}$
50 42.10±${}_{\pm\text{6.19}}$ 43.40±${}_{\pm\text{5.22}}$ 34.32±${}_{\pm\text{3.98}}$ 24.68±${}_{\pm\text{6.47}}$ 17.66±${}_{\pm\text{4.64}}$ 24.82±${}_{\pm\text{6.35}}$ 27.74±${}_{\pm\text{5.86}}$ 35.42±${}_{\pm\text{7.51}}$ 35.60±${}_{\pm\text{3.11}}$ 45.08±${}_{\pm\text{6.47}}$
100 54.84±${}_{\pm\text{2.78}}$ 52.92±${}_{\pm\text{5.69}}$ 48.22±${}_{\pm\text{3.28}}$ 36.90±${}_{\pm\text{5.15}}$ 30.36±${}_{\pm\text{3.94}}$ 42.46±${}_{\pm\text{4.13}}$ 40.28±${}_{\pm\text{3.41}}$ 48.78±${}_{\pm\text{4.36}}$ 49.80±${}_{\pm\text{1.98}}$ 54.94±${}_{\pm\text{5.72}}$
200 63.88±${}_{\pm\text{3.35}}$ 65.34±${}_{\pm\text{3.57}}$ 58.36±${}_{\pm\text{3.27}}$ 53.20±${}_{\pm\text{5.26}}$ 46.96±${}_{\pm\text{4.58}}$ 61.40±${}_{\pm\text{4.12}}$ 52.10±${}_{\pm\text{3.32}}$ 56.66±${}_{\pm\text{2.67}}$ 66.60±${}_{\pm\text{4.24}}$ 67.68±${}_{\pm\text{3.19}}$
500 74.56±${}_{\pm\text{1.97}}$ 74.18±${}_{\pm\text{2.10}}$ 68.28±${}_{\pm\text{2.82}}$ 67.98±${}_{\pm\text{2.07}}$ 61.24±${}_{\pm\text{2.35}}$ 72.78±${}_{\pm\text{2.54}}$ 67.50±${}_{\pm\text{2.57}}$ 68.28±${}_{\pm\text{3.43}}$ N/A 76.25±${}_{\pm\text{3.18}}$
biodeg 20 62.95±${}_{\pm\text{7.95}}$ 66.51±${}_{\pm\text{5.84}}$ 62.72±${}_{\pm\text{5.69}}$ 55.24±${}_{\pm\text{6.28}}$ 59.20±${}_{\pm\text{7.83}}$ 54.65±${}_{\pm\text{5.56}}$ 62.78±${}_{\pm\text{5.98}}$ 61.09±${}_{\pm\text{10.49}}$ 65.52±${}_{\pm\text{6.08}}$ 66.64±${}_{\pm\text{6.71}}$
50 67.96±${}_{\pm\text{3.45}}$ 67.69±${}_{\pm\text{4.42}}$ 66.22±${}_{\pm\text{5.70}}$ 61.64±${}_{\pm\text{6.73}}$ 60.72±${}_{\pm\text{5.73}}$ 57.48±${}_{\pm\text{8.28}}$ 69.48±${}_{\pm\text{5.35}}$ 65.93±${}_{\pm\text{4.98}}$ 67.76±${}_{\pm\text{4.90}}$ 67.90±${}_{\pm\text{3.27}}$
100 73.88±${}_{\pm\text{2.55}}$ 72.05±${}_{\pm\text{4.75}}$ 72.11±${}_{\pm\text{3.17}}$ 70.41±${}_{\pm\text{3.60}}$ 66.02±${}_{\pm\text{6.25}}$ 69.35±${}_{\pm\text{4.66}}$ 71.11±${}_{\pm\text{3.88}}$ 69.03±${}_{\pm\text{4.33}}$ 72.58±${}_{\pm\text{2.91}}$ 71.05±${}_{\pm\text{5.70}}$
200 76.38±${}_{\pm\text{4.85}}$ 74.98±${}_{\pm\text{3.15}}$ 73.93±${}_{\pm\text{3.29}}$ 75.68±${}_{\pm\text{4.15}}$ 67.82±${}_{\pm\text{3.91}}$ 72.58±${}_{\pm\text{5.07}}$ 74.74±${}_{\pm\text{2.24}}$ 73.84±${}_{\pm\text{3.82}}$ 75.85±${}_{\pm\text{1.80}}$ 76.74±${}_{\pm\text{2.44}}$
500 78.45±${}_{\pm\text{3.37}}$ 79.38±${}_{\pm\text{1.99}}$ 78.88±${}_{\pm\text{3.42}}$ 80.15±${}_{\pm\text{1.87}}$ 76.72±${}_{\pm\text{3.44}}$ 77.10±${}_{\pm\text{2.96}}$ 78.14±${}_{\pm\text{2.65}}$ 78.83±${}_{\pm\text{2.21}}$ 79.40±${}_{\pm\text{1.49}}$ 78.80±${}_{\pm\text{3.76}}$
steel 20 53.12±${}_{\pm\text{5.62}}$ 55.64±${}_{\pm\text{4.76}}$ 53.32±${}_{\pm\text{7.25}}$ 55.36±${}_{\pm\text{6.24}}$ 52.38±${}_{\pm\text{3.55}}$ 52.44±${}_{\pm\text{4.08}}$ 51.34±${}_{\pm\text{4.15}}$ 50.74±${}_{\pm\text{2.53}}$ 55.43±${}_{\pm\text{5.57}}$ 55.78±${}_{\pm\text{4.53}}$
50 66.73±${}_{\pm\text{9.11}}$ 60.79±${}_{\pm\text{5.52}}$ 59.51±${}_{\pm\text{4.15}}$ 54.82±${}_{\pm\text{4.23}}$ 54.79±${}_{\pm\text{4.69}}$ 59.71±${}_{\pm\text{6.94}}$ 57.66±${}_{\pm\text{5.19}}$ 55.89±${}_{\pm\text{4.50}}$ 63.78±${}_{\pm\text{7.20}}$ 74.18±${}_{\pm\text{13.67}}$
100 83.17±${}_{\pm\text{9.36}}$ 66.95±${}_{\pm\text{6.51}}$ 61.72±${}_{\pm\text{6.80}}$ 65.12±${}_{\pm\text{3.02}}$ 60.56±${}_{\pm\text{4.37}}$ 72.02±${}_{\pm\text{12.47}}$ 59.67±${}_{\pm\text{4.77}}$ 59.04±${}_{\pm\text{4.76}}$ 90.52±${}_{\pm\text{7.47}}$ 96.55±${}_{\pm\text{2.66}}$
200 95.94±${}_{\pm\text{2.73}}$ 81.21±${}_{\pm\text{5.01}}$ 73.14±${}_{\pm\text{5.45}}$ 70.64±${}_{\pm\text{10.67}}$ 70.26±${}_{\pm\text{9.25}}$ 74.50±${}_{\pm\text{23.57}}$ 74.57±${}_{\pm\text{9.36}}$ 65.41±${}_{\pm\text{6.70}}$ 99.14±${}_{\pm\text{1.19}}$ 99.54±${}_{\pm\text{0.62}}$
500 99.95±${}_{\pm\text{0.10}}$ 97.04±${}_{\pm\text{2.14}}$ 95.27±${}_{\pm\text{2.88}}$ 89.46±${}_{\pm\text{6.88}}$ 83.25±${}_{\pm\text{8.10}}$ 91.72±${}_{\pm\text{15.34}}$ 87.59±${}_{\pm\text{6.72}}$ 79.54±${}_{\pm\text{15.29}}$ 100.00±${}_{\pm\text{0.00}}$ 100.00±${}_{\pm\text{0.00}}$
stock 20 76.42±${}_{\pm\text{4.34}}$ 78.92±${}_{\pm\text{5.21}}$ 67.46±${}_{\pm\text{13.93}}$ 60.56±${}_{\pm\text{9.69}}$ 73.36±${}_{\pm\text{9.57}}$ 77.45±${}_{\pm\text{9.80}}$ 69.15±${}_{\pm\text{9.35}}$ 70.88±${}_{\pm\text{8.52}}$ 79.82±${}_{\pm\text{4.52}}$ 83.44±${}_{\pm\text{3.74}}$
50 83.71±${}_{\pm\text{3.40}}$ 86.23±${}_{\pm\text{2.54}}$ 84.65±${}_{\pm\text{4.44}}$ 79.31±${}_{\pm\text{6.58}}$ 76.27±${}_{\pm\text{3.89}}$ 85.70±${}_{\pm\text{3.96}}$ 81.61±${}_{\pm\text{1.97}}$ 84.98±${}_{\pm\text{4.44}}$ 87.28±${}_{\pm\text{3.65}}$ 88.21±${}_{\pm\text{3.31}}$
100 88.19±${}_{\pm\text{3.04}}$ 89.01±${}_{\pm\text{2.07}}$ 85.66±${}_{\pm\text{6.01}}$ 84.68±${}_{\pm\text{2.87}}$ 82.50±${}_{\pm\text{3.73}}$ 90.07±${}_{\pm\text{3.41}}$ 86.09±${}_{\pm\text{4.08}}$ 84.67±${}_{\pm\text{7.29}}$ 90.01±${}_{\pm\text{3.46}}$ 89.66±${}_{\pm\text{3.28}}$
200 92.32±${}_{\pm\text{1.35}}$ 92.26±${}_{\pm\text{2.33}}$ 90.94±${}_{\pm\text{1.98}}$ 89.01±${}_{\pm\text{2.53}}$ 88.92±${}_{\pm\text{2.67}}$ 91.36±${}_{\pm\text{3.79}}$ 91.04±${}_{\pm\text{1.46}}$ 91.42±${}_{\pm\text{2.66}}$ 91.72±${}_{\pm\text{2.77}}$ 92.17±${}_{\pm\text{1.51}}$
More than 10 classes energy 50 12.05±${}_{\pm\text{2.42}}$ N/A 11.60±${}_{\pm\text{3.83}}$ 14.47±${}_{\pm\text{5.32}}$ 10.95±${}_{\pm\text{4.68}}$ 10.21±${}_{\pm\text{5.11}}$ 12.81±${}_{\pm\text{2.51}}$ 12.34±${}_{\pm\text{3.55}}$ N/A 21.07±${}_{\pm\text{3.99}}$
100 29.37±${}_{\pm\text{1.72}}$ N/A 20.61±${}_{\pm\text{5.39}}$ 19.81±${}_{\pm\text{4.52}}$ 22.71±${}_{\pm\text{6.15}}$ 22.27±${}_{\pm\text{2.12}}$ 22.02±${}_{\pm\text{3.54}}$ 10.01±${}_{\pm\text{3.40}}$ N/A 27.93±${}_{\pm\text{4.16}}$
200 44.96±${}_{\pm\text{3.31}}$ N/A 36.73±${}_{\pm\text{6.03}}$ 35.92±${}_{\pm\text{8.45}}$ 33.71±${}_{\pm\text{6.54}}$ 34.73±${}_{\pm\text{5.89}}$ 37.06±${}_{\pm\text{5.26}}$ 18.81±${}_{\pm\text{7.27}}$ N/A 40.95±${}_{\pm\text{5.59}}$
collins 100 7.77±${}_{\pm\text{2.21}}$ N/A 7.76±${}_{\pm\text{0.95}}$ 6.52±${}_{\pm\text{1.16}}$ 6.11±${}_{\pm\text{1.09}}$ 8.95±${}_{\pm\text{1.90}}$ 6.21±${}_{\pm\text{1.14}}$ 5.96±${}_{\pm\text{1.07}}$ N/A 8.73±${}_{\pm\text{1.64}}$
200 10.58±${}_{\pm\text{2.57}}$ 11.46±${}_{\pm\text{2.11}}$ 9.43±${}_{\pm\text{2.20}}$ 9.84±${}_{\pm\text{1.56}}$ 8.26±${}_{\pm\text{1.75}}$ 9.80±${}_{\pm\text{1.96}}$ 8.90±${}_{\pm\text{0.83}}$ 9.79±${}_{\pm\text{0.80}}$ N/A 11.72±${}_{\pm\text{1.34}}$
texture 50 56.72±${}_{\pm\text{6.12}}$ 60.99±${}_{\pm\text{4.35}}$ 45.76±${}_{\pm\text{6.50}}$ 39.50±${}_{\pm\text{6.46}}$ 43.02±${}_{\pm\text{6.12}}$ 50.22±${}_{\pm\text{6.28}}$ 43.71±${}_{\pm\text{5.98}}$ 46.21±${}_{\pm\text{7.95}}$ N/A 69.11±${}_{\pm\text{3.27}}$
100 68.96±${}_{\pm\text{2.59}}$ 69.77±${}_{\pm\text{4.63}}$ 54.95±${}_{\pm\text{5.99}}$ 55.52±${}_{\pm\text{7.80}}$ 63.23±${}_{\pm\text{4.80}}$ 65.59±${}_{\pm\text{3.62}}$ 57.04±${}_{\pm\text{6.59}}$ 62.06±${}_{\pm\text{6.11}}$ N/A 76.35±${}_{\pm\text{2.64}}$
200 77.91±${}_{\pm\text{1.98}}$ 81.55±${}_{\pm\text{2.22}}$ 70.70±${}_{\pm\text{4.40}}$ 71.60±${}_{\pm\text{4.19}}$ 73.76±${}_{\pm\text{5.69}}$ 77.06±${}_{\pm\text{2.17}}$ 72.56±${}_{\pm\text{4.09}}$ 70.31±${}_{\pm\text{6.55}}$ N/A 82.59±${}_{\pm\text{2.15}}$
500 89.37±${}_{\pm\text{1.11}}$ 89.87±${}_{\pm\text{1.24}}$ 85.06±${}_{\pm\text{2.40}}$ 86.80±${}_{\pm\text{2.25}}$ 86.83±${}_{\pm\text{1.89}}$ 86.52±${}_{\pm\text{1.66}}$ 85.70±${}_{\pm\text{2.75}}$ 87.07±${}_{\pm\text{2.43}}$ N/A 89.69±${}_{\pm\text{1.10}}$
Average rank 3.64±${}_{\pm\text{2.09}}$ 3.45±${}_{\pm\text{1.48}}$ 6.32±${}_{\pm\text{1.86}}$ 7.33±${}_{\pm\text{2.19}}$ 8.64±${}_{\pm\text{1.82}}$ 6.30±${}_{\pm\text{2.44}}$ 6.64±${}_{\pm\text{2.18}}$ 7.62±${}_{\pm\text{1.84}}$ 3.44±${}_{\pm\text{1.54}}$ 1.62±${}_{\pm\text{1.29}}$
Table 12: Classification accuracy (%) of TabPFN, comparing data augmentation on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes that a specific generator was not applicable or the downstream predictor failed to converge, and the rank is computed with the mean balanced accuracy of other methods. We bold the highest accuracy for each dataset of different sample sizes. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ Baseline (Real data) SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 27.80±${}_{\pm\text{4.37}}$ N/A 19.21±${}_{\pm\text{3.80}}$ 20.58±${}_{\pm\text{4.63}}$ 20.80±${}_{\pm\text{4.34}}$ 18.89±${}_{\pm\text{4.37}}$ 23.97±${}_{\pm\text{3.05}}$ 10.55±${}_{\pm\text{1.61}}$ 33.42±${}_{\pm\text{5.95}}$ 34.63±${}_{\pm\text{5.78}}$
50 55.24±${}_{\pm\text{3.46}}$ 59.85±${}_{\pm\text{3.87}}$ 43.58±${}_{\pm\text{6.20}}$ 37.37±${}_{\pm\text{8.01}}$ 34.42±${}_{\pm\text{6.65}}$ 21.70±${}_{\pm\text{7.43}}$ 46.02±${}_{\pm\text{2.60}}$ 13.54±${}_{\pm\text{4.22}}$ 57.63±${}_{\pm\text{2.82}}$ 58.88±${}_{\pm\text{3.99}}$
100 74.31±${}_{\pm\text{3.49}}$ 80.05±${}_{\pm\text{3.16}}$ 68.15±${}_{\pm\text{5.54}}$ 71.10±${}_{\pm\text{2.55}}$ 57.89±${}_{\pm\text{6.13}}$ 59.28±${}_{\pm\text{8.23}}$ 65.84±${}_{\pm\text{3.03}}$ 23.69±${}_{\pm\text{10.40}}$ 77.60±${}_{\pm\text{4.03}}$ 78.26±${}_{\pm\text{3.75}}$
200 88.67±${}_{\pm\text{1.53}}$ 91.79±${}_{\pm\text{1.42}}$ 87.05±${}_{\pm\text{2.85}}$ 86.69±${}_{\pm\text{2.85}}$ 83.29±${}_{\pm\text{2.42}}$ 87.39±${}_{\pm\text{2.95}}$ 85.49±${}_{\pm\text{2.49}}$ 77.63±${}_{\pm\text{6.11}}$ 90.77±${}_{\pm\text{1.37}}$ 90.94±${}_{\pm\text{1.46}}$
500 97.31±${}_{\pm\text{0.69}}$ 97.69±${}_{\pm\text{0.77}}$ 97.51±${}_{\pm\text{0.85}}$ 97.58±${}_{\pm\text{0.85}}$ 96.89±${}_{\pm\text{0.62}}$ 97.44±${}_{\pm\text{0.85}}$ 97.40±${}_{\pm\text{0.60}}$ 97.35±${}_{\pm\text{0.61}}$ 97.24±${}_{\pm\text{0.80}}$ 97.28±${}_{\pm\text{0.62}}$
fourier 20 30.06±${}_{\pm\text{6.85}}$ N/A 22.00±${}_{\pm\text{4.62}}$ 20.10±${}_{\pm\text{4.31}}$ 14.52±${}_{\pm\text{3.96}}$ 12.22±${}_{\pm\text{2.40}}$ 21.64±${}_{\pm\text{5.91}}$ 14.64±${}_{\pm\text{3.94}}$ N/A 36.56±${}_{\pm\text{4.96}}$
50 53.62±${}_{\pm\text{4.71}}$ 53.08±${}_{\pm\text{3.34}}$ 45.82±${}_{\pm\text{4.29}}$ 37.46±${}_{\pm\text{5.82}}$ ${}_{\pm\text{2.78}}$ 22.74±${}_{\pm\text{5.11}}$ 42.14±${}_{\pm\text{3.09}}$ 11.30±${}_{\pm\text{1.50}}$ 53.15±${}_{\pm\text{3.50}}$ 53.82±${}_{\pm\text{3.92}}$
100 64.62±${}_{\pm\text{4.14}}$ 63.66±${}_{\pm\text{3.92}}$ 56.68±${}_{\pm\text{3.02}}$ 54.78±${}_{\pm\text{2.80}}$ 45.50±${}_{\pm\text{4.50}}$ 49.36±${}_{\pm\text{8.51}}$ 54.74±${}_{\pm\text{2.78}}$ 21.40±${}_{\pm\text{4.29}}$ 65.95±${}_{\pm\text{3.49}}$ 65.40±${}_{\pm\text{3.61}}$
200 71.62±${}_{\pm\text{2.59}}$ 70.56±${}_{\pm\text{3.61}}$ 66.48±${}_{\pm\text{3.82}}$ 66.14±${}_{\pm\text{4.02}}$ 62.64±${}_{\pm\text{2.60}}$ 72.12±${}_{\pm\text{2.64}}$ 65.04±${}_{\pm\text{3.20}}$ 52.18±${}_{\pm\text{7.35}}$ 69.93±${}_{\pm\text{3.91}}$ 72.48±${}_{\pm\text{3.08}}$
500 77.66±${}_{\pm\text{1.61}}$ 77.50±${}_{\pm\text{1.08}}$ 76.80±${}_{\pm\text{1.34}}$ 77.82±${}_{\pm\text{1.24}}$ 73.90±${}_{\pm\text{1.76}}$ 79.16±${}_{\pm\text{2.05}}$ 75.70±${}_{\pm\text{2.11}}$ 74.36±${}_{\pm\text{2.53}}$ 77.30±${}_{\pm\text{0.42}}$ 77.40±${}_{\pm\text{1.28}}$
biodeg 20 65.26±${}_{\pm\text{8.01}}$ 68.72±${}_{\pm\text{4.50}}$ 69.02±${}_{\pm\text{5.37}}$ 59.39±${}_{\pm\text{6.25}}$ 58.28±${}_{\pm\text{8.30}}$ 50.00±${}_{\pm\text{0.00}}$ 58.45±${}_{\pm\text{8.16}}$ 51.80±${}_{\pm\text{4.07}}$ 70.68±${}_{\pm\text{4.94}}$ 71.18±${}_{\pm\text{5.25}}$
50 75.27±${}_{\pm\text{2.63}}$ 74.65±${}_{\pm\text{3.28}}$ 73.44±${}_{\pm\text{4.02}}$ 70.21±${}_{\pm\text{3.61}}$ 55.68±${}_{\pm\text{9.27}}$ 50.00±${}_{\pm\text{0.00}}$ 72.74±${}_{\pm\text{3.74}}$ 55.75±${}_{\pm\text{7.45}}$ 75.69±${}_{\pm\text{2.44}}$ 75.56±${}_{\pm\text{3.22}}$
100 78.92±${}_{\pm\text{1.98}}$ 77.78±${}_{\pm\text{2.65}}$ 77.27±${}_{\pm\text{3.15}}$ 77.71±${}_{\pm\text{1.81}}$ 63.50±${}_{\pm\text{10.77}}$ 57.50±${}_{\pm\text{6.27}}$ 77.25±${}_{\pm\text{1.66}}$ 65.87±${}_{\pm\text{6.72}}$ 78.15±${}_{\pm\text{1.45}}$ 79.00±${}_{\pm\text{1.99}}$
200 82.59±${}_{\pm\text{1.84}}$ 81.42±${}_{\pm\text{1.27}}$ 80.48±${}_{\pm\text{1.82}}$ 80.19±${}_{\pm\text{2.48}}$ 79.16±${}_{\pm\text{2.49}}$ 80.45±${}_{\pm\text{1.48}}$ 80.88±${}_{\pm\text{1.68}}$ 80.66±${}_{\pm\text{1.49}}$ 82.56±${}_{\pm\text{1.68}}$ 82.58±${}_{\pm\text{1.90}}$
500 85.00±${}_{\pm\text{0.70}}$ 84.37±${}_{\pm\text{0.75}}$ 84.40±${}_{\pm\text{0.68}}$ 84.67±${}_{\pm\text{0.98}}$ 84.45±${}_{\pm\text{0.91}}$ 84.58±${}_{\pm\text{0.70}}$ 84.68±${}_{\pm\text{1.06}}$ 83.66±${}_{\pm\text{0.67}}$ 84.56±${}_{\pm\text{0.98}}$ 84.55±${}_{\pm\text{0.92}}$
steel 20 56.77±${}_{\pm\text{4.17}}$ 55.95±${}_{\pm\text{4.30}}$ 56.03±${}_{\pm\text{4.37}}$ 55.62±${}_{\pm\text{4.80}}$ 52.52±${}_{\pm\text{4.64}}$ 50.00±${}_{\pm\text{0.00}}$ 52.39±${}_{\pm\text{3.13}}$ 50.05±${}_{\pm\text{0.17}}$ 64.80±${}_{\pm\text{5.66}}$ 65.87±${}_{\pm\text{6.14}}$
50 82.34±${}_{\pm\text{8.38}}$ 63.42±${}_{\pm\text{3.93}}$ 62.08±${}_{\pm\text{2.69}}$ 63.98±${}_{\pm\text{4.08}}$ 52.92±${}_{\pm\text{4.72}}$ 50.64±${}_{\pm\text{2.01}}$ 61.32±${}_{\pm\text{4.55}}$ 50.36±${}_{\pm\text{1.09}}$ ${}_{\pm\text{7.84}}$ 86.30±${}_{\pm\text{6.73}}$
100 97.37±${}_{\pm\text{1.37}}$ 73.06±${}_{\pm\text{4.46}}$ 71.96±${}_{\pm\text{5.40}}$ 72.23±${}_{\pm\text{4.15}}$ 56.34±${}_{\pm\text{6.30}}$ 80.87±${}_{\pm\text{20.44}}$ 69.29±${}_{\pm\text{5.70}}$ 51.18±${}_{\pm\text{3.24}}$ 97.49±${}_{\pm\text{1.21}}$ 97.81±${}_{\pm\text{1.49}}$
200 98.84±${}_{\pm\text{0.70}}$ 82.32±${}_{\pm\text{2.88}}$ 81.78±${}_{\pm\text{3.36}}$ 83.24±${}_{\pm\text{2.68}}$ 82.92±${}_{\pm\text{6.21}}$ 99.35±${}_{\pm\text{0.70}}$ 86.40±${}_{\pm\text{4.22}}$ 64.42±${}_{\pm\text{11.35}}$ 98.80±${}_{\pm\text{0.73}}$ 98.96±${}_{\pm\text{0.71}}$
500 99.74±${}_{\pm\text{0.29}}$ 94.27±${}_{\pm\text{2.39}}$ 94.93±${}_{\pm\text{1.89}}$ 96.98±${}_{\pm\text{1.34}}$ 98.32±${}_{\pm\text{1.19}}$ 99.88±${}_{\pm\text{0.15}}$ 95.70±${}_{\pm\text{1.50}}$ 98.56±${}_{\pm\text{0.52}}$ 99.77±${}_{\pm\text{0.30}}$ 99.74±${}_{\pm\text{0.29}}$
stock 20 83.18±${}_{\pm\text{4.37}}$ 83.69±${}_{\pm\text{3.10}}$ 74.01±${}_{\pm\text{5.09}}$ 56.92±${}_{\pm\text{16.52}}$ 74.99±${}_{\pm\text{6.60}}$ 78.73±${}_{\pm\text{12.25}}$ 69.64±${}_{\pm\text{6.88}}$ 73.40±${}_{\pm\text{4.88}}$ 82.95±${}_{\pm\text{4.44}}$ 83.81±${}_{\pm\text{4.94}}$
50 90.01±${}_{\pm\text{2.07}}$ 90.01±${}_{\pm\text{2.43}}$ 82.27±${}_{\pm\text{4.30}}$ 78.91±${}_{\pm\text{4.14}}$ 78.94±${}_{\pm\text{8.78}}$ 89.68±${}_{\pm\text{1.92}}$ 83.72±${}_{\pm\text{2.50}}$ 79.00±${}_{\pm\text{6.87}}$ 89.95±${}_{\pm\text{2.08}}$ 90.15±${}_{\pm\text{1.76}}$
100 92.39±${}_{\pm\text{1.06}}$ 92.09±${}_{\pm\text{1.45}}$ 90.75±${}_{\pm\text{2.20}}$ 89.43±${}_{\pm\text{3.29}}$ 86.16±${}_{\pm\text{3.83}}$ 92.12±${}_{\pm\text{1.16}}$ 90.17±${}_{\pm\text{1.92}}$ 89.30±${}_{\pm\text{1.33}}$ 92.12±${}_{\pm\text{1.12}}$ 92.57±${}_{\pm\text{1.27}}$
200 94.16±${}_{\pm\text{0.92}}$ 93.99±${}_{\pm\text{0.70}}$ 93.57±${}_{\pm\text{1.10}}$ 93.28±${}_{\pm\text{1.59}}$ 91.92±${}_{\pm\text{2.00}}$ 94.22±${}_{\pm\text{1.10}}$ 93.05±${}_{\pm\text{1.35}}$ 92.07±${}_{\pm\text{1.76}}$ 94.17±${}_{\pm\text{0.89}}$ 94.16±${}_{\pm\text{1.07}}$
Average rank 3.08±${}_{\pm\text{1.22}}$ 4.23±${}_{\pm\text{2.32}}$ 6.12±${}_{\pm\text{1.57}}$ 6.29±${}_{\pm\text{2.07}}$ 8.42±${}_{\pm\text{1.32}}$ 6.12±${}_{\pm\text{3.38}}$ 6.54±${}_{\pm\text{1.61}}$ 8.83±${}_{\pm\text{1.46}}$ 3.12±${}_{\pm\text{1.80}}$ 2.23±${}_{\pm\text{1.83}}$
C.5.2 Results on six UCI Datasets
Table 13: Details of the six real-world tabular datasets from UCI.
Dataset UCI ID Not evaluated in TabPFN [33] # Samples ($N$) # Features ($D$) # Classes N/D # Samples per class (Min) # Samples per class (Max)
clinical 890 2,139 23 2 93 521 1,618
support2 880 9,105 42 2 217 2,904 6,201
mushroom 73 8,124 22 2 369 3,916 4,208
auction 713 2,043 7 2 292 262 1,781
abalone 1 4,153 8 19 519 14 689
statlog 144 1,000 20 2 50 300 700
Table 14: Test classification accuracy (%) aggregated over six downstream predictors, comparing data augmentation on six leakage-free UCI datasets. Note that “N/A” denotes that a specific generator was not applicable. TabEBM still achieves the best overall performance against benchmark methods.
Datasets ($N_{\text{real}}=100$) Baseline SMOTE TVAE CTGAN TabDDPM TabPFGen TabEBM (Ours)
clinical 68.63±${}_{\pm\text{5.81}}$ 71.07±${}_{\pm\text{4.67}}$ 61.80±${}_{\pm\text{2.76}}$ 65.21±${}_{\pm\text{5.77}}$ 54.03±${}_{\pm\text{5.36}}$ 69.66±${}_{\pm\text{3.65}}$ 71.20±${}_{\pm\text{3.54}}$
support2 64.23±${}_{\pm\text{1.89}}$ 65.60±${}_{\pm\text{1.52}}$ 60.70±${}_{\pm\text{0.90}}$ 59.14±${}_{\pm\text{1.88}}$ 58.31±${}_{\pm\text{1.74}}$ 64.34±${}_{\pm\text{1.19}}$ 65.28±${}_{\pm\text{1.15}}$
mushroom 95.51±${}_{\pm\text{2.48}}$ 95.84±${}_{\pm\text{1.99}}$ 93.75±${}_{\pm\text{1.18}}$ 93.26±${}_{\pm\text{2.46}}$ 79.87±${}_{\pm\text{2.29}}$ 97.05±${}_{\pm\text{1.56}}$ 96.82±${}_{\pm\text{1.51}}$
auction 51.90±${}_{\pm\text{1.91}}$ 57.35±${}_{\pm\text{1.53}}$ 53.09±${}_{\pm\text{0.91}}$ 52.35±${}_{\pm\text{1.90}}$ 51.14±${}_{\pm\text{1.76}}$ 56.82±${}_{\pm\text{1.20}}$ 57.97±${}_{\pm\text{1.16}}$
abalone 11.59±${}_{\pm\text{2.69}}$ N/A 8.49±${}_{\pm\text{1.28}}$ 7.72±${}_{\pm\text{2.67}}$ 9.95±${}_{\pm\text{2.48}}$ N/A 13.56±${}_{\pm\text{1.64}}$
statlog 56.22±${}_{\pm\text{3.20}}$ 57.30±${}_{\pm\text{2.57}}$ 53.12±${}_{\pm\text{1.52}}$ 55.55±${}_{\pm\text{3.18}}$ 53.07±${}_{\pm\text{2.95}}$ 57.65±${}_{\pm\text{2.01}}$ 57.85±${}_{\pm\text{1.95}}$
C.5.3 Results on larger sample sizes
Table 15: Test classification accuracy (%) aggregated over six downstream predictors, comparing data augmentation with increased real data availability of the “texture” dataset. Note that “N/A” denotes that a specific generator was not applicable. On larger datasets, TabEBM still outperforms other generators, but training on real data alone appears sufficient. This highlights TabEBM’s usefulness in fields with limited training samples.
$N_{\text{real}}$ Baseline SMOTE TVAE CTGAN TabDDPM TabPFGen TabEBM (Ours) Accuracy improvements by TabEBM (%)
50 72.40±${}_{\pm\text{13.07}}$ 76.40±${}_{\pm\text{10.50}}$ 55.33±${}_{\pm\text{6.20}}$ 54.80±${}_{\pm\text{12.97}}$ 62.94±${}_{\pm\text{12.06}}$ N/A 78.90±${}_{\pm\text{7.96}}$ +6.50
100 82.42±${}_{\pm\text{10.38}}$ 84.35±${}_{\pm\text{9.67}}$ 66.00±${}_{\pm\text{7.21}}$ 69.49±${}_{\pm\text{10.93}}$ 76.34±${}_{\pm\text{9.55}}$ N/A 86.01±${}_{\pm\text{7.36}}$ +3.59
200 87.54±${}_{\pm\text{7.62}}$ 89.29±${}_{\pm\text{6.20}}$ 78.37±${}_{\pm\text{6.03}}$ 82.44±${}_{\pm\text{7.15}}$ 82.53±${}_{\pm\text{7.99}}$ N/A 89.77±${}_{\pm\text{5.77}}$ +2.23
500 92.96±${}_{\pm\text{4.07}}$ 93.69±${}_{\pm\text{3.83}}$ 90.09±${}_{\pm\text{3.56}}$ 91.48±${}_{\pm\text{3.50}}$ 91.24±${}_{\pm\text{3.56}}$ N/A 93.76±${}_{\pm\text{3.64}}$ +0.80
1000 96.37±${}_{\pm\text{2.17}}$ 96.21±${}_{\pm\text{2.37}}$ 93.61±${}_{\pm\text{2.10}}$ 95.36±${}_{\pm\text{1.71}}$ 94.56±${}_{\pm\text{1.59}}$ N/A 96.30±${}_{\pm\text{2.30}}$0.07
2000 97.76±${}_{\pm\text{1.16}}$ 96.84±${}_{\pm\text{1.46}}$ 96.62±${}_{\pm\text{1.24}}$ 97.10±${}_{\pm\text{0.84}}$ 97.13±${}_{\pm\text{0.71}}$ N/A 97.83±${}_{\pm\text{1.45}}$ +0.07
3000 98.20±${}_{\pm\text{0.62}}$ 98.28±${}_{\pm\text{0.90}}$ 97.60±${}_{\pm\text{0.73}}$ 97.60±${}_{\pm\text{0.41}}$ 97.73±${}_{\pm\text{0.31}}$ N/A 98.35±${}_{\pm\text{0.91}}$ +0.15
4000 98.51±${}_{\pm\text{0.33}}$ 98.59±${}_{\pm\text{0.56}}$ 98.11±${}_{\pm\text{0.43}}$ 98.00±${}_{\pm\text{0.20}}$ 98.46±${}_{\pm\text{0.14}}$ N/A 98.55±${}_{\pm\text{0.58}}$ +0.04
C.6 Results on Statistical Fidelity
We aim to provide a fair and coherent comparison between TabEBM and existing methods and thus we follow the widely adopted evaluation process in prior studies. Specifically, we compute the three statistical fidelity metrics with the open-source implementations from the well-established benchmark, Synthcity. We note that the previous studies [88, 68] often operate under the assumption that the issues associated with multiple comparisons are less pronounced in generating low-dimensional tabular data, hence correction methods for multiple hypothesis testing are seldom employed. Following such assumptions, correction methods are not employed in this work. In addition, we would like to point out the imperfection of widely adopted univariate metrics (i.e., Inverse KL, KS test and test) in existing work. However, evaluating generators’ ability to capture the joining feature relationships remains an open research question [79]. We leave this for future work to explore.
C.6.1 Similarity between Real Train Data and Synthetic Data
Table 16: Inverse KL between real train data and synthetic data on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher fidelity. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean result of other methods. We bold the highest result for each dataset of different sample sizes. TabEBM achieves the best overall performance against benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 0.11±${}_{\pm\text{0.01}}$ ${}_{\pm\text{0.02}}$ 0.34±${}_{\pm\text{0.05}}$ 0.07±${}_{\pm\text{0.01}}$ 0.22±${}_{\pm\text{0.02}}$ 0.07±${}_{\pm\text{0.00}}$ 0.46±${}_{\pm\text{0.13}}$ 0.77±${}_{\pm\text{0.04}}$
50 0.88±${}_{\pm\text{0.01}}$ 0.80±${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.05}}$ 0.87±${}_{\pm\text{0.01}}$ 0.07±${}_{\pm\text{0.00}}$ 0.87±${}_{\pm\text{0.01}}$ 0.50±${}_{\pm\text{0.04}}$ 0.82±${}_{\pm\text{0.06}}$ 0.94±${}_{\pm\text{0.02}}$
100 0.93±${}_{\pm\text{0.01}}$ 0.79±${}_{\pm\text{0.02}}$ 0.78±${}_{\pm\text{0.03}}$ 0.90±${}_{\pm\text{0.03}}$ 0.07±${}_{\pm\text{0.00}}$ 0.91±${}_{\pm\text{0.01}}$ 0.32±${}_{\pm\text{0.05}}$ 0.92±${}_{\pm\text{0.02}}$ 0.96±${}_{\pm\text{0.01}}$
200 0.95±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.03}}$ 0.83±${}_{\pm\text{0.03}}$ 0.93±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.01}}$ 0.11±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.01}}$ 0.96±${}_{\pm\text{0.01}}$
500 0.96±${}_{\pm\text{0.00}}$ 0.70±${}_{\pm\text{0.02}}$ 0.87±${}_{\pm\text{0.03}}$ 0.94±${}_{\pm\text{0.00}}$ 0.09±${}_{\pm\text{0.00}}$ 0.95±${}_{\pm\text{0.01}}$ 0.13±${}_{\pm\text{0.01}}$ 0.96±${}_{\pm\text{0.01}}$ 0.97±${}_{\pm\text{0.01}}$
fourier 20 N/A 0.12±${}_{\pm\text{0.03}}$ 0.15±${}_{\pm\text{0.02}}$ 0.27±${}_{\pm\text{0.04}}$ 0.50±${}_{\pm\text{0.03}}$ 0.50±${}_{\pm\text{0.03}}$ 0.50±${}_{\pm\text{0.04}}$ 0.97±${}_{\pm\text{0.00}}$ 0.87±${}_{\pm\text{0.01}}$
50 0.93±${}_{\pm\text{0.01}}$ 0.79±${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.05}}$ 0.90±${}_{\pm\text{0.01}}$ 0.07±${}_{\pm\text{0.00}}$ 0.90±${}_{\pm\text{0.00}}$ 0.47±${}_{\pm\text{0.05}}$ 0.87±${}_{\pm\text{0.02}}$ 0.95±${}_{\pm\text{0.01}}$
100 0.95±${}_{\pm\text{0.01}}$ 0.76±${}_{\pm\text{0.03}}$ 0.81±${}_{\pm\text{0.03}}$ 0.93±${}_{\pm\text{0.00}}$ 0.07±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.01}}$ 0.20±${}_{\pm\text{0.05}}$ 0.94±${}_{\pm\text{0.01}}$ 0.97±${}_{\pm\text{0.01}}$
200 0.97±${}_{\pm\text{0.01}}$ 0.61±${}_{\pm\text{0.01}}$ 0.82±${}_{\pm\text{0.02}}$ 0.95±${}_{\pm\text{0.00}}$ 0.09±${}_{\pm\text{0.01}}$ 0.96±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.01}}$ 0.97±${}_{\pm\text{0.01}}$ 0.98±${}_{\pm\text{0.00}}$
500 0.97±${}_{\pm\text{0.00}}$ 0.52±${}_{\pm\text{0.03}}$ 0.90±${}_{\pm\text{0.02}}$ 0.95±${}_{\pm\text{0.01}}$ ${}_{\pm\text{0.01}}$ 0.97±${}_{\pm\text{0.00}}$ 0.09±${}_{\pm\text{0.00}}$ 0.98±${}_{\pm\text{0.00}}$ 0.98±${}_{\pm\text{0.00}}$
biodeg 20 0.47±${}_{\pm\text{0.04}}$ 0.43±${}_{\pm\text{0.04}}$ 0.43±${}_{\pm\text{0.03}}$ ${}_{\pm\text{0.05}}$ 0.34±${}_{\pm\text{0.03}}$ 0.51±${}_{\pm\text{0.04}}$ 0.37±${}_{\pm\text{0.04}}$ 0.60±${}_{\pm\text{0.07}}$ 0.87±${}_{\pm\text{0.04}}$
50 0.62±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.02}}$ 0.56±${}_{\pm\text{0.05}}$ 0.63±${}_{\pm\text{0.05}}$ 0.28±${}_{\pm\text{0.02}}$ 0.65±${}_{\pm\text{0.03}}$ 0.41±${}_{\pm\text{0.03}}$ 0.75±${}_{\pm\text{0.05}}$ 0.90±${}_{\pm\text{0.02}}$
100 0.69±${}_{\pm\text{0.05}}$ 0.66±${}_{\pm\text{0.03}}$ 0.65±${}_{\pm\text{0.05}}$ 0.67±${}_{\pm\text{0.04}}$ 0.30±${}_{\pm\text{0.04}}$ 0.69±${}_{\pm\text{0.05}}$ 0.38±${}_{\pm\text{0.04}}$ 0.76±${}_{\pm\text{0.04}}$ 0.90±${}_{\pm\text{0.04}}$
200 0.71±${}_{\pm\text{0.03}}$ 0.65±${}_{\pm\text{0.03}}$ 0.69±${}_{\pm\text{0.02}}$ 0.68±${}_{\pm\text{0.03}}$ 0.29±${}_{\pm\text{0.04}}$ 0.69±${}_{\pm\text{0.03}}$ 0.34±${}_{\pm\text{0.01}}$ 0.79±${}_{\pm\text{0.04}}$ 0.91±${}_{\pm\text{0.02}}$
500 0.80±${}_{\pm\text{0.02}}$ 0.68±${}_{\pm\text{0.02}}$ 0.73±${}_{\pm\text{0.02}}$ 0.75±${}_{\pm\text{0.02}}$ 0.26±${}_{\pm\text{0.02}}$ 0.73±${}_{\pm\text{0.02}}$ 0.37±${}_{\pm\text{0.02}}$ 0.81±${}_{\pm\text{0.04}}$ 0.92±${}_{\pm\text{0.02}}$
steel 20 0.45±${}_{\pm\text{0.05}}$ 0.37±${}_{\pm\text{0.03}}$ ${}_{\pm\text{0.04}}$ 0.47±${}_{\pm\text{0.04}}$ 0.17±${}_{\pm\text{0.03}}$ 0.43±${}_{\pm\text{0.05}}$ 0.29±${}_{\pm\text{0.05}}$ 0.53±${}_{\pm\text{0.08}}$ 0.84±${}_{\pm\text{0.03}}$
50 0.70±${}_{\pm\text{0.02}}$ 0.57±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.04}}$ 0.64±${}_{\pm\text{0.04}}$ 0.13±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.01}}$ 0.23±${}_{\pm\text{0.03}}$ 0.71±${}_{\pm\text{0.08}}$ 0.91±${}_{\pm\text{0.02}}$
100 0.71±${}_{\pm\text{0.04}}$ 0.55±${}_{\pm\text{0.02}}$ 0.63±${}_{\pm\text{0.02}}$ 0.67±${}_{\pm\text{0.02}}$ 0.13±${}_{\pm\text{0.01}}$ 0.66±${}_{\pm\text{0.02}}$ ${}_{\pm\text{0.02}}$ 0.75±${}_{\pm\text{0.05}}$ 0.92±${}_{\pm\text{0.03}}$
200 0.75±${}_{\pm\text{0.01}}$ 0.50±${}_{\pm\text{0.04}}$ 0.65±${}_{\pm\text{0.03}}$ 0.70±${}_{\pm\text{0.01}}$ 0.13±${}_{\pm\text{0.02}}$ 0.67±${}_{\pm\text{0.02}}$ 0.17±${}_{\pm\text{0.01}}$ 0.77±${}_{\pm\text{0.04}}$ 0.93±${}_{\pm\text{0.01}}$
500 0.75±${}_{\pm\text{0.01}}$ 0.51±${}_{\pm\text{0.04}}$ 0.66±${}_{\pm\text{0.04}}$ 0.70±${}_{\pm\text{0.01}}$ 0.14±${}_{\pm\text{0.01}}$ 0.68±${}_{\pm\text{0.01}}$ 0.19±${}_{\pm\text{0.01}}$ 0.80±${}_{\pm\text{0.06}}$ 0.94±${}_{\pm\text{0.02}}$
stock 20 0.55±${}_{\pm\text{0.07}}$ 0.45±${}_{\pm\text{0.07}}$ 0.43±${}_{\pm\text{0.05}}$ 0.60±${}_{\pm\text{0.05}}$ 0.24±${}_{\pm\text{0.04}}$ 0.52±${}_{\pm\text{0.06}}$ 0.35±${}_{\pm\text{0.11}}$ 0.68±${}_{\pm\text{0.12}}$ 0.89±${}_{\pm\text{0.02}}$
50 0.92±${}_{\pm\text{0.02}}$ 0.73±${}_{\pm\text{0.06}}$ 0.78±${}_{\pm\text{0.07}}$ 0.86±${}_{\pm\text{0.03}}$ 0.37±${}_{\pm\text{0.07}}$ 0.88±${}_{\pm\text{0.01}}$ 0.32±${}_{\pm\text{0.11}}$ 0.91±${}_{\pm\text{0.03}}$ 0.95±${}_{\pm\text{0.02}}$
100 0.96±${}_{\pm\text{0.02}}$ 0.67±${}_{\pm\text{0.07}}$ 0.83±${}_{\pm\text{0.05}}$ 0.91±${}_{\pm\text{0.03}}$ 0.49±${}_{\pm\text{0.10}}$ 0.93±${}_{\pm\text{0.01}}$ 0.17±${}_{\pm\text{0.04}}$ 0.95±${}_{\pm\text{0.02}}$ 0.97±${}_{\pm\text{0.01}}$
200 0.98±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.04}}$ ${}_{\pm\text{0.08}}$ 0.92±${}_{\pm\text{0.02}}$ 0.83±${}_{\pm\text{0.09}}$ 0.95±${}_{\pm\text{0.01}}$ 0.15±${}_{\pm\text{0.00}}$ 0.98±${}_{\pm\text{0.01}}$ 0.98±${}_{\pm\text{0.00}}$
More than 10 classes energy 50 N/A 0.25±${}_{\pm\text{0.06}}$ 0.34±${}_{\pm\text{0.06}}$ 0.42±${}_{\pm\text{0.06}}$ 0.24±${}_{\pm\text{0.05}}$ 0.49±${}_{\pm\text{0.10}}$ 0.24±${}_{\pm\text{0.09}}$ N/A 0.80±${}_{\pm\text{0.04}}$
100 N/A 0.28±${}_{\pm\text{0.07}}$ 0.42±${}_{\pm\text{0.09}}$ 0.44±${}_{\pm\text{0.05}}$ 0.22±${}_{\pm\text{0.07}}$ 0.41±${}_{\pm\text{0.08}}$ 0.16±${}_{\pm\text{0.04}}$ N/A 0.89±${}_{\pm\text{0.01}}$
200 N/A 0.30±${}_{\pm\text{0.08}}$ 0.43±${}_{\pm\text{0.08}}$ 0.47±${}_{\pm\text{0.09}}$ 0.25±${}_{\pm\text{0.06}}$ 0.40±${}_{\pm\text{0.08}}$ 0.12±${}_{\pm\text{0.04}}$ N/A 0.91±${}_{\pm\text{0.01}}$
collins 100 N/A 0.72±${}_{\pm\text{0.02}}$ 0.84±${}_{\pm\text{0.04}}$ 0.90±${}_{\pm\text{0.02}}$ 0.44±${}_{\pm\text{0.11}}$ 0.91±${}_{\pm\text{0.02}}$ 0.28±${}_{\pm\text{0.08}}$ N/A 0.94±${}_{\pm\text{0.01}}$
200 0.94±${}_{\pm\text{0.01}}$ 0.64±${}_{\pm\text{0.05}}$ 0.87±${}_{\pm\text{0.04}}$ 0.92±${}_{\pm\text{0.02}}$ 0.44±${}_{\pm\text{0.05}}$ 0.93±${}_{\pm\text{0.01}}$ 0.23±${}_{\pm\text{0.10}}$ N/A 0.96±${}_{\pm\text{0.01}}$
texture 50 0.89±${}_{\pm\text{0.04}}$ 0.74±${}_{\pm\text{0.04}}$ 0.71±${}_{\pm\text{0.06}}$ 0.88±${}_{\pm\text{0.03}}$ ${}_{\pm\text{0.01}}$ 0.88±${}_{\pm\text{0.02}}$ 0.45±${}_{\pm\text{0.10}}$ N/A 0.93±${}_{\pm\text{0.04}}$
100 0.96±${}_{\pm\text{0.01}}$ 0.67±${}_{\pm\text{0.05}}$ 0.81±${}_{\pm\text{0.07}}$ 0.91±${}_{\pm\text{0.02}}$ 0.11±${}_{\pm\text{0.02}}$ 0.92±${}_{\pm\text{0.02}}$ 0.22±${}_{\pm\text{0.06}}$ N/A 0.97±${}_{\pm\text{0.01}}$
200 0.96±${}_{\pm\text{0.02}}$ 0.56±${}_{\pm\text{0.04}}$ 0.80±${}_{\pm\text{0.07}}$ 0.93±${}_{\pm\text{0.01}}$ 0.12±${}_{\pm\text{0.01}}$ 0.95±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.01}}$ N/A 0.98±${}_{\pm\text{0.01}}$
500 0.97±${}_{\pm\text{0.02}}$ 0.63±${}_{\pm\text{0.06}}$ 0.84±${}_{\pm\text{0.04}}$ 0.93±${}_{\pm\text{0.02}}$ 0.14±${}_{\pm\text{0.01}}$ 0.96±${}_{\pm\text{0.01}}$ 0.11±${}_{\pm\text{0.01}}$ N/A 0.98±${}_{\pm\text{0.01}}$
Average rank 3.24±${}_{\pm\text{1.30}}$ 6.76±${}_{\pm\text{0.75}}$ 5.91±${}_{\pm\text{1.10}}$ 4.12±${}_{\pm\text{1.08}}$ 8.42±${}_{\pm\text{1.17}}$ 3.97±${}_{\pm\text{0.95}}$ 8.21±${}_{\pm\text{0.89}}$ 3.30±${}_{\pm\text{1.79}}$ 1.06±${}_{\pm\text{0.24}}$
Table 17: KS test between real train data and synthetic data on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std result and average rank across datasets. A higher rank implies higher fidelity. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean result of other methods. We bold the highest result for each dataset of different sample sizes. TabEBM achieves the best overall performance against benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 0.72±${}_{\pm\text{0.02}}$ 0.75±${}_{\pm\text{0.03}}$ 0.81±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.01}}$ 0.80±${}_{\pm\text{0.01}}$ 0.44±${}_{\pm\text{0.00}}$ 0.84±${}_{\pm\text{0.02}}$ 0.87±${}_{\pm\text{0.01}}$
50 0.90±${}_{\pm\text{0.01}}$ 0.87±${}_{\pm\text{0.00}}$ 0.87±${}_{\pm\text{0.01}}$ 0.89±${}_{\pm\text{0.01}}$ 0.62±${}_{\pm\text{0.01}}$ 0.89±${}_{\pm\text{0.00}}$ 0.73±${}_{\pm\text{0.02}}$ 0.91±${}_{\pm\text{0.00}}$ 0.93±${}_{\pm\text{0.00}}$
100 0.92±${}_{\pm\text{0.00}}$ 0.88±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.00}}$ 0.91±${}_{\pm\text{0.00}}$ 0.60±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.00}}$ 0.66±${}_{\pm\text{0.02}}$ 0.94±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.00}}$
200 0.94±${}_{\pm\text{0.00}}$ 0.87±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.00}}$ 0.93±${}_{\pm\text{0.00}}$ 0.59±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.58±${}_{\pm\text{0.01}}$ 0.95±${}_{\pm\text{0.00}}$ 0.95±${}_{\pm\text{0.00}}$
500 0.95±${}_{\pm\text{0.00}}$ 0.83±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.00}}$ 0.93±${}_{\pm\text{0.00}}$ 0.59±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.00}}$ 0.63±${}_{\pm\text{0.01}}$ 0.95±${}_{\pm\text{0.00}}$ 0.96±${}_{\pm\text{0.00}}$
fourier 20 N/A 0.71±${}_{\pm\text{0.04}}$ 0.73±${}_{\pm\text{0.02}}$ 0.79±${}_{\pm\text{0.01}}$ 0.85±${}_{\pm\text{0.00}}$ 0.85±${}_{\pm\text{0.00}}$ 0.85±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.00}}$ 0.90±${}_{\pm\text{0.00}}$
50 0.92±${}_{\pm\text{0.00}}$ 0.87±${}_{\pm\text{0.00}}$ 0.88±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.00}}$ 0.64±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.00}}$ 0.73±${}_{\pm\text{0.02}}$ 0.93±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.00}}$
100 0.94±${}_{\pm\text{0.00}}$ 0.87±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.00}}$ 0.93±${}_{\pm\text{0.00}}$ 0.63±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.62±${}_{\pm\text{0.02}}$ 0.95±${}_{\pm\text{0.00}}$ 0.95±${}_{\pm\text{0.00}}$
200 0.95±${}_{\pm\text{0.00}}$ 0.83±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.00}}$ 0.62±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.00}}$ 0.61±${}_{\pm\text{0.03}}$ 0.97±${}_{\pm\text{0.00}}$ 0.96±${}_{\pm\text{0.00}}$
500 0.96±${}_{\pm\text{0.00}}$ 0.81±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.00}}$ 0.95±${}_{\pm\text{0.00}}$ 0.62±${}_{\pm\text{0.01}}$ 0.95±${}_{\pm\text{0.00}}$ 0.63±${}_{\pm\text{0.01}}$ 0.97±${}_{\pm\text{0.00}}$ 0.97±${}_{\pm\text{0.00}}$
biodeg 20 0.63±${}_{\pm\text{0.04}}$ 0.62±${}_{\pm\text{0.04}}$ 0.64±${}_{\pm\text{0.03}}$ 0.64±${}_{\pm\text{0.04}}$ 0.56±${}_{\pm\text{0.04}}$ 0.63±${}_{\pm\text{0.04}}$ 0.65±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.02}}$ 0.70±${}_{\pm\text{0.01}}$
50 0.57±${}_{\pm\text{0.03}}$ 0.57±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.03}}$ 0.60±${}_{\pm\text{0.03}}$ 0.48±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.03}}$ 0.63±${}_{\pm\text{0.02}}$ 0.61±${}_{\pm\text{0.04}}$ 0.73±${}_{\pm\text{0.00}}$
100 0.56±${}_{\pm\text{0.03}}$ 0.55±${}_{\pm\text{0.03}}$ 0.57±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.03}}$ 0.46±${}_{\pm\text{0.03}}$ 0.58±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.02}}$ 0.73±${}_{\pm\text{0.01}}$
200 0.53±${}_{\pm\text{0.01}}$ 0.52±${}_{\pm\text{0.01}}$ 0.53±${}_{\pm\text{0.02}}$ 0.56±${}_{\pm\text{0.02}}$ 0.43±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.02}}$ 0.58±${}_{\pm\text{0.02}}$ 0.72±${}_{\pm\text{0.01}}$
500 0.53±${}_{\pm\text{0.00}}$ 0.50±${}_{\pm\text{0.01}}$ 0.52±${}_{\pm\text{0.01}}$ 0.55±${}_{\pm\text{0.01}}$ 0.43±${}_{\pm\text{0.01}}$ 0.56±${}_{\pm\text{0.01}}$ 0.56±${}_{\pm\text{0.01}}$ 0.57±${}_{\pm\text{0.01}}$ 0.72±${}_{\pm\text{0.01}}$
steel 20 0.67±${}_{\pm\text{0.02}}$ 0.63±${}_{\pm\text{0.02}}$ 0.64±${}_{\pm\text{0.03}}$ 0.66±${}_{\pm\text{0.02}}$ 0.54±${}_{\pm\text{0.02}}$ 0.65±${}_{\pm\text{0.02}}$ 0.57±${}_{\pm\text{0.03}}$ 0.67±${}_{\pm\text{0.02}}$ 0.76±${}_{\pm\text{0.01}}$
50 0.69±${}_{\pm\text{0.02}}$ 0.64±${}_{\pm\text{0.03}}$ 0.65±${}_{\pm\text{0.03}}$ 0.67±${}_{\pm\text{0.02}}$ 0.51±${}_{\pm\text{0.02}}$ 0.67±${}_{\pm\text{0.03}}$ 0.56±${}_{\pm\text{0.02}}$ 0.72±${}_{\pm\text{0.02}}$ 0.79±${}_{\pm\text{0.01}}$
100 0.69±${}_{\pm\text{0.02}}$ 0.63±${}_{\pm\text{0.02}}$ 0.65±${}_{\pm\text{0.02}}$ 0.67±${}_{\pm\text{0.01}}$ 0.50±${}_{\pm\text{0.01}}$ 0.65±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.02}}$ 0.71±${}_{\pm\text{0.03}}$ 0.79±${}_{\pm\text{0.01}}$
200 0.70±${}_{\pm\text{0.01}}$ 0.61±${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.02}}$ 0.69±${}_{\pm\text{0.01}}$ 0.50±${}_{\pm\text{0.01}}$ 0.65±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.01}}$ 0.71±${}_{\pm\text{0.02}}$ 0.79±${}_{\pm\text{0.01}}$
500 0.70±${}_{\pm\text{0.01}}$ 0.62±${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.02}}$ 0.68±${}_{\pm\text{0.01}}$ 0.49±${}_{\pm\text{0.01}}$ 0.65±${}_{\pm\text{0.01}}$ 0.58±${}_{\pm\text{0.01}}$ 0.74±${}_{\pm\text{0.02}}$ 0.80±${}_{\pm\text{0.01}}$
stock 20 0.86±${}_{\pm\text{0.02}}$ 0.82±${}_{\pm\text{0.01}}$ 0.82±${}_{\pm\text{0.02}}$ 0.86±${}_{\pm\text{0.01}}$ 0.74±${}_{\pm\text{0.03}}$ 0.86±${}_{\pm\text{0.02}}$ 0.63±${}_{\pm\text{0.07}}$ 0.89±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$
50 0.92±${}_{\pm\text{0.01}}$ 0.86±${}_{\pm\text{0.01}}$ 0.88±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.84±${}_{\pm\text{0.02}}$ 0.91±${}_{\pm\text{0.01}}$ 0.68±${}_{\pm\text{0.04}}$ 0.93±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.00}}$
100 0.94±${}_{\pm\text{0.01}}$ 0.86±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.01}}$ 0.88±${}_{\pm\text{0.02}}$ 0.93±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.02}}$ 0.95±${}_{\pm\text{0.01}}$ 0.95±${}_{\pm\text{0.01}}$
200 0.95±${}_{\pm\text{0.01}}$ 0.86±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.00}}$ 0.63±${}_{\pm\text{0.01}}$ 0.96±${}_{\pm\text{0.00}}$ 0.95±${}_{\pm\text{0.00}}$
More than 10 classes energy 50 N/A 0.70±${}_{\pm\text{0.02}}$ 0.69±${}_{\pm\text{0.04}}$ 0.73±${}_{\pm\text{0.01}}$ 0.65±${}_{\pm\text{0.03}}$ 0.72±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.03}}$ N/A 0.78±${}_{\pm\text{0.01}}$
100 N/A 0.69±${}_{\pm\text{0.02}}$ 0.74±${}_{\pm\text{0.01}}$ 0.74±${}_{\pm\text{0.01}}$ 0.69±${}_{\pm\text{0.01}}$ 0.74±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.03}}$ N/A 0.81±${}_{\pm\text{0.01}}$
200 N/A 0.71±${}_{\pm\text{0.01}}$ 0.74±${}_{\pm\text{0.02}}$ 0.75±${}_{\pm\text{0.01}}$ 0.67±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.00}}$ 0.63±${}_{\pm\text{0.02}}$ N/A 0.83±${}_{\pm\text{0.01}}$
collins 100 N/A 0.85±${}_{\pm\text{0.01}}$ 0.89±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.82±${}_{\pm\text{0.04}}$ 0.90±${}_{\pm\text{0.01}}$ 0.65±${}_{\pm\text{0.04}}$ N/A 0.93±${}_{\pm\text{0.00}}$
200 0.93±${}_{\pm\text{0.00}}$ 0.83±${}_{\pm\text{0.02}}$ 0.91±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.00}}$ 0.80±${}_{\pm\text{0.03}}$ 0.92±${}_{\pm\text{0.00}}$ 0.63±${}_{\pm\text{0.05}}$ N/A 0.94±${}_{\pm\text{0.00}}$
texture 50 0.92±${}_{\pm\text{0.01}}$ 0.86±${}_{\pm\text{0.01}}$ 0.88±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.57±${}_{\pm\text{0.03}}$ 0.90±${}_{\pm\text{0.01}}$ 0.71±${}_{\pm\text{0.05}}$ N/A 0.93±${}_{\pm\text{0.01}}$
100 0.94±${}_{\pm\text{0.01}}$ 0.86±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.00}}$ 0.61±${}_{\pm\text{0.02}}$ 0.92±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.03}}$ N/A 0.95±${}_{\pm\text{0.00}}$
200 0.96±${}_{\pm\text{0.01}}$ 0.83±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.62±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.00}}$ 0.60±${}_{\pm\text{0.01}}$ N/A 0.96±${}_{\pm\text{0.00}}$
500 0.97±${}_{\pm\text{0.00}}$ 0.85±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.61±${}_{\pm\text{0.01}}$ 0.95±${}_{\pm\text{0.00}}$ 0.64±${}_{\pm\text{0.01}}$ N/A 0.97±${}_{\pm\text{0.00}}$
Average rank 3.91±${}_{\pm\text{1.65}}$ 6.97±${}_{\pm\text{0.92}}$ 5.76±${}_{\pm\text{1.00}}$ 3.97±${}_{\pm\text{1.07}}$ 8.33±${}_{\pm\text{1.05}}$ 4.15±${}_{\pm\text{0.94}}$ 7.55±${}_{\pm\text{2.22}}$ 3.21±${}_{\pm\text{2.16}}$ 1.15±${}_{\pm\text{0.36}}$
Table 18: $\chi^{2}$ test between real train data and synthetic data on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std result and average rank across datasets. A higher rank implies higher fidelity. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean result of other methods. We bold the highest result for each dataset of different sample sizes. TabEBM achieves the best overall performance against benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 0.02±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.03}}$ 0.19±${}_{\pm\text{0.07}}$ 0.02±${}_{\pm\text{0.01}}$ 0.05±${}_{\pm\text{0.03}}$ 0.02±${}_{\pm\text{0.01}}$ 0.33±${}_{\pm\text{0.23}}$ 0.92±${}_{\pm\text{0.07}}$
50 0.95±${}_{\pm\text{0.04}}$ 0.84±${}_{\pm\text{0.05}}$ 0.50±${}_{\pm\text{0.10}}$ 0.92±${}_{\pm\text{0.04}}$ 0.01±${}_{\pm\text{0.00}}$ 0.98±${}_{\pm\text{0.02}}$ 0.53±${}_{\pm\text{0.05}}$ 0.63±${}_{\pm\text{0.14}}$ 0.96±${}_{\pm\text{0.04}}$
100 0.86±${}_{\pm\text{0.06}}$ 0.62±${}_{\pm\text{0.07}}$ 0.53±${}_{\pm\text{0.09}}$ 0.83±${}_{\pm\text{0.10}}$ 0.01±${}_{\pm\text{0.00}}$ 0.89±${}_{\pm\text{0.03}}$ 0.27±${}_{\pm\text{0.07}}$ 0.70±${}_{\pm\text{0.10}}$ 0.91±${}_{\pm\text{0.06}}$
200 0.80±${}_{\pm\text{0.05}}$ 0.46±${}_{\pm\text{0.07}}$ 0.46±${}_{\pm\text{0.09}}$ 0.77±${}_{\pm\text{0.05}}$ 0.01±${}_{\pm\text{0.00}}$ 0.76±${}_{\pm\text{0.06}}$ ${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.08}}$ 0.81±${}_{\pm\text{0.08}}$
500 0.62±${}_{\pm\text{0.06}}$ 0.25±${}_{\pm\text{0.05}}$ 0.36±${}_{\pm\text{0.06}}$ 0.61±${}_{\pm\text{0.04}}$ 0.01±${}_{\pm\text{0.00}}$ 0.57±${}_{\pm\text{0.05}}$ 0.01±${}_{\pm\text{0.00}}$ 0.65±${}_{\pm\text{0.07}}$ 0.70±${}_{\pm\text{0.05}}$
fourier 20 N/A ${}_{\pm\text{0.02}}$ 0.05±${}_{\pm\text{0.02}}$ 0.15±${}_{\pm\text{0.05}}$ 0.37±${}_{\pm\text{0.04}}$ 0.35±${}_{\pm\text{0.05}}$ 0.36±${}_{\pm\text{0.06}}$ 1.00±${}_{\pm\text{0.00}}$ 1.00±${}_{\pm\text{0.00}}$
50 0.99±${}_{\pm\text{0.01}}$ 0.87±${}_{\pm\text{0.04}}$ 0.52±${}_{\pm\text{0.10}}$ 0.97±${}_{\pm\text{0.02}}$ 0.01±${}_{\pm\text{0.00}}$ 0.99±${}_{\pm\text{0.01}}$ 0.49±${}_{\pm\text{0.06}}$ 0.74±${}_{\pm\text{0.05}}$ 0.99±${}_{\pm\text{0.01}}$
100 0.96±${}_{\pm\text{0.02}}$ 0.75±${}_{\pm\text{0.04}}$ 0.69±${}_{\pm\text{0.07}}$ 0.95±${}_{\pm\text{0.02}}$ 0.01±${}_{\pm\text{0.00}}$ 0.98±${}_{\pm\text{0.02}}$ 0.16±${}_{\pm\text{0.05}}$ 0.82±${}_{\pm\text{0.06}}$ 0.97±${}_{\pm\text{0.03}}$
200 0.92±${}_{\pm\text{0.03}}$ 0.41±${}_{\pm\text{0.07}}$ 0.59±${}_{\pm\text{0.04}}$ 0.92±${}_{\pm\text{0.03}}$ 0.01±${}_{\pm\text{0.00}}$ 0.95±${}_{\pm\text{0.02}}$ 0.02±${}_{\pm\text{0.01}}$ 0.85±${}_{\pm\text{0.05}}$ 0.95±${}_{\pm\text{0.03}}$
500 0.80±${}_{\pm\text{0.06}}$ 0.14±${}_{\pm\text{0.04}}$ 0.59±${}_{\pm\text{0.06}}$ 0.76±${}_{\pm\text{0.07}}$ 0.01±${}_{\pm\text{0.00}}$ 0.84±${}_{\pm\text{0.04}}$ 0.01±${}_{\pm\text{0.00}}$ 0.81±${}_{\pm\text{0.04}}$ 0.84±${}_{\pm\text{0.02}}$
biodeg 20 0.23±${}_{\pm\text{0.06}}$ 0.21±${}_{\pm\text{0.06}}$ 0.16±${}_{\pm\text{0.04}}$ 0.28±${}_{\pm\text{0.06}}$ 0.08±${}_{\pm\text{0.04}}$ 0.28±${}_{\pm\text{0.07}}$ 0.12±${}_{\pm\text{0.06}}$ 0.29±${}_{\pm\text{0.10}}$ 0.71±${}_{\pm\text{0.06}}$
50 0.39±${}_{\pm\text{0.02}}$ 0.33±${}_{\pm\text{0.05}}$ 0.28±${}_{\pm\text{0.06}}$ 0.41±${}_{\pm\text{0.07}}$ 0.05±${}_{\pm\text{0.02}}$ 0.45±${}_{\pm\text{0.05}}$ 0.15±${}_{\pm\text{0.03}}$ 0.44±${}_{\pm\text{0.09}}$ 0.75±${}_{\pm\text{0.06}}$
100 0.35±${}_{\pm\text{0.06}}$ 0.26±${}_{\pm\text{0.06}}$ 0.30±${}_{\pm\text{0.08}}$ 0.38±${}_{\pm\text{0.06}}$ 0.04±${}_{\pm\text{0.01}}$ 0.42±${}_{\pm\text{0.05}}$ 0.08±${}_{\pm\text{0.04}}$ 0.37±${}_{\pm\text{0.13}}$ 0.67±${}_{\pm\text{0.14}}$
200 0.25±${}_{\pm\text{0.06}}$ 0.19±${}_{\pm\text{0.05}}$ 0.24±${}_{\pm\text{0.05}}$ 0.30±${}_{\pm\text{0.07}}$ 0.03±${}_{\pm\text{0.01}}$ 0.31±${}_{\pm\text{0.06}}$ 0.02±${}_{\pm\text{0.00}}$ 0.37±${}_{\pm\text{0.06}}$ 0.59±${}_{\pm\text{0.08}}$
500 0.22±${}_{\pm\text{0.08}}$ 0.10±${}_{\pm\text{0.04}}$ 0.17±${}_{\pm\text{0.05}}$ 0.25±${}_{\pm\text{0.07}}$ 0.02±${}_{\pm\text{0.00}}$ 0.23±${}_{\pm\text{0.06}}$ 0.02±${}_{\pm\text{0.00}}$ 0.26±${}_{\pm\text{0.08}}$ 0.44±${}_{\pm\text{0.09}}$
steel 20 0.37±${}_{\pm\text{0.08}}$ 0.32±${}_{\pm\text{0.05}}$ 0.32±${}_{\pm\text{0.04}}$ 0.40±${}_{\pm\text{0.08}}$ 0.04±${}_{\pm\text{0.02}}$ 0.40±${}_{\pm\text{0.08}}$ 0.23±${}_{\pm\text{0.07}}$ 0.34±${}_{\pm\text{0.13}}$ 0.81±${}_{\pm\text{0.05}}$
50 0.68±${}_{\pm\text{0.02}}$ 0.50±${}_{\pm\text{0.06}}$ 0.51±${}_{\pm\text{0.08}}$ 0.64±${}_{\pm\text{0.05}}$ 0.03±${}_{\pm\text{0.00}}$ 0.67±${}_{\pm\text{0.03}}$ 0.10±${}_{\pm\text{0.05}}$ 0.49±${}_{\pm\text{0.15}}$ 0.85±${}_{\pm\text{0.06}}$
100 0.61±${}_{\pm\text{0.09}}$ 0.37±${}_{\pm\text{0.07}}$ 0.47±${}_{\pm\text{0.08}}$ 0.61±${}_{\pm\text{0.04}}$ 0.03±${}_{\pm\text{0.00}}$ 0.61±${}_{\pm\text{0.07}}$ 0.06±${}_{\pm\text{0.02}}$ 0.51±${}_{\pm\text{0.07}}$ 0.81±${}_{\pm\text{0.08}}$
200 0.60±${}_{\pm\text{0.04}}$ 0.23±${}_{\pm\text{0.04}}$ 0.42±${}_{\pm\text{0.06}}$ 0.60±${}_{\pm\text{0.05}}$ 0.03±${}_{\pm\text{0.00}}$ 0.53±${}_{\pm\text{0.07}}$ 0.02±${}_{\pm\text{0.00}}$ 0.52±${}_{\pm\text{0.07}}$ 0.77±${}_{\pm\text{0.06}}$
500 0.46±${}_{\pm\text{0.06}}$ 0.17±${}_{\pm\text{0.07}}$ 0.35±${}_{\pm\text{0.07}}$ 0.55±${}_{\pm\text{0.06}}$ 0.03±${}_{\pm\text{0.00}}$ 0.42±${}_{\pm\text{0.04}}$ 0.02±${}_{\pm\text{0.00}}$ 0.48±${}_{\pm\text{0.10}}$ 0.69±${}_{\pm\text{0.09}}$
stock 20 0.56±${}_{\pm\text{0.13}}$ 0.50±${}_{\pm\text{0.12}}$ 0.40±${}_{\pm\text{0.12}}$ 0.63±${}_{\pm\text{0.08}}$ 0.11±${}_{\pm\text{0.03}}$ 0.55±${}_{\pm\text{0.12}}$ 0.43±${}_{\pm\text{0.16}}$ 0.56±${}_{\pm\text{0.20}}$ 1.00±${}_{\pm\text{0.00}}$
50 0.99±${}_{\pm\text{0.03}}$ 0.85±${}_{\pm\text{0.10}}$ 0.88±${}_{\pm\text{0.13}}$ 0.99±${}_{\pm\text{0.03}}$ 0.12±${}_{\pm\text{0.06}}$ 1.00±${}_{\pm\text{0.00}}$ 0.31±${}_{\pm\text{0.14}}$ 0.87±${}_{\pm\text{0.08}}$ 0.99±${}_{\pm\text{0.03}}$
100 1.00±${}_{\pm\text{0.00}}$ 0.64±${}_{\pm\text{0.14}}$ 0.89±${}_{\pm\text{0.12}}$ 0.99±${}_{\pm\text{0.03}}$ 0.16±${}_{\pm\text{0.13}}$ 1.00±${}_{\pm\text{0.00}}$ 0.13±${}_{\pm\text{0.05}}$ 0.91±${}_{\pm\text{0.10}}$ 0.99±${}_{\pm\text{0.03}}$
200 0.99±${}_{\pm\text{0.03}}$ 0.58±${}_{\pm\text{0.09}}$ 0.76±${}_{\pm\text{0.20}}$ 0.98±${}_{\pm\text{0.04}}$ 0.71±${}_{\pm\text{0.22}}$ 1.00±${}_{\pm\text{0.00}}$ 0.10±${}_{\pm\text{0.00}}$ 0.99±${}_{\pm\text{0.03}}$ 0.99±${}_{\pm\text{0.03}}$
More than 10 classes energy 50 N/A 0.14±${}_{\pm\text{0.08}}$ 0.26±${}_{\pm\text{0.09}}$ 0.35±${}_{\pm\text{0.09}}$ 0.17±${}_{\pm\text{0.06}}$ ${}_{\pm\text{0.14}}$ 0.19±${}_{\pm\text{0.10}}$ N/A 0.80±${}_{\pm\text{0.05}}$
100 N/A 0.09±${}_{\pm\text{0.10}}$ 0.32±${}_{\pm\text{0.12}}$ 0.35±${}_{\pm\text{0.08}}$ 0.13±${}_{\pm\text{0.06}}$ 0.30±${}_{\pm\text{0.09}}$ 0.06±${}_{\pm\text{0.07}}$ N/A 0.92±${}_{\pm\text{0.01}}$
200 N/A 0.15±${}_{\pm\text{0.12}}$ 0.30±${}_{\pm\text{0.09}}$ 0.39±${}_{\pm\text{0.12}}$ 0.15±${}_{\pm\text{0.07}}$ 0.28±${}_{\pm\text{0.08}}$ 0.03±${}_{\pm\text{0.05}}$ N/A 0.96±${}_{\pm\text{0.01}}$
collins 100 N/A 0.61±${}_{\pm\text{0.08}}$ 0.73±${}_{\pm\text{0.12}}$ 0.87±${}_{\pm\text{0.08}}$ 0.10±${}_{\pm\text{0.07}}$ 0.90±${}_{\pm\text{0.08}}$ 0.20±${}_{\pm\text{0.09}}$ N/A 0.89±${}_{\pm\text{0.04}}$
200 0.75±${}_{\pm\text{0.09}}$ 0.35±${}_{\pm\text{0.10}}$ 0.62±${}_{\pm\text{0.13}}$ 0.76±${}_{\pm\text{0.10}}$ 0.07±${}_{\pm\text{0.03}}$ 0.78±${}_{\pm\text{0.06}}$ 0.05±${}_{\pm\text{0.04}}$ N/A 0.80±${}_{\pm\text{0.09}}$
texture 50 0.90±${}_{\pm\text{0.12}}$ 0.79±${}_{\pm\text{0.11}}$ 0.63±${}_{\pm\text{0.14}}$ 0.94±${}_{\pm\text{0.09}}$ 0.02±${}_{\pm\text{0.00}}$ 0.99±${}_{\pm\text{0.02}}$ 0.47±${}_{\pm\text{0.13}}$ N/A 0.93±${}_{\pm\text{0.12}}$
100 0.97±${}_{\pm\text{0.04}}$ 0.51±${}_{\pm\text{0.13}}$ 0.67±${}_{\pm\text{0.12}}$ 0.92±${}_{\pm\text{0.07}}$ 0.02±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.09}}$ 0.17±${}_{\pm\text{0.08}}$ N/A 0.97±${}_{\pm\text{0.06}}$
200 0.86±${}_{\pm\text{0.11}}$ 0.28±${}_{\pm\text{0.11}}$ 0.55±${}_{\pm\text{0.16}}$ 0.89±${}_{\pm\text{0.10}}$ 0.02±${}_{\pm\text{0.00}}$ 0.93±${}_{\pm\text{0.07}}$ 0.00±${}_{\pm\text{0.01}}$ N/A 0.91±${}_{\pm\text{0.08}}$
500 0.74±${}_{\pm\text{0.17}}$ 0.21±${}_{\pm\text{0.09}}$ 0.51±${}_{\pm\text{0.08}}$ 0.73±${}_{\pm\text{0.07}}$ 0.02±${}_{\pm\text{0.00}}$ 0.73±${}_{\pm\text{0.07}}$ ${}_{\pm\text{0.00}}$ N/A 0.82±${}_{\pm\text{0.12}}$
Average rank 3.52±${}_{\pm\text{1.08}}$ 6.88±${}_{\pm\text{1.02}}$ 6.03±${}_{\pm\text{1.10}}$ 3.55±${}_{\pm\text{1.03}}$ 8.30±${}_{\pm\text{1.02}}$ 2.82±${}_{\pm\text{1.76}}$ 8.18±${}_{\pm\text{0.88}}$ 4.27±${}_{\pm\text{1.62}}$ 1.45±${}_{\pm\text{0.71}}$
C.6.2 Similarity between Real Test Data and Synthetic Data
Table 19: Inverse KL between real test data and synthetic data on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std result and average rank across datasets. A higher rank implies higher fidelity. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean result of other methods. We bold the highest result for each dataset of different sample sizes. TabEBM achieves the best overall performance against benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 0.26±${}_{\pm\text{0.04}}$ 0.26±${}_{\pm\text{0.03}}$ 0.32±${}_{\pm\text{0.03}}$ 0.09±${}_{\pm\text{0.01}}$ 0.34±${}_{\pm\text{0.03}}$ 0.08±${}_{\pm\text{0.00}}$ 0.35±${}_{\pm\text{0.07}}$ 0.52±${}_{\pm\text{0.06}}$
50 0.78±${}_{\pm\text{0.02}}$ 0.82±${}_{\pm\text{0.02}}$ 0.63±${}_{\pm\text{0.06}}$ 0.71±${}_{\pm\text{0.04}}$ 0.08±${}_{\pm\text{0.00}}$ 0.80±${}_{\pm\text{0.03}}$ 0.52±${}_{\pm\text{0.03}}$ 0.62±${}_{\pm\text{0.04}}$ 0.75±${}_{\pm\text{0.03}}$
100 0.88±${}_{\pm\text{0.02}}$ 0.80±${}_{\pm\text{0.02}}$ 0.76±${}_{\pm\text{0.04}}$ 0.83±${}_{\pm\text{0.03}}$ 0.08±${}_{\pm\text{0.00}}$ 0.87±${}_{\pm\text{0.02}}$ 0.33±${}_{\pm\text{0.05}}$ 0.78±${}_{\pm\text{0.03}}$ 0.85±${}_{\pm\text{0.02}}$
200 0.92±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.03}}$ 0.81±${}_{\pm\text{0.03}}$ 0.91±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.00}}$ 0.91±${}_{\pm\text{0.01}}$ 0.12±${}_{\pm\text{0.01}}$ 0.89±${}_{\pm\text{0.02}}$ 0.92±${}_{\pm\text{0.01}}$
500 0.94±${}_{\pm\text{0.00}}$ 0.68±${}_{\pm\text{0.02}}$ 0.86±${}_{\pm\text{0.02}}$ 0.93±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.00}}$ 0.13±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.00}}$ 0.95±${}_{\pm\text{0.01}}$
fourier 20 N/A 0.14±${}_{\pm\text{0.02}}$ 0.18±${}_{\pm\text{0.01}}$ 0.20±${}_{\pm\text{0.01}}$ 0.21±${}_{\pm\text{0.02}}$ 0.21±${}_{\pm\text{0.01}}$ 0.22±${}_{\pm\text{0.02}}$ 0.24±${}_{\pm\text{0.01}}$ 0.48±${}_{\pm\text{0.03}}$
50 0.84±${}_{\pm\text{0.02}}$ 0.78±${}_{\pm\text{0.03}}$ 0.62±${}_{\pm\text{0.05}}$ 0.77±${}_{\pm\text{0.03}}$ 0.07±${}_{\pm\text{0.00}}$ 0.85±${}_{\pm\text{0.03}}$ 0.48±${}_{\pm\text{0.06}}$ 0.66±${}_{\pm\text{0.03}}$ 0.79±${}_{\pm\text{0.03}}$
100 0.91±${}_{\pm\text{0.02}}$ 0.72±${}_{\pm\text{0.05}}$ 0.76±${}_{\pm\text{0.03}}$ 0.88±${}_{\pm\text{0.02}}$ 0.08±${}_{\pm\text{0.00}}$ 0.90±${}_{\pm\text{0.01}}$ 0.20±${}_{\pm\text{0.05}}$ 0.81±${}_{\pm\text{0.02}}$ 0.88±${}_{\pm\text{0.02}}$
200 0.94±${}_{\pm\text{0.01}}$ 0.58±${}_{\pm\text{0.02}}$ 0.79±${}_{\pm\text{0.03}}$ 0.93±${}_{\pm\text{0.01}}$ 0.09±${}_{\pm\text{0.00}}$ 0.93±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.02}}$ 0.93±${}_{\pm\text{0.01}}$
500 0.96±${}_{\pm\text{0.00}}$ 0.50±${}_{\pm\text{0.02}}$ 0.88±${}_{\pm\text{0.02}}$ 0.93±${}_{\pm\text{0.01}}$ ${}_{\pm\text{0.01}}$ 0.95±${}_{\pm\text{0.01}}$ 0.09±${}_{\pm\text{0.00}}$ 0.95±${}_{\pm\text{0.01}}$ 0.96±${}_{\pm\text{0.00}}$
biodeg 20 0.43±${}_{\pm\text{0.03}}$ 0.44±${}_{\pm\text{0.04}}$ 0.41±${}_{\pm\text{0.03}}$ 0.41±${}_{\pm\text{0.04}}$ 0.33±${}_{\pm\text{0.03}}$ 0.44±${}_{\pm\text{0.03}}$ 0.36±${}_{\pm\text{0.03}}$ 0.45±${}_{\pm\text{0.02}}$ 0.57±${}_{\pm\text{0.03}}$
50 0.60±${}_{\pm\text{0.05}}$ 0.59±${}_{\pm\text{0.04}}$ 0.53±${}_{\pm\text{0.04}}$ 0.55±${}_{\pm\text{0.04}}$ 0.31±${}_{\pm\text{0.03}}$ 0.57±${}_{\pm\text{0.03}}$ 0.41±${}_{\pm\text{0.03}}$ 0.60±${}_{\pm\text{0.05}}$ 0.71±${}_{\pm\text{0.04}}$
100 0.65±${}_{\pm\text{0.04}}$ 0.65±${}_{\pm\text{0.02}}$ 0.62±${}_{\pm\text{0.03}}$ 0.63±${}_{\pm\text{0.03}}$ 0.31±${}_{\pm\text{0.03}}$ 0.64±${}_{\pm\text{0.03}}$ 0.37±${}_{\pm\text{0.03}}$ 0.65±${}_{\pm\text{0.03}}$ 0.77±${}_{\pm\text{0.02}}$
200 0.71±${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.03}}$ 0.66±${}_{\pm\text{0.04}}$ 0.65±${}_{\pm\text{0.03}}$ 0.31±${}_{\pm\text{0.04}}$ 0.68±${}_{\pm\text{0.02}}$ 0.33±${}_{\pm\text{0.02}}$ 0.73±${}_{\pm\text{0.04}}$ 0.83±${}_{\pm\text{0.02}}$
500 0.77±${}_{\pm\text{0.03}}$ 0.64±${}_{\pm\text{0.02}}$ 0.69±${}_{\pm\text{0.03}}$ 0.72±${}_{\pm\text{0.04}}$ 0.25±${}_{\pm\text{0.03}}$ 0.69±${}_{\pm\text{0.03}}$ 0.35±${}_{\pm\text{0.01}}$ 0.76±${}_{\pm\text{0.05}}$ 0.88±${}_{\pm\text{0.02}}$
steel 20 0.47±${}_{\pm\text{0.03}}$ 0.45±${}_{\pm\text{0.03}}$ 0.42±${}_{\pm\text{0.02}}$ 0.44±${}_{\pm\text{0.02}}$ 0.23±${}_{\pm\text{0.01}}$ 0.46±${}_{\pm\text{0.04}}$ 0.37±${}_{\pm\text{0.04}}$ 0.42±${}_{\pm\text{0.04}}$ 0.70±${}_{\pm\text{0.03}}$
50 0.65±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.03}}$ 0.58±${}_{\pm\text{0.03}}$ 0.60±${}_{\pm\text{0.04}}$ 0.21±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.03}}$ ${}_{\pm\text{0.03}}$ 0.62±${}_{\pm\text{0.05}}$ 0.83±${}_{\pm\text{0.03}}$
100 0.70±${}_{\pm\text{0.02}}$ 0.59±${}_{\pm\text{0.02}}$ 0.62±${}_{\pm\text{0.03}}$ 0.66±${}_{\pm\text{0.03}}$ 0.21±${}_{\pm\text{0.01}}$ 0.66±${}_{\pm\text{0.02}}$ 0.29±${}_{\pm\text{0.02}}$ 0.72±${}_{\pm\text{0.04}}$ 0.89±${}_{\pm\text{0.02}}$
200 0.73±${}_{\pm\text{0.01}}$ 0.55±${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.02}}$ 0.69±${}_{\pm\text{0.02}}$ 0.22±${}_{\pm\text{0.01}}$ 0.68±${}_{\pm\text{0.02}}$ 0.25±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.03}}$ 0.91±${}_{\pm\text{0.02}}$
500 0.74±${}_{\pm\text{0.01}}$ 0.55±${}_{\pm\text{0.04}}$ 0.66±${}_{\pm\text{0.02}}$ 0.69±${}_{\pm\text{0.01}}$ 0.22±${}_{\pm\text{0.01}}$ 0.70±${}_{\pm\text{0.01}}$ 0.27±${}_{\pm\text{0.01}}$ 0.78±${}_{\pm\text{0.05}}$ 0.93±${}_{\pm\text{0.02}}$
stock 20 0.50±${}_{\pm\text{0.09}}$ 0.50±${}_{\pm\text{0.09}}$ 0.41±${}_{\pm\text{0.05}}$ 0.47±${}_{\pm\text{0.08}}$ 0.25±${}_{\pm\text{0.04}}$ 0.54±${}_{\pm\text{0.09}}$ 0.40±${}_{\pm\text{0.13}}$ 0.35±${}_{\pm\text{0.05}}$ 0.65±${}_{\pm\text{0.09}}$
50 0.80±${}_{\pm\text{0.06}}$ 0.68±${}_{\pm\text{0.06}}$ 0.68±${}_{\pm\text{0.05}}$ 0.76±${}_{\pm\text{0.03}}$ 0.34±${}_{\pm\text{0.05}}$ 0.86±${}_{\pm\text{0.04}}$ 0.33±${}_{\pm\text{0.11}}$ 0.69±${}_{\pm\text{0.09}}$ 0.85±${}_{\pm\text{0.05}}$
100 0.86±${}_{\pm\text{0.04}}$ 0.61±${}_{\pm\text{0.05}}$ 0.73±${}_{\pm\text{0.06}}$ 0.85±${}_{\pm\text{0.06}}$ 0.44±${}_{\pm\text{0.07}}$ 0.90±${}_{\pm\text{0.02}}$ 0.18±${}_{\pm\text{0.04}}$ 0.84±${}_{\pm\text{0.05}}$ 0.91±${}_{\pm\text{0.04}}$
200 0.92±${}_{\pm\text{0.02}}$ 0.59±${}_{\pm\text{0.05}}$ 0.75±${}_{\pm\text{0.07}}$ 0.90±${}_{\pm\text{0.03}}$ 0.67±${}_{\pm\text{0.09}}$ 0.94±${}_{\pm\text{0.01}}$ 0.15±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.03}}$ 0.96±${}_{\pm\text{0.02}}$
More than 10 classes energy 50 N/A 0.26±${}_{\pm\text{0.06}}$ 0.33±${}_{\pm\text{0.06}}$ 0.36±${}_{\pm\text{0.08}}$ 0.23±${}_{\pm\text{0.05}}$ 0.46±${}_{\pm\text{0.10}}$ 0.22±${}_{\pm\text{0.07}}$ N/A 0.77±${}_{\pm\text{0.03}}$
100 N/A 0.27±${}_{\pm\text{0.06}}$ 0.40±${}_{\pm\text{0.08}}$ 0.43±${}_{\pm\text{0.06}}$ 0.22±${}_{\pm\text{0.06}}$ 0.39±${}_{\pm\text{0.08}}$ 0.16±${}_{\pm\text{0.05}}$ N/A 0.87±${}_{\pm\text{0.01}}$
200 N/A 0.30±${}_{\pm\text{0.07}}$ 0.41±${}_{\pm\text{0.07}}$ 0.46±${}_{\pm\text{0.09}}$ 0.24±${}_{\pm\text{0.06}}$ 0.39±${}_{\pm\text{0.08}}$ 0.12±${}_{\pm\text{0.04}}$ N/A 0.89±${}_{\pm\text{0.01}}$
collins 100 N/A 0.68±${}_{\pm\text{0.03}}$ 0.75±${}_{\pm\text{0.04}}$ 0.79±${}_{\pm\text{0.03}}$ 0.43±${}_{\pm\text{0.07}}$ 0.81±${}_{\pm\text{0.02}}$ 0.28±${}_{\pm\text{0.07}}$ N/A 0.78±${}_{\pm\text{0.02}}$
200 0.87±${}_{\pm\text{0.02}}$ 0.62±${}_{\pm\text{0.03}}$ 0.81±${}_{\pm\text{0.03}}$ 0.88±${}_{\pm\text{0.02}}$ 0.44±${}_{\pm\text{0.04}}$ 0.88±${}_{\pm\text{0.02}}$ 0.22±${}_{\pm\text{0.09}}$ N/A 0.87±${}_{\pm\text{0.01}}$
texture 50 0.82±${}_{\pm\text{0.04}}$ 0.80±${}_{\pm\text{0.04}}$ 0.70±${}_{\pm\text{0.05}}$ 0.80±${}_{\pm\text{0.07}}$ 0.11±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.01}}$ 0.48±${}_{\pm\text{0.11}}$ N/A 0.87±${}_{\pm\text{0.03}}$
100 0.89±${}_{\pm\text{0.02}}$ 0.69±${}_{\pm\text{0.03}}$ 0.79±${}_{\pm\text{0.06}}$ 0.89±${}_{\pm\text{0.02}}$ 0.13±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.23±${}_{\pm\text{0.06}}$ N/A 0.92±${}_{\pm\text{0.02}}$
200 0.93±${}_{\pm\text{0.02}}$ 0.58±${}_{\pm\text{0.05}}$ 0.79±${}_{\pm\text{0.05}}$ 0.92±${}_{\pm\text{0.01}}$ 0.14±${}_{\pm\text{0.01}}$ 0.95±${}_{\pm\text{0.01}}$ 0.10±${}_{\pm\text{0.02}}$ N/A 0.95±${}_{\pm\text{0.01}}$
500 0.96±${}_{\pm\text{0.01}}$ 0.64±${}_{\pm\text{0.06}}$ 0.85±${}_{\pm\text{0.04}}$ 0.93±${}_{\pm\text{0.01}}$ 0.15±${}_{\pm\text{0.01}}$ 0.96±${}_{\pm\text{0.00}}$ 0.12±${}_{\pm\text{0.01}}$ N/A 0.97±${}_{\pm\text{0.01}}$
Average rank 2.94±${}_{\pm\text{1.31}}$ 6.03±${}_{\pm\text{1.69}}$ 5.82±${}_{\pm\text{1.07}}$ 4.39±${}_{\pm\text{1.30}}$ 8.48±${}_{\pm\text{0.71}}$ 3.00±${}_{\pm\text{1.52}}$ 8.24±${}_{\pm\text{0.94}}$ 4.45±${}_{\pm\text{1.91}}$ 1.64±${}_{\pm\text{1.03}}$
Table 20: KS test between real test data and synthetic data on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std result and average rank across datasets. A higher rank implies higher fidelity. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean result of other methods. We bold the highest result for each dataset of different sample sizes. TabEBM achieves the best overall performance against benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 0.69±${}_{\pm\text{0.01}}$ 0.72±${}_{\pm\text{0.02}}$ 0.75±${}_{\pm\text{0.01}}$ 0.61±${}_{\pm\text{0.01}}$ 0.74±${}_{\pm\text{0.01}}$ 0.41±${}_{\pm\text{0.02}}$ 0.77±${}_{\pm\text{0.02}}$ 0.81±${}_{\pm\text{0.01}}$
50 0.88±${}_{\pm\text{0.01}}$ 0.87±${}_{\pm\text{0.01}}$ 0.87±${}_{\pm\text{0.01}}$ 0.88±${}_{\pm\text{0.01}}$ 0.61±${}_{\pm\text{0.01}}$ 0.89±${}_{\pm\text{0.01}}$ 0.73±${}_{\pm\text{0.02}}$ 0.86±${}_{\pm\text{0.01}}$ 0.88±${}_{\pm\text{0.01}}$
100 0.90±${}_{\pm\text{0.01}}$ 0.87±${}_{\pm\text{0.01}}$ 0.89±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.60±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.65±${}_{\pm\text{0.02}}$ 0.89±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$
200 0.92±${}_{\pm\text{0.01}}$ 0.86±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.59±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.01}}$ 0.58±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.01}}$
500 0.92±${}_{\pm\text{0.00}}$ 0.82±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.00}}$ 0.91±${}_{\pm\text{0.00}}$ 0.58±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.00}}$ 0.63±${}_{\pm\text{0.00}}$ 0.92±${}_{\pm\text{0.00}}$ 0.93±${}_{\pm\text{0.00}}$
fourier 20 N/A 0.67±${}_{\pm\text{0.03}}$ 0.69±${}_{\pm\text{0.02}}$ 0.73±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.01}}$ 0.76±${}_{\pm\text{0.01}}$ 0.81±${}_{\pm\text{0.01}}$
50 0.89±${}_{\pm\text{0.00}}$ 0.86±${}_{\pm\text{0.01}}$ 0.88±${}_{\pm\text{0.01}}$ 0.89±${}_{\pm\text{0.01}}$ 0.64±${}_{\pm\text{0.00}}$ 0.91±${}_{\pm\text{0.00}}$ 0.73±${}_{\pm\text{0.02}}$ 0.87±${}_{\pm\text{0.01}}$ 0.89±${}_{\pm\text{0.01}}$
100 0.92±${}_{\pm\text{0.00}}$ 0.85±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.00}}$ 0.91±${}_{\pm\text{0.00}}$ 0.62±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.00}}$ 0.61±${}_{\pm\text{0.02}}$ 0.90±${}_{\pm\text{0.00}}$ 0.91±${}_{\pm\text{0.00}}$
200 0.93±${}_{\pm\text{0.00}}$ 0.82±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.62±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.60±${}_{\pm\text{0.03}}$ 0.93±${}_{\pm\text{0.00}}$ 0.93±${}_{\pm\text{0.00}}$
500 0.94±${}_{\pm\text{0.00}}$ 0.80±${}_{\pm\text{0.00}}$ 0.92±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.62±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.00}}$ 0.63±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.00}}$ 0.94±${}_{\pm\text{0.00}}$
biodeg 20 0.61±${}_{\pm\text{0.03}}$ 0.61±${}_{\pm\text{0.03}}$ 0.63±${}_{\pm\text{0.03}}$ 0.63±${}_{\pm\text{0.03}}$ 0.55±${}_{\pm\text{0.04}}$ 0.62±${}_{\pm\text{0.03}}$ 0.64±${}_{\pm\text{0.03}}$ 0.56±${}_{\pm\text{0.03}}$ 0.67±${}_{\pm\text{0.02}}$
50 0.56±${}_{\pm\text{0.03}}$ 0.57±${}_{\pm\text{0.03}}$ 0.58±${}_{\pm\text{0.03}}$ 0.59±${}_{\pm\text{0.03}}$ 0.48±${}_{\pm\text{0.03}}$ 0.58±${}_{\pm\text{0.03}}$ 0.63±${}_{\pm\text{0.02}}$ 0.59±${}_{\pm\text{0.04}}$ 0.71±${}_{\pm\text{0.01}}$
100 0.55±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.02}}$ 0.57±${}_{\pm\text{0.03}}$ 0.58±${}_{\pm\text{0.02}}$ 0.46±${}_{\pm\text{0.03}}$ 0.58±${}_{\pm\text{0.02}}$ 0.59±${}_{\pm\text{0.03}}$ 0.58±${}_{\pm\text{0.02}}$ 0.71±${}_{\pm\text{0.01}}$
200 0.53±${}_{\pm\text{0.01}}$ 0.51±${}_{\pm\text{0.02}}$ 0.53±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.02}}$ 0.43±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.01}}$ 0.57±${}_{\pm\text{0.02}}$ 0.72±${}_{\pm\text{0.01}}$
500 0.53±${}_{\pm\text{0.00}}$ 0.49±${}_{\pm\text{0.01}}$ 0.51±${}_{\pm\text{0.01}}$ 0.54±${}_{\pm\text{0.01}}$ 0.43±${}_{\pm\text{0.01}}$ 0.55±${}_{\pm\text{0.01}}$ 0.56±${}_{\pm\text{0.01}}$ 0.57±${}_{\pm\text{0.01}}$ 0.72±${}_{\pm\text{0.00}}$
steel 20 0.65±${}_{\pm\text{0.03}}$ 0.62±${}_{\pm\text{0.03}}$ 0.63±${}_{\pm\text{0.03}}$ 0.65±${}_{\pm\text{0.02}}$ 0.54±${}_{\pm\text{0.01}}$ 0.64±${}_{\pm\text{0.02}}$ 0.56±${}_{\pm\text{0.03}}$ 0.64±${}_{\pm\text{0.02}}$ 0.73±${}_{\pm\text{0.02}}$
50 0.68±${}_{\pm\text{0.02}}$ 0.64±${}_{\pm\text{0.03}}$ 0.65±${}_{\pm\text{0.03}}$ 0.67±${}_{\pm\text{0.02}}$ 0.51±${}_{\pm\text{0.01}}$ 0.68±${}_{\pm\text{0.03}}$ 0.56±${}_{\pm\text{0.03}}$ 0.70±${}_{\pm\text{0.02}}$ 0.77±${}_{\pm\text{0.01}}$
100 0.68±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.02}}$ 0.65±${}_{\pm\text{0.01}}$ 0.66±${}_{\pm\text{0.01}}$ 0.50±${}_{\pm\text{0.01}}$ 0.66±${}_{\pm\text{0.02}}$ 0.55±${}_{\pm\text{0.02}}$ 0.70±${}_{\pm\text{0.03}}$ 0.78±${}_{\pm\text{0.01}}$
200 0.69±${}_{\pm\text{0.01}}$ 0.61±${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.02}}$ 0.68±${}_{\pm\text{0.02}}$ 0.50±${}_{\pm\text{0.01}}$ 0.65±${}_{\pm\text{0.02}}$ 0.54±${}_{\pm\text{0.01}}$ 0.71±${}_{\pm\text{0.02}}$ 0.78±${}_{\pm\text{0.01}}$
500 0.69±${}_{\pm\text{0.01}}$ 0.62±${}_{\pm\text{0.02}}$ 0.66±${}_{\pm\text{0.02}}$ 0.68±${}_{\pm\text{0.01}}$ 0.50±${}_{\pm\text{0.01}}$ 0.64±${}_{\pm\text{0.01}}$ 0.58±${}_{\pm\text{0.01}}$ 0.73±${}_{\pm\text{0.02}}$ 0.79±${}_{\pm\text{0.01}}$
stock 20 0.83±${}_{\pm\text{0.04}}$ 0.81±${}_{\pm\text{0.04}}$ 0.82±${}_{\pm\text{0.04}}$ 0.84±${}_{\pm\text{0.04}}$ 0.74±${}_{\pm\text{0.04}}$ 0.84±${}_{\pm\text{0.04}}$ 0.64±${}_{\pm\text{0.06}}$ 0.83±${}_{\pm\text{0.04}}$ 0.86±${}_{\pm\text{0.04}}$
50 0.89±${}_{\pm\text{0.02}}$ 0.84±${}_{\pm\text{0.02}}$ 0.87±${}_{\pm\text{0.03}}$ 0.88±${}_{\pm\text{0.02}}$ 0.83±${}_{\pm\text{0.02}}$ 0.89±${}_{\pm\text{0.02}}$ 0.67±${}_{\pm\text{0.05}}$ 0.88±${}_{\pm\text{0.02}}$ 0.89±${}_{\pm\text{0.02}}$
100 0.91±${}_{\pm\text{0.02}}$ 0.84±${}_{\pm\text{0.02}}$ 0.88±${}_{\pm\text{0.02}}$ 0.90±${}_{\pm\text{0.02}}$ 0.87±${}_{\pm\text{0.03}}$ 0.90±${}_{\pm\text{0.02}}$ 0.62±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.02}}$ 0.91±${}_{\pm\text{0.02}}$
200 0.92±${}_{\pm\text{0.02}}$ 0.84±${}_{\pm\text{0.02}}$ 0.88±${}_{\pm\text{0.02}}$ 0.90±${}_{\pm\text{0.02}}$ 0.90±${}_{\pm\text{0.02}}$ 0.91±${}_{\pm\text{0.02}}$ 0.63±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.02}}$ 0.92±${}_{\pm\text{0.02}}$
More than 10 classes energy 50 N/A 0.69±${}_{\pm\text{0.02}}$ 0.70±${}_{\pm\text{0.03}}$ 0.72±${}_{\pm\text{0.02}}$ 0.64±${}_{\pm\text{0.02}}$ 0.72±${}_{\pm\text{0.01}}$ 0.62±${}_{\pm\text{0.03}}$ N/A 0.76±${}_{\pm\text{0.01}}$
100 N/A 0.69±${}_{\pm\text{0.03}}$ 0.74±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.01}}$ 0.69±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.01}}$ 0.64±${}_{\pm\text{0.03}}$ N/A 0.81±${}_{\pm\text{0.01}}$
200 N/A 0.71±${}_{\pm\text{0.01}}$ 0.74±${}_{\pm\text{0.01}}$ 0.75±${}_{\pm\text{0.01}}$ 0.67±${}_{\pm\text{0.02}}$ 0.76±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.02}}$ N/A 0.83±${}_{\pm\text{0.01}}$
collins 100 N/A 0.83±${}_{\pm\text{0.01}}$ 0.88±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.81±${}_{\pm\text{0.04}}$ 0.90±${}_{\pm\text{0.01}}$ 0.65±${}_{\pm\text{0.03}}$ N/A 0.90±${}_{\pm\text{0.01}}$
200 0.91±${}_{\pm\text{0.01}}$ 0.82±${}_{\pm\text{0.02}}$ 0.90±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.80±${}_{\pm\text{0.02}}$ 0.91±${}_{\pm\text{0.01}}$ 0.63±${}_{\pm\text{0.04}}$ N/A 0.93±${}_{\pm\text{0.01}}$
texture 50 0.90±${}_{\pm\text{0.01}}$ 0.87±${}_{\pm\text{0.02}}$ 0.88±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.55±${}_{\pm\text{0.04}}$ 0.92±${}_{\pm\text{0.01}}$ 0.73±${}_{\pm\text{0.05}}$ N/A 0.92±${}_{\pm\text{0.01}}$
100 0.92±${}_{\pm\text{0.01}}$ 0.86±${}_{\pm\text{0.01}}$ 0.91±${}_{\pm\text{0.01}}$ 0.92±${}_{\pm\text{0.01}}$ 0.60±${}_{\pm\text{0.02}}$ 0.93±${}_{\pm\text{0.01}}$ 0.64±${}_{\pm\text{0.03}}$ N/A 0.94±${}_{\pm\text{0.01}}$
200 0.94±${}_{\pm\text{0.01}}$ 0.82±${}_{\pm\text{0.01}}$ 0.90±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.00}}$ 0.62±${}_{\pm\text{0.01}}$ 0.94±${}_{\pm\text{0.00}}$ 0.60±${}_{\pm\text{0.01}}$ N/A 0.94±${}_{\pm\text{0.00}}$
500 0.95±${}_{\pm\text{0.00}}$ 0.84±${}_{\pm\text{0.02}}$ 0.91±${}_{\pm\text{0.01}}$ 0.93±${}_{\pm\text{0.01}}$ 0.61±${}_{\pm\text{0.02}}$ 0.95±${}_{\pm\text{0.00}}$ 0.64±${}_{\pm\text{0.01}}$ N/A 0.95±${}_{\pm\text{0.00}}$
Average rank 3.97±${}_{\pm\text{1.73}}$ 7.09±${}_{\pm\text{0.72}}$ 5.52±${}_{\pm\text{0.97}}$ 3.76±${}_{\pm\text{1.12}}$ 8.27±${}_{\pm\text{1.21}}$ 3.27±${}_{\pm\text{1.53}}$ 7.45±${}_{\pm\text{2.36}}$ 4.39±${}_{\pm\text{1.93}}$ 1.27±${}_{\pm\text{0.72}}$
Table 21: $\chi^{2}$ test between real test data and synthetic data on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std result and average rank across datasets. A higher rank implies higher fidelity. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean result of other methods. We bold the highest result for each dataset of different sample sizes. TabEBM achieves the best overall performance against benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.02±${}_{\pm\text{0.01}}$ 0.06±${}_{\pm\text{0.03}}$
50 0.26±${}_{\pm\text{0.05}}$ 0.32±${}_{\pm\text{0.06}}$ 0.09±${}_{\pm\text{0.04}}$ 0.17±${}_{\pm\text{0.04}}$ 0.01±${}_{\pm\text{0.00}}$ 0.25±${}_{\pm\text{0.04}}$ 0.22±${}_{\pm\text{0.05}}$ 0.05±${}_{\pm\text{0.03}}$ 0.15±${}_{\pm\text{0.04}}$
100 0.39±${}_{\pm\text{0.04}}$ 0.33±${}_{\pm\text{0.05}}$ 0.18±${}_{\pm\text{0.05}}$ 0.29±${}_{\pm\text{0.09}}$ 0.01±${}_{\pm\text{0.00}}$ 0.34±${}_{\pm\text{0.04}}$ 0.13±${}_{\pm\text{0.04}}$ 0.12±${}_{\pm\text{0.04}}$ 0.26±${}_{\pm\text{0.05}}$
200 0.48±${}_{\pm\text{0.06}}$ 0.30±${}_{\pm\text{0.07}}$ 0.28±${}_{\pm\text{0.06}}$ 0.48±${}_{\pm\text{0.04}}$ 0.01±${}_{\pm\text{0.00}}$ 0.44±${}_{\pm\text{0.08}}$ 0.02±${}_{\pm\text{0.01}}$ 0.34±${}_{\pm\text{0.08}}$ 0.43±${}_{\pm\text{0.06}}$
500 0.54±${}_{\pm\text{0.05}}$ 0.24±${}_{\pm\text{0.07}}$ 0.33±${}_{\pm\text{0.06}}$ 0.60±${}_{\pm\text{0.04}}$ 0.01±${}_{\pm\text{0.00}}$ 0.55±${}_{\pm\text{0.04}}$ 0.01±${}_{\pm\text{0.00}}$ 0.53±${}_{\pm\text{0.05}}$ 0.58±${}_{\pm\text{0.06}}$
fourier 20 N/A ${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.01±${}_{\pm\text{0.00}}$ 0.10±${}_{\pm\text{0.04}}$
50 0.42±${}_{\pm\text{0.08}}$ 0.45±${}_{\pm\text{0.07}}$ 0.15±${}_{\pm\text{0.06}}$ 0.31±${}_{\pm\text{0.07}}$ 0.01±${}_{\pm\text{0.00}}$ 0.39±${}_{\pm\text{0.08}}$ 0.29±${}_{\pm\text{0.08}}$ 0.10±${}_{\pm\text{0.04}}$ 0.33±${}_{\pm\text{0.06}}$
100 0.58±${}_{\pm\text{0.08}}$ 0.44±${}_{\pm\text{0.07}}$ 0.32±${}_{\pm\text{0.07}}$ 0.52±${}_{\pm\text{0.09}}$ 0.01±${}_{\pm\text{0.00}}$ 0.48±${}_{\pm\text{0.07}}$ 0.08±${}_{\pm\text{0.04}}$ 0.26±${}_{\pm\text{0.07}}$ 0.49±${}_{\pm\text{0.07}}$
200 0.67±${}_{\pm\text{0.05}}$ 0.28±${}_{\pm\text{0.03}}$ 0.38±${}_{\pm\text{0.06}}$ 0.68±${}_{\pm\text{0.04}}$ 0.01±${}_{\pm\text{0.00}}$ 0.60±${}_{\pm\text{0.04}}$ 0.02±${}_{\pm\text{0.01}}$ 0.46±${}_{\pm\text{0.06}}$ 0.60±${}_{\pm\text{0.05}}$
500 0.76±${}_{\pm\text{0.04}}$ 0.15±${}_{\pm\text{0.06}}$ 0.55±${}_{\pm\text{0.08}}$ 0.72±${}_{\pm\text{0.07}}$ 0.01±${}_{\pm\text{0.00}}$ 0.71±${}_{\pm\text{0.05}}$ 0.01±${}_{\pm\text{0.00}}$ 0.69±${}_{\pm\text{0.04}}$ 0.74±${}_{\pm\text{0.03}}$
biodeg 20 0.05±${}_{\pm\text{0.03}}$ 0.05±${}_{\pm\text{0.04}}$ 0.04±${}_{\pm\text{0.03}}$ 0.04±${}_{\pm\text{0.02}}$ 0.03±${}_{\pm\text{0.01}}$ 0.05±${}_{\pm\text{0.03}}$ 0.03±${}_{\pm\text{0.01}}$ ${}_{\pm\text{0.01}}$ 0.13±${}_{\pm\text{0.02}}$
50 0.13±${}_{\pm\text{0.06}}$ 0.12±${}_{\pm\text{0.04}}$ 0.07±${}_{\pm\text{0.02}}$ 0.08±${}_{\pm\text{0.04}}$ 0.02±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.05}}$ 0.06±${}_{\pm\text{0.03}}$ 0.08±${}_{\pm\text{0.04}}$ 0.14±${}_{\pm\text{0.05}}$
100 0.15±${}_{\pm\text{0.04}}$ 0.14±${}_{\pm\text{0.05}}$ 0.10±${}_{\pm\text{0.03}}$ 0.12±${}_{\pm\text{0.05}}$ 0.02±${}_{\pm\text{0.00}}$ 0.12±${}_{\pm\text{0.04}}$ 0.04±${}_{\pm\text{0.02}}$ 0.07±${}_{\pm\text{0.03}}$ 0.18±${}_{\pm\text{0.04}}$
200 0.18±${}_{\pm\text{0.03}}$ 0.15±${}_{\pm\text{0.04}}$ 0.14±${}_{\pm\text{0.03}}$ 0.19±${}_{\pm\text{0.05}}$ 0.03±${}_{\pm\text{0.01}}$ 0.18±${}_{\pm\text{0.03}}$ 0.02±${}_{\pm\text{0.00}}$ 0.14±${}_{\pm\text{0.05}}$ 0.28±${}_{\pm\text{0.07}}$
500 0.25±${}_{\pm\text{0.07}}$ 0.13±${}_{\pm\text{0.05}}$ 0.21±${}_{\pm\text{0.07}}$ 0.28±${}_{\pm\text{0.05}}$ 0.03±${}_{\pm\text{0.01}}$ 0.28±${}_{\pm\text{0.07}}$ 0.02±${}_{\pm\text{0.00}}$ 0.23±${}_{\pm\text{0.07}}$ 0.38±${}_{\pm\text{0.05}}$
steel 20 0.13±${}_{\pm\text{0.04}}$ 0.16±${}_{\pm\text{0.04}}$ 0.09±${}_{\pm\text{0.03}}$ 0.10±${}_{\pm\text{0.04}}$ 0.03±${}_{\pm\text{0.00}}$ 0.14±${}_{\pm\text{0.03}}$ 0.12±${}_{\pm\text{0.05}}$ 0.07±${}_{\pm\text{0.02}}$ 0.29±${}_{\pm\text{0.06}}$
50 0.26±${}_{\pm\text{0.07}}$ 0.24±${}_{\pm\text{0.07}}$ 0.17±${}_{\pm\text{0.04}}$ 0.21±${}_{\pm\text{0.06}}$ 0.03±${}_{\pm\text{0.00}}$ 0.31±${}_{\pm\text{0.04}}$ 0.05±${}_{\pm\text{0.02}}$ 0.17±${}_{\pm\text{0.05}}$ 0.43±${}_{\pm\text{0.07}}$
100 0.32±${}_{\pm\text{0.04}}$ 0.23±${}_{\pm\text{0.06}}$ 0.24±${}_{\pm\text{0.07}}$ 0.31±${}_{\pm\text{0.06}}$ 0.03±${}_{\pm\text{0.00}}$ 0.33±${}_{\pm\text{0.06}}$ 0.04±${}_{\pm\text{0.02}}$ 0.29±${}_{\pm\text{0.06}}$ 0.49±${}_{\pm\text{0.07}}$
200 0.36±${}_{\pm\text{0.04}}$ 0.17±${}_{\pm\text{0.03}}$ 0.29±${}_{\pm\text{0.06}}$ 0.34±${}_{\pm\text{0.05}}$ 0.03±${}_{\pm\text{0.00}}$ 0.36±${}_{\pm\text{0.06}}$ 0.02±${}_{\pm\text{0.00}}$ 0.35±${}_{\pm\text{0.06}}$ 0.53±${}_{\pm\text{0.07}}$
500 0.37±${}_{\pm\text{0.04}}$ 0.15±${}_{\pm\text{0.07}}$ 0.28±${}_{\pm\text{0.05}}$ 0.36±${}_{\pm\text{0.04}}$ 0.03±${}_{\pm\text{0.00}}$ 0.36±${}_{\pm\text{0.05}}$ 0.02±${}_{\pm\text{0.00}}$ 0.38±${}_{\pm\text{0.06}}$ 0.57±${}_{\pm\text{0.06}}$
stock 20 0.31±${}_{\pm\text{0.11}}$ 0.38±${}_{\pm\text{0.13}}$ 0.22±${}_{\pm\text{0.08}}$ 0.25±${}_{\pm\text{0.08}}$ 0.10±${}_{\pm\text{0.00}}$ 0.33±${}_{\pm\text{0.11}}$ 0.42±${}_{\pm\text{0.18}}$ 0.11±${}_{\pm\text{0.03}}$ 0.41±${}_{\pm\text{0.19}}$
50 0.60±${}_{\pm\text{0.09}}$ 0.56±${}_{\pm\text{0.12}}$ 0.44±${}_{\pm\text{0.15}}$ 0.54±${}_{\pm\text{0.08}}$ 0.10±${}_{\pm\text{0.00}}$ 0.73±${}_{\pm\text{0.14}}$ 0.31±${}_{\pm\text{0.14}}$ 0.41±${}_{\pm\text{0.16}}$ 0.76±${}_{\pm\text{0.10}}$
100 0.71±${}_{\pm\text{0.11}}$ 0.40±${}_{\pm\text{0.11}}$ 0.53±${}_{\pm\text{0.15}}$ 0.68±${}_{\pm\text{0.12}}$ 0.12±${}_{\pm\text{0.06}}$ 0.85±${}_{\pm\text{0.05}}$ 0.13±${}_{\pm\text{0.05}}$ 0.65±${}_{\pm\text{0.12}}$ 0.84±${}_{\pm\text{0.07}}$
200 0.86±${}_{\pm\text{0.07}}$ 0.45±${}_{\pm\text{0.12}}$ 0.60±${}_{\pm\text{0.18}}$ 0.91±${}_{\pm\text{0.11}}$ 0.35±${}_{\pm\text{0.20}}$ 0.97±${}_{\pm\text{0.05}}$ 0.10±${}_{\pm\text{0.00}}$ 0.92±${}_{\pm\text{0.08}}$ 0.97±${}_{\pm\text{0.05}}$
More than 10 classes energy 50 N/A 0.15±${}_{\pm\text{0.08}}$ 0.25±${}_{\pm\text{0.09}}$ 0.28±${}_{\pm\text{0.10}}$ 0.17±${}_{\pm\text{0.06}}$ 0.40±${}_{\pm\text{0.14}}$ 0.15±${}_{\pm\text{0.10}}$ N/A 0.78±${}_{\pm\text{0.05}}$
100 N/A 0.10±${}_{\pm\text{0.09}}$ 0.31±${}_{\pm\text{0.12}}$ 0.34±${}_{\pm\text{0.08}}$ 0.14±${}_{\pm\text{0.06}}$ 0.30±${}_{\pm\text{0.09}}$ 0.06±${}_{\pm\text{0.07}}$ N/A 0.92±${}_{\pm\text{0.02}}$
200 N/A 0.16±${}_{\pm\text{0.11}}$ 0.30±${}_{\pm\text{0.09}}$ 0.38±${}_{\pm\text{0.12}}$ 0.15±${}_{\pm\text{0.07}}$ 0.28±${}_{\pm\text{0.08}}$ 0.03±${}_{\pm\text{0.05}}$ N/A 0.96±${}_{\pm\text{0.01}}$
collins 100 N/A 0.23±${}_{\pm\text{0.08}}$ 0.21±${}_{\pm\text{0.08}}$ 0.25±${}_{\pm\text{0.09}}$ 0.05±${}_{\pm\text{0.02}}$ 0.24±${}_{\pm\text{0.06}}$ 0.07±${}_{\pm\text{0.03}}$ N/A 0.18±${}_{\pm\text{0.07}}$
200 0.32±${}_{\pm\text{0.06}}$ 0.16±${}_{\pm\text{0.06}}$ 0.24±${}_{\pm\text{0.06}}$ 0.38±${}_{\pm\text{0.05}}$ 0.06±${}_{\pm\text{0.02}}$ 0.33±${}_{\pm\text{0.07}}$ 0.02±${}_{\pm\text{0.04}}$ N/A 0.31±${}_{\pm\text{0.07}}$
texture 50 0.33±${}_{\pm\text{0.11}}$ 0.44±${}_{\pm\text{0.10}}$ 0.21±${}_{\pm\text{0.08}}$ 0.29±${}_{\pm\text{0.10}}$ 0.02±${}_{\pm\text{0.00}}$ 0.55±${}_{\pm\text{0.15}}$ 0.27±${}_{\pm\text{0.08}}$ N/A 0.41±${}_{\pm\text{0.04}}$
100 0.44±${}_{\pm\text{0.13}}$ 0.27±${}_{\pm\text{0.06}}$ 0.27±${}_{\pm\text{0.09}}$ 0.48±${}_{\pm\text{0.15}}$ 0.02±${}_{\pm\text{0.00}}$ 0.54±${}_{\pm\text{0.06}}$ 0.07±${}_{\pm\text{0.05}}$ N/A 0.49±${}_{\pm\text{0.08}}$
200 0.51±${}_{\pm\text{0.11}}$ 0.19±${}_{\pm\text{0.10}}$ 0.34±${}_{\pm\text{0.10}}$ 0.58±${}_{\pm\text{0.11}}$ 0.02±${}_{\pm\text{0.00}}$ 0.64±${}_{\pm\text{0.09}}$ ${}_{\pm\text{0.00}}$ N/A 0.59±${}_{\pm\text{0.09}}$
500 0.63±${}_{\pm\text{0.14}}$ 0.21±${}_{\pm\text{0.09}}$ 0.49±${}_{\pm\text{0.09}}$ 0.67±${}_{\pm\text{0.08}}$ 0.02±${}_{\pm\text{0.00}}$ 0.64±${}_{\pm\text{0.08}}$ ${}_{\pm\text{0.00}}$ N/A 0.66±${}_{\pm\text{0.12}}$
Average rank 3.06±${}_{\pm\text{1.32}}$ 5.45±${}_{\pm\text{2.35}}$ 6.15±${}_{\pm\text{0.87}}$ 3.64±${}_{\pm\text{1.75}}$ 8.36±${}_{\pm\text{0.90}}$ 3.18±${}_{\pm\text{1.47}}$ 7.76±${}_{\pm\text{1.77}}$ 5.33±${}_{\pm\text{1.71}}$ 2.06±${}_{\pm\text{1.43}}$
C.7 Results on Privacy Preservation
C.7.1 Downstream Accuracy in Data Sharing
Table 22: Classification accuracy (%) aggregated over six downstream predictors, comparing data sharing on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes the inapplicability of a specific generator. Different from Table 1, “--” denotes a generator cannot satisfy the requirement of generating 500 stratified samples even after generating 10,000 synthetic samples. The results of these inapplicable or failed generators are computed with the mean results of other methods. We bold the highest result for each dataset of different sample sizes. TVAE learns the joint distribution $p({\mathbf{x}},y)$ and fails to maintain the original training label distribution. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 16.38±${}_{\pm\text{3.34}}$ 13.06±${}_{\pm\text{3.57}}$ 12.99±${}_{\pm\text{4.16}}$ 12.64±${}_{\pm\text{3.06}}$ 19.42±${}_{\pm\text{3.96}}$ 12.88±${}_{\pm\text{2.32}}$ 33.11±${}_{\pm\text{3.64}}$ 33.97±${}_{\pm\text{3.48}}$
50 54.55±${}_{\pm\text{3.94}}$ 27.29±${}_{\pm\text{3.43}}$ 17.18±${}_{\pm\text{4.87}}$ 13.20±${}_{\pm\text{3.08}}$ 14.04±${}_{\pm\text{3.29}}$ 29.34±${}_{\pm\text{4.46}}$ 13.05±${}_{\pm\text{3.34}}$ 54.82±${}_{\pm\text{3.74}}$ 55.83±${}_{\pm\text{4.35}}$
100 72.46±${}_{\pm\text{3.54}}$ 40.04±${}_{\pm\text{4.54}}$ 26.65±${}_{\pm\text{3.89}}$ 12.80±${}_{\pm\text{3.48}}$ 17.65±${}_{\pm\text{3.25}}$ 35.90±${}_{\pm\text{5.29}}$ 13.55±${}_{\pm\text{3.03}}$ 71.63±${}_{\pm\text{3.95}}$ 72.99±${}_{\pm\text{3.69}}$
200 83.12±${}_{\pm\text{2.33}}$ 45.33±${}_{\pm\text{6.01}}$ 32.52±${}_{\pm\text{5.60}}$ 14.07±${}_{\pm\text{3.94}}$ 20.79±${}_{\pm\text{4.42}}$ 41.63±${}_{\pm\text{4.24}}$ 11.68±${}_{\pm\text{3.60}}$ 84.18±${}_{\pm\text{1.90}}$ 84.29±${}_{\pm\text{2.03}}$
500 89.63±${}_{\pm\text{1.57}}$ 55.24±${}_{\pm\text{5.30}}$ 44.35±${}_{\pm\text{5.67}}$ 12.76±${}_{\pm\text{2.73}}$ 21.63±${}_{\pm\text{4.90}}$ 54.26±${}_{\pm\text{3.47}}$ 11.09±${}_{\pm\text{3.19}}$ 91.19±${}_{\pm\text{1.48}}$ 90.99±${}_{\pm\text{1.53}}$
fourier 20 N/A -- 13.30±${}_{\pm\text{3.14}}$ 10.72±${}_{\pm\text{3.03}}$ 10.03±${}_{\pm\text{2.42}}$ 17.14±${}_{\pm\text{4.25}}$ 10.29±${}_{\pm\text{2.81}}$ 33.87±${}_{\pm\text{4.93}}$ 37.06±${}_{\pm\text{4.80}}$
50 55.11±${}_{\pm\text{3.56}}$- N/A 9.88±${}_{\pm\text{2.51}}$ 11.96±${}_{\pm\text{2.30}}$ 28.63±${}_{\pm\text{3.71}}$ 11.52±${}_{\pm\text{2.46}}$ 56.11±${}_{\pm\text{3.28}}$ 56.88±${}_{\pm\text{2.90}}$
100 64.06±${}_{\pm\text{3.34}}$ 34.80±${}_{\pm\text{4.60}}$ 23.35±${}_{\pm\text{4.20}}$ 10.85±${}_{\pm\text{3.02}}$ 17.68±${}_{\pm\text{3.42}}$ 32.59±${}_{\pm\text{4.02}}$ 9.82±${}_{\pm\text{2.84}}$ 64.42±${}_{\pm\text{2.96}}$ 64.93±${}_{\pm\text{3.06}}$
200 70.78±${}_{\pm\text{2.99}}$ 40.49±${}_{\pm\text{3.43}}$ 32.98±${}_{\pm\text{5.20}}$ 10.28±${}_{\pm\text{3.74}}$ 28.63±${}_{\pm\text{4.22}}$ 39.34±${}_{\pm\text{3.41}}$ 9.55±${}_{\pm\text{2.06}}$ 71.37±${}_{\pm\text{2.56}}$ 72.01±${}_{\pm\text{2.67}}$
500 74.84±${}_{\pm\text{1.47}}$ 46.83±${}_{\pm\text{3.74}}$ 46.13±${}_{\pm\text{4.33}}$ 11.13±${}_{\pm\text{2.47}}$ 27.25±${}_{\pm\text{6.04}}$ 47.19±${}_{\pm\text{2.89}}$ 10.31±${}_{\pm\text{2.19}}$ 75.86±${}_{\pm\text{1.91}}$ 76.58±${}_{\pm\text{1.54}}$
biodeg 20 68.75±${}_{\pm\text{4.96}}$ 64.07±${}_{\pm\text{7.39}}$ 53.60±${}_{\pm\text{7.49}}$ 54.58±${}_{\pm\text{8.02}}$ 48.39±${}_{\pm\text{2.73}}$ 58.69±${}_{\pm\text{6.75}}$ 48.65±${}_{\pm\text{9.50}}$ 69.14±${}_{\pm\text{5.08}}$ 69.66±${}_{\pm\text{5.18}}$
50 72.08±${}_{\pm\text{3.01}}$ 67.11±${}_{\pm\text{5.03}}$ 62.21±${}_{\pm\text{7.04}}$ 55.12±${}_{\pm\text{8.81}}$ 47.86±${}_{\pm\text{3.58}}$ 70.14±${}_{\pm\text{3.92}}$ 52.14±${}_{\pm\text{7.45}}$ 73.53±${}_{\pm\text{3.39}}$ 73.96±${}_{\pm\text{3.13}}$
100 75.61±${}_{\pm\text{2.48}}$ 70.90±${}_{\pm\text{4.96}}$ 70.33±${}_{\pm\text{3.28}}$ 58.08±${}_{\pm\text{6.71}}$ 48.78±${}_{\pm\text{2.92}}$ 70.79±${}_{\pm\text{2.88}}$ 49.15±${}_{\pm\text{7.71}}$ 76.65±${}_{\pm\text{2.03}}$ 76.56±${}_{\pm\text{2.29}}$
200 78.97±${}_{\pm\text{1.46}}$ 71.54±${}_{\pm\text{4.84}}$ 71.42±${}_{\pm\text{4.30}}$ 55.78±${}_{\pm\text{7.76}}$ 47.63±${}_{\pm\text{3.98}}$ 72.74±${}_{\pm\text{3.46}}$ 50.75±${}_{\pm\text{6.70}}$ 79.66±${}_{\pm\text{2.49}}$ 79.80±${}_{\pm\text{2.15}}$
500 81.26±${}_{\pm\text{1.42}}$ 74.57±${}_{\pm\text{3.49}}$ 76.32±${}_{\pm\text{2.93}}$ 52.78±${}_{\pm\text{4.17}}$ 47.31±${}_{\pm\text{2.99}}$ 75.67±${}_{\pm\text{2.27}}$ 48.05±${}_{\pm\text{7.50}}$ 81.36±${}_{\pm\text{1.69}}$ 81.10±${}_{\pm\text{1.52}}$
steel 20 57.55±${}_{\pm\text{4.83}}$ 54.39±${}_{\pm\text{7.65}}$ 52.13±${}_{\pm\text{5.78}}$ 51.21±${}_{\pm\text{6.68}}$ 51.28±${}_{\pm\text{4.58}}$ 52.03±${}_{\pm\text{4.88}}$ 49.25±${}_{\pm\text{4.36}}$ 63.27±${}_{\pm\text{5.66}}$ 63.30±${}_{\pm\text{5.47}}$
50 64.84±${}_{\pm\text{4.07}}$ 59.06±${}_{\pm\text{3.85}}$ 56.70±${}_{\pm\text{4.96}}$ 53.24±${}_{\pm\text{4.77}}$ 49.18±${}_{\pm\text{4.67}}$ 57.16±${}_{\pm\text{4.35}}$ 49.90±${}_{\pm\text{3.96}}$ 78.55±${}_{\pm\text{5.17}}$ 79.99±${}_{\pm\text{6.37}}$
100 72.48±${}_{\pm\text{4.51}}$ 61.74±${}_{\pm\text{4.16}}$ 59.37±${}_{\pm\text{5.08}}$ 52.79±${}_{\pm\text{4.06}}$ 45.01±${}_{\pm\text{5.84}}$ 57.26±${}_{\pm\text{4.46}}$ 49.66±${}_{\pm\text{3.56}}$ 90.12±${}_{\pm\text{4.20}}$ 92.33±${}_{\pm\text{2.57}}$
200 77.85±${}_{\pm\text{3.50}}$ 65.45±${}_{\pm\text{4.07}}$ 63.31±${}_{\pm\text{4.63}}$ 51.02±${}_{\pm\text{2.94}}$ 43.71±${}_{\pm\text{7.51}}$ 61.41±${}_{\pm\text{4.21}}$ 50.20±${}_{\pm\text{3.70}}$ 94.61±${}_{\pm\text{1.75}}$ 95.54±${}_{\pm\text{1.45}}$
500 84.21±${}_{\pm\text{3.52}}$ 70.26±${}_{\pm\text{4.93}}$ 70.62±${}_{\pm\text{5.09}}$ 49.97±${}_{\pm\text{4.05}}$ 46.98±${}_{\pm\text{4.30}}$ 66.33±${}_{\pm\text{5.72}}$ 51.60±${}_{\pm\text{3.43}}$ 96.31±${}_{\pm\text{1.25}}$ 97.04±${}_{\pm\text{1.17}}$
stock 20 81.62±${}_{\pm\text{4.62}}$ 69.35±${}_{\pm\text{8.30}}$ 51.71±${}_{\pm\text{12.09}}$ 67.05±${}_{\pm\text{11.69}}$ 75.94±${}_{\pm\text{14.43}}$ 64.16±${}_{\pm\text{9.18}}$ 48.55±${}_{\pm\text{12.54}}$ 82.55±${}_{\pm\text{4.42}}$ 83.60±${}_{\pm\text{4.09}}$
50 87.43±${}_{\pm\text{2.47}}$ 76.07±${}_{\pm\text{5.02}}$ 69.12±${}_{\pm\text{5.00}}$ 70.61±${}_{\pm\text{8.50}}$ 85.99±${}_{\pm\text{2.82}}$ 78.38±${}_{\pm\text{3.56}}$ ${}_{\pm\text{9.74}}$ 88.09±${}_{\pm\text{2.37}}$ 88.49±${}_{\pm\text{2.34}}$
100 89.63±${}_{\pm\text{1.30}}$ 81.18±${}_{\pm\text{4.24}}$ 78.44±${}_{\pm\text{3.91}}$ 72.05±${}_{\pm\text{5.43}}$ 88.27±${}_{\pm\text{2.34}}$ 84.02±${}_{\pm\text{2.86}}$ 51.07±${}_{\pm\text{10.98}}$ 90.13±${}_{\pm\text{1.57}}$ 90.58±${}_{\pm\text{1.34}}$
200 91.11±${}_{\pm\text{1.29}}$ 84.05±${}_{\pm\text{2.63}}$ 82.12±${}_{\pm\text{2.93}}$ 75.44±${}_{\pm\text{3.54}}$ 89.92±${}_{\pm\text{1.56}}$ 85.66±${}_{\pm\text{2.26}}$ 49.74±${}_{\pm\text{12.06}}$ 91.09±${}_{\pm\text{1.52}}$ 91.07±${}_{\pm\text{1.07}}$
More than 10 classes energy 50 N/A 7.17±${}_{\pm\text{1.81}}$ 5.20±${}_{\pm\text{2.04}}$ 5.33±${}_{\pm\text{1.61}}$ 4.21±${}_{\pm\text{1.69}}$ 7.26±${}_{\pm\text{1.89}}$ 4.52±${}_{\pm\text{0.57}}$ N/A 23.80±${}_{\pm\text{2.60}}$
100 N/A -- 7.66±${}_{\pm\text{1.96}}$ 5.52±${}_{\pm\text{1.49}}$ 4.08±${}_{\pm\text{1.40}}$ 9.87±${}_{\pm\text{2.01}}$ 4.01±${}_{\pm\text{1.12}}$ N/A 30.15±${}_{\pm\text{3.21}}$
200 N/A -- 7.57±${}_{\pm\text{2.01}}$ 6.92±${}_{\pm\text{1.96}}$ 3.42±${}_{\pm\text{1.25}}$ 11.85±${}_{\pm\text{2.39}}$ 4.18±${}_{\pm\text{0.91}}$ N/A 35.74±${}_{\pm\text{3.59}}$
collins 100 N/A -- 5.51±${}_{\pm\text{0.87}}$ 4.58±${}_{\pm\text{0.95}}$ 11.74±${}_{\pm\text{1.73}}$ 5.34±${}_{\pm\text{1.48}}$ 3.77±${}_{\pm\text{0.63}}$ N/A 13.12±${}_{\pm\text{1.75}}$
200 17.60±${}_{\pm\text{1.83}}$- 5.55±${}_{\pm\text{1.34}}$ 4.70±${}_{\pm\text{0.83}}$ 13.64±${}_{\pm\text{1.54}}$ 5.46±${}_{\pm\text{0.90}}$ 3.92±${}_{\pm\text{0.92}}$ N/A 16.80±${}_{\pm\text{1.55}}$
texture 50 75.50±${}_{\pm\text{2.93}}$ 23.86±${}_{\pm\text{8.05}}$ 12.00±${}_{\pm\text{4.19}}$ 12.16±${}_{\pm\text{3.20}}$ 17.18±${}_{\pm\text{5.34}}$ 16.82±${}_{\pm\text{6.16}}$ 9.63±${}_{\pm\text{2.75}}$ N/A 78.84±${}_{\pm\text{3.35}}$
100 83.77±${}_{\pm\text{2.52}}$ 26.87±${}_{\pm\text{4.19}}$ 14.05±${}_{\pm\text{6.20}}$ 13.95±${}_{\pm\text{5.32}}$ 20.08±${}_{\pm\text{5.52}}$ 23.82±${}_{\pm\text{5.96}}$ 10.21±${}_{\pm\text{2.69}}$ N/A 85.88±${}_{\pm\text{1.87}}$
200 87.96±${}_{\pm\text{1.79}}$ 21.59±${}_{\pm\text{8.99}}$ 15.36±${}_{\pm\text{4.43}}$ 12.17±${}_{\pm\text{4.57}}$ 20.59±${}_{\pm\text{5.96}}$ 42.28±${}_{\pm\text{5.81}}$ 9.52±${}_{\pm\text{2.91}}$ N/A 88.84±${}_{\pm\text{1.68}}$
500 91.65±${}_{\pm\text{1.14}}$ 34.48±${}_{\pm\text{9.28}}$ 26.45±${}_{\pm\text{8.77}}$ 12.37±${}_{\pm\text{4.63}}$ 20.07±${}_{\pm\text{5.56}}$ 57.94±${}_{\pm\text{4.77}}$ 10.22±${}_{\pm\text{3.69}}$ N/A 91.36±${}_{\pm\text{1.12}}$
Average rank 2.76±${}_{\pm\text{0.76}}$ 4.52±${}_{\pm\text{0.77}}$ 6.18±${}_{\pm\text{0.95}}$ 7.48±${}_{\pm\text{0.76}}$ 7.03±${}_{\pm\text{2.07}}$ 4.91±${}_{\pm\text{1.16}}$ 8.52±${}_{\pm\text{0.57}}$ 2.36±${}_{\pm\text{0.89}}$ 1.24±${}_{\pm\text{0.56}}$
Table 23: Classification accuracy (%) of LR, comparing data sharing on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes the inapplicability of a specific generator. Different from Table 1, “--” denotes a generator cannot satisfy the requirement of generating 500 stratified samples even after generating 10,000 synthetic samples. The results of these inapplicable or failed generators are computed with the mean results of other methods. We bold the highest result for each dataset of different sample sizes. TVAE learns the joint distribution $p({\mathbf{x}},y)$ and fails to maintain the original training label distribution. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 14.51±${}_{\pm\text{3.57}}$ 13.00±${}_{\pm\text{5.74}}$ 12.08±${}_{\pm\text{4.67}}$ 13.83±${}_{\pm\text{3.50}}$ 22.46±${}_{\pm\text{5.01}}$ 12.52±${}_{\pm\text{2.66}}$ 38.00±${}_{\pm\text{2.16}}$ 38.04±${}_{\pm\text{2.34}}$
50 61.15±${}_{\pm\text{4.22}}$ 24.43±${}_{\pm\text{3.31}}$ 16.81±${}_{\pm\text{5.60}}$ 12.55±${}_{\pm\text{2.78}}$ 14.74±${}_{\pm\text{4.36}}$ 33.77±${}_{\pm\text{5.53}}$ 12.26±${}_{\pm\text{4.83}}$ 63.01±${}_{\pm\text{3.66}}$ 62.94±${}_{\pm\text{4.10}}$
100 78.16±${}_{\pm\text{3.53}}$ 40.22±${}_{\pm\text{4.23}}$ 27.14±${}_{\pm\text{5.20}}$ 12.40±${}_{\pm\text{5.31}}$ 19.34±${}_{\pm\text{3.12}}$ 40.99±${}_{\pm\text{5.14}}$ 12.33±${}_{\pm\text{6.06}}$ 80.54±${}_{\pm\text{3.22}}$ 80.18±${}_{\pm\text{2.97}}$
200 88.75±${}_{\pm\text{1.37}}$ 45.70±${}_{\pm\text{7.29}}$ 34.17±${}_{\pm\text{6.11}}$ 13.87±${}_{\pm\text{4.39}}$ 22.01±${}_{\pm\text{5.64}}$ 44.76±${}_{\pm\text{4.63}}$ 12.78±${}_{\pm\text{5.20}}$ 90.89±${}_{\pm\text{1.53}}$ 90.09±${}_{\pm\text{1.86}}$
500 94.29±${}_{\pm\text{1.17}}$ 60.39±${}_{\pm\text{2.81}}$ 47.87±${}_{\pm\text{7.19}}$ 12.58±${}_{\pm\text{3.76}}$ 24.69±${}_{\pm\text{5.87}}$ 58.91±${}_{\pm\text{2.80}}$ 9.79±${}_{\pm\text{3.65}}$ 95.86±${}_{\pm\text{1.59}}$ 95.38±${}_{\pm\text{1.40}}$
fourier 20 N/A -- 11.64±${}_{\pm\text{4.35}}$ 10.16±${}_{\pm\text{2.99}}$ 8.48±${}_{\pm\text{2.86}}$ 18.75±${}_{\pm\text{4.29}}$ 11.38±${}_{\pm\text{4.85}}$ 42.98±${}_{\pm\text{5.14}}$ 43.04±${}_{\pm\text{5.10}}$
50 58.24±${}_{\pm\text{2.05}}$- N/A 9.82±${}_{\pm\text{4.25}}$ 13.48±${}_{\pm\text{2.85}}$ 33.97±${}_{\pm\text{3.71}}$ 12.00±${}_{\pm\text{2.92}}$ 60.30±${}_{\pm\text{1.58}}$ 60.28±${}_{\pm\text{1.63}}$
100 65.32±${}_{\pm\text{2.38}}$ 28.60±${}_{\pm\text{2.38}}$ 21.26±${}_{\pm\text{2.88}}$ 11.24±${}_{\pm\text{3.65}}$ 19.72±${}_{\pm\text{3.78}}$ 37.64±${}_{\pm\text{3.66}}$ 11.34±${}_{\pm\text{2.54}}$ 67.36±${}_{\pm\text{2.21}}$ 67.38±${}_{\pm\text{2.34}}$
200 70.74±${}_{\pm\text{2.13}}$ 37.50±${}_{\pm\text{4.11}}$ 28.86±${}_{\pm\text{4.32}}$ 10.12±${}_{\pm\text{4.56}}$ 33.58±${}_{\pm\text{3.65}}$ 42.75±${}_{\pm\text{2.57}}$ 7.52±${}_{\pm\text{2.05}}$ 72.20±${}_{\pm\text{3.02}}$ 72.18±${}_{\pm\text{2.94}}$
500 73.96±${}_{\pm\text{0.94}}$ 46.84±${}_{\pm\text{2.99}}$ 41.18±${}_{\pm\text{4.18}}$ 10.60±${}_{\pm\text{3.33}}$ 31.88±${}_{\pm\text{9.69}}$ 48.67±${}_{\pm\text{1.58}}$ 10.56±${}_{\pm\text{3.12}}$ 75.58±${}_{\pm\text{1.99}}$ 76.00±${}_{\pm\text{2.93}}$
biodeg 20 69.93±${}_{\pm\text{5.59}}$ 66.92±${}_{\pm\text{7.16}}$ 50.10±${}_{\pm\text{10.26}}$ 56.08±${}_{\pm\text{10.54}}$ 46.15±${}_{\pm\text{4.22}}$ 61.15±${}_{\pm\text{7.27}}$ 45.12±${}_{\pm\text{9.23}}$ 70.78±${}_{\pm\text{3.95}}$ 71.24±${}_{\pm\text{4.85}}$
50 74.71±${}_{\pm\text{3.39}}$ 70.19±${}_{\pm\text{4.89}}$ 62.90±${}_{\pm\text{6.28}}$ 56.85±${}_{\pm\text{10.24}}$ 42.74±${}_{\pm\text{2.82}}$ 73.38±${}_{\pm\text{2.61}}$ 54.52±${}_{\pm\text{12.44}}$ 75.67±${}_{\pm\text{2.36}}$ 76.47±${}_{\pm\text{2.99}}$
100 77.48±${}_{\pm\text{1.83}}$ 71.77±${}_{\pm\text{5.56}}$ 71.42±${}_{\pm\text{3.27}}$ 59.17±${}_{\pm\text{7.35}}$ 46.24±${}_{\pm\text{3.89}}$ 74.48±${}_{\pm\text{2.84}}$ 48.03±${}_{\pm\text{12.56}}$ 77.94±${}_{\pm\text{2.45}}$ 78.34±${}_{\pm\text{2.11}}$
200 80.56±${}_{\pm\text{1.77}}$ 75.37±${}_{\pm\text{4.47}}$ 72.61±${}_{\pm\text{5.99}}$ 55.71±${}_{\pm\text{7.21}}$ 44.16±${}_{\pm\text{4.71}}$ 75.09±${}_{\pm\text{2.93}}$ 52.94±${}_{\pm\text{11.50}}$ 81.21±${}_{\pm\text{2.06}}$ 81.50±${}_{\pm\text{1.84}}$
500 82.68±${}_{\pm\text{1.01}}$ 77.92±${}_{\pm\text{2.73}}$ 77.55±${}_{\pm\text{2.55}}$ 54.90±${}_{\pm\text{5.52}}$ 46.53±${}_{\pm\text{2.91}}$ 77.87±${}_{\pm\text{2.10}}$ 46.40±${}_{\pm\text{12.51}}$ 82.38±${}_{\pm\text{1.61}}$ 82.14±${}_{\pm\text{1.30}}$
steel 20 56.80±${}_{\pm\text{5.46}}$ 54.06±${}_{\pm\text{11.24}}$ 54.41±${}_{\pm\text{5.57}}$ 52.14±${}_{\pm\text{7.31}}$ 51.40±${}_{\pm\text{6.83}}$ 54.38±${}_{\pm\text{4.84}}$ 48.32±${}_{\pm\text{6.98}}$ 66.75±${}_{\pm\text{9.74}}$ 67.05±${}_{\pm\text{9.39}}$
50 66.80±${}_{\pm\text{6.52}}$ 61.11±${}_{\pm\text{3.73}}$ 58.25±${}_{\pm\text{6.67}}$ 52.95±${}_{\pm\text{6.68}}$ 49.07±${}_{\pm\text{7.58}}$ 61.31±${}_{\pm\text{4.45}}$ 48.86±${}_{\pm\text{7.10}}$ 93.51±${}_{\pm\text{4.94}}$ 92.13±${}_{\pm\text{4.90}}$
100 79.38±${}_{\pm\text{4.02}}$ 64.64±${}_{\pm\text{3.99}}$ 60.57±${}_{\pm\text{5.90}}$ 54.40±${}_{\pm\text{5.67}}$ 42.51±${}_{\pm\text{5.11}}$ 58.86±${}_{\pm\text{3.66}}$ 50.95±${}_{\pm\text{3.77}}$ 99.21±${}_{\pm\text{0.86}}$ 99.17±${}_{\pm\text{0.93}}$
200 82.66±${}_{\pm\text{3.91}}$ 71.78±${}_{\pm\text{3.68}}$ 65.44±${}_{\pm\text{5.02}}$ 51.72±${}_{\pm\text{4.16}}$ 38.19±${}_{\pm\text{8.24}}$ 63.31±${}_{\pm\text{5.98}}$ 53.46±${}_{\pm\text{2.96}}$ 99.45±${}_{\pm\text{0.69}}$ 99.51±${}_{\pm\text{0.69}}$
500 90.21±${}_{\pm\text{3.80}}$ 76.49±${}_{\pm\text{5.04}}$ 75.41±${}_{\pm\text{6.20}}$ 48.71±${}_{\pm\text{6.15}}$ 46.00±${}_{\pm\text{4.77}}$ 72.11±${}_{\pm\text{7.56}}$ 52.55±${}_{\pm\text{3.50}}$ 99.70±${}_{\pm\text{0.22}}$ 99.72±${}_{\pm\text{0.20}}$
stock 20 80.31±${}_{\pm\text{4.03}}$ 71.68±${}_{\pm\text{8.45}}$ 54.58±${}_{\pm\text{13.04}}$ 70.64±${}_{\pm\text{9.77}}$ 71.05±${}_{\pm\text{11.98}}$ 68.03±${}_{\pm\text{7.38}}$ 49.84±${}_{\pm\text{19.78}}$ 79.58±${}_{\pm\text{4.43}}$ 80.33±${}_{\pm\text{3.52}}$
50 81.28±${}_{\pm\text{2.87}}$ 75.75±${}_{\pm\text{3.84}}$ 69.67±${}_{\pm\text{3.14}}$ 72.99±${}_{\pm\text{6.89}}$ 77.72±${}_{\pm\text{4.99}}$ 76.64±${}_{\pm\text{2.54}}$ 49.01±${}_{\pm\text{17.32}}$ 82.37±${}_{\pm\text{3.13}}$ 82.06±${}_{\pm\text{2.58}}$
100 83.90±${}_{\pm\text{1.88}}$ 78.79±${}_{\pm\text{3.25}}$ 77.37±${}_{\pm\text{3.85}}$ 75.64±${}_{\pm\text{3.63}}$ 80.43±${}_{\pm\text{3.70}}$ 78.22±${}_{\pm\text{2.48}}$ 49.60±${}_{\pm\text{17.93}}$ 83.65±${}_{\pm\text{1.65}}$ 83.56±${}_{\pm\text{1.81}}$
200 83.61±${}_{\pm\text{1.32}}$ 79.31±${}_{\pm\text{2.69}}$ 80.20±${}_{\pm\text{2.16}}$ 76.37±${}_{\pm\text{1.92}}$ 79.32±${}_{\pm\text{2.34}}$ 78.89±${}_{\pm\text{2.42}}$ 48.18±${}_{\pm\text{17.04}}$ 83.69±${}_{\pm\text{1.41}}$ 83.82±${}_{\pm\text{1.34}}$
More than 10 classes energy 50 N/A 7.51±${}_{\pm\text{1.98}}$ 5.79±${}_{\pm\text{2.56}}$ 5.89±${}_{\pm\text{1.99}}$ 3.42±${}_{\pm\text{2.02}}$ 8.72±${}_{\pm\text{1.70}}$ 4.52±${}_{\pm\text{1.16}}$ N/A 21.56±${}_{\pm\text{1.66}}$
100 N/A -- 7.16±${}_{\pm\text{1.69}}$ 6.15±${}_{\pm\text{1.09}}$ 4.88±${}_{\pm\text{1.99}}$ 10.67±${}_{\pm\text{2.27}}$ 4.03±${}_{\pm\text{1.33}}$ N/A 27.59±${}_{\pm\text{2.14}}$
200 N/A -- 7.82±${}_{\pm\text{2.19}}$ 6.63±${}_{\pm\text{2.11}}$ 3.48±${}_{\pm\text{1.64}}$ 12.96±${}_{\pm\text{2.22}}$ 4.34±${}_{\pm\text{1.04}}$ N/A 33.01±${}_{\pm\text{2.44}}$
collins 100 N/A -- 5.33±${}_{\pm\text{0.98}}$ 5.49±${}_{\pm\text{1.10}}$ 12.30±${}_{\pm\text{2.14}}$ 5.98±${}_{\pm\text{1.61}}$ 3.61±${}_{\pm\text{1.35}}$ N/A 14.02±${}_{\pm\text{2.45}}$
200 18.91±${}_{\pm\text{1.64}}$- 5.36±${}_{\pm\text{1.46}}$ 5.13±${}_{\pm\text{0.92}}$ 14.70±${}_{\pm\text{1.70}}$ 6.04±${}_{\pm\text{0.86}}$ 3.28±${}_{\pm\text{1.59}}$ N/A 19.11±${}_{\pm\text{1.47}}$
texture 50 86.20±${}_{\pm\text{2.75}}$ 25.72±${}_{\pm\text{6.89}}$ 13.32±${}_{\pm\text{4.63}}$ 13.86±${}_{\pm\text{4.99}}$ 23.90±${}_{\pm\text{6.56}}$ 18.17±${}_{\pm\text{8.82}}$ 10.11±${}_{\pm\text{3.94}}$ N/A 88.46±${}_{\pm\text{2.81}}$
100 93.17±${}_{\pm\text{2.01}}$ 26.04±${}_{\pm\text{2.24}}$ 12.34±${}_{\pm\text{5.07}}$ 12.44±${}_{\pm\text{5.67}}$ 27.67±${}_{\pm\text{6.90}}$ 22.55±${}_{\pm\text{5.85}}$ 12.00±${}_{\pm\text{5.29}}$ N/A 94.23±${}_{\pm\text{1.30}}$
200 95.70±${}_{\pm\text{1.24}}$ 16.65±${}_{\pm\text{9.20}}$ 17.15±${}_{\pm\text{5.04}}$ 13.14±${}_{\pm\text{6.40}}$ 29.67±${}_{\pm\text{6.59}}$ 43.54±${}_{\pm\text{7.07}}$ 11.12±${}_{\pm\text{5.24}}$ N/A 95.99±${}_{\pm\text{1.14}}$
500 97.17±${}_{\pm\text{0.38}}$ 39.69±${}_{\pm\text{10.43}}$ 27.24±${}_{\pm\text{8.08}}$ 11.68±${}_{\pm\text{4.59}}$ 27.75±${}_{\pm\text{7.52}}$ 60.15±${}_{\pm\text{5.45}}$ 10.95±${}_{\pm\text{5.25}}$ N/A 96.53±${}_{\pm\text{0.54}}$
Average rank 2.82±${}_{\pm\text{0.80}}$ 4.70±${}_{\pm\text{0.78}}$ 6.39±${}_{\pm\text{1.14}}$ 7.48±${}_{\pm\text{0.76}}$ 6.79±${}_{\pm\text{2.00}}$ 4.64±${}_{\pm\text{1.37}}$ 8.52±${}_{\pm\text{0.71}}$ 2.24±${}_{\pm\text{1.10}}$ 1.42±${}_{\pm\text{0.61}}$
Table 24: Classification accuracy (%) of KNN, comparing data sharing on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes the inapplicability of a specific generator. Different from Table 1, “--” denotes a generator cannot satisfy the requirement of generating 500 stratified samples even after generating 10,000 synthetic samples. The results of these inapplicable or failed generators are computed with the mean results of other methods. We bold the highest result for each dataset of different sample sizes. TVAE learns the joint distribution $p({\mathbf{x}},y)$ and fails to maintain the original training label distribution. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 16.84±${}_{\pm\text{3.19}}$ 12.84±${}_{\pm\text{3.39}}$ 11.26±${}_{\pm\text{3.06}}$ 11.63±${}_{\pm\text{1.92}}$ 16.53±${}_{\pm\text{3.52}}$ 11.76±${}_{\pm\text{2.51}}$ 35.75±${}_{\pm\text{4.48}}$ 35.76±${}_{\pm\text{4.39}}$
50 55.10±${}_{\pm\text{3.65}}$ 27.34±${}_{\pm\text{2.43}}$ 15.55±${}_{\pm\text{4.09}}$ 13.25±${}_{\pm\text{4.00}}$ 12.03±${}_{\pm\text{3.72}}$ 24.09±${}_{\pm\text{4.43}}$ 12.53±${}_{\pm\text{2.50}}$ 53.42±${}_{\pm\text{3.59}}$ 53.44±${}_{\pm\text{3.32}}$
100 69.44±${}_{\pm\text{2.83}}$ 39.23±${}_{\pm\text{3.86}}$ 22.23±${}_{\pm\text{2.89}}$ 12.64±${}_{\pm\text{3.23}}$ 16.71±${}_{\pm\text{3.38}}$ 30.64±${}_{\pm\text{4.80}}$ 13.66±${}_{\pm\text{2.29}}$ 67.23±${}_{\pm\text{2.65}}$ 67.35±${}_{\pm\text{2.69}}$
200 77.12±${}_{\pm\text{2.73}}$ 41.50±${}_{\pm\text{4.16}}$ 27.68±${}_{\pm\text{5.17}}$ 14.16±${}_{\pm\text{4.10}}$ 20.29±${}_{\pm\text{3.05}}$ 36.01±${}_{\pm\text{2.74}}$ 11.91±${}_{\pm\text{4.20}}$ 75.56±${}_{\pm\text{2.00}}$ 76.17±${}_{\pm\text{2.04}}$
500 83.82±${}_{\pm\text{1.86}}$ 50.56±${}_{\pm\text{5.32}}$ 35.07±${}_{\pm\text{4.52}}$ 13.44±${}_{\pm\text{2.78}}$ 19.90±${}_{\pm\text{3.06}}$ 46.40±${}_{\pm\text{3.68}}$ 11.99±${}_{\pm\text{2.62}}$ 83.63±${}_{\pm\text{1.63}}$ 84.35±${}_{\pm\text{1.33}}$
fourier 20 N/A -- 19.60±${}_{\pm\text{4.25}}$ 11.06±${}_{\pm\text{3.51}}$ 9.16±${}_{\pm\text{1.73}}$ 15.58±${}_{\pm\text{3.19}}$ 9.78±${}_{\pm\text{2.29}}$ 42.66±${}_{\pm\text{5.78}}$ 42.78±${}_{\pm\text{5.83}}$
50 60.16±${}_{\pm\text{1.62}}$- N/A 9.74±${}_{\pm\text{2.10}}$ 11.66±${}_{\pm\text{2.39}}$ 24.32±${}_{\pm\text{4.59}}$ 12.86±${}_{\pm\text{2.44}}$ 58.52±${}_{\pm\text{1.71}}$ 58.56±${}_{\pm\text{1.99}}$
100 66.54±${}_{\pm\text{2.75}}$ 38.18±${}_{\pm\text{5.70}}$ 19.86±${}_{\pm\text{3.82}}$ 10.68±${}_{\pm\text{3.49}}$ 18.42±${}_{\pm\text{3.40}}$ 27.64±${}_{\pm\text{4.26}}$ 9.48±${}_{\pm\text{3.68}}$ 64.64±${}_{\pm\text{2.26}}$ 64.98±${}_{\pm\text{2.58}}$
200 71.08±${}_{\pm\text{2.07}}$ 42.72±${}_{\pm\text{2.73}}$ 31.68±${}_{\pm\text{4.44}}$ 9.70±${}_{\pm\text{2.10}}$ 26.96±${}_{\pm\text{2.98}}$ 34.32±${}_{\pm\text{4.26}}$ 9.76±${}_{\pm\text{2.16}}$ 70.00±${}_{\pm\text{1.77}}$ 70.50±${}_{\pm\text{1.86}}$
500 75.06±${}_{\pm\text{1.95}}$ 47.58±${}_{\pm\text{3.46}}$ 42.68±${}_{\pm\text{3.18}}$ 10.00±${}_{\pm\text{1.07}}$ 26.18±${}_{\pm\text{3.18}}$ 41.87±${}_{\pm\text{3.25}}$ 10.88±${}_{\pm\text{1.70}}$ 73.46±${}_{\pm\text{1.80}}$ 73.95±${}_{\pm\text{1.45}}$
biodeg 20 69.11±${}_{\pm\text{3.28}}$ 65.06±${}_{\pm\text{6.01}}$ 54.27±${}_{\pm\text{6.17}}$ 52.02±${}_{\pm\text{5.98}}$ 47.87±${}_{\pm\text{4.28}}$ 55.35±${}_{\pm\text{6.64}}$ 47.41±${}_{\pm\text{12.04}}$ 67.89±${}_{\pm\text{4.68}}$ 69.62±${}_{\pm\text{4.52}}$
50 72.84±${}_{\pm\text{2.33}}$ 66.10±${}_{\pm\text{4.38}}$ 61.89±${}_{\pm\text{7.87}}$ 52.26±${}_{\pm\text{8.81}}$ 50.78±${}_{\pm\text{5.21}}$ 67.72±${}_{\pm\text{3.53}}$ 53.48±${}_{\pm\text{10.48}}$ 72.01±${}_{\pm\text{4.14}}$ 73.77±${}_{\pm\text{3.46}}$
100 75.82±${}_{\pm\text{2.03}}$ 69.86±${}_{\pm\text{3.93}}$ 70.02±${}_{\pm\text{4.13}}$ 58.95±${}_{\pm\text{5.48}}$ 48.67±${}_{\pm\text{4.05}}$ 69.39±${}_{\pm\text{2.95}}$ 51.11±${}_{\pm\text{13.28}}$ 74.77±${}_{\pm\text{2.03}}$ 75.15±${}_{\pm\text{1.93}}$
200 79.67±${}_{\pm\text{0.95}}$ 70.25±${}_{\pm\text{5.58}}$ 71.89±${}_{\pm\text{3.71}}$ 53.54±${}_{\pm\text{7.72}}$ 51.40±${}_{\pm\text{7.02}}$ 70.21±${}_{\pm\text{3.87}}$ 49.42±${}_{\pm\text{6.95}}$ 77.78±${}_{\pm\text{2.42}}$ 78.23±${}_{\pm\text{2.08}}$
500 82.51±${}_{\pm\text{1.48}}$ 72.72±${}_{\pm\text{3.89}}$ 75.53±${}_{\pm\text{3.97}}$ 52.92±${}_{\pm\text{5.87}}$ 46.14±${}_{\pm\text{4.19}}$ 73.57±${}_{\pm\text{3.14}}$ 50.03±${}_{\pm\text{3.77}}$ 80.55±${}_{\pm\text{1.56}}$ 81.15±${}_{\pm\text{1.54}}$
steel 20 62.72±${}_{\pm\text{5.46}}$ 56.75±${}_{\pm\text{7.62}}$ 49.84±${}_{\pm\text{5.79}}$ 49.97±${}_{\pm\text{6.59}}$ 52.29±${}_{\pm\text{4.71}}$ 52.67±${}_{\pm\text{5.56}}$ 49.23±${}_{\pm\text{4.34}}$ 70.71±${}_{\pm\text{3.87}}$ 69.36±${}_{\pm\text{4.02}}$
50 69.72±${}_{\pm\text{3.48}}$ 60.43±${}_{\pm\text{5.12}}$ 57.65±${}_{\pm\text{4.62}}$ 54.41±${}_{\pm\text{4.04}}$ 48.61±${}_{\pm\text{6.80}}$ 57.24±${}_{\pm\text{5.82}}$ 49.66±${}_{\pm\text{1.98}}$ 82.03±${}_{\pm\text{3.08}}$ 80.57±${}_{\pm\text{3.35}}$
100 76.33±${}_{\pm\text{4.19}}$ 63.34±${}_{\pm\text{2.92}}$ 60.89±${}_{\pm\text{5.83}}$ 52.24±${}_{\pm\text{4.96}}$ 42.38±${}_{\pm\text{6.73}}$ 58.01±${}_{\pm\text{6.58}}$ 51.17±${}_{\pm\text{1.99}}$ 87.79±${}_{\pm\text{3.06}}$ 87.76±${}_{\pm\text{3.25}}$
200 79.90±${}_{\pm\text{1.97}}$ 68.00±${}_{\pm\text{2.85}}$ 66.31±${}_{\pm\text{4.27}}$ 50.19±${}_{\pm\text{4.02}}$ 43.82±${}_{\pm\text{7.44}}$ 63.29±${}_{\pm\text{4.68}}$ 50.99±${}_{\pm\text{2.53}}$ 90.90±${}_{\pm\text{1.80}}$ 91.07±${}_{\pm\text{1.75}}$
500 85.06±${}_{\pm\text{1.90}}$ 73.24±${}_{\pm\text{3.16}}$ 72.49±${}_{\pm\text{2.82}}$ 49.19±${}_{\pm\text{4.79}}$ 45.77±${}_{\pm\text{8.19}}$ 68.81±${}_{\pm\text{4.42}}$ 52.45±${}_{\pm\text{3.59}}$ 92.81±${}_{\pm\text{1.15}}$ 92.83±${}_{\pm\text{1.35}}$
stock 20 84.27±${}_{\pm\text{5.28}}$ 66.05±${}_{\pm\text{9.50}}$ 52.15±${}_{\pm\text{10.98}}$ 58.69±${}_{\pm\text{15.22}}$ 77.28±${}_{\pm\text{18.40}}$ 58.03±${}_{\pm\text{12.41}}$ 45.13±${}_{\pm\text{15.32}}$ 84.45±${}_{\pm\text{4.02}}$ 84.69±${}_{\pm\text{4.16}}$
50 89.64±${}_{\pm\text{2.35}}$ 75.40±${}_{\pm\text{3.27}}$ 67.64±${}_{\pm\text{6.93}}$ 64.81±${}_{\pm\text{9.27}}$ 88.94±${}_{\pm\text{1.62}}$ 77.67±${}_{\pm\text{4.49}}$ 48.55±${}_{\pm\text{11.95}}$ 89.69±${}_{\pm\text{1.89}}$ 89.70±${}_{\pm\text{1.90}}$
100 91.93±${}_{\pm\text{0.77}}$ 80.87±${}_{\pm\text{2.69}}$ 78.37±${}_{\pm\text{3.93}}$ 67.33±${}_{\pm\text{6.34}}$ 91.03±${}_{\pm\text{1.15}}$ 84.53±${}_{\pm\text{3.34}}$ 53.19±${}_{\pm\text{11.60}}$ 91.98±${}_{\pm\text{0.63}}$ 92.35±${}_{\pm\text{0.75}}$
200 93.46±${}_{\pm\text{0.93}}$ 84.65±${}_{\pm\text{1.61}}$ 81.64±${}_{\pm\text{3.69}}$ 70.56±${}_{\pm\text{3.41}}$ 91.81±${}_{\pm\text{0.69}}$ 85.58±${}_{\pm\text{2.64}}$ 49.98±${}_{\pm\text{11.39}}$ 92.53±${}_{\pm\text{1.09}}$ 92.94±${}_{\pm\text{0.99}}$
More than 10 classes energy 50 N/A 6.75±${}_{\pm\text{1.25}}$ 4.88±${}_{\pm\text{1.00}}$ 5.21±${}_{\pm\text{1.39}}$ 4.71±${}_{\pm\text{1.44}}$ 6.06±${}_{\pm\text{1.55}}$ 4.63±${}_{\pm\text{0.47}}$ N/A 25.36±${}_{\pm\text{2.29}}$
100 N/A -- 7.53±${}_{\pm\text{1.81}}$ 5.33±${}_{\pm\text{1.43}}$ 3.68±${}_{\pm\text{1.03}}$ 9.05±${}_{\pm\text{2.14}}$ 4.37±${}_{\pm\text{1.22}}$ N/A 29.63±${}_{\pm\text{2.43}}$
200 N/A -- 7.22±${}_{\pm\text{0.99}}$ 6.01±${}_{\pm\text{1.31}}$ 3.30±${}_{\pm\text{0.91}}$ 10.56±${}_{\pm\text{1.57}}$ 3.99±${}_{\pm\text{0.73}}$ N/A 33.85±${}_{\pm\text{3.11}}$
collins 100 N/A -- 5.61±${}_{\pm\text{0.95}}$ 4.32±${}_{\pm\text{0.72}}$ 13.46±${}_{\pm\text{1.68}}$ 4.71±${}_{\pm\text{1.10}}$ 3.80±${}_{\pm\text{0.09}}$ N/A 15.20±${}_{\pm\text{2.00}}$
200 19.94±${}_{\pm\text{2.15}}$- 5.24±${}_{\pm\text{0.54}}$ 4.49±${}_{\pm\text{0.68}}$ 15.29±${}_{\pm\text{1.82}}$ 4.58±${}_{\pm\text{0.70}}$ 3.89±${}_{\pm\text{0.28}}$ N/A 17.66±${}_{\pm\text{1.80}}$
texture 50 78.67±${}_{\pm\text{2.72}}$ 22.15±${}_{\pm\text{10.01}}$ 12.36±${}_{\pm\text{3.58}}$ 11.19±${}_{\pm\text{3.12}}$ 15.85±${}_{\pm\text{5.14}}$ 13.67±${}_{\pm\text{6.44}}$ 9.05±${}_{\pm\text{0.13}}$ N/A 75.57±${}_{\pm\text{2.61}}$
100 85.31±${}_{\pm\text{2.44}}$ 25.80±${}_{\pm\text{5.61}}$ 13.82±${}_{\pm\text{7.03}}$ 11.67±${}_{\pm\text{4.14}}$ 19.32±${}_{\pm\text{4.92}}$ 25.54±${}_{\pm\text{5.86}}$ 9.09±${}_{\pm\text{0.10}}$ N/A 84.69±${}_{\pm\text{1.70}}$
200 88.17±${}_{\pm\text{1.71}}$ 18.33±${}_{\pm\text{7.83}}$ 13.85±${}_{\pm\text{3.95}}$ 10.55±${}_{\pm\text{4.55}}$ 23.17±${}_{\pm\text{7.57}}$ 42.92±${}_{\pm\text{6.67}}$ 9.07±${}_{\pm\text{0.56}}$ N/A 89.16±${}_{\pm\text{1.75}}$
500 90.78±${}_{\pm\text{1.16}}$ 29.29±${}_{\pm\text{7.03}}$ 24.91±${}_{\pm\text{9.99}}$ 10.23±${}_{\pm\text{3.48}}$ 23.85±${}_{\pm\text{9.99}}$ 57.58±${}_{\pm\text{4.98}}$ 10.12±${}_{\pm\text{2.55}}$ N/A 91.23±${}_{\pm\text{0.95}}$
Average rank 2.03±${}_{\pm\text{0.99}}$ 4.39±${}_{\pm\text{0.84}}$ 6.03±${}_{\pm\text{1.07}}$ 7.70±${}_{\pm\text{0.81}}$ 6.97±${}_{\pm\text{2.05}}$ 5.30±${}_{\pm\text{0.92}}$ 8.33±${}_{\pm\text{0.82}}$ 2.73±${}_{\pm\text{0.83}}$ 1.52±${}_{\pm\text{0.51}}$
Table 25: Classification accuracy (%) of MLP, comparing data sharing on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes the inapplicability of a specific generator. Different from Table 1, “--” denotes a generator cannot satisfy the requirement of generating 500 stratified samples even after generating 10,000 synthetic samples. The results of these inapplicable or failed generators are computed with the mean results of other methods. We bold the highest result for each dataset of different sample sizes. TVAE learns the joint distribution $p({\mathbf{x}},y)$ and fails to maintain the original training label distribution. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 15.81±${}_{\pm\text{3.26}}$ 12.08±${}_{\pm\text{4.27}}$ 13.06±${}_{\pm\text{5.22}}$ 12.93±${}_{\pm\text{3.80}}$ 20.17±${}_{\pm\text{4.44}}$ 12.98±${}_{\pm\text{2.50}}$ 36.01±${}_{\pm\text{2.68}}$ 36.23±${}_{\pm\text{2.49}}$
50 56.85±${}_{\pm\text{4.56}}$ 26.83±${}_{\pm\text{4.86}}$ 14.19±${}_{\pm\text{4.81}}$ 13.68±${}_{\pm\text{1.80}}$ 14.92±${}_{\pm\text{3.38}}$ 31.36±${}_{\pm\text{4.84}}$ 13.78±${}_{\pm\text{3.91}}$ 58.69±${}_{\pm\text{4.22}}$ 58.83±${}_{\pm\text{4.59}}$
100 75.86±${}_{\pm\text{3.03}}$ 41.60±${}_{\pm\text{4.43}}$ 25.78±${}_{\pm\text{4.77}}$ 11.34±${}_{\pm\text{3.38}}$ 20.82±${}_{\pm\text{4.55}}$ 40.31±${}_{\pm\text{5.63}}$ 14.54±${}_{\pm\text{3.04}}$ 77.47±${}_{\pm\text{3.39}}$ 77.56±${}_{\pm\text{3.65}}$
200 87.85±${}_{\pm\text{1.99}}$ 50.87±${}_{\pm\text{6.26}}$ 32.49±${}_{\pm\text{6.76}}$ 12.90±${}_{\pm\text{3.77}}$ 23.85±${}_{\pm\text{5.06}}$ 45.40±${}_{\pm\text{5.31}}$ 10.24±${}_{\pm\text{3.54}}$ 90.01±${}_{\pm\text{2.00}}$ 89.48±${}_{\pm\text{2.00}}$
500 94.37±${}_{\pm\text{1.66}}$ 62.81±${}_{\pm\text{4.29}}$ 47.22±${}_{\pm\text{6.43}}$ 12.06±${}_{\pm\text{2.49}}$ 24.51±${}_{\pm\text{7.67}}$ 60.31±${}_{\pm\text{4.02}}$ 10.41±${}_{\pm\text{2.60}}$ 96.26±${}_{\pm\text{1.35}}$ 96.03±${}_{\pm\text{1.56}}$
fourier 20 N/A -- 12.16±${}_{\pm\text{2.77}}$ 10.56±${}_{\pm\text{3.94}}$ 10.38±${}_{\pm\text{2.81}}$ 19.66±${}_{\pm\text{4.97}}$ 11.04±${}_{\pm\text{3.61}}$ 34.40±${}_{\pm\text{3.85}}$ 35.04±${}_{\pm\text{3.63}}$
50 52.72±${}_{\pm\text{2.11}}$- N/A 9.42±${}_{\pm\text{3.23}}$ 13.78±${}_{\pm\text{3.07}}$ 33.94±${}_{\pm\text{2.46}}$ 10.88±${}_{\pm\text{3.14}}$ 55.12±${}_{\pm\text{1.81}}$ 55.30±${}_{\pm\text{1.65}}$
100 60.82±${}_{\pm\text{2.85}}$ 33.16±${}_{\pm\text{3.91}}$ 21.44±${}_{\pm\text{3.15}}$ 10.88±${}_{\pm\text{4.21}}$ 21.96±${}_{\pm\text{2.34}}$ 39.68±${}_{\pm\text{4.04}}$ 8.12±${}_{\pm\text{3.25}}$ 63.54±${}_{\pm\text{1.81}}$ 64.08±${}_{\pm\text{1.87}}$
200 68.64±${}_{\pm\text{2.78}}$ 42.18±${}_{\pm\text{3.95}}$ 30.56±${}_{\pm\text{5.91}}$ 10.94±${}_{\pm\text{4.87}}$ 33.80±${}_{\pm\text{5.10}}$ 43.92±${}_{\pm\text{2.17}}$ 8.46±${}_{\pm\text{2.68}}$ 71.76±${}_{\pm\text{1.62}}$ 71.36±${}_{\pm\text{1.94}}$
500 73.78±${}_{\pm\text{1.12}}$ 50.48±${}_{\pm\text{4.03}}$ 45.94±${}_{\pm\text{3.55}}$ 13.10±${}_{\pm\text{4.09}}$ 35.18±${}_{\pm\text{5.84}}$ 51.10±${}_{\pm\text{3.53}}$ 9.66±${}_{\pm\text{2.62}}$ 76.50±${}_{\pm\text{1.87}}$ 77.65±${}_{\pm\text{1.81}}$
biodeg 20 68.59±${}_{\pm\text{4.24}}$ 62.24±${}_{\pm\text{8.67}}$ 53.56±${}_{\pm\text{5.65}}$ 55.15±${}_{\pm\text{9.09}}$ 45.87±${}_{\pm\text{3.29}}$ 57.76±${}_{\pm\text{4.42}}$ 45.98±${}_{\pm\text{11.23}}$ 72.04±${}_{\pm\text{4.95}}$ 72.10±${}_{\pm\text{4.69}}$
50 73.48±${}_{\pm\text{2.67}}$ 65.96±${}_{\pm\text{6.24}}$ 62.15±${}_{\pm\text{7.11}}$ 55.84±${}_{\pm\text{11.19}}$ 45.55±${}_{\pm\text{8.63}}$ 72.40±${}_{\pm\text{3.40}}$ 49.95±${}_{\pm\text{6.48}}$ 77.20±${}_{\pm\text{2.84}}$ 77.24±${}_{\pm\text{3.18}}$
100 77.20±${}_{\pm\text{1.72}}$ 70.20±${}_{\pm\text{5.67}}$ 71.90±${}_{\pm\text{3.27}}$ 57.99±${}_{\pm\text{6.21}}$ 49.10±${}_{\pm\text{5.76}}$ 71.42±${}_{\pm\text{2.59}}$ 45.52±${}_{\pm\text{6.61}}$ 78.58±${}_{\pm\text{2.25}}$ 78.93±${}_{\pm\text{2.11}}$
200 81.30±${}_{\pm\text{0.91}}$ 72.26±${}_{\pm\text{5.86}}$ 71.59±${}_{\pm\text{4.85}}$ 55.75±${}_{\pm\text{7.93}}$ 44.37±${}_{\pm\text{7.27}}$ 73.82±${}_{\pm\text{4.38}}$ 48.45±${}_{\pm\text{8.83}}$ 82.06±${}_{\pm\text{1.67}}$ 82.25±${}_{\pm\text{1.69}}$
500 83.55±${}_{\pm\text{0.62}}$ 74.36±${}_{\pm\text{4.32}}$ 77.55±${}_{\pm\text{2.54}}$ 53.37±${}_{\pm\text{4.84}}$ 43.97±${}_{\pm\text{4.60}}$ 76.51±${}_{\pm\text{2.60}}$ 47.04±${}_{\pm\text{9.41}}$ 83.63±${}_{\pm\text{1.19}}$ 83.33±${}_{\pm\text{0.86}}$
steel 20 57.49±${}_{\pm\text{5.60}}$ 53.28±${}_{\pm\text{11.47}}$ 50.57±${}_{\pm\text{7.01}}$ 52.01±${}_{\pm\text{8.53}}$ 52.76±${}_{\pm\text{9.11}}$ 51.25±${}_{\pm\text{4.91}}$ 47.98±${}_{\pm\text{5.99}}$ 64.34±${}_{\pm\text{6.21}}$ 64.33±${}_{\pm\text{5.70}}$
50 66.70±${}_{\pm\text{3.60}}$ 62.02±${}_{\pm\text{5.08}}$ 57.41±${}_{\pm\text{5.18}}$ 54.51±${}_{\pm\text{6.50}}$ 50.58±${}_{\pm\text{7.20}}$ 59.85±${}_{\pm\text{5.16}}$ 50.78±${}_{\pm\text{6.19}}$ 83.06±${}_{\pm\text{6.05}}$ 82.38±${}_{\pm\text{5.79}}$
100 76.17±${}_{\pm\text{4.30}}$ 63.05±${}_{\pm\text{4.11}}$ 60.62±${}_{\pm\text{4.71}}$ 54.81±${}_{\pm\text{5.06}}$ 44.64±${}_{\pm\text{9.29}}$ 59.60±${}_{\pm\text{4.83}}$ 48.93±${}_{\pm\text{2.99}}$ 95.36±${}_{\pm\text{3.13}}$ 95.45±${}_{\pm\text{3.29}}$
200 80.23±${}_{\pm\text{2.17}}$ 68.45±${}_{\pm\text{3.46}}$ 64.73±${}_{\pm\text{5.71}}$ 50.83±${}_{\pm\text{4.76}}$ 39.73±${}_{\pm\text{12.23}}$ 62.36±${}_{\pm\text{3.96}}$ 47.89±${}_{\pm\text{7.71}}$ 98.34±${}_{\pm\text{1.12}}$ 98.37±${}_{\pm\text{0.81}}$
500 85.76±${}_{\pm\text{3.61}}$ 73.89±${}_{\pm\text{4.48}}$ 71.48±${}_{\pm\text{4.21}}$ 49.32±${}_{\pm\text{6.51}}$ 43.65±${}_{\pm\text{5.66}}$ 68.12±${}_{\pm\text{4.63}}$ 53.56±${}_{\pm\text{4.70}}$ 99.56±${}_{\pm\text{0.34}}$ 99.45±${}_{\pm\text{0.44}}$
stock 20 82.93±${}_{\pm\text{4.66}}$ 71.78±${}_{\pm\text{7.33}}$ 55.09±${}_{\pm\text{13.94}}$ 68.46±${}_{\pm\text{8.87}}$ 76.71±${}_{\pm\text{15.45}}$ 62.13±${}_{\pm\text{10.43}}$ 52.68±${}_{\pm\text{12.66}}$ 83.83±${}_{\pm\text{3.86}}$ 83.91±${}_{\pm\text{3.99}}$
50 89.48±${}_{\pm\text{2.12}}$ 75.93±${}_{\pm\text{3.72}}$ 67.04±${}_{\pm\text{4.78}}$ 73.18±${}_{\pm\text{9.70}}$ 87.47±${}_{\pm\text{2.35}}$ 78.81±${}_{\pm\text{3.68}}$ 48.79±${}_{\pm\text{5.17}}$ 90.25±${}_{\pm\text{1.96}}$ 90.36±${}_{\pm\text{2.20}}$
100 90.85±${}_{\pm\text{0.72}}$ 81.61±${}_{\pm\text{3.37}}$ 79.76±${}_{\pm\text{3.51}}$ 71.03±${}_{\pm\text{5.88}}$ 90.17±${}_{\pm\text{1.66}}$ 85.83±${}_{\pm\text{2.70}}$ 50.41±${}_{\pm\text{14.72}}$ 91.60±${}_{\pm\text{1.00}}$ 91.72±${}_{\pm\text{0.97}}$
200 92.03±${}_{\pm\text{0.79}}$ 84.67±${}_{\pm\text{2.45}}$ 82.92±${}_{\pm\text{2.01}}$ 74.97±${}_{\pm\text{4.41}}$ 91.21±${}_{\pm\text{1.00}}$ 87.06±${}_{\pm\text{2.18}}$ 44.90±${}_{\pm\text{17.08}}$ 92.24±${}_{\pm\text{0.91}}$ 92.00±${}_{\pm\text{0.93}}$
More than 10 classes energy 50 N/A 7.77±${}_{\pm\text{1.39}}$ 5.64±${}_{\pm\text{2.40}}$ 5.75±${}_{\pm\text{1.72}}$ 4.30±${}_{\pm\text{2.07}}$ 8.28±${}_{\pm\text{1.91}}$ 4.46±${}_{\pm\text{0.25}}$ N/A 23.91±${}_{\pm\text{1.49}}$
100 N/A -- 7.44±${}_{\pm\text{2.38}}$ 5.34±${}_{\pm\text{1.60}}$ 3.68±${}_{\pm\text{1.68}}$ 10.38±${}_{\pm\text{1.67}}$ 4.29±${}_{\pm\text{0.71}}$ N/A 29.24±${}_{\pm\text{2.45}}$
200 N/A -- 8.74±${}_{\pm\text{2.17}}$ 7.38±${}_{\pm\text{2.33}}$ 3.84±${}_{\pm\text{1.29}}$ 12.88±${}_{\pm\text{2.09}}$ 4.08±${}_{\pm\text{1.12}}$ N/A 38.27±${}_{\pm\text{3.50}}$
collins 100 N/A -- 6.06±${}_{\pm\text{0.96}}$ 5.04±${}_{\pm\text{1.41}}$ 12.80±${}_{\pm\text{1.48}}$ 5.69±${}_{\pm\text{1.80}}$ 3.67±${}_{\pm\text{0.43}}$ N/A 13.66±${}_{\pm\text{1.86}}$
200 18.91±${}_{\pm\text{1.62}}$- 5.78±${}_{\pm\text{1.75}}$ ${}_{\pm\text{0.94}}$ 15.94±${}_{\pm\text{1.74}}$ 5.84±${}_{\pm\text{1.19}}$ 4.42±${}_{\pm\text{1.18}}$ N/A 19.49±${}_{\pm\text{1.56}}$
texture 50 83.45±${}_{\pm\text{3.05}}$ 28.11±${}_{\pm\text{7.00}}$ 14.57±${}_{\pm\text{6.49}}$ 12.56±${}_{\pm\text{2.92}}$ 22.50±${}_{\pm\text{6.71}}$ 18.07±${}_{\pm\text{4.88}}$ 9.88±${}_{\pm\text{4.02}}$ N/A 85.41±${}_{\pm\text{2.73}}$
100 90.73±${}_{\pm\text{1.80}}$ 33.20±${}_{\pm\text{4.80}}$ 15.18±${}_{\pm\text{7.46}}$ 13.72±${}_{\pm\text{6.03}}$ 29.36±${}_{\pm\text{8.98}}$ 25.80±${}_{\pm\text{5.24}}$ 10.77±${}_{\pm\text{5.02}}$ N/A 91.87±${}_{\pm\text{1.42}}$
200 93.31±${}_{\pm\text{1.19}}$ 26.95±${}_{\pm\text{15.32}}$ 18.10±${}_{\pm\text{5.37}}$ 10.39±${}_{\pm\text{4.68}}$ 28.79±${}_{\pm\text{10.29}}$ 44.44±${}_{\pm\text{5.46}}$ 7.70±${}_{\pm\text{2.74}}$ N/A 93.78±${}_{\pm\text{1.41}}$
500 95.61±${}_{\pm\text{1.10}}$ 39.39±${}_{\pm\text{9.19}}$ 29.60±${}_{\pm\text{8.79}}$ 13.15±${}_{\pm\text{4.59}}$ 27.70±${}_{\pm\text{6.21}}$ 62.88±${}_{\pm\text{5.56}}$ ${}_{\pm\text{3.10}}$ N/A 95.06±${}_{\pm\text{0.96}}$
Average rank 2.91±${}_{\pm\text{0.71}}$ 4.73±${}_{\pm\text{0.75}}$ 6.45±${}_{\pm\text{1.12}}$ 7.48±${}_{\pm\text{0.87}}$ 6.79±${}_{\pm\text{2.13}}$ 4.70±${}_{\pm\text{1.33}}$ 8.39±${}_{\pm\text{0.70}}$ 2.21±${}_{\pm\text{1.05}}$ 1.33±${}_{\pm\text{0.60}}$
Table 26: Classification accuracy (%) of RF, comparing data sharing on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes the inapplicability of a specific generator. Different from Table 1, “--” denotes a generator cannot satisfy the requirement of generating 500 stratified samples even after generating 10,000 synthetic samples. The results of these inapplicable or failed generators are computed with the mean results of other methods. We bold the highest result for each dataset of different sample sizes. TVAE learns the joint distribution $p({\mathbf{x}},y)$ and fails to maintain the original training label distribution. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 19.29±${}_{\pm\text{3.85}}$ 13.63±${}_{\pm\text{2.95}}$ 12.71±${}_{\pm\text{2.84}}$ 12.75±${}_{\pm\text{2.09}}$ 20.47±${}_{\pm\text{4.29}}$ 13.25±${}_{\pm\text{2.90}}$ 31.86±${}_{\pm\text{2.24}}$ 34.05±${}_{\pm\text{2.24}}$
50 56.09±${}_{\pm\text{2.77}}$ 32.29±${}_{\pm\text{3.05}}$ 19.37±${}_{\pm\text{5.89}}$ 12.62±${}_{\pm\text{2.05}}$ 14.87±${}_{\pm\text{2.75}}$ 32.60±${}_{\pm\text{4.15}}$ 13.29±${}_{\pm\text{2.50}}$ 54.25±${}_{\pm\text{1.99}}$ 56.96±${}_{\pm\text{3.32}}$
100 71.14±${}_{\pm\text{2.65}}$ 44.50±${}_{\pm\text{3.08}}$ 29.72±${}_{\pm\text{3.69}}$ ${}_{\pm\text{3.12}}$ 17.71±${}_{\pm\text{3.51}}$ 38.22±${}_{\pm\text{4.85}}$ 13.67±${}_{\pm\text{1.00}}$ 70.68±${}_{\pm\text{3.24}}$ 72.74±${}_{\pm\text{2.97}}$
200 81.22±${}_{\pm\text{2.51}}$ 50.10±${}_{\pm\text{3.89}}$ 37.08±${}_{\pm\text{5.17}}$ 14.34±${}_{\pm\text{3.15}}$ 22.23±${}_{\pm\text{3.43}}$ 43.48±${}_{\pm\text{3.74}}$ 11.76±${}_{\pm\text{3.27}}$ 81.93±${}_{\pm\text{2.47}}$ 82.35±${}_{\pm\text{2.09}}$
500 87.93±${}_{\pm\text{1.52}}$ 58.17±${}_{\pm\text{5.02}}$ 49.44±${}_{\pm\text{5.57}}$ 12.03±${}_{\pm\text{2.40}}$ 22.88±${}_{\pm\text{2.67}}$ 53.67±${}_{\pm\text{2.94}}$ 10.59±${}_{\pm\text{3.14}}$ 90.81±${}_{\pm\text{1.37}}$ 89.78±${}_{\pm\text{1.73}}$
fourier 20 N/A -- 11.28±${}_{\pm\text{1.63}}$ 11.52±${}_{\pm\text{2.65}}$ 9.24±${}_{\pm\text{2.16}}$ 17.76±${}_{\pm\text{3.96}}$ 9.78±${}_{\pm\text{1.33}}$ 31.94±${}_{\pm\text{3.25}}$ 38.16±${}_{\pm\text{4.48}}$
50 65.32±${}_{\pm\text{3.86}}$- N/A 9.98±${}_{\pm\text{1.98}}$ 12.12±${}_{\pm\text{2.56}}$ 31.76±${}_{\pm\text{3.66}}$ 11.24±${}_{\pm\text{2.70}}$ 65.64±${}_{\pm\text{3.73}}$ 66.04±${}_{\pm\text{2.91}}$
100 73.82±${}_{\pm\text{2.70}}$ 44.84±${}_{\pm\text{5.88}}$ 29.68±${}_{\pm\text{5.97}}$ 11.04±${}_{\pm\text{2.51}}$ 21.90±${}_{\pm\text{4.68}}$ 35.60±${}_{\pm\text{4.77}}$ 9.30±${}_{\pm\text{2.87}}$ 74.28±${}_{\pm\text{2.81}}$ 74.92±${}_{\pm\text{3.12}}$
200 78.66±${}_{\pm\text{1.64}}$ 48.56±${}_{\pm\text{3.52}}$ 41.94±${}_{\pm\text{5.54}}$ 10.24±${}_{\pm\text{2.93}}$ 31.36±${}_{\pm\text{4.10}}$ 44.80±${}_{\pm\text{3.88}}$ 9.88±${}_{\pm\text{2.32}}$ 79.48±${}_{\pm\text{2.18}}$ 79.52±${}_{\pm\text{2.03}}$
500 79.64±${}_{\pm\text{1.39}}$ 54.80±${}_{\pm\text{4.51}}$ 57.06±${}_{\pm\text{4.59}}$ 11.94±${}_{\pm\text{2.26}}$ 25.08±${}_{\pm\text{4.65}}$ 54.04±${}_{\pm\text{3.10}}$ 10.66±${}_{\pm\text{2.71}}$ 81.64±${}_{\pm\text{1.61}}$ 81.85±${}_{\pm\text{0.53}}$
biodeg 20 69.34±${}_{\pm\text{6.13}}$ 64.52±${}_{\pm\text{7.42}}$ 52.83±${}_{\pm\text{7.85}}$ 53.88±${}_{\pm\text{7.31}}$ 49.80±${}_{\pm\text{0.45}}$ 60.95±${}_{\pm\text{6.69}}$ 50.47±${}_{\pm\text{9.67}}$ 67.37±${}_{\pm\text{5.45}}$ 67.06±${}_{\pm\text{4.88}}$
50 70.17±${}_{\pm\text{3.69}}$ 68.14±${}_{\pm\text{3.98}}$ 62.50±${}_{\pm\text{6.74}}$ 54.84±${}_{\pm\text{7.67}}$ 49.94±${}_{\pm\text{1.03}}$ 67.21±${}_{\pm\text{4.25}}$ 51.99±${}_{\pm\text{6.12}}$ 71.94±${}_{\pm\text{4.16}}$ 71.91±${}_{\pm\text{2.77}}$
100 74.80±${}_{\pm\text{2.96}}$ 71.82±${}_{\pm\text{3.70}}$ 68.29±${}_{\pm\text{2.67}}$ 56.44±${}_{\pm\text{5.66}}$ 49.62±${}_{\pm\text{0.68}}$ 67.68±${}_{\pm\text{2.88}}$ 50.71±${}_{\pm\text{5.41}}$ 75.68±${}_{\pm\text{2.01}}$ 74.93±${}_{\pm\text{2.08}}$
200 77.49±${}_{\pm\text{2.13}}$ 69.22±${}_{\pm\text{3.80}}$ 69.63±${}_{\pm\text{2.86}}$ 55.47±${}_{\pm\text{7.86}}$ 48.70±${}_{\pm\text{1.53}}$ 70.40±${}_{\pm\text{3.01}}$ 48.04±${}_{\pm\text{5.01}}$ 78.64±${}_{\pm\text{2.29}}$ 76.96±${}_{\pm\text{3.02}}$
500 80.31±${}_{\pm\text{1.32}}$ 74.83±${}_{\pm\text{1.76}}$ 74.44±${}_{\pm\text{2.90}}$ 51.30±${}_{\pm\text{3.00}}$ 49.68±${}_{\pm\text{0.61}}$ 73.88±${}_{\pm\text{1.59}}$ 45.73±${}_{\pm\text{8.13}}$ 79.56±${}_{\pm\text{1.95}}$ 78.66±${}_{\pm\text{1.76}}$
steel 20 55.71±${}_{\pm\text{3.47}}$ 55.23±${}_{\pm\text{4.31}}$ 52.14±${}_{\pm\text{3.64}}$ 50.93±${}_{\pm\text{6.29}}$ 49.93±${}_{\pm\text{0.56}}$ 52.02±${}_{\pm\text{5.80}}$ 48.88±${}_{\pm\text{4.70}}$ 57.20±${}_{\pm\text{2.92}}$ 58.27±${}_{\pm\text{2.40}}$
50 62.30±${}_{\pm\text{2.94}}$ 56.05±${}_{\pm\text{3.07}}$ 55.40±${}_{\pm\text{4.01}}$ 52.16±${}_{\pm\text{3.88}}$ 49.98±${}_{\pm\text{0.29}}$ 54.55±${}_{\pm\text{2.89}}$ 50.65±${}_{\pm\text{1.92}}$ 64.26±${}_{\pm\text{2.55}}$ 66.98±${}_{\pm\text{3.30}}$
100 67.63±${}_{\pm\text{3.75}}$ 58.93±${}_{\pm\text{4.82}}$ 56.65±${}_{\pm\text{4.04}}$ 51.86±${}_{\pm\text{2.52}}$ 47.84±${}_{\pm\text{2.76}}$ 56.13±${}_{\pm\text{2.99}}$ 49.33±${}_{\pm\text{2.96}}$ 72.47±${}_{\pm\text{4.01}}$ 77.42±${}_{\pm\text{3.85}}$
200 71.41±${}_{\pm\text{3.46}}$ 62.15±${}_{\pm\text{4.45}}$ 59.89±${}_{\pm\text{4.05}}$ 51.44±${}_{\pm\text{1.39}}$ 48.94±${}_{\pm\text{1.59}}$ 58.23±${}_{\pm\text{2.57}}$ 49.25±${}_{\pm\text{3.05}}$ 81.56±${}_{\pm\text{4.46}}$ 86.14±${}_{\pm\text{4.04}}$
500 77.17±${}_{\pm\text{2.01}}$ 67.42±${}_{\pm\text{3.15}}$ 65.34±${}_{\pm\text{4.17}}$ 50.86±${}_{\pm\text{3.85}}$ 49.83±${}_{\pm\text{0.48}}$ 60.96±${}_{\pm\text{4.42}}$ 51.04±${}_{\pm\text{3.62}}$ 86.49±${}_{\pm\text{4.38}}$ 90.60±${}_{\pm\text{4.65}}$
stock 20 80.81±${}_{\pm\text{4.71}}$ 71.58±${}_{\pm\text{6.48}}$ 48.00±${}_{\pm\text{11.14}}$ 66.07±${}_{\pm\text{12.69}}$ 77.42±${}_{\pm\text{15.77}}$ 61.80±${}_{\pm\text{9.34}}$ 46.62±${}_{\pm\text{11.62}}$ 84.30±${}_{\pm\text{4.96}}$ 84.65±${}_{\pm\text{3.38}}$
50 89.45±${}_{\pm\text{2.09}}$ 74.30±${}_{\pm\text{8.64}}$ 66.75±${}_{\pm\text{7.07}}$ 69.95±${}_{\pm\text{10.51}}$ 87.73±${}_{\pm\text{2.21}}$ 79.90±${}_{\pm\text{4.82}}$ 47.28±${}_{\pm\text{12.21}}$ 89.27±${}_{\pm\text{1.89}}$ 90.04±${}_{\pm\text{2.09}}$
100 91.38±${}_{\pm\text{1.80}}$ 84.09±${}_{\pm\text{3.75}}$ 79.16±${}_{\pm\text{4.14}}$ 72.00±${}_{\pm\text{4.77}}$ 90.96±${}_{\pm\text{2.04}}$ 86.38±${}_{\pm\text{2.97}}$ 47.78±${}_{\pm\text{10.15}}$ 92.07±${}_{\pm\text{1.56}}$ 92.22±${}_{\pm\text{1.36}}$
200 93.46±${}_{\pm\text{0.79}}$ 86.85±${}_{\pm\text{2.22}}$ 83.47±${}_{\pm\text{3.47}}$ 78.62±${}_{\pm\text{3.19}}$ 92.72±${}_{\pm\text{0.81}}$ 87.66±${}_{\pm\text{0.89}}$ 49.91±${}_{\pm\text{13.88}}$ 93.35±${}_{\pm\text{1.32}}$ 93.07±${}_{\pm\text{0.99}}$
More than 10 classes energy 50 N/A 6.42±${}_{\pm\text{1.49}}$ 4.52±${}_{\pm\text{2.22}}$ 4.88±${}_{\pm\text{1.49}}$ 4.67±${}_{\pm\text{1.31}}$ 6.62±${}_{\pm\text{2.47}}$ 4.47±${}_{\pm\text{0.38}}$ N/A 28.79±${}_{\pm\text{3.70}}$
100 N/A -- 7.92±${}_{\pm\text{1.14}}$ 5.40±${}_{\pm\text{1.11}}$ 4.21±${}_{\pm\text{0.94}}$ 9.03±${}_{\pm\text{1.53}}$ 3.37±${}_{\pm\text{1.21}}$ N/A 39.20±${}_{\pm\text{3.05}}$
200 N/A -- 7.12±${}_{\pm\text{1.77}}$ 6.01±${}_{\pm\text{1.38}}$ 3.26±${}_{\pm\text{0.91}}$ 11.78±${}_{\pm\text{3.10}}$ 4.32±${}_{\pm\text{0.75}}$ N/A 42.55±${}_{\pm\text{3.93}}$
collins 100 N/A -- 6.10±${}_{\pm\text{0.63}}$ 4.09±${}_{\pm\text{0.82}}$ 11.95±${}_{\pm\text{1.92}}$ 4.95±${}_{\pm\text{1.23}}$ 3.99±${}_{\pm\text{0.64}}$ N/A 13.67±${}_{\pm\text{1.31}}$
200 17.74±${}_{\pm\text{1.75}}$- 5.86±${}_{\pm\text{1.39}}$ 4.56±${}_{\pm\text{0.98}}$ 13.22±${}_{\pm\text{1.22}}$ 5.23±${}_{\pm\text{0.63}}$ 4.08±${}_{\pm\text{0.62}}$ N/A 16.24±${}_{\pm\text{1.45}}$
texture 50 71.18±${}_{\pm\text{2.80}}$ 22.41±${}_{\pm\text{7.88}}$ 10.33±${}_{\pm\text{1.47}}$ 10.55±${}_{\pm\text{1.87}}$ 12.35±${}_{\pm\text{4.09}}$ 19.83±${}_{\pm\text{5.72}}$ 9.50±${}_{\pm\text{2.91}}$ N/A 75.77±${}_{\pm\text{3.60}}$
100 79.85±${}_{\pm\text{2.02}}$ 26.34±${}_{\pm\text{1.17}}$ 14.84±${}_{\pm\text{6.52}}$ 15.77±${}_{\pm\text{5.11}}$ 14.16±${}_{\pm\text{4.42}}$ 26.07±${}_{\pm\text{6.44}}$ 8.97±${}_{\pm\text{0.33}}$ N/A 81.99±${}_{\pm\text{2.39}}$
200 84.17±${}_{\pm\text{2.07}}$ 22.70±${}_{\pm\text{5.57}}$ 12.59±${}_{\pm\text{4.81}}$ 11.61±${}_{\pm\text{3.40}}$ 11.68±${}_{\pm\text{3.73}}$ 46.83±${}_{\pm\text{4.78}}$ 10.17±${}_{\pm\text{3.08}}$ N/A 84.90±${}_{\pm\text{1.98}}$
500 88.31±${}_{\pm\text{1.36}}$ 32.92±${}_{\pm\text{10.58}}$ 27.62±${}_{\pm\text{9.73}}$ 13.33±${}_{\pm\text{4.79}}$ 10.84±${}_{\pm\text{2.37}}$ 62.04±${}_{\pm\text{4.01}}$ 9.50±${}_{\pm\text{3.84}}$ N/A 88.14±${}_{\pm\text{1.50}}$
Average rank 2.55±${}_{\pm\text{0.84}}$ 4.39±${}_{\pm\text{0.80}}$ 6.09±${}_{\pm\text{1.01}}$ 7.42±${}_{\pm\text{0.90}}$ 7.09±${}_{\pm\text{1.94}}$ 5.15±${}_{\pm\text{1.03}}$ 8.55±${}_{\pm\text{0.62}}$ 2.36±${}_{\pm\text{0.89}}$ 1.39±${}_{\pm\text{0.70}}$
Table 27: Classification accuracy (%) of XGBoost, comparing data sharing on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes the inapplicability of a specific generator. Different from Table 1, “--” denotes a generator cannot satisfy the requirement of generating 500 stratified samples even after generating 10,000 synthetic samples. The results of these inapplicable or failed generators are computed with the mean results of other methods. We bold the highest result for each dataset of different sample sizes. TVAE learns the joint distribution $p({\mathbf{x}},y)$ and fails to maintain the original training label distribution. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 15.47±${}_{\pm\text{2.77}}$ 13.60±${}_{\pm\text{3.62}}$ 15.03±${}_{\pm\text{4.60}}$ 11.77±${}_{\pm\text{3.36}}$ 18.32±${}_{\pm\text{3.84}}$ 14.61±${}_{\pm\text{2.21}}$ 23.54±${}_{\pm\text{4.33}}$ 25.22±${}_{\pm\text{3.56}}$
50 38.51±${}_{\pm\text{4.83}}$ 25.36±${}_{\pm\text{3.19}}$ 18.88±${}_{\pm\text{4.31}}$ 13.84±${}_{\pm\text{4.50}}$ 12.72±${}_{\pm\text{1.99}}$ 24.39±${}_{\pm\text{4.09}}$ 13.13±${}_{\pm\text{4.44}}$ 41.88±${}_{\pm\text{6.20}}$ 43.79±${}_{\pm\text{6.81}}$
100 60.71±${}_{\pm\text{5.70}}$ 33.23±${}_{\pm\text{4.65}}$ 23.54±${}_{\pm\text{3.74}}$ 14.36±${}_{\pm\text{3.09}}$ 14.08±${}_{\pm\text{1.84}}$ 28.49±${}_{\pm\text{6.26}}$ 14.48±${}_{\pm\text{4.96}}$ 54.91±${}_{\pm\text{6.96}}$ ${}_{\pm\text{6.11}}$
200 73.31±${}_{\pm\text{3.15}}$ 35.50±${}_{\pm\text{6.72}}$ 26.74±${}_{\pm\text{5.01}}$ 15.45±${}_{\pm\text{4.45}}$ 14.71±${}_{\pm\text{3.38}}$ 36.61±${}_{\pm\text{3.61}}$ 10.41±${}_{\pm\text{2.52}}$ 75.20±${}_{\pm\text{2.13}}$ 76.96±${}_{\pm\text{2.90}}$
500 83.11±${}_{\pm\text{1.82}}$ 41.18±${}_{\pm\text{6.31}}$ 37.95±${}_{\pm\text{5.62}}$ 13.56±${}_{\pm\text{2.40}}$ 14.47±${}_{\pm\text{2.40}}$ 48.92±${}_{\pm\text{4.09}}$ 11.25±${}_{\pm\text{3.92}}$ 85.30±${}_{\pm\text{1.56}}$ 85.39±${}_{\pm\text{1.90}}$
fourier 20 N/A -- 11.16±${}_{\pm\text{2.78}}$ 10.36±${}_{\pm\text{2.58}}$ 12.38±${}_{\pm\text{3.08}}$ 14.64±${}_{\pm\text{4.90}}$ 9.18±${}_{\pm\text{1.60}}$ 21.80±${}_{\pm\text{4.86}}$ 26.80±${}_{\pm\text{4.82}}$
50 41.70±${}_{\pm\text{8.00}}$- N/A 10.06±${}_{\pm\text{1.96}}$ 10.00±${}_{\pm\text{1.41}}$ 19.96±${}_{\pm\text{4.44}}$ 12.24±${}_{\pm\text{3.23}}$ 43.10±${}_{\pm\text{6.37}}$ 47.16±${}_{\pm\text{5.42}}$
100 54.52±${}_{\pm\text{5.80}}$ 27.78±${}_{\pm\text{5.49}}$ 19.40±${}_{\pm\text{4.93}}$ 10.98±${}_{\pm\text{2.29}}$ 12.22±${}_{\pm\text{3.24}}$ 24.24±${}_{\pm\text{4.70}}$ 10.34±${}_{\pm\text{2.64}}$ 51.40±${}_{\pm\text{5.42}}$ 52.62±${}_{\pm\text{5.05}}$
200 65.96±${}_{\pm\text{4.60}}$ 31.60±${}_{\pm\text{2.49}}$ 24.76±${}_{\pm\text{5.83}}$ 11.84±${}_{\pm\text{3.99}}$ 13.12±${}_{\pm\text{3.24}}$ 31.74±${}_{\pm\text{3.75}}$ 11.36±${}_{\pm\text{2.39}}$ 62.68±${}_{\pm\text{4.32}}$ 66.26±${}_{\pm\text{4.52}}$
500 71.44±${}_{\pm\text{1.57}}$ 34.72±${}_{\pm\text{1.86}}$ 36.56±${}_{\pm\text{5.98}}$ 11.64±${}_{\pm\text{2.58}}$ 12.72±${}_{\pm\text{3.83}}$ 37.86±${}_{\pm\text{3.31}}$ 9.42±${}_{\pm\text{1.66}}$ 72.00±${}_{\pm\text{2.94}}$ 74.45±${}_{\pm\text{1.66}}$
biodeg 20 66.52±${}_{\pm\text{5.96}}$ 60.35±${}_{\pm\text{7.63}}$ 58.55±${}_{\pm\text{6.57}}$ 54.65±${}_{\pm\text{8.42}}$ 50.67±${}_{\pm\text{4.16}}$ 59.80±${}_{\pm\text{7.09}}$ 52.58±${}_{\pm\text{13.73}}$ 65.99±${}_{\pm\text{6.54}}$ 66.60±${}_{\pm\text{6.87}}$
50 68.01±${}_{\pm\text{2.67}}$ 63.62±${}_{\pm\text{4.85}}$ 60.53±${}_{\pm\text{7.91}}$ 56.10±${}_{\pm\text{6.94}}$ 48.13±${}_{\pm\text{3.77}}$ 68.90±${}_{\pm\text{4.66}}$ 52.89±${}_{\pm\text{9.18}}$ 68.60±${}_{\pm\text{4.27}}$ 68.82±${}_{\pm\text{3.18}}$
100 71.98±${}_{\pm\text{4.21}}$ 68.71±${}_{\pm\text{4.56}}$ 68.58±${}_{\pm\text{3.57}}$ 60.67±${}_{\pm\text{9.00}}$ 49.08±${}_{\pm\text{3.11}}$ 68.74±${}_{\pm\text{3.58}}$ 49.51±${}_{\pm\text{8.42}}$ 74.73±${}_{\pm\text{1.88}}$ 72.69±${}_{\pm\text{3.36}}$
200 74.29±${}_{\pm\text{2.14}}$ 68.94±${}_{\pm\text{4.03}}$ 69.87±${}_{\pm\text{3.15}}$ 59.26±${}_{\pm\text{7.34}}$ 47.26±${}_{\pm\text{2.99}}$ 70.89±${}_{\pm\text{3.66}}$ 55.64±${}_{\pm\text{7.90}}$ 76.03±${}_{\pm\text{4.77}}$ 77.15±${}_{\pm\text{2.53}}$
500 76.04±${}_{\pm\text{3.09}}$ 70.53±${}_{\pm\text{5.19}}$ 73.65±${}_{\pm\text{3.54}}$ 53.73±${}_{\pm\text{4.78}}$ 47.86±${}_{\pm\text{4.57}}$ 74.39±${}_{\pm\text{2.84}}$ 49.08±${}_{\pm\text{11.19}}$ 78.99±${}_{\pm\text{2.89}}$ 77.61±${}_{\pm\text{2.83}}$
steel 20 55.76±${}_{\pm\text{4.40}}$ 53.17±${}_{\pm\text{6.28}}$ 53.71±${}_{\pm\text{9.38}}$ 51.36±${}_{\pm\text{5.69}}$ 51.32±${}_{\pm\text{6.29}}$ 51.36±${}_{\pm\text{6.40}}$ 51.11±${}_{\pm\text{4.18}}$ 55.71±${}_{\pm\text{5.46}}$ 54.95±${}_{\pm\text{5.04}}$
50 61.07±${}_{\pm\text{3.58}}$ 56.97±${}_{\pm\text{3.93}}$ 54.38±${}_{\pm\text{4.00}}$ 53.54±${}_{\pm\text{3.82}}$ 46.83±${}_{\pm\text{6.18}}$ 53.60±${}_{\pm\text{4.17}}$ 49.46±${}_{\pm\text{6.57}}$ 63.64±${}_{\pm\text{6.52}}$ 71.58±${}_{\pm\text{14.28}}$
100 64.12±${}_{\pm\text{5.76}}$ 59.03±${}_{\pm\text{5.74}}$ 57.01±${}_{\pm\text{5.40}}$ 52.86±${}_{\pm\text{3.06}}$ 44.22±${}_{\pm\text{7.97}}$ 55.21±${}_{\pm\text{4.00}}$ 46.92±${}_{\pm\text{7.56}}$ 88.61±${}_{\pm\text{12.72}}$ 96.34±${}_{\pm\text{2.73}}$
200 74.96±${}_{\pm\text{6.42}}$ 58.42±${}_{\pm\text{6.66}}$ 59.25±${}_{\pm\text{3.48}}$ 51.94±${}_{\pm\text{3.30}}$ 43.96±${}_{\pm\text{10.64}}$ 58.92±${}_{\pm\text{3.56}}$ 49.58±${}_{\pm\text{5.96}}$ 98.85±${}_{\pm\text{1.71}}$ 99.28±${}_{\pm\text{0.72}}$
500 81.79±${}_{\pm\text{5.79}}$ 61.11±${}_{\pm\text{8.40}}$ 65.70±${}_{\pm\text{7.71}}$ 51.68±${}_{\pm\text{2.71}}$ 46.63±${}_{\pm\text{6.69}}$ 59.19±${}_{\pm\text{7.43}}$ 49.98±${}_{\pm\text{5.15}}$ 99.62±${}_{\pm\text{1.11}}$ 99.92±${}_{\pm\text{0.13}}$
stock 20 77.64±${}_{\pm\text{6.07}}$ 64.61±${}_{\pm\text{10.04}}$ 50.93±${}_{\pm\text{6.99}}$ 67.70±${}_{\pm\text{12.94}}$ 74.63±${}_{\pm\text{13.41}}$ 68.29±${}_{\pm\text{7.28}}$ 45.59±${}_{\pm\text{10.96}}$ 80.14±${}_{\pm\text{4.76}}$ 84.22±${}_{\pm\text{4.50}}$
50 84.83±${}_{\pm\text{2.98}}$ 77.89±${}_{\pm\text{7.53}}$ 73.57±${}_{\pm\text{5.76}}$ 70.29±${}_{\pm\text{6.08}}$ 84.75±${}_{\pm\text{4.01}}$ 79.19±${}_{\pm\text{2.51}}$ 54.82±${}_{\pm\text{11.81}}$ 86.97±${}_{\pm\text{3.25}}$ 88.59±${}_{\pm\text{3.54}}$
100 87.69±${}_{\pm\text{1.26}}$ 80.12±${}_{\pm\text{8.75}}$ 77.23±${}_{\pm\text{4.37}}$ 71.03±${}_{\pm\text{8.08}}$ 85.72±${}_{\pm\text{4.00}}$ 83.91±${}_{\pm\text{3.27}}$ 55.60±${}_{\pm\text{10.94}}$ 89.21±${}_{\pm\text{3.39}}$ 91.14±${}_{\pm\text{1.97}}$
200 90.58±${}_{\pm\text{3.02}}$ 83.34±${}_{\pm\text{3.80}}$ 81.04±${}_{\pm\text{3.24}}$ 76.60±${}_{\pm\text{5.86}}$ 91.52±${}_{\pm\text{3.49}}$ 87.49±${}_{\pm\text{3.41}}$ 55.43±${}_{\pm\text{12.95}}$ 90.78±${}_{\pm\text{3.56}}$ 90.94±${}_{\pm\text{1.18}}$
More than 10 classes energy 50 N/A 7.37±${}_{\pm\text{2.92}}$ 5.16±${}_{\pm\text{2.03}}$ 4.90±${}_{\pm\text{1.44}}$ 3.95±${}_{\pm\text{1.60}}$ 6.62±${}_{\pm\text{1.84}}$ N/A N/A 19.38±${}_{\pm\text{3.84}}$
100 N/A -- 8.25±${}_{\pm\text{2.78}}$ 5.40±${}_{\pm\text{2.24}}$ 3.95±${}_{\pm\text{1.37}}$ 10.21±${}_{\pm\text{2.45}}$ N/A N/A 25.08±${}_{\pm\text{5.99}}$
200 N/A -- 6.97±${}_{\pm\text{2.92}}$ 8.56±${}_{\pm\text{2.68}}$ 3.21±${}_{\pm\text{1.53}}$ 11.08±${}_{\pm\text{2.96}}$ N/A N/A 31.03±${}_{\pm\text{4.94}}$
collins 100 N/A -- 4.45±${}_{\pm\text{0.82}}$ 3.98±${}_{\pm\text{0.70}}$ 8.18±${}_{\pm\text{1.43}}$ 5.39±${}_{\pm\text{1.68}}$ N/A N/A 9.03±${}_{\pm\text{1.14}}$
200 12.52±${}_{\pm\text{1.98}}$- 5.52±${}_{\pm\text{1.57}}$ 4.43±${}_{\pm\text{0.63}}$ 9.06±${}_{\pm\text{1.21}}$ 5.61±${}_{\pm\text{1.10}}$ N/A N/A 11.49±${}_{\pm\text{1.45}}$
texture 50 57.99±${}_{\pm\text{3.35}}$ 20.94±${}_{\pm\text{8.45}}$ 9.44±${}_{\pm\text{4.78}}$ 12.63±${}_{\pm\text{3.07}}$ 11.32±${}_{\pm\text{4.19}}$ 14.35±${}_{\pm\text{4.95}}$ N/A N/A 68.97±${}_{\pm\text{4.98}}$
100 69.80±${}_{\pm\text{4.34}}$ 23.00±${}_{\pm\text{7.10}}$ 14.06±${}_{\pm\text{4.92}}$ 16.18±${}_{\pm\text{5.64}}$ 9.90±${}_{\pm\text{2.39}}$ 19.13±${}_{\pm\text{6.41}}$ N/A N/A 76.62±${}_{\pm\text{2.55}}$
200 78.46±${}_{\pm\text{2.72}}$ 23.32±${}_{\pm\text{7.04}}$ 15.09±${}_{\pm\text{2.96}}$ 15.15±${}_{\pm\text{3.79}}$ 9.66±${}_{\pm\text{1.64}}$ 33.67±${}_{\pm\text{5.07}}$ N/A N/A 80.37±${}_{\pm\text{2.13}}$
500 86.39±${}_{\pm\text{1.70}}$ 31.13±${}_{\pm\text{9.16}}$ 22.88±${}_{\pm\text{7.24}}$ 13.46±${}_{\pm\text{5.72}}$ 10.22±${}_{\pm\text{1.70}}$ 47.06±${}_{\pm\text{3.83}}$ N/A N/A 85.83±${}_{\pm\text{1.68}}$
Average rank 2.76±${}_{\pm\text{0.90}}$ 4.97±${}_{\pm\text{0.92}}$ 6.39±${}_{\pm\text{1.32}}$ 7.42±${}_{\pm\text{0.75}}$ 7.48±${}_{\pm\text{2.35}}$ 4.97±${}_{\pm\text{1.26}}$ 7.11±${}_{\pm\text{2.15}}$ 2.62±${}_{\pm\text{0.95}}$ 1.27±${}_{\pm\text{0.52}}$
Table 28: Classification accuracy (%) of TabPFN, comparing data sharing on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std balanced accuracy and average accuracy rank across datasets. A higher rank implies higher accuracy. Note that “N/A” denotes the inapplicability of a specific generator. Different from Table 1, “--” denotes a generator cannot satisfy the requirement of generating 500 stratified samples even after generating 10,000 synthetic samples. The results of these inapplicable or failed generators are computed with the mean results of other methods. We bold the highest result for each dataset of different sample sizes. TVAE learns the joint distribution $p({\mathbf{x}},y)$ and fails to maintain the original training label distribution. TabEBM achieves the best overall performance against Baseline and benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF GOGGLE TabPFGen TabEBM
At most 10 classes protein 20 N/A 16.36±${}_{\pm\text{3.39}}$ 13.22±${}_{\pm\text{1.42}}$ 13.80±${}_{\pm\text{4.56}}$ 12.93±${}_{\pm\text{3.67}}$ 18.55±${}_{\pm\text{2.64}}$ 12.14±${}_{\pm\text{1.16}}$ 33.46±${}_{\pm\text{5.96}}$ 34.50±${}_{\pm\text{5.86}}$
50 59.60±${}_{\pm\text{3.63}}$ 27.51±${}_{\pm\text{3.72}}$ 18.26±${}_{\pm\text{4.54}}$ 13.28±${}_{\pm\text{3.34}}$ 14.95±${}_{\pm\text{3.52}}$ 29.82±${}_{\pm\text{3.75}}$ 13.30±${}_{\pm\text{1.88}}$ 57.68±${}_{\pm\text{2.79}}$ 59.00±${}_{\pm\text{3.97}}$
100 79.47±${}_{\pm\text{3.48}}$ 41.49±${}_{\pm\text{6.98}}$ 31.50±${}_{\pm\text{3.05}}$ 13.69±${}_{\pm\text{2.75}}$ 17.21±${}_{\pm\text{3.12}}$ 36.73±${}_{\pm\text{5.05}}$ 12.61±${}_{\pm\text{0.81}}$ 78.93±${}_{\pm\text{4.23}}$ 79.02±${}_{\pm\text{3.74}}$
200 90.49±${}_{\pm\text{2.23}}$ 48.30±${}_{\pm\text{7.73}}$ 36.99±${}_{\pm\text{5.39}}$ 13.71±${}_{\pm\text{3.79}}$ 21.67±${}_{\pm\text{5.92}}$ 43.53±${}_{\pm\text{5.42}}$ 12.96±${}_{\pm\text{2.86}}$ 91.47±${}_{\pm\text{1.28}}$ 90.71±${}_{\pm\text{1.30}}$
500 94.28±${}_{\pm\text{1.36}}$ 58.31±${}_{\pm\text{8.05}}$ 48.52±${}_{\pm\text{4.70}}$ 12.91±${}_{\pm\text{2.56}}$ 23.35±${}_{\pm\text{7.76}}$ 57.37±${}_{\pm\text{3.28}}$ 12.51±${}_{\pm\text{3.23}}$ 95.26±${}_{\pm\text{1.37}}$ 95.00±${}_{\pm\text{1.28}}$
fourier 20 N/A -- 13.98±${}_{\pm\text{3.05}}$ 10.64±${}_{\pm\text{2.52}}$ 10.56±${}_{\pm\text{1.86}}$ 16.44±${}_{\pm\text{4.20}}$ 10.56±${}_{\pm\text{3.20}}$ 29.42±${}_{\pm\text{6.72}}$ 36.56±${}_{\pm\text{4.95}}$
50 52.52±${}_{\pm\text{3.71}}$- N/A 10.24±${}_{\pm\text{1.53}}$ 10.72±${}_{\pm\text{1.53}}$ 27.82±${}_{\pm\text{3.41}}$ 9.90±${}_{\pm\text{0.33}}$ 53.98±${}_{\pm\text{4.48}}$ 53.92±${}_{\pm\text{3.82}}$
100 63.36±${}_{\pm\text{3.56}}$ 36.24±${}_{\pm\text{4.25}}$ 28.44±${}_{\pm\text{4.45}}$ 10.30±${}_{\pm\text{1.94}}$ 11.84±${}_{\pm\text{3.10}}$ 30.76±${}_{\pm\text{2.70}}$ 10.32±${}_{\pm\text{2.04}}$ 65.32±${}_{\pm\text{3.26}}$ 65.58±${}_{\pm\text{3.38}}$
200 69.62±${}_{\pm\text{4.71}}$ 40.36±${}_{\pm\text{3.77}}$ 40.08±${}_{\pm\text{5.17}}$ 8.82±${}_{\pm\text{3.98}}$ 32.98±${}_{\pm\text{6.25}}$ 38.50±${}_{\pm\text{3.82}}$ 10.30±${}_{\pm\text{0.76}}$ 72.08±${}_{\pm\text{2.45}}$ 72.26±${}_{\pm\text{2.70}}$
500 75.14±${}_{\pm\text{1.83}}$ 46.56±${}_{\pm\text{5.61}}$ 53.38±${}_{\pm\text{4.52}}$ 9.52±${}_{\pm\text{1.46}}$ 32.46±${}_{\pm\text{9.02}}$ 49.62±${}_{\pm\text{2.60}}$ 10.68±${}_{\pm\text{1.36}}$ 75.98±${}_{\pm\text{1.25}}$ 75.60±${}_{\pm\text{0.85}}$
biodeg 20 69.01±${}_{\pm\text{4.56}}$ 65.35±${}_{\pm\text{7.46}}$ 52.29±${}_{\pm\text{8.45}}$ 55.71±${}_{\pm\text{6.77}}$ 50.00±${}_{\pm\text{0.00}}$ 57.11±${}_{\pm\text{8.37}}$ 50.35±${}_{\pm\text{1.11}}$ 70.77±${}_{\pm\text{4.90}}$ 71.36±${}_{\pm\text{5.30}}$
50 73.27±${}_{\pm\text{3.31}}$ 68.69±${}_{\pm\text{5.88}}$ 63.31±${}_{\pm\text{6.35}}$ 54.85±${}_{\pm\text{8.00}}$ 50.00±${}_{\pm\text{0.00}}$ 71.23±${}_{\pm\text{5.04}}$ 50.00±${}_{\pm\text{0.00}}$ 75.78±${}_{\pm\text{2.59}}$ 75.56±${}_{\pm\text{3.19}}$
100 76.39±${}_{\pm\text{2.12}}$ 73.06±${}_{\pm\text{6.33}}$ 71.75±${}_{\pm\text{2.75}}$ 55.26±${}_{\pm\text{6.54}}$ 50.00±${}_{\pm\text{0.00}}$ 73.02±${}_{\pm\text{2.42}}$ 50.00±${}_{\pm\text{0.00}}$ 78.22±${}_{\pm\text{1.55}}$ 79.28±${}_{\pm\text{2.16}}$
200 ${}_{\pm\text{0.85}}$ 73.19±${}_{\pm\text{5.28}}$ 72.93±${}_{\pm\text{5.24}}$ 54.94±${}_{\pm\text{8.51}}$ 49.88±${}_{\pm\text{0.38}}$ 76.03±${}_{\pm\text{2.88}}$ 50.00±${}_{\pm\text{0.00}}$ 82.22±${}_{\pm\text{1.72}}$ 82.70±${}_{\pm\text{1.73}}$
500 82.47±${}_{\pm\text{1.02}}$ 77.06±${}_{\pm\text{3.03}}$ 79.18±${}_{\pm\text{2.08}}$ 50.47±${}_{\pm\text{1.03}}$ 49.67±${}_{\pm\text{1.05}}$ 77.79±${}_{\pm\text{1.36}}$ 50.00±${}_{\pm\text{0.00}}$ 83.03±${}_{\pm\text{0.96}}$ 83.67±${}_{\pm\text{0.83}}$
steel 20 56.80±${}_{\pm\text{4.60}}$ 53.88±${}_{\pm\text{5.00}}$ 52.12±${}_{\pm\text{3.28}}$ 50.86±${}_{\pm\text{5.69}}$ 50.00±${}_{\pm\text{0.00}}$ 50.47±${}_{\pm\text{1.75}}$ 50.00±${}_{\pm\text{0.00}}$ 64.92±${}_{\pm\text{5.75}}$ 65.82±${}_{\pm\text{6.29}}$
50 62.45±${}_{\pm\text{4.31}}$ 57.80±${}_{\pm\text{2.19}}$ 57.11±${}_{\pm\text{5.27}}$ 51.84±${}_{\pm\text{3.72}}$ 50.00±${}_{\pm\text{0.00}}$ 56.40±${}_{\pm\text{3.61}}$ 50.00±${}_{\pm\text{0.00}}$ 84.81±${}_{\pm\text{7.86}}$ 86.32±${}_{\pm\text{6.58}}$
100 71.25±${}_{\pm\text{5.06}}$ 61.44±${}_{\pm\text{3.38}}$ 60.45±${}_{\pm\text{4.62}}$ 50.59±${}_{\pm\text{3.10}}$ 48.48±${}_{\pm\text{3.17}}$ 55.74±${}_{\pm\text{4.69}}$ 50.66±${}_{\pm\text{2.08}}$ 97.29±${}_{\pm\text{1.40}}$ 97.85±${}_{\pm\text{1.38}}$
200 77.96±${}_{\pm\text{3.06}}$ 63.92±${}_{\pm\text{3.33}}$ 64.22±${}_{\pm\text{5.27}}$ 50.00±${}_{\pm\text{0.00}}$ 47.65±${}_{\pm\text{4.95}}$ 62.33±${}_{\pm\text{4.50}}$ 50.00±${}_{\pm\text{0.01}}$ 98.58±${}_{\pm\text{0.72}}$ 98.89±${}_{\pm\text{0.69}}$
500 85.29±${}_{\pm\text{4.02}}$ 69.41±${}_{\pm\text{5.33}}$ 73.34±${}_{\pm\text{5.43}}$ 50.09±${}_{\pm\text{0.27}}$ 50.00±${}_{\pm\text{0.00}}$ 68.78±${}_{\pm\text{5.84}}$ 50.00±${}_{\pm\text{0.00}}$ 99.71±${}_{\pm\text{0.30}}$ 99.71±${}_{\pm\text{0.27}}$
stock 20 83.79±${}_{\pm\text{3.01}}$ 70.41±${}_{\pm\text{7.99}}$ 49.52±${}_{\pm\text{16.46}}$ 70.75±${}_{\pm\text{10.65}}$ 78.58±${}_{\pm\text{11.59}}$ 66.66±${}_{\pm\text{8.26}}$ 51.45±${}_{\pm\text{4.91}}$ 82.98±${}_{\pm\text{4.47}}$ 83.79±${}_{\pm\text{4.96}}$
50 89.91±${}_{\pm\text{2.44}}$ 77.18±${}_{\pm\text{3.13}}$ 70.04±${}_{\pm\text{2.32}}$ 72.42±${}_{\pm\text{8.52}}$ 89.35±${}_{\pm\text{1.75}}$ 78.06±${}_{\pm\text{3.32}}$ 50.00±${}_{\pm\text{0.00}}$ 89.97±${}_{\pm\text{2.12}}$ 90.17±${}_{\pm\text{1.73}}$
100 92.05±${}_{\pm\text{1.34}}$ 81.60±${}_{\pm\text{3.66}}$ 78.73±${}_{\pm\text{3.67}}$ 75.24±${}_{\pm\text{3.91}}$ 91.32±${}_{\pm\text{1.47}}$ 85.24±${}_{\pm\text{2.39}}$ 49.83±${}_{\pm\text{0.54}}$ 92.25±${}_{\pm\text{1.21}}$ 92.46±${}_{\pm\text{1.18}}$
200 93.50±${}_{\pm\text{0.91}}$ 85.46±${}_{\pm\text{3.04}}$ 83.47±${}_{\pm\text{3.02}}$ 75.51±${}_{\pm\text{2.44}}$ 92.95±${}_{\pm\text{1.04}}$ 87.25±${}_{\pm\text{2.03}}$ 50.00±${}_{\pm\text{0.00}}$ 93.93±${}_{\pm\text{0.85}}$ 93.65±${}_{\pm\text{1.01}}$
Average rank 2.77±${}_{\pm\text{0.69}}$ 4.73±${}_{\pm\text{0.87}}$ 5.88±${}_{\pm\text{1.26}}$ 7.42±${}_{\pm\text{1.06}}$ 7.33±${}_{\pm\text{1.72}}$ 5.21±${}_{\pm\text{0.83}}$ 8.42±${}_{\pm\text{0.52}}$ 1.85±${}_{\pm\text{0.62}}$ 1.40±${}_{\pm\text{0.49}}$
C.7.2 DCR Evaluation
Table 29: DCR between real train data and synthetic data on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std result and average rank across datasets. A higher rank implies better privacy preservation. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean result of other methods. We bold the highest result for each dataset of different sample sizes. Even though ARF and NFLOW show high DCR, our experiments demonstrate that they do not learn the data distribution well, leading to poor downstream accuracy. TabEBM achieves competitive overall DCR against benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF TabPFGen TabEBM
At most 10 classes protein 20 N/A 0.24±${}_{\pm\text{0.06}}$ 0.36±${}_{\pm\text{0.10}}$ 0.29±${}_{\pm\text{0.07}}$ ${}_{\pm\text{0.06}}$ 0.49±${}_{\pm\text{0.03}}$ 0.24±${}_{\pm\text{0.11}}$ 0.39±${}_{\pm\text{0.05}}$
50 0.20±${}_{\pm\text{0.03}}$ 0.34±${}_{\pm\text{0.09}}$ 0.42±${}_{\pm\text{0.06}}$ 0.21±${}_{\pm\text{0.06}}$ 0.62±${}_{\pm\text{0.01}}$ 0.47±${}_{\pm\text{0.08}}$ 0.21±${}_{\pm\text{0.11}}$ 0.37±${}_{\pm\text{0.11}}$
100 0.20±${}_{\pm\text{0.03}}$ 0.33±${}_{\pm\text{0.07}}$ 0.37±${}_{\pm\text{0.05}}$ 0.26±${}_{\pm\text{0.10}}$ 0.54±${}_{\pm\text{0.03}}$ 0.46±${}_{\pm\text{0.06}}$ 0.12±${}_{\pm\text{0.07}}$ 0.27±${}_{\pm\text{0.15}}$
200 0.19±${}_{\pm\text{0.03}}$ 0.31±${}_{\pm\text{0.05}}$ 0.35±${}_{\pm\text{0.04}}$ 0.31±${}_{\pm\text{0.06}}$ 0.51±${}_{\pm\text{0.04}}$ 0.44±${}_{\pm\text{0.05}}$ 0.10±${}_{\pm\text{0.03}}$ 0.26±${}_{\pm\text{0.06}}$
500 0.19±${}_{\pm\text{0.02}}$ 0.32±${}_{\pm\text{0.06}}$ 0.30±${}_{\pm\text{0.05}}$ 0.33±${}_{\pm\text{0.02}}$ 0.48±${}_{\pm\text{0.03}}$ 0.43±${}_{\pm\text{0.05}}$ 0.09±${}_{\pm\text{0.06}}$ 0.23±${}_{\pm\text{0.13}}$
fourier 20 N/A 0.19±${}_{\pm\text{0.17}}$ 0.61±${}_{\pm\text{0.04}}$ 0.52±${}_{\pm\text{0.08}}$ 0.56±${}_{\pm\text{0.05}}$ 0.57±${}_{\pm\text{0.04}}$ 0.48±${}_{\pm\text{0.00}}$ 0.40±${}_{\pm\text{0.00}}$
50 ${}_{\pm\text{0.02}}$ 0.29±${}_{\pm\text{0.26}}$ 0.48±${}_{\pm\text{0.17}}$ 0.33±${}_{\pm\text{0.10}}$ 0.67±${}_{\pm\text{0.05}}$ 0.54±${}_{\pm\text{0.07}}$ 0.31±${}_{\pm\text{0.07}}$ 0.43±${}_{\pm\text{0.03}}$
100 0.23±${}_{\pm\text{0.02}}$ 0.53±${}_{\pm\text{0.06}}$ 0.50±${}_{\pm\text{0.04}}$ 0.37±${}_{\pm\text{0.09}}$ 0.60±${}_{\pm\text{0.08}}$ 0.59±${}_{\pm\text{0.06}}$ 0.31±${}_{\pm\text{0.07}}$ 0.44±${}_{\pm\text{0.04}}$
200 0.22±${}_{\pm\text{0.02}}$ 0.56±${}_{\pm\text{0.06}}$ 0.53±${}_{\pm\text{0.05}}$ 0.37±${}_{\pm\text{0.08}}$ 0.58±${}_{\pm\text{0.02}}$ 0.56±${}_{\pm\text{0.04}}$ ${}_{\pm\text{0.03}}$ 0.46±${}_{\pm\text{0.04}}$
500 0.25±${}_{\pm\text{0.03}}$ 0.67±${}_{\pm\text{0.04}}$ 0.54±${}_{\pm\text{0.06}}$ 0.40±${}_{\pm\text{0.06}}$ 0.62±${}_{\pm\text{0.04}}$ 0.61±${}_{\pm\text{0.05}}$ 0.28±${}_{\pm\text{0.03}}$ 0.45±${}_{\pm\text{0.05}}$
biodeg 20 0.29±${}_{\pm\text{0.05}}$ 0.19±${}_{\pm\text{0.08}}$ 0.26±${}_{\pm\text{0.07}}$ 0.33±${}_{\pm\text{0.08}}$ 0.26±${}_{\pm\text{0.15}}$ 0.46±${}_{\pm\text{0.05}}$ 0.38±${}_{\pm\text{0.03}}$ 0.39±${}_{\pm\text{0.04}}$
50 0.18±${}_{\pm\text{0.05}}$ 0.17±${}_{\pm\text{0.06}}$ 0.16±${}_{\pm\text{0.04}}$ 0.24±${}_{\pm\text{0.04}}$ 0.14±${}_{\pm\text{0.05}}$ 0.31±${}_{\pm\text{0.05}}$ 0.31±${}_{\pm\text{0.07}}$ 0.30±${}_{\pm\text{0.07}}$
100 0.11±${}_{\pm\text{0.04}}$ 0.17±${}_{\pm\text{0.04}}$ 0.17±${}_{\pm\text{0.04}}$ 0.22±${}_{\pm\text{0.06}}$ 0.10±${}_{\pm\text{0.02}}$ 0.21±${}_{\pm\text{0.02}}$ 0.24±${}_{\pm\text{0.07}}$ 0.21±${}_{\pm\text{0.08}}$
200 0.08±${}_{\pm\text{0.02}}$ 0.14±${}_{\pm\text{0.02}}$ 0.14±${}_{\pm\text{0.03}}$ 0.20±${}_{\pm\text{0.04}}$ 0.11±${}_{\pm\text{0.04}}$ 0.20±${}_{\pm\text{0.04}}$ 0.13±${}_{\pm\text{0.08}}$ 0.15±${}_{\pm\text{0.07}}$
500 0.08±${}_{\pm\text{0.03}}$ 0.16±${}_{\pm\text{0.03}}$ 0.13±${}_{\pm\text{0.03}}$ 0.18±${}_{\pm\text{0.05}}$ 0.10±${}_{\pm\text{0.04}}$ 0.18±${}_{\pm\text{0.03}}$ 0.05±${}_{\pm\text{0.03}}$ 0.09±${}_{\pm\text{0.04}}$
steel 20 0.38±${}_{\pm\text{0.09}}$ 0.21±${}_{\pm\text{0.10}}$ 0.25±${}_{\pm\text{0.12}}$ 0.33±${}_{\pm\text{0.07}}$ 0.48±${}_{\pm\text{0.15}}$ 0.43±${}_{\pm\text{0.08}}$ 0.23±${}_{\pm\text{0.05}}$ 0.21±${}_{\pm\text{0.03}}$
50 0.27±${}_{\pm\text{0.11}}$ 0.27±${}_{\pm\text{0.09}}$ 0.22±${}_{\pm\text{0.09}}$ 0.34±${}_{\pm\text{0.05}}$ 0.45±${}_{\pm\text{0.15}}$ 0.40±${}_{\pm\text{0.06}}$ 0.15±${}_{\pm\text{0.06}}$ 0.24±${}_{\pm\text{0.08}}$
100 0.20±${}_{\pm\text{0.11}}$ 0.28±${}_{\pm\text{0.07}}$ 0.22±${}_{\pm\text{0.08}}$ 0.30±${}_{\pm\text{0.08}}$ 0.37±${}_{\pm\text{0.19}}$ 0.41±${}_{\pm\text{0.07}}$ 0.15±${}_{\pm\text{0.09}}$ 0.25±${}_{\pm\text{0.10}}$
200 0.19±${}_{\pm\text{0.08}}$ 0.28±${}_{\pm\text{0.04}}$ 0.22±${}_{\pm\text{0.04}}$ 0.32±${}_{\pm\text{0.06}}$ 0.31±${}_{\pm\text{0.11}}$ ${}_{\pm\text{0.04}}$ 0.14±${}_{\pm\text{0.06}}$ 0.30±${}_{\pm\text{0.09}}$
500 0.17±${}_{\pm\text{0.04}}$ 0.29±${}_{\pm\text{0.07}}$ 0.24±${}_{\pm\text{0.07}}$ 0.32±${}_{\pm\text{0.05}}$ 0.21±${}_{\pm\text{0.07}}$ 0.37±${}_{\pm\text{0.05}}$ 0.10±${}_{\pm\text{0.05}}$ 0.25±${}_{\pm\text{0.09}}$
stock 20 0.24±${}_{\pm\text{0.05}}$ 0.37±${}_{\pm\text{0.08}}$ 0.42±${}_{\pm\text{0.07}}$ 0.46±${}_{\pm\text{0.05}}$ 0.45±${}_{\pm\text{0.12}}$ ${}_{\pm\text{0.05}}$ 0.41±${}_{\pm\text{0.06}}$ 0.46±${}_{\pm\text{0.03}}$
50 0.16±${}_{\pm\text{0.03}}$ 0.41±${}_{\pm\text{0.08}}$ 0.34±${}_{\pm\text{0.04}}$ 0.43±${}_{\pm\text{0.06}}$ 0.28±${}_{\pm\text{0.07}}$ 0.39±${}_{\pm\text{0.03}}$ 0.37±${}_{\pm\text{0.14}}$ 0.46±${}_{\pm\text{0.02}}$
100 0.15±${}_{\pm\text{0.04}}$ 0.39±${}_{\pm\text{0.04}}$ 0.33±${}_{\pm\text{0.05}}$ 0.46±${}_{\pm\text{0.04}}$ 0.17±${}_{\pm\text{0.02}}$ 0.33±${}_{\pm\text{0.03}}$ 0.34±${}_{\pm\text{0.09}}$ 0.44±${}_{\pm\text{0.03}}$
200 0.14±${}_{\pm\text{0.02}}$ 0.38±${}_{\pm\text{0.04}}$ 0.28±${}_{\pm\text{0.03}}$ 0.45±${}_{\pm\text{0.05}}$ 0.11±${}_{\pm\text{0.01}}$ 0.32±${}_{\pm\text{0.03}}$ 0.39±${}_{\pm\text{0.11}}$ 0.46±${}_{\pm\text{0.04}}$
More than 10 classes energy 50 N/A 0.18±${}_{\pm\text{0.20}}$ 0.36±${}_{\pm\text{0.08}}$ 0.48±${}_{\pm\text{0.07}}$ ${}_{\pm\text{0.00}}$ 0.46±${}_{\pm\text{0.06}}$ N/A 0.44±${}_{\pm\text{0.02}}$
100 N/A 0.08±${}_{\pm\text{0.17}}$ 0.40±${}_{\pm\text{0.03}}$ 0.46±${}_{\pm\text{0.05}}$ ${}_{\pm\text{0.00}}$ ${}_{\pm\text{0.04}}$ N/A ${}_{\pm\text{0.04}}$
200 N/A 0.04±${}_{\pm\text{0.11}}$ 0.30±${}_{\pm\text{0.16}}$ 0.45±${}_{\pm\text{0.05}}$ ${}_{\pm\text{0.00}}$ 0.38±${}_{\pm\text{0.03}}$ N/A 0.42±${}_{\pm\text{0.04}}$
collins 100 N/A ${}_{\pm\text{0.00}}$ 0.30±${}_{\pm\text{0.05}}$ 0.33±${}_{\pm\text{0.05}}$ 0.26±${}_{\pm\text{0.05}}$ 0.33±${}_{\pm\text{0.06}}$ N/A 0.38±${}_{\pm\text{0.11}}$
200 0.18±${}_{\pm\text{0.03}}$ ${}_{\pm\text{0.00}}$ 0.23±${}_{\pm\text{0.06}}$ 0.29±${}_{\pm\text{0.08}}$ 0.19±${}_{\pm\text{0.03}}$ 0.33±${}_{\pm\text{0.09}}$ N/A 0.36±${}_{\pm\text{0.12}}$
texture 50 0.21±${}_{\pm\text{0.05}}$ ${}_{\pm\text{0.00}}$ 0.17±${}_{\pm\text{0.19}}$ 0.26±${}_{\pm\text{0.08}}$ 0.26±${}_{\pm\text{0.13}}$ 0.42±${}_{\pm\text{0.07}}$ N/A 0.40±${}_{\pm\text{0.13}}$
100 0.16±${}_{\pm\text{0.04}}$ 0.08±${}_{\pm\text{0.14}}$ 0.35±${}_{\pm\text{0.06}}$ 0.26±${}_{\pm\text{0.03}}$ 0.37±${}_{\pm\text{0.09}}$ 0.46±${}_{\pm\text{0.08}}$ N/A 0.42±${}_{\pm\text{0.10}}$
200 0.13±${}_{\pm\text{0.03}}$ 0.24±${}_{\pm\text{0.18}}$ 0.29±${}_{\pm\text{0.12}}$ 0.32±${}_{\pm\text{0.05}}$ 0.40±${}_{\pm\text{0.08}}$ 0.37±${}_{\pm\text{0.05}}$ N/A 0.42±${}_{\pm\text{0.07}}$
500 0.13±${}_{\pm\text{0.02}}$ 0.04±${}_{\pm\text{0.11}}$ 0.15±${}_{\pm\text{0.11}}$ 0.34±${}_{\pm\text{0.05}}$ 0.37±${}_{\pm\text{0.07}}$ 0.31±${}_{\pm\text{0.03}}$ N/A 0.44±${}_{\pm\text{0.05}}$
Average rank 6.64±${}_{\pm\text{1.35}}$ 5.48±${}_{\pm\text{2.08}}$ 4.82±${}_{\pm\text{1.42}}$ 3.45±${}_{\pm\text{1.75}}$ 4.03±${}_{\pm\text{2.82}}$ 2.21±${}_{\pm\text{1.17}}$ 5.85±${}_{\pm\text{1.85}}$ 3.52±${}_{\pm\text{1.95}}$
C.7.3 Delta-presence Evaluation
Table 30: $\delta$-presence between real train data and synthetic data on eight real-world tabular datasets with varied real data availability. We report the mean $\pm$ std result and average rank across datasets. A higher rank implies better privacy preservation. Note that “N/A” denotes that a specific generator was not applicable, and the rank is computed with the mean result of other methods. We bold the best result for each dataset of different sample sizes. TabEBM achieves the best overall performance against benchmark generators.
Datasets $N_{\text{real}}$ SMOTE TVAE CTGAN NFLOW TabDDPM ARF TabPFGen TabEBM
At most 10 classes protein 20 N/A 0.03±${}_{\pm\text{0.00}}$ 0.33±${}_{\pm\text{0.94}}$ 0.13±${}_{\pm\text{0.23}}$ 0.07±${}_{\pm\text{0.11}}$ 0.05±${}_{\pm\text{0.04}}$ 0.03±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$
50 0.07±${}_{\pm\text{0.00}}$ 0.07±${}_{\pm\text{0.00}}$ 0.07±${}_{\pm\text{0.00}}$ 0.39±${}_{\pm\text{0.92}}$ 0.11±${}_{\pm\text{0.10}}$ 0.09±${}_{\pm\text{0.03}}$ 0.07±${}_{\pm\text{0.00}}$ 0.07±${}_{\pm\text{0.00}}$
100 0.19±${}_{\pm\text{0.02}}$ 0.36±${}_{\pm\text{0.19}}$ 0.50±${}_{\pm\text{0.88}}$ 2.45±${}_{\pm\text{3.23}}$ 0.23±${}_{\pm\text{0.03}}$ 0.36±${}_{\pm\text{0.14}}$ 0.17±${}_{\pm\text{0.01}}$ 0.17±${}_{\pm\text{0.01}}$
200 0.62±${}_{\pm\text{0.23}}$ 2.84±${}_{\pm\text{2.07}}$ 2.03±${}_{\pm\text{1.05}}$ 11.55±${}_{\pm\text{7.74}}$ 1.71±${}_{\pm\text{0.72}}$ 2.73±${}_{\pm\text{1.46}}$ 0.57±${}_{\pm\text{0.17}}$ 0.57±${}_{\pm\text{0.17}}$
500 1.15±${}_{\pm\text{0.12}}$ 4.62±${}_{\pm\text{2.61}}$ 2.41±${}_{\pm\text{1.32}}$ 26.85±${}_{\pm\text{9.98}}$ 6.51±${}_{\pm\text{1.41}}$ 2.26±${}_{\pm\text{0.53}}$ 1.20±${}_{\pm\text{0.32}}$ 1.20±${}_{\pm\text{0.32}}$
fourier 20 N/A 0.02±${}_{\pm\text{0.00}}$ 0.02±${}_{\pm\text{0.00}}$ 0.06±${}_{\pm\text{0.07}}$ 0.04±${}_{\pm\text{0.03}}$ ${}_{\pm\text{0.04}}$ 0.02±${}_{\pm\text{0.00}}$ 0.02±${}_{\pm\text{0.00}}$
50 0.08±${}_{\pm\text{0.00}}$ ${}_{\pm\text{0.01}}$ 0.09±${}_{\pm\text{0.01}}$ 0.09±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$
100 0.18±${}_{\pm\text{0.01}}$ 0.31±${}_{\pm\text{0.10}}$ 0.26±${}_{\pm\text{0.12}}$ 2.74±${}_{\pm\text{2.76}}$ 0.30±${}_{\pm\text{0.07}}$ 0.53±${}_{\pm\text{0.16}}$ 0.17±${}_{\pm\text{0.01}}$ 0.17±${}_{\pm\text{0.01}}$
200 0.73±${}_{\pm\text{0.48}}$ 1.56±${}_{\pm\text{0.63}}$ 3.52±${}_{\pm\text{1.67}}$ 10.60±${}_{\pm\text{4.58}}$ 1.52±${}_{\pm\text{0.67}}$ 2.52±${}_{\pm\text{1.14}}$ 0.45±${}_{\pm\text{0.05}}$ 0.44±${}_{\pm\text{0.05}}$
500 1.42±${}_{\pm\text{0.31}}$ 6.06±${}_{\pm\text{3.50}}$ 5.65±${}_{\pm\text{4.16}}$ 25.63±${}_{\pm\text{12.53}}$ 3.39±${}_{\pm\text{1.53}}$ 2.99±${}_{\pm\text{0.80}}$ 1.18±${}_{\pm\text{0.20}}$ 1.16±${}_{\pm\text{0.18}}$
biodeg 20 0.03±${}_{\pm\text{0.00}}$ 0.09±${}_{\pm\text{0.10}}$ 0.12±${}_{\pm\text{0.15}}$ 0.22±${}_{\pm\text{0.32}}$ 0.03±${}_{\pm\text{0.00}}$ 0.14±${}_{\pm\text{0.30}}$ 0.03±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$
50 0.08±${}_{\pm\text{0.01}}$ 0.10±${}_{\pm\text{0.04}}$ 0.08±${}_{\pm\text{0.01}}$ 0.12±${}_{\pm\text{0.04}}$ 0.10±${}_{\pm\text{0.04}}$ 0.10±${}_{\pm\text{0.04}}$ 0.08±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.00}}$
100 0.35±${}_{\pm\text{0.25}}$ 0.52±${}_{\pm\text{0.29}}$ 0.98±${}_{\pm\text{0.92}}$ 1.14±${}_{\pm\text{0.89}}$ 0.59±${}_{\pm\text{0.41}}$ 0.51±${}_{\pm\text{0.30}}$ 0.24±${}_{\pm\text{0.06}}$ 0.24±${}_{\pm\text{0.06}}$
200 2.35±${}_{\pm\text{1.76}}$ 2.27±${}_{\pm\text{1.01}}$ 2.80±${}_{\pm\text{1.74}}$ 4.18±${}_{\pm\text{1.74}}$ 5.10±${}_{\pm\text{4.44}}$ 1.70±${}_{\pm\text{0.82}}$ 0.74±${}_{\pm\text{0.23}}$ 0.74±${}_{\pm\text{0.23}}$
500 2.91±${}_{\pm\text{1.43}}$ 3.95±${}_{\pm\text{1.08}}$ 3.58±${}_{\pm\text{1.71}}$ 8.85±${}_{\pm\text{3.07}}$ 11.92±${}_{\pm\text{8.19}}$ 2.37±${}_{\pm\text{1.04}}$ 1.57±${}_{\pm\text{0.40}}$ 1.61±${}_{\pm\text{0.42}}$
steel 20 0.04±${}_{\pm\text{0.01}}$ ${}_{\pm\text{0.08}}$ 0.10±${}_{\pm\text{0.12}}$ 0.08±${}_{\pm\text{0.06}}$ 0.06±${}_{\pm\text{0.02}}$ 0.15±${}_{\pm\text{0.30}}$ 0.03±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$
50 0.08±${}_{\pm\text{0.01}}$ 0.12±${}_{\pm\text{0.03}}$ 0.09±${}_{\pm\text{0.02}}$ 0.17±${}_{\pm\text{0.08}}$ 1.04±${}_{\pm\text{0.98}}$ 0.09±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.01}}$
100 0.22±${}_{\pm\text{0.03}}$ 0.39±${}_{\pm\text{0.12}}$ 0.44±${}_{\pm\text{0.40}}$ 0.73±${}_{\pm\text{0.30}}$ 2.17±${}_{\pm\text{2.47}}$ 0.27±${}_{\pm\text{0.06}}$ 0.20±${}_{\pm\text{0.03}}$ 0.20±${}_{\pm\text{0.03}}$
200 1.39±${}_{\pm\text{0.78}}$ 2.05±${}_{\pm\text{0.99}}$ 3.39±${}_{\pm\text{2.50}}$ 5.55±${}_{\pm\text{3.58}}$ 6.68±${}_{\pm\text{4.55}}$ 1.29±${}_{\pm\text{0.69}}$ 0.49±${}_{\pm\text{0.07}}$ 0.48±${}_{\pm\text{0.06}}$
500 1.93±${}_{\pm\text{0.60}}$ 4.39±${}_{\pm\text{3.30}}$ 5.23±${}_{\pm\text{3.90}}$ 10.52±${}_{\pm\text{7.22}}$ 33.92±${}_{\pm\text{25.31}}$ 2.39±${}_{\pm\text{0.88}}$ 1.68±${}_{\pm\text{0.88}}$ 1.70±${}_{\pm\text{0.89}}$
stock 20 0.03±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$ 0.04±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$ 0.03±${}_{\pm\text{0.00}}$
50 0.08±${}_{\pm\text{0.00}}$ 0.09±${}_{\pm\text{0.00}}$ 0.09±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.09±${}_{\pm\text{0.01}}$ 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$
100 0.19±${}_{\pm\text{0.02}}$ 0.23±${}_{\pm\text{0.04}}$ 0.26±${}_{\pm\text{0.12}}$ 0.23±${}_{\pm\text{0.04}}$ 0.22±${}_{\pm\text{0.04}}$ ${}_{\pm\text{0.02}}$ 0.18±${}_{\pm\text{0.01}}$ 0.18±${}_{\pm\text{0.01}}$
200 0.51±${}_{\pm\text{0.08}}$ 0.97±${}_{\pm\text{0.30}}$ 1.83±${}_{\pm\text{1.14}}$ 0.90±${}_{\pm\text{0.32}}$ 0.57±${}_{\pm\text{0.07}}$ 0.60±${}_{\pm\text{0.16}}$ 0.49±${}_{\pm\text{0.07}}$ 0.48±${}_{\pm\text{0.08}}$
More than 10 classes energy 50 N/A 0.09±${}_{\pm\text{0.02}}$ 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.06±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ N/A 0.08±${}_{\pm\text{0.00}}$
100 N/A 1.56±${}_{\pm\text{1.51}}$ 0.19±${}_{\pm\text{0.03}}$ 0.19±${}_{\pm\text{0.03}}$ 0.99±${}_{\pm\text{2.46}}$ 0.16±${}_{\pm\text{0.01}}$ N/A 0.16±${}_{\pm\text{0.00}}$
200 N/A 4.15±${}_{\pm\text{3.04}}$ 1.67±${}_{\pm\text{0.84}}$ 0.72±${}_{\pm\text{0.16}}$ 13.28±${}_{\pm\text{9.21}}$ 0.44±${}_{\pm\text{0.04}}$ N/A 0.38±${}_{\pm\text{0.03}}$
collins 100 N/A 1.79±${}_{\pm\text{1.23}}$ 0.18±${}_{\pm\text{0.02}}$ 0.18±${}_{\pm\text{0.03}}$ 0.17±${}_{\pm\text{0.02}}$ 0.17±${}_{\pm\text{0.01}}$ N/A 0.16±${}_{\pm\text{0.01}}$
200 0.43±${}_{\pm\text{0.08}}$ 4.20±${}_{\pm\text{2.55}}$ 0.76±${}_{\pm\text{0.26}}$ 1.73±${}_{\pm\text{1.09}}$ 0.94±${}_{\pm\text{0.59}}$ 0.95±${}_{\pm\text{0.47}}$ N/A 0.39±${}_{\pm\text{0.05}}$
texture 50 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.08±${}_{\pm\text{0.00}}$ 0.10±${}_{\pm\text{0.02}}$ 0.08±${}_{\pm\text{0.00}}$ N/A 0.08±${}_{\pm\text{0.00}}$
100 0.17±${}_{\pm\text{0.01}}$ 0.42±${}_{\pm\text{0.22}}$ 0.19±${}_{\pm\text{0.02}}$ 0.41±${}_{\pm\text{0.16}}$ 0.44±${}_{\pm\text{0.32}}$ 0.25±${}_{\pm\text{0.05}}$ N/A 0.17±${}_{\pm\text{0.01}}$
200 0.57±${}_{\pm\text{0.18}}$ 2.34±${}_{\pm\text{1.14}}$ 1.76±${}_{\pm\text{1.05}}$ 5.78±${}_{\pm\text{2.82}}$ 7.02±${}_{\pm\text{4.14}}$ 1.54±${}_{\pm\text{0.94}}$ N/A 0.45±${}_{\pm\text{0.08}}$
500 1.33±${}_{\pm\text{0.61}}$ 2.88±${}_{\pm\text{1.09}}$ 2.42±${}_{\pm\text{1.35}}$ 23.40±${}_{\pm\text{10.74}}$ 16.64±${}_{\pm\text{12.35}}$ 2.36±${}_{\pm\text{1.06}}$ N/A 1.03±${}_{\pm\text{0.14}}$
Average rank 3.30±${}_{\pm\text{1.37}}$ 5.91±${}_{\pm\text{1.68}}$ 5.30±${}_{\pm\text{1.74}}$ 6.45±${}_{\pm\text{1.86}}$ 5.82±${}_{\pm\text{2.11}}$ 4.45±${}_{\pm\text{1.66}}$ 3.00±${}_{\pm\text{1.90}}$ 1.76±${}_{\pm\text{1.05}}$

### v2

@@
 When used for data augmentation, our synthetic data consistently improves the classification performance across diverse datasets of various sizes, especially small ones.
+Code is available at https://github.com/andreimargeloiu/TabEBM.
 1 Introduction
 Figure 1: Evaluation of TabEBM and other state-of-the-art tabular generative methods across six key metrics (larger area indicates better performance).
-The results demonstrate that TabEBM excels in data augmentation (utility), leading to downstream classification improvements by a larger margin than all other methods.
-In scientific fields such as medicine, physics, and chemistry, collecting tabular data is often challenging due to the experimental nature of data acquisition [5, 50, 4, 32, 77, 12].
-Due to the small size of such datasets [5, 46], training machine learning models that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], often suffer from poor performance [75, 52, 37].
-In the case of vision and language tasks, which focus on image and text data, a standard remedy to address model performance due to data scarcity is employing data augmentation techniques [73, 74, 60, 72] that generate additional synthetic samples from existing data.
-This increases the diversity of the training set and often leads to improved model generalisation [73].
-For instance, techniques such as rotating, flipping, or altering the colour of images [73], or generating paraphrased sentences for text [74], have proven effective for many downstream tasks.
-However, applying data augmentation to tabular data remains understudied, as tabular datasets are very diverse and lack explicit symmetries [8], such as rotations or translations seen in images.
-Additionally, feature correlations in tabular data are weaker compared to the strong spatial or semantic relationships in image or text data [8].
+The results demonstrate that TabEBM excels in data augmentation (utility), with a larger area than all other methods.
+Many scientific domains within medicine, physics, and chemistry often rely on intricate and challenging data acquisition procedures [5, 50, 4, 32, 76, 12], that typically render small-size tabular datasets [5, 46].
+Using these to train machine learning models, that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], can lead to poor performance [74, 52, 37].
+In the case of learning tasks which leverage image and text data, a standard remedy to address performance issues due to data scarcity is employing data augmentation techniques [72, 73, 60, 71] that generate additional synthetic samples from existing data.
+However, applying data augmentation to tabular data introduces additional challenges, as tabular datasets are often very diverse and lack explicit symmetries [8].
 Consequently, existing tabular data augmentation methods often yield mixed results and can even degrade model performance [51, 71, 48], hindering their widespread adoption.
 Tabular augmentation typically involves training generative models to approximate either the joint distribution $p({\mathbf{x}},y)$ [85, 24] or the class-conditional distribution $p({\mathbf{x}}|y)$ [85, 42, 83, 47, 48].
-A key challenge of joint distribution methods is maintaining the original training label distribution, as sampling from such generators can produce label distributions that deviate from the original and even fail to generate data for specific classes (see Appendix B for an example).
+A key challenge of joint distribution methods is maintaining the original training label distribution, as sampling from such generators can produce label distributions that deviate from the original and even fail to generate data for specific classes (see Appendix C for an example).
 These issues compromise the effectiveness of data augmentation [51] by undermining the label accuracy and distribution.
@@
 Our analysis compares TabEBM with eight leading tabular generative models across various datasets, demonstrating that TabEBM consistently improves data augmentation performance on small datasets, while our generated data demonstrates better statistical fidelity and privacy-preserving properties (Figure 1).
-  • Library: We release TabEBM as an open-source library (included as a zip file with this submission and to be made publicly available after publication, details in Section A.5).
+  • Library: We release TabEBM as an open-source library, available at https://github.com/andreimargeloiu/TabEBM.
 Our library enables off-the-shelf data generation and data augmentation on any tabular dataset without requiring training.
+Further details are available in Section B.5.
 2 TabEBM
@@
 For example, in $R^{3}$, a negative sample might have coordinates $[\alpha^{\text{neg}}_{\text{dist}}\sigma_{1},\alpha^{\text{neg}}_{\text{dist}}\sigma_{2},-\alpha^{\text{neg}}_{\text{dist}}\sigma_{3}]$.
-Placing the negative samples at the corners of a hypercube ensures they are easily distinguishable from the real data, which is crucial for an accurate energy function (see Section C.1.1).
-This placement is also robust to variations in the number and distance of the negative samples (see Sections C.1.2 and C.1.3).
+Placing the negative samples at the corners of a hypercube ensures they are easily distinguishable from the real data, which is crucial for an accurate energy function (see Section D.1.1).
+This placement is also robust to variations in the number and distance of the negative samples (see Sections D.1.2 and D.1.3).
 We create the combined dataset $\mathcal{D}_{c}$ for the surrogate binary classification task by labelling $\mathcal{X}_{c}$ as 1 and $\mathcal{X}_{c}^{\text{neg}}$ as 0:
@@
 In practice, the step size and the noise standard deviations are often chosen separately, resulting in a biased sampler that allows for faster training.
-Section C.2 further shows that TabEBM is stable to hyperparameters for the sampling process.
+Section D.2 further shows that TabEBM is stable to hyperparameters for the sampling process.
 In our method, SGLD performs iterative augmentation.
@@
 Output: Set of synthetic samples for class $c$
-1:Initialise the surrogate binary classification task and train the model
+1:Initialise a surrogate binary classification task and train the model
 2:Assign new labels to the samples $\mathcal{X}_{c}$ from class $c$, setting them to class 1
@@
 Datasets.
-We utilise 14 open-source tabular datasets–eight from OpenML [7] and six from UCI [22] – across five domains: Medicine, Chemistry, Engineering, Language and Economics.
+We utilise eight open-source datasets from OpenML [7] – across five domains: Medicine, Chemistry, Engineering, Language and Economics.
+We note that TabPFN almost drains the small-size OpenML datasets for its meta-validation [33], which might lead to data leakage when evaluating TabEBM.
+To provide fair comparisons, we further select six leakage-free datasets from UCI [22].
 These diverse datasets contain 7 to 77 features and 698 to 5500 samples across 2 to 26 classes.
 Five datasets contain both numerical and categorical features, while the remaining are numerical only.
-To provide comprehensive and fine-grained conclusions on model performance, we further enlarge the evaluation scope by varying the degrees of data availability (i.e., $N_{\text{real}}$), leading up to 33 different test cases for the eight OpenML datasets.
-Section A.1 provide detailed dataset descriptions.
+We further enlarge the evaluation scope by varying the degrees of data availability (i.e., $N_{\text{real}}$), leading up to 33 different test cases for the eight OpenML datasets.
+Section B.1 provide detailed descriptions.
 Benchmark generators.
@@
 Furthermore, we also include a “Baseline” model, where no data augmentation is applied (i.e., only real data is used to train downstream predictors).
-In Section A.6, we detail the settings used for TabEBM and all other generators.
-We have attached the implementation code for TabEBM and will make it publicly available post-publication.
+In Section B.6, we detail the settings used for TabEBM and all other generators.
 Downstream predictors.
@@
 We split each subset into stratified training and validation sets with a ratio of 4:1.
-We provide detailed descriptions of data splitting in Section A.2 and preprocessing in Section A.3.
+We provide detailed descriptions of data splitting in Section B.2 and preprocessing in Section B.3.
 We repeat the splitting ten times, summing up to 10 runs per subset size.
@@
 When aggregating results across datasets, we use the average distance to the minimum (ADTM) metric via affine renormalisation between the top-performing and worse-performing models [30, 54].
-We provide the evaluation results averaged over six downstream predictors for a general conclusion, and the fine-grained numerical results for each predictor are in Appendix C.
+We provide the evaluation results averaged over six downstream predictors for a general conclusion, and the fine-grained numerical results for each predictor are in Appendix D.
 Data augmentation setup.
@@
 Therefore, we perform data augmentation with a large synthetic set ($N_{\text{syn}}=500$) across all splits, and the synthetic data has the same class distribution as the real training data.
-We provide an illustrative figure of the data splitting setup in Section A.2.
+We provide an illustrative figure of the data splitting setup in Section B.2.
 Table 1: Classification accuracy (%) aggregated over six downstream predictors, comparing data augmentation on eight real-world tabular datasets with varied real data availability.
@@
 A key observation is that most modern benchmark generators even underperform the Baseline, indicating poor approximated distributions in the low-sample-size regime.
-Moreover, TabEBM achieves the largest overall performance improvement on six additional UCI datasets, further supporting its effectiveness (see Section C.5.2 for detailed results).
+Moreover, TabEBM achieves the largest overall performance improvement on six leakage-free UCI datasets, further supporting its effectiveness (see Section D.5.2 for details).
 Furthermore, TabEBM is the most widely applicable method among the top three competitive generators on the considered datasets: (i) SMOTE requires at least two samples per class for interpolation, and thus it is not applicable for some datasets, such as the “protein” dataset ($N_{\text{real}}=20$); (ii) TabPFGen cannot scale up to more than ten classes, such as the “collins” dataset.
@@
 We evaluate the similarity between real and synthetic data via (i) average inverse of the Kullback–Leibler Divergence (inverse KL) [17], (ii) p-value of Kolmogorov–Smirnov test (KS test) [39] and (iii) p-value of Chi-squared test ($\chi^{2}$ test) [55].
-The full results, including $\chi^{2}$ test and the similarity between real test data and synthetic data, are in Section C.6.
+The full results, including $\chi^{2}$ test and the similarity between real test data and synthetic data, are in Section D.6.
 For all three metrics, a bigger value denotes that synthetic data is more likely to have the same distribution as real data.
 In Figure 7 (a1&a2), TabEBM consistently achieves the highest accuracy and distribution similarity between real train data and synthetic data, showing that TabEBM learns the distributions of real train data better than benchmark generators.
-In Section C.6, we further show that TabEBM remains the most competitive method in similarity between real test data and synthetic data.
+In Section D.6, we further show that TabEBM remains the most competitive method in similarity between real test data and synthetic data.
 This indicates that TabEBM can extrapolate beyond real train data and thus generate synthetic data from a more general distribution that aligns with both train and test data.
@@
 Specifically, we evaluate synthetic data via three metrics: (i) balanced accuracy of downstream predictors trained with only synthetic data (i.e., train-on-synthetic, test-on-real [85, 42, 87]); (ii) median Distance to Closest Record (DCR) [88], where a greater DCR denotes synthetic data is less likely to be copied from real data; and (iii) ${\delta}$-presense [62], where a smaller value denotes a lower re-identification risk for real data from synthetic data.
-Full numerical results are in Section C.7.
+Full numerical results are in Section D.7.
 Figure 7 (b1&b2) shows that TabEBM consistently finds a better trade-off between accuracy and privacy preservation.
@@
 Note that “accuracy improvement” is for data augmentation, and “accuracy change” is for data sharing.
-The figures with with standard deviations are in Section C.4.
+The figures with with standard deviations are in Section D.4.
 TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
@@
 Since SGLD sampling performs gradient ascent on the density, the TabEBM-generated samples will be close to the real data.
-These findings are consistent across datasets (see Section C.3), where TabPFN’s logits remain positive, with similar ranges and a relatively constant sum as distance increases, warranting further investigation.
+These findings are consistent across datasets (see Section D.3), where TabPFN’s logits remain positive, with similar ranges and a relatively constant sum as distance increases, warranting further investigation.
 Overall, TabPFN’s distance-based uncertainty is useful for inferring accurate energy functions within our TabEBM framework.
@@
 For instance, TableGAN employs a convolutional neural network to optimise the label quality [63], and TVAE is introduced in [85] as a variant of VAE for tabular data.
-However, these methods learn the joint distribution and thus cannot preserve the stratification of the original data (Appendix B).
+However, these methods learn the joint distribution and thus cannot preserve the stratification of the original data (Appendix C).
 CTGAN [85] refines the generation to be class-conditional.
@@
 For instance, another promising generative approach uses pre-trained models like Prior-Data Fitted Networks (PFNs), and the recent TabPFGen [48] adapts such models into one shared class-conditional generator.
-However, TabPFGen’s shared generator can lead to inaccurate density estimates, particularly in high-noise and class-imbalance situations (see examples in Appendix B).
+However, TabPFGen’s shared generator can lead to inaccurate density estimates, particularly in high-noise and class-imbalance situations (see examples in Appendix C).
 As noise increases, TabPFGen’s inferred densities fluctuate significantly and diverge from the true data distributions.
-In contrast, TabEBM uses class-specific EBMs to model each class’s marginal distributions, and the results in Appendix B reveal that our design choice reduces the impact of noise and data imbalance.
+In contrast, TabEBM uses class-specific EBMs to model each class’s marginal distributions, and the results in Appendix C reveal that our design choice reduces the impact of noise and data imbalance.
 TabEBM focuses on approximating and generating for one class at a time, remaining unaffected by noise from other classes.
@@
 TabEBM can handle datasets with over 1000 samples, overcoming TabPFN’s limitation, as it processes one class at a time.
-In Section C.5.3, we show that TabEBM outperforms other generators on larger datasets, though the performance gains decrease as the sample size increases.
-Although TabPFN has a 10-class limit, TabEBM can handle an unlimited number of classes by training TabPFN on binary surrogate tasks instead of the original multi-class problem.
-Our results also show that TabEBM can handle categorical data by encoding categorical features with leave-one-out target statistics.
-We stress that TabEBM is compatible with any classifier that can be adapted into EBMs, as described in Section 2.
+In Section D.5.3, we show that TabEBM outperforms other generators on larger datasets, though the performance gains decrease as the sample size increases.
+Although we implement TabEBM with TabPFN in this paper, we stress that TabEBM is compatible with any classifier that can be adapted into EBMs, as described in Section 2.
 As foundational models for tabular data evolve [81], new models capable of handling more features and samples are expected.
 Integrating them into TabEBM will enhance its ability to manage high-dimensional datasets, increasing its versatility and utility.
+In addition, we would like to point out the imperfection of widely adopted univariate statistical fidelity metrics (i.e., Inverse KL, KS test and $\chi^{2}$ test).
+However, evaluating generators’ ability to capture the joining feature relationships remains an open research question [78].
+We leave this for future work to explore.
 5 Conclusion
@@
 We release our method as an open-source library, allowing users to generate data immediately without additional training.
-Broader Impact Statement
+Acknowledgments and Disclosure of Funding
+The authors would like to thank Francisco Vargas, Randall Balestriero, and Otilia Stretcu for their insightful discussions and valuable input early in the project.
+NS acknowledges the support of the U.S.
+Army Medical Research and Development Command of the Department of Defense; through the FY22 Breast Cancer Research Program of the Congressionally Directed Medical Research Programs, Clinical Research Extension Award GRANT13769713.
+Opinions, interpretations, conclusions, and recommendations are those of the authors and are not necessarily endorsed by the Department of Defense.
+Appendix: TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models
+[…]
+[…]
+Appendix A Broader Impact Statement
 This paper introduces a novel data augmentation approach, TabEBM, that aims to advance the field of machine learning by addressing challenges in the low-sample-size regime.
@@
 Therefore, this risk, together with other data privacy risks during downstream deployment, must be carefully managed.
-Appendix: TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models
-[…]
-[…]
-Appendix A Reproducibility
-A.1 Datasets
+Appendix B Reproducibility
+B.1 Datasets
 All eight datasets are publicly available on OpenML [7], and their details are listed in Table 3.
@@
 texture 40499 5,500 40 11 137.5 500 500
-A.2 Data Splitting
+B.2 Data Splitting
 Figure 9 shows the data splitting setup used across all datasets.
@@
 Figure 9: Data splitting strategies for data augmentation for all datasets.
-A.3 Data Preprocessing
+B.3 Data Preprocessing
 Following the procedures presented in prior work [54, 30], we perform preprocessing in two steps.
@@
 Finally, we apply the same transformation to the validation and test data before conducting evaluations.
-A.4 Software and Computing Resources
+B.4 Software and Computing Resources
 Software implementation.
@@
 This ensured that TabEBM and all benchmark models were trained and evaluated on the same set of samples.
-We also attach our code to this submission and will release it under the MIT licence upon publication.
-The experimental environment settings, including library dependencies, are specified in the associated code for reference and reproduction purposes.
+The experimental environment settings, including library dependencies, are specified in the open-source library for reference and reproduction purposes.
 Computing Resources.
@@
 The operating system was Ubuntu 20.4.4 LTS.
-A.5 TabEBM open-source library
-We implemented TabEBM as an extensible library, included as a zip file in the Supplementary material.
-The library will be made publicly available after publication.
+B.5 TabEBM open-source library
+We implemented TabEBM as an extensible library, and the code is available on https://github.com/andreimargeloiu/TabEBM.
 For practitioners, it offers an easy-to-use, domain-agnostic tool that requires no training, making it particularly suitable for data augmentation, especially in small datasets.
@@
     Generate synthetic data: The library can generate data for augmentation.
-    from TabEBM import TabEBM
+    from tabebm.TabEBM import TabEBM
     tabebm = TabEBM()
@@
     Compute and visualise the energy function: The library allows computation of TabEBM’s energy function and the unnormalised data density.
-The demo notebook, TabEBM_approximated_density.ipynb, shows the TabEBM-inferred densities under conditions of data noise and class imbalance (thus recreating the plots from Appendix B).
-A.6 Implementation of Generators
+The demo notebook, TabEBM_approximated_density.ipynb, shows the TabEBM-inferred densities under conditions of data noise and class imbalance (thus recreating the plots from Appendix C).
+B.6 Implementation of Generators
 TabEBM.
@@
 We use four surrogate negative samples, $\mathcal{X}_{c}^{\text{neg}}$, positioned at $\alpha^{\text{neg}}_{\text{dist}}=5$ standard deviations from zero, in random corners of a hypercube in ${\mathbb{R}}^{D}$ (as explained in Section 2.2), distant from any real data.
-In Section C.1, we show that TabEBM is robust to the distribution of the negative samples.
+In Section D.1, we show that TabEBM is robust to the distribution of the negative samples.
 We use SGLD [84] for sampling from TabEBM, where the starting points ${\mathbf{x}}_{0}^{\text{synth}}$ are initialised by adding Gaussian noise with zero mean and standard deviation $\sigma_{\text{start}}=0.01$ to a randomly selected sample of the specific class, i.e., ${\mathbf{x}}_{0}^{\text{synth}}\sim\mathcal{N}(\mathcal{X}_{c},\sigma_{\text{start}}^{2}\mathbf{I})$.
 For SGLD, we used the following parameters: step size $\alpha_{\text{step}}=0.1$, noise scale $\alpha_{\text{noise}}=0.01$ and number of steps $T=200$.
-We found TabEBM to be robust to the SGLD settings (see Section C.2).
-We attach the implementation code for TabEBM and will release it as a public open-source library available after publication (see Section A.5 for details on our library).
+We found TabEBM to be robust to the SGLD settings (see Section D.2).
 TabPFGen.
@@
 For the other six benchmark generators, we use their open-source implementations in Synthcity [67]. Following prior studies [87, 80, 71, 48], we use the default settings for all generators.
-A.7 Implementation of Downstream Predictors
+B.7 Implementation of Downstream Predictors
 We implemented TabPFN with its official implementation [33] and the other five downstream predictors with the scikit-learn library [65].
 Following prior studies [80, 71], we use the default settings for all downstream predictors.
-Appendix B Limitations of Existing Generative Methods
+Appendix C Limitations of Existing Generative Methods
 We showcase three limitations of current generative models: (1) Figure 10 shows that models approximating the joint distribution $p({\mathbf{x}},y)$ may fail to preserve the stratification of the real data and even fail to generate samples from specific classes.
@@
 This results in very robust inferred conditional distributions even under heavy class imbalance (e.g., see that $p({\mathbf{x}}|y=1)$ remains relatively constant).
-Appendix C Extended Experimental Results
-C.1 Ablations on the distribution of the surrogate negative samples
-C.1.1 Ablations on placing the negative samples
+Appendix D Extended Experimental Results
+D.1 Ablations on the distribution of the surrogate negative samples
+D.1.1 Ablations on placing the negative samples
 Figure 13: TabEBM energy $E_{c}(x)$ for different choices of negative samples.
@@
 This occurs because TabPFN is uncertain when points of different classes are close, affecting its logits magnitude and making them unsuitable for density estimation.
-C.1.2 Varying the number of negative samples
+D.1.2 Varying the number of negative samples
 We evaluate the impact of the ratio $|\mathcal{X}^{\text{neg}}_{c}|:|\mathcal{X}_{c}|$ between the negative samples $\mathcal{X}^{\text{neg}}_{c}$ and the real samples $|\mathcal{X}_{c}|$.
@@
 Average accuracy 64.99 64.95 64.94 64.98 64.97 62.07
-C.1.3 Varying the distance of the negative samples
+D.1.3 Varying the distance of the negative samples
 We assess the effect of varying the distance of negative samples.
@@
 Average accuracy 65.25 65.05 65.17 65.07 64.92 64.97 62.07
-C.2 Ablations on the sensitivity to the hyperparameters of SGLD sampling
+D.2 Ablations on the sensitivity to the hyperparameters of SGLD sampling
 We vary two key hyperparameters of SGLD on the “biodeg” binary dataset with $N_{\text{real}}=100$: the step size $\alpha_{\text{step}}$ and the noise scale $\alpha_{\text{noise}}$.
@@
 0.05 75.93 75.89 75.94 75.70
-C.3 Distribution of Logits and Unnormalized Density in TabEBM
+D.3 Distribution of Logits and Unnormalized Density in TabEBM
 Figure 14: Additional results for Section 3.4.
@@
 Because sampling using SGLD perform gradient ascent on the density, the TabEBM-generated samples will be similar when using one or both logits.
-C.4 Complete Trade-off Figures with Error Bars
+D.4 Complete Trade-off Figures with Error Bars
 Figure 15: (a1&a2): Median inverse KL and KS test vs.
@@
 TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
-C.5 Results on Data Augmentation
-C.5.1 Results on eight OpenML datasets.
+D.5 Results on Data Augmentation
+D.5.1 Results on eight OpenML datasets.
 Table 7: Classification accuracy (%) of LR, comparing data augmentation on eight real-world tabular datasets with varied real data availability.
@@
 Average rank 3.08±${}_{\pm\text{1.22}}$ 4.23±${}_{\pm\text{2.32}}$ 6.12±${}_{\pm\text{1.57}}$ 6.29±${}_{\pm\text{2.07}}$ 8.42±${}_{\pm\text{1.32}}$ 6.12±${}_{\pm\text{3.38}}$ 6.54±${}_{\pm\text{1.61}}$ 8.83±${}_{\pm\text{1.46}}$ 3.12±${}_{\pm\text{1.80}}$ 2.23±${}_{\pm\text{1.83}}$
-C.5.2 Results on six UCI Datasets
+D.5.2 Results on six UCI Datasets
 Table 13: Details of the six real-world tabular datasets from UCI.
@@
 statlog 56.22±${}_{\pm\text{3.20}}$ 57.30±${}_{\pm\text{2.57}}$ 53.12±${}_{\pm\text{1.52}}$ 55.55±${}_{\pm\text{3.18}}$ 53.07±${}_{\pm\text{2.95}}$ 57.65±${}_{\pm\text{2.01}}$ 57.85±${}_{\pm\text{1.95}}$
-C.5.3 Results on larger sample sizes
+D.5.3 Results on larger sample sizes
 Table 15: Test classification accuracy (%) aggregated over six downstream predictors, comparing data augmentation with increased real data availability of the “texture” dataset.
@@
 4000 98.51±${}_{\pm\text{0.33}}$ 98.59±${}_{\pm\text{0.56}}$ 98.11±${}_{\pm\text{0.43}}$ 98.00±${}_{\pm\text{0.20}}$ 98.46±${}_{\pm\text{0.14}}$ N/A 98.55±${}_{\pm\text{0.58}}$ +0.04
-C.6 Results on Statistical Fidelity
+D.6 Results on Statistical Fidelity
 We aim to provide a fair and coherent comparison between TabEBM and existing methods and thus we follow the widely adopted evaluation process in prior studies.
@@
 We leave this for future work to explore.
-C.6.1 Similarity between Real Train Data and Synthetic Data
+D.6.1 Similarity between Real Train Data and Synthetic Data
 Table 16: Inverse KL between real train data and synthetic data on eight real-world tabular datasets with varied real data availability.
@@
 Average rank 3.52±${}_{\pm\text{1.08}}$ 6.88±${}_{\pm\text{1.02}}$ 6.03±${}_{\pm\text{1.10}}$ 3.55±${}_{\pm\text{1.03}}$ 8.30±${}_{\pm\text{1.02}}$ 2.82±${}_{\pm\text{1.76}}$ 8.18±${}_{\pm\text{0.88}}$ 4.27±${}_{\pm\text{1.62}}$ 1.45±${}_{\pm\text{0.71}}$
-C.6.2 Similarity between Real Test Data and Synthetic Data
+D.6.2 Similarity between Real Test Data and Synthetic Data
 Table 19: Inverse KL between real test data and synthetic data on eight real-world tabular datasets with varied real data availability.
@@
 Average rank 3.06±${}_{\pm\text{1.32}}$ 5.45±${}_{\pm\text{2.35}}$ 6.15±${}_{\pm\text{0.87}}$ 3.64±${}_{\pm\text{1.75}}$ 8.36±${}_{\pm\text{0.90}}$ 3.18±${}_{\pm\text{1.47}}$ 7.76±${}_{\pm\text{1.77}}$ 5.33±${}_{\pm\text{1.71}}$ 2.06±${}_{\pm\text{1.43}}$
-C.7 Results on Privacy Preservation
-C.7.1 Downstream Accuracy in Data Sharing
+D.7 Results on Privacy Preservation
+D.7.1 Downstream Accuracy in Data Sharing
 Table 22: Classification accuracy (%) aggregated over six downstream predictors, comparing data sharing on eight real-world tabular datasets with varied real data availability.
@@
 Average rank 2.77±${}_{\pm\text{0.69}}$ 4.73±${}_{\pm\text{0.87}}$ 5.88±${}_{\pm\text{1.26}}$ 7.42±${}_{\pm\text{1.06}}$ 7.33±${}_{\pm\text{1.72}}$ 5.21±${}_{\pm\text{0.83}}$ 8.42±${}_{\pm\text{0.52}}$ 1.85±${}_{\pm\text{0.62}}$ 1.40±${}_{\pm\text{0.49}}$
-C.7.2 DCR Evaluation
+D.7.2 DCR Evaluation
 Table 29: DCR between real train data and synthetic data on eight real-world tabular datasets with varied real data availability.
@@
 Average rank 6.64±${}_{\pm\text{1.35}}$ 5.48±${}_{\pm\text{2.08}}$ 4.82±${}_{\pm\text{1.42}}$ 3.45±${}_{\pm\text{1.75}}$ 4.03±${}_{\pm\text{2.82}}$ 2.21±${}_{\pm\text{1.17}}$ 5.85±${}_{\pm\text{1.85}}$ 3.52±${}_{\pm\text{1.95}}$
-C.7.3 Delta-presence Evaluation
+D.7.3 Delta-presence Evaluation
 Table 30: $\delta$-presence between real train data and synthetic data on eight real-world tabular datasets with varied real data availability.
@@
 Average rank 3.30±${}_{\pm\text{1.37}}$ 5.91±${}_{\pm\text{1.68}}$ 5.30±${}_{\pm\text{1.74}}$ 6.45±${}_{\pm\text{1.86}}$ 5.82±${}_{\pm\text{2.11}}$ 4.45±${}_{\pm\text{1.66}}$ 3.00±${}_{\pm\text{1.90}}$ 1.76±${}_{\pm\text{1.05}}$
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: Please refer to Section 1, where we state our research topic and objective and then connect them to our contributions.
+We also provide an open-source library of the proposed method (https://github.com/andreimargeloiu/TabEBM).
+  5. 2.
+    Limitations
+  6. Question: Does the paper discuss the limitations of the work performed by the authors?
+  7. Answer: [Yes]
+  8. Justification: Please refer to Section 4, where we analyse the limitations of the proposed model and promising directions for future research.
+  9. 3.
+    Theory Assumptions and Proofs
+  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  11. Answer: [Yes]
+  12. Justification: Please refer to Section 2, where we provide full theoretical derivations of our proposed method.
+  13. 4.
+    Experimental Result Reproducibility
+  14.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  15. Answer: [Yes]
+  16.
+Justification: Please refer to Appendix B, where we provide full details on reproducing the results in the paper.
+Furthermore, we have also provided the open-source implementation of our proposed method.
+  17. 5.
+    Open access to data and code
+  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  19. Answer: [Yes]
+  20.
+Justification: Please refer to Appendix B.
+All datasets used in this paper are publicly available, and the implementations of benchmark generators are open-source.
+In addition, we have also provided the implementation of our proposed method.
+  21. 6.
+    Experimental Setting/Details
+  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  23. Answer: [Yes]
+  24. Justification: Please refer to Appendix B, where we provide full descriptions of the experimental setup to reproduce the results in the paper.
+  25. 7.
+    Experiment Statistical Significance
+  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  27. Answer: [Yes]
+  28.
+Justification: Please refer to Section 3, where we provide standard deviations for all tables.
+Figure 6 and Figure 4 (Right) contain error bars.
+Due to the page limit, the error bars for all other figures are available in Appendix D.
+In Section 3.2 and Section D.6, we show statistical significance tests of the similarity between real data and synthetic data.
+  29. 8.
+    Experiments Compute Resources
+  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  31. Answer: [Yes]
+  32. Justification: Please refer to Section B.4, where we provide full details on the computation resources used in the paper.
+  33. 9.
+    Code Of Ethics
+  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  35. Answer: [Yes]
+  36. Justification: We carefully check the NeurIPS Code of Ethics, and we confirm that our work follows the Code in every respect.
+  37. 10.
+    Broader Impacts
+  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  39. Answer: [Yes]
+  40. Justification: Please refer to Appendix A, where we include the societal impacts of our work.
+  41. 11.
+    Safeguards
+  42.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  43. Answer: [N/A]
+  44. Justification: [N/A]
+  45. 12.
+    Licenses for existing assets
+  46.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  47. Answer: [Yes]
+  48. Justification: Please refer to Appendix B, where we provide the open-source licenses followed by the creators or original owners of assets.
+  49. 13.
+    New Assets
+  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  51. Answer: [Yes]
+  52. Justification: We provide the implementation of our method as a python library attached to this submission. We will make it publicly available post-publication.
+  53. 14.
+    Crowdsourcing and Research with Human Subjects
+  54.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  55. Answer: [N/A]
+  56. Justification: [N/A]
+  57. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  58.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  59. Answer: [No]
+  60. Justification:[N/A]

### v3

@@
 Abstract
-Data collection is often difficult in critical fields such as medicine, physics, and chemistry.
-As a result, classification methods usually perform poorly with these small datasets, leading to weak predictive performance.
-Increasing the training set with additional synthetic data, similar to data augmentation in images, is commonly believed to improve downstream classification performance.
+Data collection is often difficult in critical fields such as medicine, physics, and chemistry, yielding typically only small tabular datasets.
+However, classification methods tend to struggle with these small datasets, leading to poor predictive performance.
+Increasing the training set with additional synthetic data, similar to data augmentation in images, is commonly believed to improve downstream tabular classification performance.
 However, current tabular generative methods that learn either the joint distribution $p(\mathbf{x},y)$ or the class-conditional distribution $p(\mathbf{x}\mid y)$ often overfit on small datasets, resulting in poor-quality synthetic data, usually worsening classification performance compared to using real data alone.
 To solve these challenges, we introduce TabEBM, a novel class-conditional generative method using Energy-Based Models (EBMs).
-Unlike existing methods that use a shared model to approximate all class-conditional densities, our key innovation is to create distinct EBM generative models for each class, each modelling its class-specific data distribution individually.
+Unlike existing tabular methods that use a shared model to approximate all class-conditional densities, our key innovation is to create distinct EBM generative models for each class, each modelling its class-specific data distribution individually.
 This approach creates robust energy landscapes, even in ambiguous class distributions.
 Our experiments show that TabEBM generates synthetic data with higher quality and better statistical fidelity than existing methods.
-When used for data augmentation, our synthetic data consistently improves the classification performance across diverse datasets of various sizes, especially small ones.
+When used for data augmentation, our synthetic data consistently leads to improved classification performance across diverse datasets of various sizes, especially small ones.
 Code is available at https://github.com/andreimargeloiu/TabEBM.
@@
 The results demonstrate that TabEBM excels in data augmentation (utility), with a larger area than all other methods.
-Many scientific domains within medicine, physics, and chemistry often rely on intricate and challenging data acquisition procedures [5, 50, 4, 32, 76, 12], that typically render small-size tabular datasets [5, 46].
-Using these to train machine learning models, that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], can lead to poor performance [74, 52, 37].
+Many scientific domains within medicine, physics, and chemistry often rely on intricate and challenging data acquisition procedures [5, 50, 4, 32, 76, 12] that typically render small-size tabular datasets [5, 46].
+Using these to train machine learning models that can aid in tasks such as disease diagnosis [52, 37], material property prediction [35], and chemical compound classification [11], can lead to poor performance [74, 52, 37].
 In the case of learning tasks which leverage image and text data, a standard remedy to address performance issues due to data scarcity is employing data augmentation techniques [72, 73, 60, 71] that generate additional synthetic samples from existing data.
-However, applying data augmentation to tabular data introduces additional challenges, as tabular datasets are often very diverse and lack explicit symmetries [8].
+However, applying data augmentation to tabular data introduces additional challenges, as tabular datasets are often very diverse and lack explicit symmetries [8], such as rotations or translations seen in images.
 Consequently, existing tabular data augmentation methods often yield mixed results and can even degrade model performance [51, 71, 48], hindering their widespread adoption.
@@
 These issues compromise the effectiveness of data augmentation [51] by undermining the label accuracy and distribution.
-On the other hand, while class-conditional models learning $p({\mathbf{x}}|y)$ preserve the stratification of the original data, they often employ a shared model to represent all class-conditional densities.
+On the other hand, while class-conditional models that learn $p({\mathbf{x}}|y)$ preserve the stratification of the original data, they often employ a shared model to represent all class-conditional densities.
 This, however, can lead to overfitting, particularly in imbalanced datasets where the model may prioritise more frequent classes [21], ignoring unique features needed for generating label-invariant samples.
 Additionally, in datasets with limited data, this can lead to mode collapse [68, 70], where the model does not effectively capture the diversity of each class [70], and thus tends to perform poorly in a multi-class setting.
-Figure 2: An overview of the proposed TabEBM.
+Figure 2: An overview of TabEBM.
 We learn distinct class-specific Energy-Based Models (EBMs) $E_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}blue}}}({\mathbf{x}})$ and $E_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{red}}}({\mathbf{x}})$ exclusively on the points of their respective class.
@@
 This is achieved by setting the energy function to $E({\mathbf{x}},y)=-f_{\theta}({\mathbf{x}})$.
-Furthermore, the energy function for the marginal distribution $p({\mathbf{x}})$ is obtained by marginalising over $p({\mathbf{x}},y)$, resulting in $E({\mathbf{x}})=-\text{LogSumExp}_{y^{\prime}}f_{\theta}({\mathbf{x}})[y^{\prime}]$.
+Furthermore, the energy function for the marginal distribution $p({\mathbf{x}})$ is obtained by marginalising over $p({\mathbf{x}},y)$, resulting in $E({\mathbf{x}})=-\mathrm{{LogSumExp}}_{y^{\prime}}f_{\theta}({\mathbf{x}})[y^{\prime}]$.
 Such an energy-based model, trained with EBM-specific protocols on multiple classes, is typically used as a classifier, as demonstrated on several computer vision tasks in [29].
@@
 Note that TabPFN is intended for inference only, with no updates to its parameters (see Section 4 for more details about TabPFN).
-In this context, “training” the TabPFN classifier is analogous to the K Nearest Neighbour algorithm, which simply performs inference based on a training dataset provided to the model.
+In this context, “training” the TabPFN classifier is analogous to the K-Nearest Neighbour algorithm, which simply performs inference based on a training dataset provided to the model.
 We apply TabPFN multiple times on separate datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{C}\}$ to obtain multiple classifiers $\{f_{\theta}^{1},f_{\theta}^{2},\ldots,f_{\theta}^{C}\}$.
@@
 Datasets.
-We utilise eight open-source datasets from OpenML [7] – across five domains: Medicine, Chemistry, Engineering, Language and Economics.
-We note that TabPFN almost drains the small-size OpenML datasets for its meta-validation [33], which might lead to data leakage when evaluating TabEBM.
-To provide fair comparisons, we further select six leakage-free datasets from UCI [22].
+We utilise eight open-source tabular datasets from OpenML [7] across five domains: Medicine, Chemistry, Engineering, Language and Economics.
+As TabPFN utilises many small-size OpenML datasets in its meta-validation [33], it can lead to data leakage when evaluating TabEBM.
+Therefore, to provide fair comparisons, we select six additional leakage-free datasets from UCI [22].
 These diverse datasets contain 7 to 77 features and 698 to 5500 samples across 2 to 26 classes.
@@
 We further enlarge the evaluation scope by varying the degrees of data availability (i.e., $N_{\text{real}}$), leading up to 33 different test cases for the eight OpenML datasets.
-Section B.1 provide detailed descriptions.
+Section B.1 provides detailed descriptions.
 Benchmark generators.
@@
 The optimal $N_{\text{syn}}$ remains an open problem for tabular data [51, 71, 31].
-Prior work [47, 48] mainly use synthetic sets with equivalent sizes to the real sets (i.e., $N_{\text{real}}=N_{\text{syn}}$).
-However, we observe that $N_{\text{real}}=N_{\text{syn}}$ can lead to highly unstable results, especially on small datasets we investigate.
+Prior works [47, 48] mainly use synthetic sets with equivalent sizes to the real sets (i.e., $N_{\text{real}}=N_{\text{syn}}$).
+However, we observe that $N_{\text{real}}=N_{\text{syn}}$ can lead to highly unstable results, especially on small datasets that we investigate.
 Recent work has used different $N_{\text{syn}}$ for various generators, such as by applying post-processing [31, 71].
-In this work, we want to provide a head-to-head comparison of the effect of data augmentation across subsampled datasets of varying size $N_{\text{real}}\in\{20,50,100,200,500\}$.
+In this work, we want to provide a head-to-head comparison of the effect of data augmentation across subsampled datasets of varying sizes $N_{\text{real}}\in\{20,50,100,200,500\}$.
 Therefore, we perform data augmentation with a large synthetic set ($N_{\text{syn}}=500$) across all splits, and the synthetic data has the same class distribution as the real training data.
 We provide an illustrative figure of the data splitting setup in Section B.2.
+3.1 Data Augmentation Improvement (Q1)
+We evaluate the effect of using synthetic data for data augmentation by comparing the balanced accuracy of downstream predictors before and after augmentation.
+Typically, higher classification accuracy (i.e., $\text{ACC}_{\text{Generator}})$ and accuracy improvements (i.e., $\text{ACC}_{\text{Generator}}-\text{ACC}_{\text{Baseline}}>0$) demonstrate the effectiveness of the synthetic data for data augmentation.
+TabEBM effectively improves downstream performance across sample sizes, especially for very low-sample-size regimes.
+Table 1 and Figure 4 (Left) show that TabEBM exhibits competitive performance in data augmentation, generally achieving the highest downstream accuracy and average rank across most datasets and sample sizes.
+Notably, TabEBM is the only generator that consistently improves performance across sample sizes.
+A key observation is that most modern benchmark generators underperform even the Baseline, indicating poor approximated distributions in the low-sample-size regime.
+Moreover, TabEBM achieves the largest overall performance improvement on six leakage-free UCI datasets, further supporting its effectiveness (see Section D.5.2 for details).
 Table 1: Classification accuracy (%) aggregated over six downstream predictors, comparing data augmentation on eight real-world tabular datasets with varied real data availability.
@@
 TabEBM generally outperforms benchmark generators across varying sample sizes and number of classes.
-3.1 Data Augmentation Improvement (Q1)
-We evaluate the effect of using synthetic data for data augmentation by comparing the balanced accuracy of downstream predictors before and after augmentation.
-Typically, higher classification accuracy (i.e., $\text{ACC}_{\text{Generator}})$ and accuracy improvements (i.e., $\text{ACC}_{\text{Generator}}-\text{ACC}_{\text{Baseline}}>0$) demonstrate the effectiveness of the synthetic data for data augmentation.
-TabEBM effectively improves downstream performance across sample sizes, especially for very low-sample-size regimes.
-Table 1 and Figure 4 (Left) show that TabEBM exhibits competitive performance in data augmentation, generally achieving the highest downstream accuracy and average rank across most datasets and sample sizes.
-Notably, TabEBM is the only generator that consistently improves performance across sample sizes.
-A key observation is that most modern benchmark generators even underperform the Baseline, indicating poor approximated distributions in the low-sample-size regime.
-Moreover, TabEBM achieves the largest overall performance improvement on six leakage-free UCI datasets, further supporting its effectiveness (see Section D.5.2 for details).
 Furthermore, TabEBM is the most widely applicable method among the top three competitive generators on the considered datasets: (i) SMOTE requires at least two samples per class for interpolation, and thus it is not applicable for some datasets, such as the “protein” dataset ($N_{\text{real}}=20$); (ii) TabPFGen cannot scale up to more than ten classes, such as the “collins” dataset.
 In addition, TabEBM can stabilise downstream performance, especially when real data is very scarce ($N_{\text{real}}=20$): TabEBM leads to smaller standard deviations than Baseline on seven out of eight datasets.
-TabEBM effectively improves downstream performance across the number of classes, especially for more than ten classes.
+TabEBM effectively improves downstream performance across any number of classes, especially for more than ten classes.
 Figure 4 (Right) shows that TabEBM consistently outperforms the Baseline with notable improvements, particularly in datasets with more than ten classes.
 In contrast, an increased number of classes tends to cause a performance degradation in the benchmark generators.
-Figure 5: Mean normalised balanced accuracy improvement (%) on imbalanced datasets. TabEBM consistently outperforms the Baseline and other generators across different levels of data imbalance.
-Figure 6: Median data augmentation time vs. mean normalised balanced accuracy. TabEBM achieves higher downstream accuracy while typically operating 3-30 times faster than most other methods.
 TabEBM is robust on imbalanced datasets.
@@
 The results show that TabEBM is practical, as it achieves higher downstream accuracy with lower time costs.
+Figure 5: Mean normalised balanced accuracy improvement (%) on imbalanced datasets. TabEBM consistently outperforms the Baseline and other generators across different levels of data imbalance.
+Figure 6: Median data augmentation time vs. mean normalised balanced accuracy. TabEBM achieves higher downstream accuracy while typically operating 3-30 times faster than most other methods.
 3.2 Statistical Fidelity (Q2)
-We evaluate the fidelity of synthetic data by focusing on two aspects of synthetic data: similarity to real train data and real test data.
-We evaluate the similarity between real and synthetic data via (i) average inverse of the Kullback–Leibler Divergence (inverse KL) [17], (ii) p-value of Kolmogorov–Smirnov test (KS test) [39] and (iii) p-value of Chi-squared test ($\chi^{2}$ test) [55].
-The full results, including $\chi^{2}$ test and the similarity between real test data and synthetic data, are in Section D.6.
+We evaluate the fidelity of synthetic data by measuring the similarity of synthetic data to real train data and to real test data (Figure 7).
+We evaluate this similarity via (i) average inverse of the Kullback–Leibler Divergence (inverse KL) [17], (ii) p-value of Kolmogorov–Smirnov test (KS test) [39] and (iii) p-value of Chi-squared test ($\chi^{2}$ test) [55].
+For full numerical results, including $\chi^{2}$ test, see Section D.6.
 For all three metrics, a bigger value denotes that synthetic data is more likely to have the same distribution as real data.
-In Figure 7 (a1&a2), TabEBM consistently achieves the highest accuracy and distribution similarity between real train data and synthetic data, showing that TabEBM learns the distributions of real train data better than benchmark generators.
+In Figure 7 (a1&a2), TabEBM consistently exhibits the highest accuracy and distribution similarity between real train data and synthetic data, indicating that TabEBM learns the distributions of real train data better than benchmark generators.
 In Section D.6, we further show that TabEBM remains the most competitive method in similarity between real test data and synthetic data.
@@
 More broadly, data privacy is a critical concern for organisations and governments handling sensitive data [75].
-Privacy-preserving synthetic data also allows researchers and practitioners to bypass ethical and logistical issues while enabling model training and testing [38].
+Privacy-preserving synthetic data allows researchers and practitioners to bypass ethical and logistical issues while enabling model training and testing [38].
 We further explore the use of TabEBM-generated data for data sharing, where only synthetic data is accessible for downstream users [75, 89, 86, 23, 42].
@@
 The relatively high DCR for TabEBM indicates that it can extrapolate beyond real train data, aligning with the finding that TabEBM’s synthetic data is statistically similar to real test data (Section 3.2).
-These results further suggest that TabEBM learns the general distribution of real data, and TabEBM can generate high-quality synthetic data suitable for various purposes, including privacy preservation.
+These results further suggest that TabEBM learns the general distribution of real data, and can generate high-quality synthetic data suitable for various purposes, including privacy preservation.
 Figure 7: (a1&a2): Median inverse KL and KS test vs.
@@
 Note that “accuracy improvement” is for data augmentation, and “accuracy change” is for data sharing.
-The figures with with standard deviations are in Section D.4.
+Complete results with standard deviations are in Section D.4.
 TabEBM generates high-fidelity synthetic data that can also be used for privacy preservation.
@@
 Figure 8 shows the logit distribution of TabPFN trained on surrogate binary tasks and the corresponding energy function of TabEBM (with TabPFN as the binary classifier) as the Euclidean distance from the real data increases.
-We found it essential to place the negative samples far from the real data, since TabPFN, which pre-trained to approximate Bayesian inference [33], has its confidence influenced by the distance from the training data [53].
+We found it essential to place the negative samples far from the real data, since TabPFN, which is pre-trained to approximate Bayesian inference [33], has its confidence influenced by the distance from the training data [53].
 Figure 8 (left) shows that TabPFN outputs high logit values near the real data.
 As the distance from the real data increases, the logit $f({\mathbf{x}})[1]$ decreases smoothly until the two logits become similar, making the classifier uncertain (because the class probabilities become equal).
-The right figure shows that TabEBM’s inferred density drops significantly as the maximum logit decreases, because $p_{c}({\mathbf{x}})\propto(\exp(f({\mathbf{x}})[0])+\exp(f({\mathbf{x}})[1]))$ from Equation 3.
+Figure 8 (right) shows that TabEBM’s inferred density drops significantly as the maximum logit decreases, because $p_{c}({\mathbf{x}})\propto(\exp(f({\mathbf{x}})[0])+\exp(f({\mathbf{x}})[1]))$ from Equation 3.
 Since SGLD sampling performs gradient ascent on the density, the TabEBM-generated samples will be close to the real data.
@@
 TabDDPM [42] Diffusion $p({\mathbf{x}}\mid y)$
-ARF [83] Random Forest $p({\mathbf{x}}\mid y)$
-GOGGLE [47] VAE $p({\mathbf{x}}\mid y)$
+ARF [83] Random Forest $p({\mathbf{x}},y)$
+GOGGLE [47] GNN $p({\mathbf{x}}\mid y)$
 TabPFGen [48] PFN $p({\mathbf{x}}\mid y)$
@@
 Integrating them into TabEBM will enhance its ability to manage high-dimensional datasets, increasing its versatility and utility.
-In addition, we would like to point out the imperfection of widely adopted univariate statistical fidelity metrics (i.e., Inverse KL, KS test and $\chi^{2}$ test).
-However, evaluating generators’ ability to capture the joining feature relationships remains an open research question [78].
-We leave this for future work to explore.
+Finally, note that, generators that are limited in modelling multivariate distributions may still perform well on univariate fidelity metrics, which is a standard approach to evaluating such models.
+However, evaluating their ability to learn more complex, high-order, relationships between features remains an open research question [78], which we leave for future work.
 5 Conclusion
-We introduce TabEBM, the first tabular data augmentation method that creates class-specific EBM generators, learning the marginal distribution for each class separately.
+We introduced TabEBM, the first tabular data augmentation method that creates class-specific EBM generators, learning the marginal distribution for each class separately.
 We also provide the first comprehensive analysis of tabular data augmentation across various dataset sizes.
@@
 The authors would like to thank Francisco Vargas, Randall Balestriero, and Otilia Stretcu for their insightful discussions and valuable input early in the project.
-NS acknowledges the support of the U.S.
+NS and MJ acknowledge the support of the U.S.
 Army Medical Research and Development Command of the Department of Defense; through the FY22 Breast Cancer Research Program of the Congressionally Directed Medical Research Programs, Clinical Research Extension Award GRANT13769713.
 Opinions, interpretations, conclusions, and recommendations are those of the authors and are not necessarily endorsed by the Department of Defense.
-Appendix: TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models
-[…]
+Appendix
 […]
@@
   3. Answer: [Yes]
-  4.
-Justification: Please refer to Section 1, where we state our research topic and objective and then connect them to our contributions.
-We also provide an open-source library of the proposed method (https://github.com/andreimargeloiu/TabEBM).
+  4. Justification: Section 1 details our research objectives and highlights our contributions.
   5. 2.
@@
   7. Answer: [Yes]
-  8. Justification: Please refer to Section 4, where we analyse the limitations of the proposed model and promising directions for future research.
+  8. Justification: Presented in Section 4.
   9. 3.
@@
   11. Answer: [Yes]
-  12. Justification: Please refer to Section 2, where we provide full theoretical derivations of our proposed method.
+  12. Justification: Section 2 presents the theoretical results of our proposed method.
   13. 4.
@@
   15. Answer: [Yes]
-  16.
-Justification: Please refer to Appendix B, where we provide full details on reproducing the results in the paper.
-Furthermore, we have also provided the open-source implementation of our proposed method.
+  16. Justification: Refer to Appendix B, where we provide full details on reproducing the results in the paper. We provide an open-source library of the proposed method.
   17. 5.
@@
   20.
-Justification: Please refer to Appendix B.
+Justification: Refer to Appendix B.
 All datasets used in this paper are publicly available, and the implementations of benchmark generators are open-source.
-In addition, we have also provided the implementation of our proposed method.
+We also provide an open-source library https://github.com/andreimargeloiu/TabEBM .
   21. 6.
@@
   23. Answer: [Yes]
-  24. Justification: Please refer to Appendix B, where we provide full descriptions of the experimental setup to reproduce the results in the paper.
+  24. Justification: Appendix B provides full descriptions of the experimental setup.
   25. 7.
@@
   28.
-Justification: Please refer to Section 3, where we provide standard deviations for all tables.
+Justification: Refer to Section 3, where we provide standard deviations for all tables.
 Figure 6 and Figure 4 (Right) contain error bars.
@@
   31. Answer: [Yes]
-  32. Justification: Please refer to Section B.4, where we provide full details on the computation resources used in the paper.
+  32. Justification: Refer to Section B.4, where we provide full details on the computation resources used in the paper.
   33. 9.
@@
   39. Answer: [Yes]
-  40. Justification: Please refer to Appendix A, where we include the societal impacts of our work.
+  40. Justification: Refer to Appendix A, where we include the societal impacts of our work.
   41. 11.
@@
   47. Answer: [Yes]
-  48. Justification: Please refer to Appendix B, where we provide the open-source licenses followed by the creators or original owners of assets.
+  48. Justification: Refer to Appendix B, where we provide the open-source licenses followed by the creators or original owners of assets.
   49. 13.